A Complete Bibliography of Publications in Computer Networks (Amsterdam, Netherlands: 2020–2029)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 April 2024 Version 1.33

Title word cross-reference

1 [1684]. 1 + 1 [239, 1104]. 2 [1559, 1609, 1989]. 3 [602, 934, 1145, 1273, 1328, 1669, 1971]. 5 [8, 15, 16, 18, 44, 88, 115, 126, 141, 145, 166, 167, 172, 179, 201, 204, 206, 207, 217, 234]. ^{2} [1627]. 3 [1309]. _{16} [2033]. _{2} [2052]. _{K} [340]. EON [489]. K [718, 1065, 1641, 2053]. κ [1916]. m [1904]. M_{K} [340]. H^{2} [476]. μ [1879]. P^{2} [597]. Q [1780, 1793].

* [771]. *-flow [771].

-anonymity [1065, 2053]. -based [418, 883, 1898]. -Coverage [1916]. -D [1273, 1328, 1684]. -dimensional [1609].

-learning [1793]. -learning-enabled [1780]. -reconstruction [1898]. -SHARP [597]. -V2X [1898].

/1/M [340]. /Hyper [340].

100 [665]. 107082 [361]. 107186 [296]. 107699 [547]. 109133 [1588]. 109529 [1765]. 109750 [1764]. 11 [27, 450, 930]. 14 [72, 502, 973]. 15 [1, 423, 913]. 169 [361]. 173 [296]. 185 [547]. 19 [168, 209, 382, 615, 1077, 1377, 1512].

2.0 [291, 2029, 2086]. 20 [202, 666, 1111]. 2019 [3]. 2020 [1, 27, 50, 72, 89, 107, 127, 148, 168, 186, 202, 222, 245, 271, 297, 325, 362, 398]. 2021 [423, 450, 481, 502, 524, 548, 571, 593, 615, 637, 666, 689, 711, 742, 778, 814, 849, 880].

2022 [913, 930, 953, 973, 1001, 1025, 1045, 1077, 1093, 1111, 1138, 1170, 1205, 1222, 1263, 1306, 1344]. 2023 [1381, 1417, 1442, 1466, 1483, 1521, 1551, 1566, 1589, 1615, 1636, 1651, 1675, 1695, 1727, 1767, 1809, 1844]. 2024 [1906, 1927, 1945, 1968, 2003, 2043, 2079, 2106]. 214 [1588]. 22 [107, 148, 548, 593, 1025]. 24 [297, 398, 778, 880, 1222, 1344]. 256 [753]. 26 [50, 481, 953]. 26th [1187]. 2IoT [1476].

3 [1707]. 360-degree [600]. 3GPP [58, 87, 2134]. 3MRP [1148]. 3P [330]. 3R [1874].

4 [222, 245, 689, 711, 1138, 1170, 2041]. 4.0 [1885]. 40/100 [1715]. 40802.11 [135]. 4G [494, 677, 787, 904, 963, 1574, 1863]. 4G/5G [494, 677, 904, 963].

5 [186, 637, 696, 1067, 1093]. 5.0 [1092, 2034]. 5G [185, 230, 265, 278, 283, 301, 323, 328, 332, 360, 361, 388, 390, 396, 408, 417, 479, 485, 494, 526, 527, 536, 566, 570, 575, 585, 619, 655, 670, 677, 693, 710, 744, 770, 777, 786, 796, 800, 803, 824, 837, 838, 889, 898, 906, 951, 963, 991, 1031, 1033, 1034, 1051, 1073, 1079, 1090, 1097, 1129, 1160, 1197, 1215, 1219, 1220, 1241, 1297, 1303, 1320, 1366, 1433, 1438, 1455, 1456, 1544, 1584, 1587, 1606, 1718, 1745, 1746, 1774, 1813, 1849, 1853, 1861, 1863, 1873, 1885, 1887, 1961, 1964, 1975, 1998, 2005, 2099, 2134]. 5G [248, 274, 318, 393, 394, 442, 478, 480, 497, 562, 671, 723, 851, 859, 899, 904, 950, 1256, 1298, 1325, 1337, 1397, 1445, 1448, 1450, 1678, 1817, 1900, 1943, 1958, 2073, 2101, 2102, 2140]. 5G-air-simulator [230]. 5G-AKA [786, 1325]. 5G-and-Beyond [332, 1584]. 5G-based [1337]. 5G-Enabled [185, 480, 710, 1160]. 5G-Flow [803]. 5G-MEC [1958]. 5G-NB-IoT [837]. 5G-RAN [1220]. 5G-satellite [1197]. 5G-VIOS [1975]. 5G/6G [1097, 1303, 1438, 1455]. 5G/B5G [388, 1861]. 5GMEC [1337]. 5GMEC-DP [1337]. 5GUK [248].

6 [332, 1438]. 60 [473, 741]. 64-antenna [332]. 6G [416, 583, 1097, 1145, 1176, 1183, 1303, 1305, 1401, 1438, 1455, 1534, 1588, 1591, 1601, 1610, 1724, 1902, 1925, 2022, 2032, 2093, 2135]. 6G-IoT [2093]. 6G-Tactile [1925]. 6Graph [940]. 6Search [1789]. 6TiSCH [308].

7 [89, 524, 1001, 1937].

8 [127, 571, 1045]. 802.11 [1086]. 802.11ac [329, 511, 759]. 802.11ad [1640]. 802.11ah [199, 1054, 2010]. 802.11ax [370]. 802.11p [380]. 802.15.4a [894]. 802.15.4e [1657]. 802.15.4e-based [1657]. 802.1X [2086].

9

[271, 325, 362, 742, 814, 849, 1205, 1263, 1306].

A-DECS [663]. AAE [1819].

AAE-DSVDD [1819]. AAPC [241]. ABAC [1153]. ABE [598]. abnormality [908]. above [665]. abstraction [436, 1603]. abundance [768]. accelerated [400, 1079, 1346, 2075]. Accelerating [1399, 1450, 1926, 2001]. acceleration [1182, 1519]. Accelerator [241, 1716, 1848]. Accelerator-aware [241]. acceptance [697]. Access [35, 58, 68, 147, 170, 171, 213, 218, 237, 261, 299, 301, 313, 335, 360, 374, 379, 395, 408, 413, 419, 462, 497, 528, 534, 560, 596, 626, 696, 724, 744, 756, 847, 920, 943, 955, 972, 976, 1004, 1008, 1029, 1034, 1079, 1129, 1133, 1149, 1167, 1197, 1256, 1271, 1319, 1350, 1359, 1372, 1444, 1481, 1500, 1506, 1543, 1584, 1591, 1592, 1597, 1624, 1653, 1661, 1670, 1738, 1742, 1799, 1856, 1951, 1954, 1974, 1987, 2010, 2022, 2073, 2076, 2134]. accessible [2033]. according [989]. accountability [1671]. accountable [598, 1635]. Accounting [622, 1608, 1931, 2059]. accuracy [74, 212, 539, 1355]. Accurate [265, 269, 683, 797, 890, 1270, 1595, 2082]. Achievable [1684, 1990]. Achieve [765, 1445]. Achieving [511, 746, 2053]. ACHO [309]. ACK [110]. ACL [445]. ACM [545]. ACoPE [6]. Acoustic [124, 153, 356, 716, 775, 1127, 1269, 1994, 2013, 2040, 2070]. Acoustic-based [356]. acquisition [220, 1475]. across [86, 500, 1282]. action [1265, 1790, 2033, 2049]. activations [1005]. active [872, 1530, 1807]. activeness [1560]. activities [346]. Activity [1377, 2082]. activity-level [1377]. actor [1833]. actors [1833]. Ad [9, 125, 146, 465, 534, 617, 668, 1098, 1148, 1492, 1575, 1580, 1599, 1620, 1685, 1805, 1827, 1984, 2091]. ad-hoc [9, 668, 1492, 1620, 1685, 1827, 2091]. ADA [1569]. ADA-PC [1569]. Adapt [1779]. adaptable [343, 938]. adaptation [96, 867, 1339, 1503, 1632, 1840, 2006, 2084, 2130]. adaption [1780]. Adaptive [6, 77, 174, 191, 194, 253, 275, 343, 383, 558, 566, 576, 663, 789, 869, 883, 919, 984, 1009, 1056, 1082, 1084, 1115, 1130, 1242, 1317, 1320, 1331, 1333, 1415, 1424, 1448, 1546, 1575, 1677, 1691, 1698, 1709, 1734, 1779, 1923, 1954, 1964, 2074]. Adaptively [827, 1686]. added [1154]. addition [247]. additional [1989]. additive [2132]. Address [801, 938, 940, 994, 1136, 1324, 1449, 1983]. addresses [657]. Addressing [938, 1437]. AdFeed [587]. adjacency [841]. adjustable [103]. adjusted [1915]. adjusting [1769]. adjustment [682, 1468, 1599, 2036]. Admission [545, 683]. Adoption [1873, 2067]. ads [460, 1131]. ADS-B [460]. ADT [1389]. advance [1938]. Advanced [563, 1425, 1514, 1586, 2118]. advancements [388, 1942]. Advances [947, 955, 1109, 1167, 1330, 1993]. Adversarial [359, 590, 686, 877, 1172, 1253, 1333, 2060]. advertising [587]. advisor [836]. Aerial [48, 111, 172, 321, 488, 518, 654, 739, 998, 1095, 1109, 1401, 1492, 1548, 2122]. Aerial/Ground [1109]. AERPAW [672]. AES [813]. affinity [1027]. Against [232, 317, 324, 372, 590, 611, 922, 988, 1061, 1422, 1425, 1435, 1568, 1665, 1754, 1765, 1877, 2016, 2053, 2132]. age [1553]. age-sensitive [1553]. Agent [231, 318, 707, 808, 859, 1062, 1281, 1408, 1432, 1464, 1668, 1874, 2012, 2021]. agents [290, 2074]. aggregate [460, 557]. aggregated [268]. Aggregation [193, 216, 238, 264, 480, 824, 827, 1446, 1875]. aggressiveness [1006]. agnostic [1961]. agree [865]. agreement [405, 1238, 1492]. agricultural [518, 1036]. agriculture [68, 132]. Ahead [390, 474]. AHP [219, 1780]. AI [283, 397, 487, 497, 498, 708, 951, 1090, 1097, 1262, 1297, 1389, 1395, 1407, 1477, 1611, 1900, 1961, 2071]. AI-based [283, 497, 708, 2071]. AI-Driven [498, 1297, 1389, 1961]. AI-enabled [397, 1611]. AI/ML [1090, 1407, 1900]. aided [4, 714, 831, 959, 1038, 1124, 1244, 1260, 1265, 1326, 1490, 1583, 1605, 1743, 1800, 2022]. aimed [112]. air [166, 230, 257, 664, 904, 923, 1072, 1321, 1605, 1971, 2066]. air-to-ground [1605]. airtime [13]. AKA [328, 786, 889, 1325, 1606]. Alarm [620, 1396]. Alexa [995]. Algorithm [5, 9, 30, 35, 45, 79, 103, 124, 140, 142, 150, 151, 205, 208, 232, 235, 250, 273, 308, 334, 383, 464, 480, 489, 558, 576, 590, 621, 642, 671, 687, 693, 722, 739, 761, 793, 817, 840, 1019, 1047, 1082, 1155, 1174, 1314, 1367, 1401, 1468, 1477, 1537, 1561, 1569, 1609, 1619, 1626, 1657, 1741, 1794, 1796, 1848, 1867, 1872, 1896, 1909, 1979, 1988, 2012, 2036, 2039, 2047, 2083, 2122, 2129]. Algorithms [152, 178, 492, 679, 808, 812, 896, 1036, 1178, 1186, 1539, 1703, 1823, 1929, 1950, 2111]. align [1452]. all-I [1350]. alleviation [1447]. allocate [1833]. Allocating [1572]. Allocation [13, 15, 35, 40, 109, 138, 150, 193, 217, 235, 242, 244, 263, 299, 339, 342, 367, 381, 413, 426, 472, 479, 487, 489, 527, 536, 553, 573, 577, 580, 589, 601, 619, 645, 650, 706, 723, 724, 732, 738, 800, 893, 939, 950, 975, 978, 1034, 1038, 1157, 1159, 1217, 1219, 1231, 1233, 1240, 1272, 1281, 1282, 1284, 1311, 1315, 1328, 1329, 1339, 1357, 1391, 1430, 1432, 1444, 1456, 1472, 1477, 1490, 1505, 1583, 1584, 1654, 1657, 1661, 1668–1670, 1691, 1692, 1718, 1729, 1733, 1775, 1829, 1861, 1908, 1912, 1933, 1934, 2024, 2026, 2084, 2116]. ALOHA [820, 1744, 1803]. ALOHA-NUI [1803]. Alone [331, 825]. alternate [1954]. alternative [220, 437, 879]. alternatives [622]. altitude [123, 1855]. Amazon [995]. Ambient [552, 1258, 1343, 1942, 2119]. amid [2050]. AMLFN [1575]. AMLFN-AD [1575]. among [64, 251, 1039]. amplifier [961]. AMQP [1525]. Analyses [1574, 2022]. Analysis [14, 46, 59, 98, 101, 105, 121, 195, 243, 261, 282, 291, 304, 307, 313, 319, 364, 369, 380, 428, 442, 447, 463, 467, 506, 519, 554, 668, 680, 725, 740, 759, 768, 780, 792, 812, 867, 926, 957, 965, 999, 1007, 1041, 1055, 1067, 1086, 1155, 1179, 1291, 1318, 1334, 1357, 1396, 1412, 1465, 1480, 1495, 1539, 1612, 1688, 1714, 1744, 1749, 1755, 1761, 1762, 1798, 1799, 1801, 1826, 1828, 1863, 1899, 1938, 1989, 1993, 1998, 2015, 2054, 2067, 2135, 2138]. analysis-based [1155]. Analytical [135, 219, 265, 315, 1013, 1679]. Analytics [68, 214, 809, 991, 1317, 1779, 2001, 2084]. analyze [2099]. Analyzing [897, 1012, 1285, 1473, 1771]. Anchor [1571, 1934]. Anchor-based [1571]. Android [39, 57, 447, 1286]. angle [1406]. anisotropic [289]. ANN [152]. announcements [531]. anomalies [531]. Anomaly [32, 75, 317, 448, 463, 582, 633, 861, 944, 999, 1076, 1185, 1248, 1379, 1556, 1766, 1932, 2019]. anonymity [328, 581, 1065, 1434, 1465, 2053]. anonymity-preserving [1434]. Anonymization [1843, 1879]. anonymous [405, 460, 829, 1469, 1504, 1671, 1894]. answering [1820]. Ant [457, 761, 1802, 1872]. antagonistic [855]. antenna [143, 332, 805, 1231, 1775, 1955]. antennas [1405, 1685]. Anti [142, 427, 505, 1181, 1486, 1739, 1898]. anti-collision [142]. anti-cybercrime [1739]. Anti-jamming [427, 505, 1181]. anti-LPB [1898]. Anti-quantum [1486]. anticipated [2140]. AoI [2119]. AP [1937]. APGNN [1396]. APIs [39]. App [288, 878, 911, 1248, 1270, 1596]. applicability [279]. applicable [813]. Application [69, 395, 414, 568, 645, 745, 879, 948, 1028, 1065, 1268, 1481, 1516]. Applications [132, 185, 287, 306, 355, 356, 373, 513, 557, 584, 610, 662, 717, 829, 891, 921, 1016, 1040, 1071, 1117, 1132, 1146, 1227, 1258, 1277, 1369, 1493, 1511, 1512, 1539, 1542, 1577, 1585, 1624, 1640, 1671, 1679, 1692, 1772, 1841, 1860, 1940, 1942, 1956, 2022, 2048, 2094, 2102]. applied [838]. Applying [1749]. Approach [6, 18, 93, 105, 134, 137, 139, 188, 228, 263, 269, 292, 305, 311, 329, 345, 368, 374, 381, 407, 420, 435, 437, 454, 462, 491, 507, 550, 555, 579, 598, 607, 617, 635, 693, 700, 729, 757, 806, 808, 835, 844, 852, 888, 899, 940, 985, 1018, 1029, 1043, 1116, 1125, 1140, 1203, 1217, 1252, 1289, 1335, 1339, 1343, 1350, 1355, 1380, 1384, 1408, 1428, 1432, 1471, 1493, 1499, 1529, 1554, 1559, 1560, 1600, 1643, 1659, 1677, 1733, 1735, 1751, 1783, 1787, 1789, 1793, 1821, 1833, 1879, 1882, 1889, 1970, 1984, 2013, 2018, 2070, 2074, 2075, 2094, 2099, 2121, 2127]. approach [312, 723, 1108, 1508, 1527, 1966]. Approaches [104, 387, 494, 774, 823, 838, 890, 1067, 1273, 1370, 1726, 1939, 2015, 2018, 2034, 2045]. Approver [821]. Approximate [729, 934, 1724, 1960]. approximation [307, 923, 1887]. apps [1377, 2054]. April [89, 107, 524, 548, 1001, 1025, 1483, 1521, 2003, 2043]. AQM [192, 2098]. AQRS [1486]. Aquila [1691]. AR-GAIL [1333]. Arbiter [688]. Arbiter-based [688]. Architecting [851]. architectural [504]. Architecture [25, 88, 120, 167, 195, 318, 348, 361, 499, 522, 533, 627, 630, 719, 792, 803, 899, 1066, 1096, 1128, 1163, 1302, 1303, 1362, 1419, 1672, 1714, 1741, 1841, 1873, 1877, 1919, 2049]. architecture-based [2049]. architecture-level [348]. Architectures [44, 55, 281, 417, 610, 645, 897, 1016, 1075, 1157, 1367]. Area [150, 227, 240, 289, 618, 1102, 1279, 1504, 1610, 1646, 1674, 1918, 1979]. area-scalable [227]. areas [322, 1880]. Arena [332]. ARES [1331]. ARIMA [646, 2045]. ARIMA-BP [646]. ARM [813]. ARM-based [813]. armed [299, 1641, 2013]. arrangement [411]. arrival [212]. Art [390, 416, 474, 543, 632, 653, 1021, 1510, 1550, 2032]. artifacts [1012]. Artificial [389, 416, 422, 448, 568, 705, 1021, 1249, 1378, 1511, 1587, 2032, 2045]. AS-level [1465]. ASIRIA [1099]. Aspects [1585]. ASRPH [745]. assess [1332]. Assessing [160, 282, 599, 1804]. Assessment [155, 220, 296, 564, 677, 1040, 1145, 1542, 1593, 1600, 1604, 1842, 1852, 2097, 2100, 2118, 2127]. asset [805, 875]. assets [927, 1334, 1995]. Assigning [720]. Assignment [129, 133, 178, 263, 484, 713, 749, 1024, 1057, 1123, 1309, 1479, 1488, 1689, 1722, 1737, 1991]. assist [2124]. assistance [631]. Assisted [52, 87, 196, 368, 413, 415, 479, 556, 613, 738, 822, 859, 975, 1035, 1036, 1069, 1100, 1210, 1249, 1348, 1366, 1408, 1410, 1492, 1523, 1534, 1629, 1655, 1669, 1674, 1753, 1787, 1795, 1798, 1814, 1862, 1870, 1896, 1955, 2021, 2084, 2094, 2119, 2126, 2128]. Assisting [236, 1961]. association [199, 244, 374, 619, 744, 757, 1054, 1124, 1276, 1668, 2124]. associations [559, 1723]. assumptions [160]. assurance [1758]. asymmetric [645]. asynchronous [819, 840, 1149, 1318, 1569]. AT-GCN [1836]. Atlas [165]. ATSSS [2005]. ATT&CK [2118]. ATT&CK-based [2118]. Attack [76, 152, 190, 405, 501, 512, 535, 538, 590, 785, 799, 802, 813, 957, 1095, 1163, 1188, 1207, 1250, 1368, 1463, 1515, 1575, 1578, 1593, 1632, 1667, 1793, 1822, 1836, 2103, 2120]. attack-resilience [405]. attacker [1349]. Attackers [1314]. Attacks [24, 46, 110, 158, 160, 164, 304, 305, 317, 492, 505, 555, 603, 688, 772, 811, 825, 842, 854, 904, 997, 1042, 1054, 1061, 1155, 1252, 1435, 1455, 1527, 1693, 1704, 1754, 1755, 1765, 1838, 1957, 1983, 1992, 2118, 2132]. Attention [1185, 1549, 1568, 1775, 1788, 1797, 2023, 2078, 2082]. Attention-based [2078]. attentive [727]. Attribute [698, 882, 1153, 1308, 1686, 1799, 1832, 1894]. Attribute-based [698, 882, 1686, 1832, 1894]. attributes [932, 1452]. Auction [78, 237, 301, 402, 551, 1282, 1462, 1548]. auction-based [1462]. auditing [704, 1245, 1592, 1830, 2108]. augmentation [1604]. augmented [700, 842, 867]. Augmenting [918]. August [222, 689, 1138, 1675]. AURA [619]. AURA-5G [619]. Aurora [61]. Authentication [104, 119, 240, 247, 268, 327, 328, 401, 405, 460, 477, 517, 530, 597, 644, 726, 753, 783, 828, 829, 834, 882, 889, 1015, 1017, 1030, 1113, 1173, 1198, 1238, 1283, 1296, 1304, 1327, 1354, 1404, 1434, 1476, 1492, 1502, 1504, 1534, 1544, 1563, 1606, 1639, 1671, 1834, 1917, 1922, 1925, 1931, 1987, 1995, 2020, 2070, 2086, 2093]. authority [448, 698, 1277, 1671, 1686, 1807]. Authorization [1597, 1931]. Auto [329, 330, 344, 812, 1188, 1537, 1556, 2025]. Auto-3P [330]. auto-encoder [1188, 1556]. auto-provisioning [2025]. auto-scaling [344]. Auto-tune [812, 1537]. autoencoder [740, 1043]. autoencoder-based [740]. Autoencoders [818]. Automated [1341, 1664, 1723, 2063]. Automatic [270, 910, 915, 1166, 2104]. automation [2068]. automotive [644]. Autonomous [324, 330, 478, 498, 519, 631, 851, 1109, 1362, 1411, 1455, 1537]. AUTOSAR [1313]. AUTOSAR-based [1313]. Autoscaling [1288]. availability [468, 1047]. available [61]. average [841, 1491]. avoidance [141, 486, 775, 1115, 1570]. avoided [2015]. Aware [10, 20, 64, 65, 94, 109, 133, 136, 189, 197, 221, 241, 256, 258, 278, 281, 314, 339, 341, 360, 379, 407, 422, 427, 430, 506, 550, 565, 587, 618, 642, 643, 674, 678, 701, 707, 721, 730, 781, 789, 802, 817, 831, 853, 867, 933, 976, 977, 1029, 1056, 1057, 1070, 1089, 1125, 1137, 1148, 1162, 1166, 1180, 1201, 1230, 1240, 1247, 1255, 1310, 1315, 1331, 1415, 1460, 1493, 1499, 1503, 1524, 1530, 1535, 1607, 1610, 1655, 1669, 1697, 1708, 1784, 1797, 1826, 1829, 1869, 1882, 1897, 1933, 1947, 1948, 1950, 1956, 1970, 1974, 1987, 2015, 2026, 2056, 2061, 2124]. aware [47, 87, 125, 129, 173, 178, 201, 220, 392, 406, 455, 671, 851, 979, 1027, 1152, 1305, 1340, 1436, 1732, 1748, 1815, 1920, 1999, 2047]. Awareness [909, 915, 1050, 1103, 1827, 1883]. AWS [428]. Azure [428].

B [460]. B2H [1092]. B5G [388, 535, 966, 998, 1260, 1276, 1724, 1861]. B5G-driven [966]. BaaS [368]. back [1468]. backbone [1210]. backbone-assisted [1210]. backend [2033]. backhaul [413, 782, 961, 990, 1031, 1034, 1433, 1444, 1631]. backhauling [562, 1298]. backoff [528, 2010]. backscatter [196, 552, 1175, 1258, 1490, 1683, 2031, 2119]. backscatter-aided [1490]. backscatter-assisted [196, 2119]. backstage [1568]. backstepping [192]. backtracking [489]. backup [484, 532, 1933]. backward [1899]. bag [1956]. bag-of-tasks [1956]. Balanar [1785]. Balanced [559, 939, 1103, 1700]. Balancing [77, 105, 231, 249, 285, 360, 425, 545, 599, 791, 981, 1006, 1039, 1080, 1184, 1209, 1316, 1526, 1579, 1619, 1785, 1882, 1962]. Band [65, 311, 395, 458, 495, 696, 1107, 1229, 1232, 1319, 1341, 1470, 1828, 1831, 1973, 2038, 2098]. bandit [1641, 2013]. bandits [299]. bands [1988, 2051]. Bandwidth [6, 96, 235, 308, 386, 459, 623, 764, 792, 869, 963, 1006, 1262, 1317, 1470, 1524, 1657, 1692, 1698, 1953, 2065, 2116]. Bandwidth-aware [1524]. bandwidth-constrained [1953]. Bandwidth-efficient [1262, 1317]. bandwidth-optimal [623]. bargaining [242]. barrier [1609]. Bartolomeu [86]. Base [111, 123, 234, 626, 820, 1688, 1774, 1836, 1951, 2075]. baseband [2140]. baseband-function [2140]. Based [16, 58, 117, 218, 228, 239, 266, 281, 283, 289, 305, 328, 330, 332, 340, 345, 356, 357, 389, 457, 459, 462, 529, 576, 580, 581, 596, 609, 682, 697, 716, 717, 721, 726, 731, 738, 763, 772, 773, 810, 828, 942, 949, 970, 991, 996, 1017, 1040, 1051, 1068, 1084, 1123, 1156, 1157, 1166, 1173, 1181, 1190, 1207, 1277, 1304, 1313, 1335, 1341, 1363, 1384, 1441, 1459, 1468, 1477, 1482, 1502, 1505, 1539, 1561, 1593, 1620, 1667, 1672, 1686, 1702, 1721, 1730, 1777, 1806, 1841, 1861, 1865, 1898, 1911, 1932, 1964, 1976, 2016, 2062, 2070, 2089, 2099]. based [5, 26, 61, 67, 174, 177, 181, 208, 219, 227, 242, 244, 259, 306, 313, 317, 373, 419, 420, 445, 458, 465, 537, 542, 555, 590, 595, 598, 614, 636, 688, 708, 713, 715, 730, 751, 753, 775, 813, 916, 987, 1067, 1073, 1086, 1098, 1102, 1105, 1128, 1133, 1162, 1188, 1192, 1196, 1199, 1219, 1255, 1287, 1311, 1320, 1338, 1349, 1359, 1398, 1454, 1461, 1471, 1495, 1499, 1507, 1520, 1532, 1535, 1548, 1571, 1583, 1601, 1628, 1629, 1640, 1645, 1694, 1793, 1797, 1816, 1832, 1842, 1874, 1887, 1910, 1915, 1922, 1930, 1959, 1981, 1998, 2010, 2054, 2059, 2061, 2105, 2116, 2122, 2133]. based [23, 48, 92, 101, 192, 254, 268, 288, 295, 303, 322, 368, 403, 412, 449, 452, 476, 490, 523, 545, 551, 570, 591, 607, 639, 650, 702, 776, 785, 786, 834, 841, 863, 896, 898, 929, 934, 975, 980, 1000, 1009, 1020, 1064, 1075, 1082, 1120, 1129, 1175, 1182, 1209, 1240, 1246, 1250, 1283, 1284, 1286, 1294, 1299, 1314, 1352, 1367, 1392, 1446, 1460, 1529, 1576, 1579, 1581, 1627, 1639, 1657, 1660, 1669, 1671, 1677, 1710, 1713, 1716, 1717, 1722, 1770, 1775, 1784, 1789, 1805, 1820, 1825, 1830, 1835, 1836, 1850, 1858, 1867, 1871, 1894, 1921, 1925, 1939, 1949, 1979, 1991, 2024, 2029, 2047– 2049, 2066, 2071, 2120]. based [12, 43, 150, 154, 157, 180, 284, 304, 321, 358, 385, 387, 411, 418, 453, 464, 470, 487, 497, 512, 517, 633, 640, 643, 668, 698, 704, 719, 725, 739, 740, 771, 790, 800, 802, 817, 822, 829, 861, 862, 870, 876, 879, 911, 924, 932, 998, 1007, 1029, 1050, 1071, 1072, 1101, 1150, 1155, 1158, 1179, 1217, 1218, 1228, 1236, 1238, 1242, 1257, 1261, 1269, 1271, 1293, 1300, 1322, 1327, 1329, 1375, 1379, 1388, 1394, 1409, 1419, 1424, 1470, 1494, 1500, 1531, 1643, 1662, 1703, 1723, 1751, 1752, 1757, 1837, 1847, 1854, 1873, 1880, 1891, 1929, 1950, 1963, 1982, 2091, 2103]. based [11, 30, 45, 63, 75, 119, 120, 143, 146, 159, 176, 226, 235, 253, 260, 261, 318, 334, 352, 432, 435, 446, 461, 480, 520, 538, 558, 597, 600, 631, 646, 652, 657, 661, 675, 685, 691, 723, 724, 754, 794, 856, 882, 909, 950, 952, 978, 981, 999, 1006, 1015, 1027, 1063, 1081, 1088, 1143, 1184, 1213, 1237, 1266, 1279, 1281, 1324, 1325, 1337, 1361, 1401, 1450, 1458, 1462, 1527, 1580, 1592, 1658, 1681, 1687, 1732, 1738, 1748, 1788, 1794, 1796, 1799, 1808, 1815, 1823, 1827, 1860, 1864, 1866, 1872, 1884, 1912, 1914, 1917, 1952, 1954, 1999, 2006, 2012, 2020, 2023, 2028, 2037, 2087, 2100, 2102, 2118, 2126, 2129]. based [66, 287, 312, 755, 788, 883, 1153, 1204, 1812, 1831, 1862, 1957, 2078, 2083]. based-Context [643]. Basic [1313]. basics [1520]. BASM [821]. batch [1991]. batch-based [1991]. batched [2085]. Battery [302, 1459, 1999]. battery-constricted [1999]. Battery-less [302]. Battlefield [1857]. Bayesian [45, 386, 1226]. BBR [1365, 1445]. BBRv2 [1006]. BBU [234, 360, 1209]. BC [1354]. BC-enabled [1354]. BCSA [1561]. be [760, 1978]. Beacon [1068]. Beam [804, 1069, 1186, 1294, 1405, 1955]. beam-width [1955]. beamformer [1659]. beamforming [515, 741, 756, 757, 982, 1611, 1664]. beat [140]. bee [321, 422]. BEE-DRONES [321]. behavior [1177, 1473, 1835, 1837, 1899]. behavioral [1334]. behaviors [833, 1579]. bell [704]. benchmarking [543]. Benders [2075]. benefits [122, 195, 1107]. BER [133]. Best [140, 821, 1699, 1839]. best-known [140]. better [94, 1362]. between [276, 753, 925, 1256]. Beyond [179, 332, 394, 583, 693, 744, 777, 800, 803, 811, 898, 906, 950, 951, 1033, 1073, 1090, 1176, 1297, 1320, 1414, 1584, 1587, 1588, 1611, 1718, 1873, 1900, 1902, 1937, 1943, 1947, 1961, 2073, 2134]. BFLS [1496]. BFT [971, 1356]. BGP [531, 922, 1003, 1185, 1300, 1402, 1965, 2019]. BGPattacks [1645]. BGSD [1398]. bi [350]. bi-level [350]. biased [899]. bidirectional [1722]. big [68, 71, 595, 622, 936]. BigFooT [971]. Binary [997, 1561]. bins [322]. bipartite [582]. bistatic [1683]. Bit [36]. BiTCN [2048]. Bitcoin [130, 985, 1014, 1364]. bitmap [542]. bitmap-based [542]. bitrate [1130]. bivariate [1236]. Black [324, 2140]. Black-box [324, 2140]. BLE [1164, 1461]. Blending [1536]. block [134, 901, 1014]. Blockage [782]. blockages [1664]. Blockchain [177, 266, 287, 421, 461, 595, 598, 608, 610, 710, 715, 719, 753, 774, 810, 821, 841, 856, 859, 862, 863, 865, 875, 876, 897, 924, 927, 949, 950, 952, 966, 970, 971, 1000, 1015, 1044, 1092, 1102, 1149, 1150, 1174, 1175, 1196, 1213, 1239, 1277, 1300, 1322, 1349, 1361, 1394, 1419, 1451, 1459, 1479, 1486, 1492, 1496, 1502, 1511, 1585, 1608, 1635, 1645, 1687, 1706, 1710, 1749, 1750, 1778, 1795, 1799, 1830, 1850, 1857, 1858, 1893, 1910, 1939, 1963, 1985, 2020, 2028, 2059, 2063, 2087, 2093, 2095, 2108, 2117, 2125, 2128]. Blockchain-assisted [1795, 2128]. Blockchain-based [177, 595, 598, 715, 810, 856, 862, 863, 924, 949, 952, 970, 1000, 1015, 1102, 1150, 1175, 1300, 1322, 1349, 1459, 1502, 1710, 1830, 1850, 1858, 1963, 2020, 2059]. blockchain-empowered [1857]. Blockchain-Enabled [421, 821, 865, 950, 966, 2093].

Blockchain-enhanced [1479]. Blockchain-native [927]. blockchain-oriented [710]. blockchain-powered [774]. Blockchained [967, 2136]. blockchains [603, 912, 1105]. blocking [1180]. blocks [848]. Bloom [736, 1662]. Blueprint [575]. Bluetooth [343, 692, 992, 1067, 1106]. BMS [1524]. Board [2, 28, 51, 73, 90, 108, 128, 149, 169, 187, 203, 223, 246, 272, 298, 326, 363, 399, 424, 451, 482, 503, 525, 549, 572, 594, 616, 638, 667, 690, 712, 743, 779, 815, 850, 881, 914, 931, 954, 974, 1002, 1026, 1046, 1060, 1078, 1094, 1112, 1139, 1171, 1206, 1223, 1264, 1307, 1345, 1382, 1418, 1443, 1467, 1484, 1522, 1552, 1567, 1590, 1616, 1637, 1652, 1676, 1696, 1728, 1768, 1810, 1845, 1907, 1928, 1946, 1969, 2004, 2044, 2080, 2107]. boards [2041]. Body [121, 240, 1102, 1504, 1577]. bonding [720]. boosted [1703]. Boosting [1568]. bot [1598]. both [142, 605]. botnet [863, 1299, 1439, 1788]. botnets [372]. bottleneck [662, 1837, 1859]. bound [1210, 1656]. Boundary [1916]. bounded [1684]. box [324, 2140]. BP [143, 646, 1236]. BP-MGKM [1236]. brain [116]. branch [732, 2049]. Breadth [679]. Breadth-first [679]. Broadband [225, 286, 440, 1049]. Broadcast [368, 377, 1101, 1686]. Broadcasting [255, 1163]. broker [1514, 1981]. Brownian [1368]. browser [432]. browser-based [432]. browsing [221, 809, 1414]. BRP [235]. BS [853]. BSS [299]. BTSM [821]. budget [1478, 1540]. budgeted [1970]. Buffer [4, 567, 831]. Buffer-aided [4, 831]. buffering [43]. Building [67, 176, 912, 1292, 1575]. buildings [822]. Bulwark [2030]. Burn [654, 875]. Burn-to-Claim [875]. bursty [315, 439, 531]. Business [1235]. Butterworth [1959]. BwShare [96]. Bytes [675]. Bytes-based [675]. ByteSGAN [877]. Byzantine [840, 1451, 1877].

C [219, 315, 318, 1107, 1209, 1440]. C-RAN [315, 318, 1209]. C-V2X [1440]. CA [528, 717]. CAB [546]. Cable [1450]. cables [2009]. Cache [26, 122, 307, 510, 642, 813, 832, 879, 1062, 1741, 1793, 1948]. Cache-based [879]. cache-enabled [1948]. CacheNet [956]. caches [250, 348, 895, 1153]. Caching [138, 173, 189, 295, 366, 433, 546, 560, 626, 714, 769, 833, 864, 1130, 1143, 1158, 1160, 1254, 1380, 1554, 1612, 1644, 1711, 1726, 1782, 1846, 1871, 1892, 1948, 1970, 1976, 2012, 2083, 2085, 2129]. CAI [643]. CAIN [1610]. Calculation [1266]. calendaring [823]. Calibration [741]. Call [409]. camouflaged [751]. Campus [98, 209, 1855, 2069, 2089]. can [140, 1511]. cancellation [1954]. candidate [1780]. capabilities [484, 2073]. capability [552]. capacitated [342, 763]. Capacity [265, 410, 982, 1431, 1683, 1813, 1847, 1984, 2114]. CapRadar [1698]. capsule [754]. CAPTAIN [124]. carbon [2059]. cardinality [1868]. care [240, 517, 745, 1509]. Carlo [1924]. carrying [2081]. CART [36]. cascading [232]. case [49, 113, 141, 428, 578, 787, 851, 1105, 1378, 1387, 1716, 2042, 2113]. cases [230]. cash [255]. Catchers [677]. categorization [1921]. category [2042]. causal [1128]. cause [1396]. CBF [1721]. CBNet [1152]. CC [1198]. CCI [1761]. CDN [976, 1294]. ceiling [332]. Cell [19, 138, 207, 267, 407, 408, 900, 1393, 1401, 1473, 1733, 1846, 2089, 2124]. CellOS [300]. cells [562, 1018]. Cellular [10, 32, 58, 105, 139, 189, 263, 300, 352, 353, 388, 413, 426, 442, 467, 494, 508, 580, 764, 838, 853, 1051, 1154, 1191, 1274, 1278, 1281, 1310, 1411, 1498, 1506, 1518, 1669, 1713, 1905, 1934, 1955, 2018]. cellular-connected [388]. cellular/WiFi [764]. Center [118, 191, 280, 316, 473, 483, 641, 771, 1081, 1218, 1232, 1431, 1720, 1790, 1950, 2036, 2104]. centers [314, 1503]. central [1249]. Centrality [795, 2016]. Centralized [823, 1541, 1714]. Centric [19, 105, 114, 295, 467, 533, 592, 611, 937, 1143, 1782, 1985]. certificate [596]. Certificateless [188, 460, 824, 1101, 1245, 1795]. Certifying [1459]. CETAnalytics [214]. CFmMIMOs [1629]. Chain [8, 91, 194, 256, 313, 351, 504, 693, 893, 917, 939, 1108, 1142, 1255, 1282, 1330, 1373, 1428, 1458, 1462, 1478, 1985, 2007, 2036]. chained [331, 1656]. Chaining [241, 292, 375, 678, 835, 1056, 1115, 1120, 1125, 1159, 1215, 1228, 1437, 1849, 2002]. Chains [785, 865, 1198, 1272, 1426, 1485, 1501, 1701, 1786, 2061]. Challenges [44, 69, 81, 197, 274, 369, 388, 416, 493, 653, 709, 809, 810, 941, 955, 1022, 1126, 1134, 1167, 1221, 1257, 1258, 1303, 1330, 1369, 1370, 1437, 1463, 1510, 1512, 1520, 1623, 1725, 1726, 1758, 1799, 1839, 1840, 1902, 1942, 2032]. changing [1641]. Channel [65, 103, 109, 121, 123, 134, 299, 427, 463, 552, 580, 696, 720, 740, 751, 811, 812, 893, 894, 997, 1042, 1048, 1073, 1181, 1224, 1234, 1329, 1537, 1657, 1684, 1687, 1733, 1748, 1831, 1865, 1934, 1951, 2010, 2040]. channel-aware [1748]. channeled [1420]. channels [279, 401, 612, 644, 720, 869, 1352, 1761, 1831, 2040, 2050]. chaos [205]. characteristics [1855]. Characterization [428, 879, 1656]. Characterizing [347]. Charge [2055]. charger [100]. Charging [888, 1103, 1211]. chemical [730]. China [85]. chip [934, 1989]. choices [113]. Choose [301, 760]. CI GRU [1788]. Cipher [292]. CIPO [2123]. circles [558]. circuit [771, 1982]. circular [1369]. circulation [927]. Citadel [613]. Cities [211]. city [496, 719, 826, 1140, 1191, 1387]. city-scale [1191]. Claim [875]. clarification [651]. class [70, 625, 1179, 1819, 2062, 2121]. classes [918, 1116]. Classification [36, 120, 156, 159, 214, 264, 345, 346, 375, 447, 537, 591, 716, 727, 839, 846, 877, 919, 962, 1116, 1142, 1273, 1286, 1374, 1377, 1413, 1439, 1489, 1507, 1547, 1549, 1602, 1781, 1788, 1811, 1819, 1890, 1903, 1921, 1923, 1953, 2048, 2060, 2115, 2131]. classifier [176, 969, 1142, 1207, 1299]. classify [163]. Classifying [1334]. clean [1380]. clean-slate [1380]. Client [12, 327, 1073, 1124, 1888, 2065]. client-based [12]. client-server [327]. clients [581, 1751]. clipping [961]. clock [7, 747]. clogging [477]. clone [957]. cloning [555]. closed [1455]. CLoSER [1846]. Cloud [17, 61, 96, 118, 144, 175, 198, 207, 210, 233, 234, 237, 266, 281, 283, 290, 320, 335, 344, 360, 368, 374, 428, 448, 449, 490, 512, 517, 557, 704– 706, 730, 783, 798, 816, 870, 885, 887, 891, 892, 952, 984, 1007, 1120, 1128, 1157, 1169, 1207, 1245, 1304, 1308, 1335, 1367, 1430, 1475, 1492, 1503, 1519, 1528, 1531, 1592, 1594, 1666, 1670, 1729, 1732, 1752, 1757, 1777, 1830, 1919, 1926, 1977, 2020, 2025, 2075, 2137]. cloud-assisted [1492]. Cloud-Based [517, 1335]. cloud-edge [1128, 1475, 1594, 1732, 1757, 2137]. cloud-edge-end [1752]. Cloud-Enabled [449]. cloud-Fog [281, 2020]. cloud-native [1977]. Cloud-RAN [1729]. Cloud-RANs [144, 2075]. cloud-to-things [783]. cloud-to-user [428]. Cloudlets [1039]. CloudPilot [1519]. clouds [337, 936, 1214, 1536, 1708, 2138]. Cluster [11, 105, 465, 490, 523, 739, 959, 1409, 1632]. cluster-aided [959]. Cluster-based [11, 523]. Clustered [367, 553, 717, 1329, 1385, 1706, 1955, 2088]. Clustering [9, 30, 219, 259, 284, 319, 374, 516, 868, 999, 1027, 1037, 1137, 1328, 1336, 1473, 1805]. clustering-based [999]. CN [126]. CNCAA [142]. CNN [67, 70, 907, 1182, 1959, 2048]. CNN-based [67]. CNN-LSTM [1959]. CNN-SSDI [907]. CNNParted [1634]. CNNs [1005]. CO [2052]. coalitional [1261]. CoAP [767]. Cobalt [1920]. code [1244, 1283, 1591]. code-based [1283]. Coded [36, 714, 900, 959, 1014, 1385, 2088]. Codes [762, 2040]. Coding [99, 174, 193, 514, 713, 1424]. Coercion [177]. Coercion-free [177]. coexistence [696, 1438, 1638, 1988]. coflows [429, 1201]. Cognition [1625, 1712]. Cognitive [11, 40, 80, 103, 262, 311, 323, 396, 427, 438, 552, 565, 701, 728, 943, 960, 1051, 1175, 1181, 1329, 1500, 1761]. Collaboration [1015, 1128, 1395, 1582, 1594, 1625, 1741, 1752, 1757]. collaboration-driven [1395]. Collaborative [146, 193, 660, 663, 707, 715, 808, 885, 1098, 1158, 1182, 1340, 1527, 1658, 1664, 1870, 1911, 1944, 1963, 2014, 2020, 2035, 2047, 2056, 2083]. Collection [38, 124, 432, 710, 826, 1038, 1407, 1409, 1410, 1416, 1454]. collective [999]. collector [1573]. collided [142]. Collision [87, 141, 142, 486, 817, 1019, 1383, 1803, 1828]. collision-aware [817]. Collision-free [1019, 1803]. colluding [1480]. collusion [1879]. Colony [422, 457, 1802, 1872]. Colony-inspired [1802]. column [723]. CoMap [1261]. combination [152]. combinatorial [1641, 2013]. Combine [2131]. combined [570, 1267, 2056]. Combining [1193, 1290, 1324, 1625]. commerce [419]. commercial [602, 685, 796]. commercial-off-the-shelf [602]. commodity [249, 2072]. Common [345, 865, 1731]. Communication [47, 57, 109, 143, 145, 151, 263, 268, 349, 358, 394, 395, 422, 426, 539, 581, 584, 620, 629, 634, 650, 665, 685, 751, 753, 766, 793, 819, 894, 934, 980, 1070, 1072, 1083, 1097, 1141, 1244, 1256, 1377, 1378, 1383, 1506, 1600, 1622, 1630, 1631, 1638, 1680, 1684, 1690, 1717, 1745, 1764, 1776, 1795, 1808, 1865, 1885, 1888, 1898, 1909, 1986, 1989, 2031, 2034, 2139]. communication-channel-based [751]. Communication-efficient [819, 1888]. Communications [18, 87, 116, 153, 163, 179, 206, 286, 417, 464, 583, 728, 831, 851, 874, 946, 1100, 1187, 1257, 1258, 1260, 1278, 1297, 1524, 1577, 1605, 1743, 1800, 1842, 1869, 1883, 2006, 2018, 2021, 2029, 2042, 2055, 2110]. communities [79]. community [62, 150, 1294, 1740, 2120]. community-based [1294]. CoMP [66, 1870]. CoMP-Based [66]. Comparative [220, 447, 540, 1041, 1172, 1799, 1801, 1826, 1989, 2045]. Comparison [69, 284, 410, 628, 2040]. comparisons [1385]. Compatible [889, 1150, 1618]. compensated [7]. Compensation [66, 1384]. competition [1217]. Competitive [304, 567, 1160, 1557]. compilation [132]. compiler [2033]. completion [347, 1087]. complex [6, 23, 79, 582, 927, 1593, 1600, 1778, 1820, 1972]. complex-policy [6]. Complexity [1210, 1645, 1876]. compliant [937]. COMPNW [1588, 1765]. Component [1354]. Component-wise [1354]. components [230, 836]. Composite [535]. composition [14, 20, 1091]. Comprehensive [161, 214, 291, 632, 686, 752, 771, 837, 838, 1023, 1055, 1072, 1134, 1144, 1168, 1350, 1370, 1520, 1725, 1772, 1773, 1798, 1863, 1902, 1939, 1943, 1961, 2067, 2096, 2138]. compressed [54]. compression [1420]. compressive [63]. Computation [228, 260, 385, 387, 560, 724, 769, 844, 950, 958, 978, 1190, 1233, 1386, 1471, 1548, 1658, 1911, 1935, 1958, 2137]. computational [807, 972]. computationally [64]. compute [557]. compute-aggregate [557]. Computer [296, 361, 547, 993, 1593, 1650, 1764, 1903]. computers [1187]. Computing [20, 147, 206, 211, 347, 353, 387, 397, 400, 402, 448, 477, 512, 551, 553, 557, 577, 592, 609, 639, 670, 681, 687, 700, 703, 705, 706, 708, 781, 784, 807, 816, 819, 823, 828, 830, 958, 975, 996, 1013, 1016, 1029, 1061, 1070, 1072, 1075, 1080, 1083, 1087, 1096, 1132, 1178, 1182, 1190, 1207, 1221, 1230, 1240, 1284, 1288, 1304, 1328, 1358, 1366, 1371, 1376, 1388, 1403, 1419, 1430, 1454, 1471, 1477, 1481, 1485, 1512, 1558, 1561, 1562, 1582, 1584, 1628, 1682, 1686, 1709, 1724, 1729, 1730, 1747, 1753, 1813, 1829, 1908, 1911, 1940, 1942, 2026, 2034, 2035, 2066, 2110, 2122, 2137]. computing [260, 366, 415, 439, 560, 663, 698, 724, 783, 794, 800, 864, 885, 966, 972, 978, 1035, 1108, 1118, 1213, 1262, 1326, 1337, 1340, 1386, 1399, 1425, 1472, 1474, 1527, 1655, 1692, 1732, 1796, 1862, 1987, 2037, 2064, 2073]. computing-based [639]. CoMSeC [834]. concatenated [1244]. concave [293]. concept [375, 998]. Concepts [1900]. conceptual [2135]. Concurrent [15, 78, 299, 685, 1591]. Condition [65]. conditions [894]. conference [1649]. Conferencing [1734]. confidence [969]. configurable [1772]. Configuration [329, 867, 1603, 1935, 2046]. configurations [236]. confirmation [130]. conflict [1363]. confusable [1155]. Congestion [192, 425, 494, 628, 872, 1028, 1098, 1115, 1192, 1415, 1660, 1697, 1713, 1714, 1824, 1859, 1867, 1929, 1962, 2111, 2124]. Congestion-aware [1697, 2124]. congestion-induced [1962]. Conjunctive [1449]. Connected [257, 306, 388, 1109, 1198, 1225, 1321]. Connecting [1940]. Connection [343, 1662]. Connectivity [444, 680, 729, 1036, 1071, 1085, 1145, 1360, 1364, 1383, 1569, 1685]. ConPoolUBF [1662]. conquer [811]. consensus [287, 521, 840, 841, 865, 912, 971, 1044, 1174, 1277, 1451, 1749, 1952, 2100]. consequences [190]. Conservative [1285]. considering [151, 484, 1619, 1886, 1933]. Consistency [159, 1128, 1426, 1607, 2030]. Consistent [647]. consolidation [175, 234, 1760, 1889]. Consortium [1451, 2028]. Constant [923]. constellation [1015, 1692, 1909]. constellations [1533]. constrained [7, 125, 320, 402, 819, 891, 1010, 1028, 1121, 1358, 1376, 1383, 1390, 1394, 1524, 1953]. constraints [249, 429, 567, 1085, 1468, 1786]. constricted [1999]. Construction [739, 986, 1599]. consumer [533, 1776]. consumption [276, 692, 1035, 1255, 1397, 1424, 1541, 1679, 1729, 1860, 2099, 2138].

Contact [642]. containerized [1935, 2138]. containment [1710]. contemporary [471, 547]. Content [94, 224, 255, 295, 335, 490, 510, 769, 833, 867, 886, 895, 996, 1065, 1074, 1116, 1324, 1497, 1642, 1670, 1694, 1711, 1726, 1846, 1871, 1892, 1948]. Content-aware [867]. content-based [1324, 1694]. Contention [903, 1447, 1721]. contention-based [1721]. contention-free [903]. contents [433]. Context [21, 220, 643, 1303]. Context-aware [220]. contexts [1482]. Contextual [1377]. Continuity [840]. continuous [768, 859]. continuum [2025]. contract [1267, 1499, 1645, 1925, 2023]. contract-based [1925]. ContractBox [1635]. contracts [1105]. contrastive [1964]. ConTrib [622]. Control [16, 110, 116, 192, 211, 343, 379, 494, 498, 506, 523, 534, 545, 561, 565, 573, 579, 596, 628, 644, 705, 731, 860, 872, 883, 915, 1028, 1098, 1124, 1133, 1192, 1228, 1305, 1338, 1359, 1372, 1415, 1486, 1526, 1592, 1597, 1660, 1700, 1713, 1799, 1824, 1856, 1859, 1867, 1904, 1905, 1924, 1929, 2010, 2016, 2071]. control-theory-based [1338]. controlled [1129]. controller [231, 355, 455, 624, 639, 759, 763, 787, 976, 1037, 1155, 1700, 1732, 1918, 2011]. controllers [456, 500, 599, 791, 1347, 1931]. controlling [1620]. conventional [466]. convergence [722, 1965]. converging [1952]. conversion [721]. convex [1248]. Convolution [907, 1507]. Convolutional [120, 226, 269, 496, 675, 944, 1342, 1634, 1775]. Cooperation [11, 19, 864, 900, 943]. Cooperative [4, 115, 153, 253, 290, 461, 560, 642, 756, 831, 1039, 1062, 1227, 1269, 1408, 1441, 1454, 1472, 1480, 1506, 1743, 1823, 1827, 1872, 1892, 1904, 1954, 2012, 2050, 2119]. CoopFL [1399]. Coordinated [19, 1249, 1311, 1444, 1930]. coordination [19, 521, 735, 750, 1678, 1828, 1937]. cope [1751]. copper [171]. copper-to-fiber [171]. Core [1256, 1462, 2039, 2099].

Core-selecting [1462]. correct [760]. Correcting [762]. correction [1244, 2040]. Correlated [1978]. Correlation [99, 925, 1604, 1720]. correlations [117, 600]. Corridor [270]. Corrigendum [296, 361, 547, 1588, 1764, 1765, 1903]. cosine [1162]. COSMOS [847, 1742]. Cost [53, 136, 333, 335, 456, 707, 716, 763, 860, 977, 1013, 1070, 1125, 1296, 1557, 1613, 1661, 1670, 1735, 1956, 1974, 1981, 2005, 2036, 2037, 2062]. cost-aware [707, 1070, 1125, 1956, 1974]. Cost-efficient [2005]. cost-energy-QoE [136]. cost-latency [1013]. COTS [22]. Count [1285, 1685]. counter [528]. countermeasure [1188]. countermeasures [603, 1646]. counters [1377]. counting [1324]. Coupling [175, 1883]. Coupon [1595]. covariance [1368]. cover [679, 1350]. Coverage [201, 467, 806, 1069, 1071, 1085, 1279, 1406, 1544, 1557, 1609, 1610, 1688, 1802, 1863, 1916]. coverage-based [1279]. coverage-centric [467]. covert [401, 612, 644, 980, 1663, 1687]. COVID [209, 382, 1377, 1512]. COVID-19 [209, 382, 1377, 1512]. CPS [1663]. CPU [1656]. CPU-bound [1656]. CR [756, 1691]. CR-NOMA [1691]. CRAN [2116]. CRC [1244]. CRC-aided [1244]. Credential [1017, 1894]. credibility [651]. credit [641, 1192, 1963, 2063]. credit-based [1192]. criteria [230, 464]. critical [352, 1169]. CRLM [1823]. CRN [2119]. CRNs [702]. Cross [16, 24, 324, 368, 504, 617, 1066, 1330, 1368, 1479, 1540, 1578, 1678, 1750, 1841, 1856, 1921, 2102]. cross-chain [1330]. Cross-domain [504, 1368, 1578, 1750, 1856]. cross-ISP [1479]. Cross-layer [368, 617, 1066, 1540, 1841, 2102]. cross-link [1678]. cross-producting [1921]. Cross-site [24]. cross-task [324]. crosschain [874]. crossing [1234]. crosstalk [339, 1436, 1886, 2015]. crosstalk-avoided [2015]. crosstalk-aware [1436, 2015]. crosstalk-derived [1886]. crowd [664, 1103, 1225, 1454, 1737, 1739]. crowd-sourced [1225, 1739]. crowded [58]. Crowdsensing [119, 129, 404, 1149, 1446, 1625, 1712, 1858, 2128]. crowdsourced [255, 1191, 1208]. Crowdsourcing [78, 114, 183, 420, 430, 737, 1213, 1488, 1722, 1991]. cruising [843]. crypto [611]. crypto-ransomware [611]. cryptocurrency [1364]. cryptographically [505]. cryptosystem [327]. CSI [515, 604, 757, 1480, 1743]. CSMA [528, 717, 725, 1243]. CSMA/CA [528, 717]. CTAP [780]. CTI [1739]. CTS [893]. CUBIC [1518]. Cuckoo [1561, 1884]. CUPS [465]. CUPS-based [465]. Current [653, 810, 1577, 1623, 1646]. CUSTOM [252]. Customized [400]. CUSUM [1295]. CWT [1354]. CWT-DPA [1354]. CYBELE [68]. Cyber [82, 469, 540, 613, 658, 1042, 1053, 1332, 1336, 1496, 1546, 1604, 1648, 1822]. Cyber-Physical [469, 1042, 1332, 1336, 1546]. Cyber-security [82, 658]. cybercrime [1739]. cybersecurity [855, 1126, 1853]. Cybersecurity-vs-QoS [1853]. cycle [412, 1064, 1540]. cycles [2110]. cycling [2046].

D [587, 602, 934, 1145, 1273, 1328, 1559, 1669, 1678, 1684, 1971, 1989]. D-AdFeed [587]. D-TDD [1678]. D2D [10, 80, 87, 136, 189, 263, 268, 413, 422, 426, 650, 707, 1070, 1083, 1090, 1097, 1100, 1274, 1278, 1423, 1456, 1506, 1544, 1626, 1869]. D2D-assisted [413]. D2D-enabled [10, 189, 1070]. D2DLive [1626]. DA-GNN [2023]. DACP [1856]. DADCNF [1449]. DAG [784, 891, 1792]. DAG-SFCs [1792]. damages [117]. Danger [845]. DAP [683]. DAP-Sketch [683]. dark [1055]. DARVAN [1469]. DASH [99, 1806]. Data [10, 13, 25, 32, 38, 49, 68, 106, 124, 157, 158, 238, 264, 276, 314, 316, 336, 343, 354, 359, 419, 433, 459, 468, 476, 480, 483, 555, 569, 595, 611, 641, 703, 708, 710, 717, 736, 758, 764, 776, 826, 832, 901, 902, 908, 946, 951, 965, 1000, 1043, 1052, 1067, 1081, 1102, 1113, 1132, 1195, 1218, 1228, 1275, 1348, 1354, 1357, 1407, 1416, 1429, 1446, 1452–1454, 1461, 1499, 1503, 1515, 1526, 1547, 1549, 1562, 1579, 1595, 1620, 1635, 1647, 1669, 1684, 1720, 1771, 1772, 1776, 1787, 1793, 1830, 1875, 1878, 1892, 1938, 1950, 1970, 1978, 1985, 2104, 2108, 2117, 2128]. data [71, 118, 155, 170, 173, 191, 193, 225, 249, 280, 296, 366, 443, 454, 473, 516, 631, 663, 695, 704, 771, 848, 856, 936, 952, 959, 1038, 1063, 1140, 1143, 1165, 1232, 1337, 1409, 1410, 1424, 1445, 1475, 1622, 1662, 1760, 1778, 1790, 1808, 1851, 1864, 1868, 2011, 2019, 2036, 2058]. data-aware [1669]. data-blocks [848]. Data-centric [611]. Data-driven [155, 296, 758]. data-efficient [901]. data-sharing [1102]. data-smart [708]. databases [1022]. Datacenter [140, 939, 1152, 1201, 1785, 1824]. datacenters [92, 338, 730, 1192, 1282, 1947, 2111]. dataplanes [436]. Dataset [182, 226, 654, 772, 839, 925, 993, 994, 1259, 1439, 1565, 1586, 1673, 2069, 2097, 2139]. Datasets [322, 391, 540, 737, 1739]. day [305]. DC [1745]. DDoS [492, 501, 535, 818, 854, 1061, 1084, 1095, 1188, 1207, 1239, 1295, 1527, 1725, 1836, 1838, 1957, 2014, 2103]. DDoS-MQTT-IoT [1673]. DE-GNN [2115]. de-randomization [994, 1324]. deadline [1201, 1785]. deadline-aware [1201]. deadlines [179, 2037]. debris [1880]. December [3, 362, 398, 849, 880, 1306, 1344, 1844]. decentralization [841, 2095, 2136]. decentralize [1760]. Decentralized [299, 521, 622, 715, 823, 882, 903, 933, 1034, 1091, 1149, 1150, 1201, 1469, 1503, 1642, 1645, 1714, 1815, 1827, 1857, 1944]. Decision [552, 1508, 1561, 1597, 1967]. Decision-Making [1561, 1597]. DEcisions [436]. decoding [183]. decomposition [31, 861, 1735, 2075]. decoupled [1167, 1481]. DECS [663]. Dedicated [1202]. deduplication [2128]. Deep [55, 116, 163, 280, 288, 292, 317, 364, 379, 406, 470, 478, 491, 570, 579, 588, 608, 636, 654, 660, 700, 727, 737, 738, 744, 757, 776, 798, 844, 872, 878, 883, 908, 929, 951, 962, 969, 972, 978, 1020, 1029, 1037, 1050, 1061, 1062, 1081, 1089, 1126, 1177, 1200, 1202, 1217, 1219, 1233, 1253, 1265, 1273, 1274, 1281, 1338, 1339, 1341, 1342, 1370, 1377, 1401, 1414, 1471, 1477, 1495, 1500, 1501, 1527, 1547, 1556, 1568, 1614, 1628, 1629, 1659, 1668, 1677, 1711, 1713, 1720, 1732, 1733, 1754, 1790, 1796, 1835, 1839, 1862, 1877, 1949, 1957, 1980, 1981, 2006, 2014, 2021, 2049, 2091, 2098, 2101, 2116, 2128]. DeepLoad [439]. deeply [751]. DeepMetricCorr [1720]. Defending [1155, 1754]. Defense [283, 613, 1053, 1061, 1239, 1425, 1508, 1648, 1662, 1693, 1725]. defenses [110, 1054, 1992]. Defined [76, 84, 95, 164, 216, 231, 349, 355, 443, 446, 455, 475, 512, 520, 523, 545, 606, 613, 624, 630, 658, 673, 706, 746, 790, 858, 978, 1037, 1066, 1096, 1102, 1120, 1163, 1184, 1246, 1535, 1607, 1641, 1693, 1742, 1782, 1796, 1876, 1897, 1918, 1926, 1977, 1996, 2103, 2137]. defining [71]. Defragmentation [239]. degradation [56]. degree [600, 841, 988]. DEHO [1207]. Delay [7, 78, 151, 304, 320, 341, 379, 385, 402, 494, 511, 536, 558, 639, 759, 781, 928, 958, 1006, 1008, 1088, 1092, 1121, 1210, 1310, 1340, 1358, 1376, 1405, 1493, 1518, 1524, 1591, 1697, 1755, 1806, 1817, 1837, 1974, 2047, 2061, 2102]. Delay-aware [379, 781, 2047]. delay-based [1006, 1806]. Delay-constrained [320, 1121, 1358, 1376, 1524]. delay-guaranteed [385, 1697]. delay-tolerant [1092]. delayed [1644]. Delegation [824, 1491, 1894]. Delivering [380]. Delivery [116, 276, 352, 510, 710, 769, 970, 996, 1176, 1497, 1588, 1734, 1871, 1892, 1948]. delta [1470]. delta-based [1470]. Demand [26, 316, 339, 380, 1152, 1533, 1736, 1947, 2005, 2036, 2056]. Demand-aware [1152, 1947]. demands [960, 1672]. Democratizing [204]. Dendrimer [1645]. deniable [1922]. Denial [160, 164, 317, 512, 1084, 1252, 1455, 1463]. denial-of-service [164]. Dense [35, 208, 413, 559, 562, 580, 738, 906, 1323, 1386, 1393, 1733]. Density [197, 1124, 1744]. Density-aware [197]. departure [212]. dependent [1468, 1476, 1665]. deployed [243, 1003, 1280, 1291]. Deploying [1121, 1792]. deployment [91, 105, 111, 562, 763, 789, 816, 853, 888, 987, 1169, 1230, 1255, 1373, 1392, 1538, 1629, 1665, 1669, 1742, 1753, 1843, 1936, 2047, 2061]. deployments [1224, 1738, 1863]. depth [1357]. derived [1886]. description [694]. Design [29, 33, 84, 113, 229, 230, 247, 307, 310, 388, 412, 427, 515, 519, 575, 623, 674, 681, 701, 714, 759, 773, 792, 917, 956, 976, 982, 1028, 1066, 1086, 1169, 1181, 1246, 1449, 1464, 1465, 1514, 1583, 1640, 1672, 1699, 1723, 1741, 1742, 1791, 1831, 1841, 1847, 1881, 1996, 1997, 2051]. designated [1245]. Designing [351, 781, 1313]. designs [1014]. DESiRED [2098]. DESLR [1748]. destructive [1289]. Detail [409]. detect [802, 1207, 1645]. Detecting [39, 79, 531, 612, 751, 1063, 1195, 1663, 1835, 1957, 2054, 2057]. Detection [29, 32, 55, 59, 74, 137, 152, 283, 305, 317, 414, 420, 441, 448, 449, 463, 469, 491, 492, 535, 538, 540, 555, 570, 582, 611, 614, 625, 633, 654, 675, 676, 715, 718, 772, 801, 845, 854, 861, 901, 907, 908, 916, 929, 941, 957, 967, 968, 983, 985, 1058, 1076, 1084, 1099, 1172, 1185, 1188, 1199, 1225, 1234, 1242, 1248, 1250, 1252, 1287, 1295, 1299, 1335, 1336, 1379, 1396, 1440, 1449, 1496, 1515, 1529, 1556, 1598, 1689, 1704, 1740, 1744, 1755, 1762, 1771, 1812, 1821, 1828, 1839, 1853, 1859, 1932, 1939, 1941, 1994, 2014, 2019, 2090, 2103, 2130, 2133]. detection [70, 75, 146, 176, 226, 312, 446, 661, 752, 754, 798, 890, 944, 999, 1204, 1267, 1363, 1439, 1527, 1531, 1575, 1578, 1594, 1766, 1783, 1866, 1920, 1938, 2023]. detector [67, 269, 863, 1703]. detectors [324]. deterioration [676]. Determining [276]. deterministic [683, 840, 1203, 1653, 1914]. DetNet [1104]. detouring [1571]. development [1407]. deviation [1226]. Device [47, 87, 182, 607, 634, 889, 994, 1238, 1324, 1524, 1847, 1883, 1924, 1936, 2060]. Device-to-Device [47, 634, 889, 1883]. Devices [57, 91, 104, 251, 270, 485, 602, 605, 685, 726, 733, 813, 935, 968, 1042, 1122, 1133, 1283, 1679, 1834, 1841, 1999, 2070, 2109]. DFS [1622]. DGA [1439, 1788]. diagnose [1395]. Diagnoser [1449]. diagnosis [793]. DID [2105]. DID-based [2105]. different [604, 857, 904, 1665]. differential [235, 855, 1495, 1508, 1978, 2111]. Differentially [1337]. differentiated [2111]. Differentiation [118, 852]. differentiations [989]. DiffServ [918]. DIFFUSE [1091]. diffusion [304]. Digital [441, 598, 741, 777, 1292, 1369, 1509, 1655, 1840, 1984, 2042, 2096]. Digital-care [1509]. dimension [1084]. dimensional [770, 1251, 1571, 1609, 1794, 1875, 2082]. dimensioning [1433]. direction [319, 1667]. directional [1685, 1971]. directions [584, 1438, 1509, 1577, 1801]. disaggregated [1791]. disagreement [737]. disaster [294, 314, 434, 1523]. disaster-aware [314]. disciplines [113]. Discontinuous [1164]. Discovering [563, 582]. Discovery [87, 102, 181, 251, 605, 694, 858, 1398, 1789, 1971]. Discrimination [258, 1295]. Discrimination-aware [258]. disposable [435]. disruptive [437]. Dissecting [1005]. Dissemination [13, 83, 393, 886, 915, 1247, 1620, 1827]. Distance [869, 1368]. distinct [514]. Distinguishing [896]. distribute [2027]. Distributed [42, 48, 92, 109, 146, 247, 253, 287, 308, 314, 317, 374, 421, 457, 490, 521, 566, 648, 659, 744, 791, 808, 825, 844, 870, 879, 882, 891, 920, 959, 978, 1004, 1033, 1056, 1061, 1080, 1084, 1090, 1091, 1097, 1125, 1128, 1157, 1182, 1202, 1216, 1252, 1277, 1282, 1329, 1367, 1371, 1398, 1455, 1463, 1464, 1477, 1514, 1527, 1562, 1569, 1580, 1619, 1622, 1660, 1704, 1717, 1769, 1773, 1828, 1877, 1879, 1971, 2035, 2049, 2052, 2053, 2071, 2132]. distribution [188, 208, 490, 687, 812, 1018, 1179, 1670, 1868, 1983, 2069]. distributions [988]. DIVA [2105]. dive [1177]. Diver [2139]. Diver-robot [2139]. diverse [516, 1679, 2131]. Diversity [134, 179, 200, 433, 587, 1272, 2039]. diversity-aware [587]. Diversity-improved [433]. divide [811]. divide-and-conquer [811]. Divisible [1403]. division [178, 339, 1057, 1315, 1390, 1701, 2092]. DL [883, 1384, 1831]. DL-based [1384, 1831]. DLT [1403]. DM [373]. DM-GKM [373]. DMatrix [797]. DMO [1164]. DMZ [2011]. DNA [1287]. DNN [898, 1399, 1658, 1682, 1826]. DNN-based [898]. DNS [752, 825, 1063, 1177, 1598, 1879]. DNSxP [695]. DOCUS [1295]. DOCUS-DDoS [1295]. Does [2052]. DoH [1131, 1762]. Domain [248, 256, 282, 435, 471, 504, 537, 547, 605, 661, 824, 1125, 1327, 1368, 1419, 1532, 1578, 1597, 1750, 1840, 1856, 1975, 2014, 2130]. domain-based [1419]. domains [500]. Dominant [39, 1557]. donations [255]. DONES [1716]. DoS/DDoS [1673]. DoS/DDoS-MQTT-IoT [1673]. Double [55, 551, 868, 1244, 1281, 1288]. double-stage [868]. down [959]. Downlink [367, 696, 725, 759, 1130, 1167, 1316, 1481, 2022, 2126]. downlink-uplink [1167]. downlink/uplink [367]. DP [1337]. DPA [1354]. DPoS [841]. drift [1547, 2076]. drifted [1596]. Driven [8, 155, 207, 296, 491, 498, 681, 758, 951, 966, 1176, 1271, 1297, 1323, 1389, 1395, 1411, 1556, 1588, 1699, 1756, 1883, 1932, 1961, 2087, 2111, 2131]. driver [53]. Driving [324, 631]. DRL [1190, 1896, 2087]. DRL-driven [2087]. DROI [5]. Drone [59, 808, 834, 1366, 1498, 1534, 1905]. drone-assisted [1366, 1534]. Drone-enabled [834]. drones [321, 596, 970, 1017, 1802, 1917, 1940]. Drops [1195, 1367]. drug [116]. drug-delivery [116]. DSMLB [1184]. DSO [1367]. DSSS [505]. DSVDD [1819]. DTLShps [1237]. Dual [143, 467, 754, 860, 981, 1032, 1334, 1360, 1455, 1592, 1970, 1987, 2023, 2115]. dual-antenna [143]. dual-grained [1334]. dual-layer [1455]. dual-link [860]. dual-radio [1032]. dual-slope [467]. dual-weight [981]. due [445]. Duplex [4, 134, 550, 847, 1086, 1274, 1828]. Duplicate [801, 1449]. duplication [1745]. duration [642]. duration-aware [642]. during [209, 431, 1377, 1512, 2057]. duty [1540]. DV [1068, 1794]. DV-Hop [1068, 1794]. DVB [1831]. DVB-S2X [1831]. DWT [1188]. Dynamic [5, 17, 19, 35, 37, 109, 174, 216, 247, 273, 295, 334, 373, 375, 379, 404, 459, 489, 510, 527, 532, 534, 562, 588, 636, 657, 671, 678, 682, 728, 919, 971, 975, 993, 1037, 1047, 1147, 1180, 1184, 1198, 1209, 1215, 1262, 1281, 1322, 1416, 1423, 1461, 1478, 1500, 1517, 1528, 1582, 1593, 1654, 1661, 1698, 1709, 1732, 1745, 1763, 1797, 1805, 1813, 1818, 1856, 1861, 1872, 1947, 1974, 1977, 1982, 2013, 2098, 2116]. Dynamically [918, 1822]. dynamics [304, 855, 1473]. DynamicTuple [919]. DynGATT [1461]. DySO [1516].

e-commerce [419]. E-health [266, 952]. e-Learning [209]. EagerCC [1824]. EAP [2086]. EAP-OAUTH [2086]. Early [314, 449, 676, 866, 1731, 1828, 1941]. earth [1692]. eavesdroppers [1480, 1743]. Eavesdropping [2046]. eBPF [476, 1450, 2072]. EC [1507]. EC-GCN [1507]. ECC [1494]. ECC-based [1494]. ECCPWS [1494]. ECN [760]. Economic [512, 1039]. economics [62]. economy [122, 1278, 1369]. ecosystem [897, 1055]. ecosystems [1008]. ED [1817]. ED-VNE [1817]. Edge [48, 49, 195, 210, 270, 353, 400, 402, 439, 511, 551, 553, 577, 609, 639, 670, 681, 687, 693, 700, 705, 769, 781, 784, 800, 807, 816, 819, 823, 859, 877, 933, 935, 958, 975, 977, 1016, 1029, 1061, 1067, 1075, 1080, 1087, 1120, 1128, 1130, 1132, 1160, 1182, 1190, 1194, 1239, 1284, 1288, 1328, 1346, 1349, 1358, 1366, 1373, 1376, 1420, 1429, 1430, 1454, 1471, 1477, 1482, 1485, 1554, 1558, 1561, 1572, 1584, 1591, 1628, 1635, 1644, 1655, 1708, 1709, 1711, 1730, 1747, 1750, 1792, 1829, 1846, 1875, 1911, 1951, 1970, 1987, 2025, 2026, 2034, 2035, 2047, 2056, 2066, 2083, 2084, 2110, 2137]. edge [91, 206, 207, 260, 267, 366, 387, 392, 397, 415, 538, 560, 626, 642, 648, 663, 724, 864, 870, 909, 936, 944, 966, 972, 978, 1035, 1062, 1108, 1114, 1118, 1147, 1213, 1262, 1326, 1337, 1340, 1371, 1386, 1388, 1399, 1421, 1472, 1474, 1475, 1481, 1527, 1536, 1582, 1592, 1594, 1597, 1658, 1732, 1752, 1753, 1757, 1796, 1862, 2012, 2037, 2039, 2064, 2073, 2102]. edge-assisted [2084]. edge-based [1067, 1482]. edge-cloud [210, 2025]. edge-cloud-based [1592]. edge-computing [781, 2035]. edge-edge [663]. edge-end [1582]. Edge-enhanced [1875]. edge-to-core [2039]. EdgeML [1346]. edges [91, 984, 1935, 2074]. EDiPSo [858]. Editorial [2, 3, 28, 51, 73, 90, 108, 128, 149, 169, 185, 187, 203, 223, 246, 272, 298, 326, 363, 399, 424, 451, 482, 503, 525, 549, 572, 594, 616, 638, 658, 667, 690, 708, 712, 743, 779, 815, 850, 881, 914, 931, 947, 954, 974, 1002, 1026, 1044, 1046, 1060, 1078, 1094, 1109, 1112, 1139, 1171, 1187, 1196, 1206, 1223, 1264, 1307, 1345, 1382, 1418, 1443, 1467, 1484, 1522, 1552, 1567, 1590, 1616, 1637, 1652, 1676, 1696, 1728, 1763, 1768, 1810, 1845, 1907, 1928, 1946, 1969, 2004, 2044, 2080]. Editorial [2107]. EDTP [928]. education [1901]. EEM [1694]. Effect [205, 961, 1406, 1446, 1916]. effect-based [1446]. Effective [59, 92, 167, 171, 214, 226, 372, 683, 771, 897, 1505, 1667, 1704]. effectiveness [160]. effects [1685]. Efficiency [144, 175, 215, 280, 323, 553, 556, 626, 820, 857, 1311, 1516, 1787, 1861, 1893, 1986, 2017, 2126]. Efficient [5, 9, 37, 47, 53, 54, 59, 102, 177, 234, 237, 293, 317, 341, 385, 406, 414, 466, 468, 535, 546, 570, 574, 580, 591, 635, 702, 730, 739, 761, 774, 775, 781, 795, 811, 812, 819, 858, 868, 901, 934, 938, 969, 1017, 1020, 1028, 1108, 1141, 1165, 1173, 1174, 1189, 1200, 1207, 1231, 1236, 1260, 1261, 1274, 1276, 1294, 1296, 1300, 1304, 1314, 1317, 1339, 1420, 1428, 1432, 1451, 1495, 1503, 1524, 1532, 1537, 1594, 1607, 1627, 1634, 1678, 1680, 1688, 1701, 1706, 1715, 1717, 1733, 1734, 1747, 1789, 1859, 1888, 1908, 1930, 1980, 2000, 2005, 2034, 2058, 2061, 2092, 2109, 2123]. efficient [96, 176, 253, 415, 648, 652, 728, 765, 830, 869, 873, 885, 900, 979, 1018, 1081, 1151, 1254, 1262, 1308, 1368, 1401, 1410, 1536, 1622, 1748, 1788, 1790, 1909, 1934, 1936]. effort [1699]. efforts [282]. egress [574]. EH [2119]. EH-CRN [2119]. eHDDP [605]. ehealth [955]. EHF [731]. EID [1247]. Elastic [42, 77, 133, 144, 178, 239, 339, 506, 541, 592, 721, 748, 749, 869, 939, 1180, 1251, 1309, 1315, 1390, 1391, 1535, 1694, 1701, 1886, 1896, 2015, 2062, 2081, 2092]. Elastically [1926]. electromagnetic [181]. Electrosense [183]. elimination [414]. emails [1020]. eMBB [976]. embedded [449, 1066, 1634, 1834, 1890]. Embedding [5, 273, 420, 472, 697, 1235, 1338, 1390, 1400, 1532, 1701, 1735, 1930, 2115]. Emergency [393, 434, 1156, 1383, 1752, 1911]. emerging [585, 1072, 1350, 2096]. EMF [508]. emission [2059]. emissions [2052]. Empirical [604, 1547, 1848, 1998, 2018]. Employing [1416]. empowered [1274, 1650, 1857, 1967, 2032, 2073, 2093]. Empowering [87, 912]. emulation [213, 672, 1075]. emulators [1901]. Enable [741, 1700]. Enabled [8, 10, 16, 52, 68, 71, 111, 154, 167, 185, 189, 211, 283, 367, 393, 397, 416, 421, 449, 480, 496, 533, 553, 710, 821, 834, 865, 869, 917, 935, 950, 966, 979, 1008, 1011, 1018, 1030, 1070, 1089, 1141, 1160, 1247, 1310, 1328, 1354, 1373, 1386, 1456, 1472, 1493, 1578, 1607, 1611, 1631, 1732, 1747, 1780, 1835, 1840, 1866, 1948, 1950, 1985, 2088, 2093]. enablers [179, 1900]. Enabling [179, 231, 388, 824, 1091, 1092, 1375, 1791, 1878, 1902, 1980]. eNCache [1892]. encapsulation [43]. encoder [1188, 1556]. encounter [334]. encrypted [214, 288, 432, 591, 846, 877, 911, 1116, 1142, 1146, 1489, 1507, 1547, 1549, 1596, 1602, 1781, 1811, 1899, 1903, 1920, 2054, 2115, 2121, 2131]. Encryption [384, 698, 876, 882, 1283, 1308, 1414, 1686, 1851, 2117, 2132]. End [112, 233, 248, 494, 1145, 1541, 1546, 1582, 1653, 1658, 1752, 1758, 1831]. End-to-edge [1658]. End-to-end [233, 248, 494, 1145, 1541, 1546, 1758, 1831]. endogenous [673]. endpoints [1012]. Energy [9, 38, 40, 47, 64, 112, 134, 136, 144, 175, 215, 221, 234, 275, 276, 323, 341, 343, 353, 367, 377, 385, 402, 431, 556, 626, 640, 692, 702, 707, 739, 761, 775, 781, 820, 830, 857, 868, 885, 928, 934, 958, 960, 992, 1017, 1018, 1029, 1035, 1067, 1089, 1106, 1137, 1141, 1165, 1174, 1189, 1231, 1240, 1254, 1260, 1274, 1276, 1304, 1305, 1311, 1314, 1339, 1424, 1493, 1500, 1503, 1524, 1610, 1627, 1629, 1679, 1680, 1688, 1715, 1729, 1733, 1747, 1748, 1787, 1833, 1841, 1869, 1870, 1908, 1913, 1930, 1934, 1950, 1979, 1986, 1994, 2017, 2058, 2061, 2092]. energy [5, 253, 415, 648, 652, 1027, 1081, 1410, 1790, 1817, 1954, 1999, 2126]. Energy-aware [47, 64, 1027, 1089, 1240, 1610, 1999]. energy-efficiency [626]. Energy-Efficient [5, 234, 253, 341, 385, 415, 648, 652, 775, 781, 868, 885, 1017, 1081, 1141, 1165, 1174, 1189, 1254, 1274, 1276, 1304, 1314, 1339, 1503, 1524, 1747, 1748, 1908, 1930, 1934]. energy-harvesting [38, 367, 707, 960]. energy-optimal [1833]. energy-saving [1979]. energy-spectral [1311]. ENERO [1200]. enforcement [6, 649, 1053, 1166, 1271]. Enforcing [1856]. engine [1364, 2001]. Engineering [466, 574, 579, 1049, 1218, 1643, 1980]. enhance [600]. Enhanced [31, 45, 76, 478, 605, 663, 786, 813, 980, 1068, 1313, 1479, 1520, 1602, 1630, 1875, 1910, 2029, 2091, 2098]. Enhancement [943, 1036, 1199, 1457, 1606, 1750]. enhancements [1122]. Enhancing [25, 695, 1243, 1310, 1516, 1983, 1992, 2017, 2086]. eNodeB [1791]. eNodeBs [578]. enrichment [733]. Ensemble [23, 75, 176, 969, 1355]. Ensemble-model-based [23]. ensure [1988]. ensuring [351]. enterprise [772, 1334, 1363, 1704]. entity [1770]. Entropy [988, 1848, 2103]. entry [501]. Environment [219, 260, 270, 490, 512, 530, 570, 710, 802, 928, 996, 1030, 1087, 1132, 1169, 1198, 1207, 1240, 1327, 1497, 1572, 1597, 1772, 1778, 1855, 1956, 2101]. environmental [2071]. environments [20, 58, 68, 312, 612, 757, 767, 805, 870, 1091, 1096, 1204, 1308, 1343, 1374, 1419, 1477, 1684, 1777, 1856, 1885, 1889, 1922, 1923, 2041]. envisioned [1799]. EON [178, 2049]. EONs [1436]. ephemeral [277]. epidemic [1486]. epidemiological [304]. EPIoT [1910]. Equalizing [1523]. equation [1837]. Equilibrium [381, 1441]. equipped [1088]. equivalence [60]. era [71, 204, 274, 1134, 1414, 1534, 2135]. erasure [959]. erasure-coded [959]. ERRANT [213]. Error [762, 1244, 1384, 2040]. ESN [1664]. ESRAS [1296]. estimate [265, 898, 1848]. estimating [1666, 1807, 1868, 1984]. Estimation [430, 729, 812, 887, 902, 1073, 1349, 1364, 1831, 1865]. Ethereum [1899]. Ethernet [235, 1715, 1716]. Ethernet-based [1716]. ETSI [1721]. Europe [336]. evaluated [578]. Evaluating [348, 941, 1673, 2136]. Evaluation [12, 29, 34, 41, 101, 220, 254, 370, 384, 396, 403, 436, 528, 580, 604, 627, 672, 686, 697, 767, 836, 936, 1053, 1119, 1251, 1293, 1318, 1319, 1322, 1347, 1412, 1612, 1623, 1749, 1843, 1929, 1963, 2041, 2076]. evaluations [740]. Event [411, 656, 1247, 1694]. event-participant [411]. events [808, 1289]. everything [206]. eviction [250, 1644]. evidence [2018]. evil [1568]. EvoIoT [1374]. Evolution [632, 868, 1744, 1837, 2067]. evolutionary [79, 152, 276, 456, 1374, 2122]. Evolving [55, 1816]. Exact [307]. Examples [359]. excessive [1620]. Exchange [248, 291, 659, 726, 2113]. execution [490, 539, 574, 707, 1477]. exemplar [1489]. exfiltration [695, 1063]. existing [904]. Expediting [1192]. experience [221, 224, 1806, 1905]. experiment [575]. Experimental [34, 388, 436, 665, 758, 804, 947, 1056, 1119, 1539, 1804, 1855, 2038, 2076, 2100]. experimentation [68, 440, 991, 1742, 1791]. Explainable [1378, 2063]. explaining [1252]. exploit [182]. exploitation [131, 1996]. Exploiting [91, 99, 600, 886, 1105, 1129]. exploration [26, 1338]. Exploring [1088, 1452, 1541, 1685, 1801]. exponentially [761]. exponentially-ant [761]. exposure [508, 1241]. expression [54]. Extended [1861, 1953]. Extending [279, 1847]. extensible [871]. extensions [58, 282]. extensive [925]. extract [1289]. extracting [1203]. extraction [785, 910, 1250, 1526]. extremely [279, 1470].

F [1812]. F-NIDS [1812]. F2DC [1286]. F4Tele [483]. fabric [792]. face [1547]. facial [2097]. Facilitating [507]. facilities [1716]. factor [240, 383, 829, 1476, 1534]. factories [1752]. factorization [729, 887]. fading [134, 1761, 1831, 1904, 2050]. Failure [445, 623, 784, 1052, 1246, 1298, 1378, 1458, 1665]. Failure-resilient [784]. failures [232, 431, 624, 676, 1177, 1422, 1877, 1918, 1933]. Fair [691, 720, 1048, 1505, 1553, 1558, 1893, 1960, 1988, 2021]. Fairness [11, 125, 486, 622, 924, 1006, 1360, 1468, 1498, 1785]. Fairness-Aware [125]. falsification [46]. family [447]. FANET [788, 2016]. FANETs [788, 1333, 1376]. farming [48, 68, 69, 71, 131, 822]. Fast [60, 157, 199, 265, 269, 647, 794, 797, 902, 1054, 1064, 1246, 1361, 1517, 1720, 1781, 1864, 1890, 1903, 2047]. FastCredit [1192]. Faster [70]. FastTraffic [1781, 1903]. FASUS [199]. fat [1218]. fat-tree [1218]. Fault [17, 95, 191, 541, 793, 840, 1137, 1259, 1272, 1301, 1336, 1396, 1608, 1777, 1864, 1908]. Fault-tolerant [1272, 1608, 1777]. FCNR [647]. FDMA [1190]. FDMA-based [1190]. FDR [1480]. feasibility [1681]. feasible [916]. Feature [137, 159, 176, 447, 463, 854, 890, 901, 910, 1199, 1250, 1299, 1379, 1531, 1581, 1887, 1890]. feature-embedded [1890]. Features [563, 591, 929, 1302, 1953, 2048, 2097, 2131]. featuring [1746]. February [27, 50, 450, 481, 930, 953, 1417, 1442, 1765, 1927, 1945]. FedAAA [1931]. FedAAA-SDN [1931]. Federated [400, 500, 635, 776, 819, 827, 924, 941, 966– 968, 998, 1039, 1076, 1125, 1135, 1154, 1193, 1213, 1346, 1376, 1399, 1420, 1475, 1482, 1496, 1541, 1550, 1558, 1672, 1751, 1756, 1766, 1774, 1812, 1821, 1888, 1931, 1999, 2018, 2063– 2065, 2090, 2109, 2132, 2135, 2136]. Federated-Learning [2090]. Federation [1073, 1749]. FedPA [827]. FedSBS [2090]. FedSSC [1888]. fee [1175]. Feedback [1401, 1805]. Feedback-based [1401, 1805]. fees [130]. femto [580]. FemtoCaching [1408]. femtocell [66]. Few [799, 1368, 1602, 1632]. Few-shot [799, 1368, 1602, 1632]. FFRLI [1864]. Fi [98, 507, 696, 787, 994, 1445, 1617, 1937, 1988]. fiber [171, 961]. fiber-optic [961]. Fidelity [1351]. Field [345, 358, 1066, 1182, 1808, 1847]. field-based [1182]. fieldable [1066]. fields [289]. file [1179, 1830]. Filter [158, 736, 1005, 1595, 1662, 1884, 1959]. Filtering [576, 1020, 2047]. Filtering-based [2047]. Finding [354, 435, 1702, 1884]. Fine [120, 534, 1191, 1340, 1413, 1592, 1672, 1707, 1899, 2115]. Fine-grained [534, 1191, 1340, 1413, 1592, 1672, 1707, 1899, 2115]. fine-tuned [120]. fingerprint [1915]. fingerprint-based [1915]. Fingerprinting [391, 799, 825, 866, 1005, 1270, 1368, 1370, 1576, 1598, 1623, 1632, 1667, 1693, 1811, 1983]. fingerprints [911, 1324]. FINT [1232]. fire [449]. firefly [808]. firewall [1216]. First [165, 679, 1243, 1487, 1667, 1905, 1951]. first-person-view [1905]. fitness [568]. fitting [578]. five [808]. five-step [808]. FIWARE [716]. Fixed [841, 843]. FL [1193, 1672]. FL-based [1672]. flagging [376]. FLAME [654]. flash [1295]. FLchain [1318]. fledged [1996]. Flex [1198]. Flex-CC [1198]. flexgrid [342]. Flexible [20, 309, 453, 635, 873, 926, 938, 965, 1132, 1198, 1219, 1232, 1237, 1433, 1445, 1488]. flight [415, 1038]. Flood [496, 501, 1662]. flooding [1957]. Flow [33, 39, 140, 158, 191, 193, 216, 389, 406, 445, 466, 509, 523, 591, 601, 682, 771, 803, 887, 1011, 1270, 1352, 1519, 1530, 1705, 1714, 1720, 1771, 1868, 1926, 1953, 2131]. Flow-level [191, 466, 1270]. FlowFight [718]. flows [403, 501, 1104, 1310, 1428, 1641, 1653, 1792, 1848, 1884]. fluctuations [1955]. flying [123, 1492, 1599, 1659, 2091]. flying-beamformer [1659]. FlyingLoRa [1410]. focus [1760]. Fog [211, 224, 281, 312, 347, 385, 477, 691, 698, 703, 714, 828, 830, 949, 996, 1008, 1013, 1030, 1083, 1096, 1143, 1153, 1178, 1204, 1221, 1230, 1240, 1293, 1403, 1419, 1425, 1493, 1531, 1633, 1686, 1777, 1882, 1908, 1956, 2020]. fog-aided [714]. Fog-based [312, 385, 1143, 1153, 1293]. Fog-cloud [1531]. fog-enabled [1008, 1030, 1493]. Fog-oriented [949]. fog/cloud [1777]. footprints [1702]. Forecasting [52, 496, 646, 942, 1191, 1528, 1774, 1813, 2018, 2045, 2077]. forest [802].

Form [1449]. formal [184, 291, 957]. Formalization [252, 557]. format [284, 1436]. formation [64, 2075]. formatting [1622]. formulation [250]. formulations [412]. Formullar [453]. forward [1173, 1893]. forward-secure [1173, 1893]. forwarder [1422]. Forwarding [26, 60, 157, 472, 755, 832, 1031, 1380, 1721, 1815, 1979]. fostering [68]. FoT [1721]. four [194]. four-stage [194]. FPGA [453, 1079]. FPGA-accelerated [1079]. FPGA-based [453]. FPGAs [36]. fractal [1084]. fractional [1990]. Fragmentation [133, 178, 339, 1251, 1886]. fragmentation-aware [133, 339]. Frame [497, 1291, 1734, 2051]. Framed [1351]. frames [1678]. Framework [42, 163, 250, 251, 267, 278, 290, 309, 315, 330, 355, 409, 411, 488, 504, 513, 518, 535, 570, 587, 589, 619, 620, 657, 660, 715, 789, 821, 836, 862, 871, 949, 965, 1009, 1039, 1053, 1075, 1076, 1090, 1097, 1102, 1150, 1185, 1201, 1229, 1265, 1332, 1351, 1376, 1397, 1398, 1420, 1450, 1459, 1464, 1470, 1479, 1507, 1515, 1528, 1542, 1561, 1575, 1595, 1634, 1645, 1682, 1694, 1710, 1719, 1757, 1766, 1782, 1832, 1850, 1857, 1897, 1930, 1944, 2076, 2087, 2093]. frameworks [131]. fraudsters [1568]. free [177, 425, 868, 903, 1019, 1448, 1503, 1508, 1580, 1624, 1803, 1922, 1965]. freed [760]. frequencies [286]. Frequency [902, 1370, 1510, 1702, 1861, 1915]. Friendship [692, 905]. fronthaul [2116]. fruit [70]. FSO [483, 1406]. FT [61]. FT-Aurora [61]. FTTx [674]. Full [4, 43, 550, 847, 1086, 1274, 1828, 1996]. Full-duplex [4, 550, 847, 1274, 1828]. full-fledged [1996]. fully [374, 598, 741, 1469]. fully-digital [741]. Function [8, 194, 256, 375, 418, 504, 536, 645, 671, 678, 693, 697, 857, 899, 979, 1091, 1108, 1157, 1159, 1214, 1220, 1255, 1272, 1280, 1373, 1422, 1426, 1437, 1458, 1462, 1478, 1485, 1501, 1665, 1701, 1786, 1828, 1849, 1936, 2000, 2005, 2007, 2036, 2061, 2062, 2140]. function-based [2062]. functionality [1114]. Functions [14, 264, 309, 333, 694, 942, 1367, 1639, 1656]. fusing [2058]. Fusion [386, 631, 717, 785, 929, 1132, 1135, 1199, 1293, 1515, 1575, 1581, 1712, 2048]. fusion-based [2048]. Future [44, 233, 273, 295, 655, 734, 1134, 1305, 1376, 1411, 1509, 1520, 1550, 1577, 1610, 1724, 1801, 1865, 2096]. fuzziness [841]. Fuzzy [492, 520, 739, 830, 1013, 1165, 1987]. fuzzy-based [520].

G [8, 15, 16, 18, 44, 88, 115, 126, 141, 145, 166, 167, 172, 179, 201, 204, 206, 207, 217, 219, 234, 1148]. G-3MRP [1148]. G-air-simulator [166]. G-enabled [167]. GA [1320, 1474]. GA-based [1320]. GAIL [1333]. Gait [1617]. gambling [2054]. Game [35, 210, 228, 242, 260, 407, 452, 617, 640, 650, 800, 855, 870, 958, 1095, 1148, 1224, 1454, 1508, 1529, 1550, 1641, 1737, 1858, 1872, 2066, 2122]. Game-based [228, 870]. game-theoretic [210, 800]. game-theoretical [1148]. games [1118]. gaming [984, 1666, 2101]. GAN [1009]. GAN-based [1009]. GANs [686]. GAT [1398]. GAT-based [1398]. gated [1556]. Gateway [162, 877, 1457, 1524, 1951]. gateway-to-device [1524]. gateways [383, 1429, 1878, 1924, 1926]. Gathering [106, 1787, 2058]. GATT [1461]. GATT-based [1461]. Gaussian [226]. Gbps [1715]. GCACS [596]. GCACS-IoD [596]. GCN [1507, 1836]. General [119, 1470, 1542, 1621]. generalisation [2130]. Generalized [381, 854, 1253]. Generation [182, 514, 519, 606, 659, 661, 723, 807, 821, 1009, 1053, 1372, 1509, 1614, 1619, 1822, 1975, 2068]. Generative [686, 877, 1333, 1489, 2060]. Generator [1433]. generic [596, 1332, 1868, 2127]. Genetic [219, 621, 642, 687, 693]. genetical [1301]. geo [314, 1128, 1279, 1282]. geo-distributed [314, 1282]. geo-indistinguishability [1279]. geo-replication [1128]. geofenced [2113]. Geographic [884, 1400]. geographical [652, 986, 1230]. geographical-aware [1230]. geographically [92, 490, 516]. geography [2052]. geolocation [1770]. geosynchronous [1929]. gesture [993, 2139]. GFDA [1250]. GFDA-WGAN [1250]. GHz [332, 473, 665, 696, 741, 804, 1885]. GKM [373]. GLADS [1549]. GlareShell [2133]. global [664, 1434, 1549, 2053]. global-local [1549]. GNN [1579, 1770, 2023, 2115]. GNN-based [1770]. gNodeB [1885]. goal [583]. goal-oriented [583]. Going [1611]. Good [544]. gossiping [1193]. Gov [1719]. governance [1719]. GPIO [1613]. GPS [322, 1225]. GR [143]. Grace [1613]. Gradient [1199, 1420]. gradual [171]. GRAIN [1142]. grained [534, 1191, 1334, 1340, 1413, 1592, 1672, 1707, 1899, 2115]. GramMatch [910]. Grano [432]. Grano-GT [432]. Grant [1448, 1624]. Grant-free [1448, 1624]. Granular [432, 1142]. Graph [338, 406, 472, 537, 797, 917, 940, 1007, 1185, 1270, 1292, 1312, 1329, 1342, 1396, 1507, 1568, 1593, 1777, 1797, 1820, 1822, 1860, 1932, 2023, 2054, 2078, 2115, 2133]. Graph-aware [406]. Graph-based [1007, 1777, 1860, 2054]. graph-theoretic [940]. Graphical [1280]. graphlet [79]. GraphNEI [1770]. GraphPowerNet [1860]. grasshopper [205]. GRE [1400]. Greedy [1400, 1441]. Green [371, 513, 822, 837, 1174, 1305, 1500, 1601, 1718]. Green-PoW [1174]. greenhouses [49]. Grid [82, 332, 584, 590, 656, 1071, 1801]. grids [618, 1624]. grooming [77, 489]. Ground [321, 432, 529, 631, 1015, 1109, 1605, 1855, 2066]. ground-to-air [1605]. Group [218, 268, 352, 373, 487, 783, 889, 1173, 1236, 1248, 1446, 1717]. Group-based [218, 352, 487]. grouping [565, 2010, 2024]. GT [432]. guarantee [96, 650, 1785, 1870].

Guaranteed [385, 455, 475, 551, 1591, 1697, 2008]. guaranteeing [2002]. guarantees [344, 1661]. guard [2051]. Guest [658, 1044, 1109, 1187]. guide [122]. guided [1388]. Guifi.net [62]. Gym [1407, 1464].

H [1376]. H-HOME [1376]. half [134]. half-duplex [134]. hand [993, 2139]. Handling [95, 625, 897, 1043, 2076]. handoff [701, 963]. Handover [208, 220, 365, 408, 464, 731, 1018, 1197, 1291, 1628, 2029, 2089]. handovers [1574]. handshake [1237, 2055]. HAP [2017]. HAP-LEO [2017]. hard [179]. hardening [1047]. hardware [43, 546, 676, 798, 836, 1798, 1848, 2072]. hardware-assisted [1798]. harnessing [1512]. harvesting [38, 40, 134, 275, 367, 707, 958, 960, 1500, 1627, 1629, 1841, 1954, 1986]. Hash [157, 597]. HashXor [485]. hay [354]. HDPoA [1277]. head [13]. healing [1301]. Health [240, 266, 517, 581, 745, 774, 952, 1985]. health-care [240, 517]. HealthBlock [856]. Healthcare [281, 405, 609, 856, 859, 952, 1000, 1092, 1113, 1196, 1512, 1925]. HealthRec [1985]. HealthRec-Chain [1985]. HECC [1238]. HELAD [75]. herd [845]. Het-MEC [1070]. Heterogeneous [30, 66, 75, 105, 163, 217, 242, 289, 306, 313, 320, 323, 340, 350, 365, 378, 400, 420, 455, 467, 478, 621, 787, 806, 833, 961, 1085, 1087, 1132, 1212, 1291, 1323, 1352, 1393, 1397– 1399, 1408, 1481, 1578, 1738, 1757, 1859, 1913, 1924, 1956, 2054, 2124]. HeteroTiC [1352]. HetNets [42, 80, 136, 220, 244, 407, 1276, 1619]. heuristic [687]. Hidden [243, 1011, 1058]. hide [2054]. Hiding [1876]. HIDRA [1419]. Hierarchal [219]. Hierarchical [460, 859, 901, 1116, 1756, 1780, 1890, 2053, 2060, 2122]. hierarchy [1013]. High [6, 7, 410, 462, 494, 511, 626, 641, 699, 718, 746, 759, 792, 794, 919, 1032, 1067, 1124, 1151, 1310, 1445, 1618, 1663, 1683, 1868, 1884, 1953, 2040, 2091]. high-bandwidth [6, 792]. High-Capacity [410]. high-density [1124]. high-performance [919, 1151, 1618]. high-precision [7]. high-rate [759]. high-speed [641, 699, 1663, 1868, 1884, 1953]. high-throughput [1445, 2091]. Higher [421, 697, 1893, 1972]. Higher-Level [421]. higher-order [1972]. highly [61, 1244]. highway [1721]. hijacking [190, 922]. History [653]. hit [864]. HITLinQ [2114]. hoc [9, 125, 146, 465, 534, 617, 668, 1098, 1148, 1492, 1580, 1599, 1620, 1685, 1805, 1827, 1984, 2091]. hole [254]. Holistic [852]. home [162, 1030, 1238, 1376, 1592, 2020]. Honesty [1277]. Honesty-based [1277]. Honey [1648]. Honey-X [1648]. Hop [7, 65, 1019, 1031, 1034, 1068, 1151, 1569, 1660, 1661, 1685, 1794, 1947, 1951, 2013, 2114]. hopping [103, 427, 985, 1181]. Horizontal [1288]. hospital [805]. host [1095, 1465]. host- [1465]. hotspots [1510]. HPC [68]. HSA [943]. HSA-SPC [943]. HSGAN [2060]. HSGAN-IoT [2060]. HSWF [1983]. HTTP [419, 1707]. HTTP-level [419]. HTTP/3 [1707]. HTTPS [1063, 1920]. hub [350]. hub-spoke [350]. hull [1248]. Human [22, 121, 227, 993, 1183, 1241, 1379, 2082]. human-based [227]. human-computer [993]. Human-in-the-loop [1183]. Hybrid [29, 139, 152, 153, 218, 311, 312, 337, 445, 468, 476, 523, 545, 561, 586, 605, 606, 632, 709, 749, 780, 790, 835, 854, 859, 869, 896, 943, 1283, 1556, 1559, 1562, 1648, 1659, 1790, 1957, 2014, 2037, 2045, 2077, 2089, 2114, 2126]. Hyper [340]. hypergraph [92, 1255]. hyperparameter [854]. Hypervisor [892].

I/O [892]. I/Q [391]. I2P [680]. IaaS [61]. IAB [723, 1316]. IBDA [1922]. IBTA [1163]. IBTA-IoV [1163]. ICAN [1793]. ICMP [512]. ICN [514, 1065, 1726]. ICN-IoT [1726]. ICNIRP [508]. ICSDN [1380]. ICT [837]. ICV [1572]. ID [1761]. Identification [22, 117, 277, 333, 907, 910, 911, 1312, 1378, 1596, 1617, 1770, 1819, 1899, 1915, 2048]. identifiers [1642]. identifies [948]. identify [1641]. Identifying [46, 282, 1920]. Identity [485, 726, 842, 1304, 1359, 1832, 1922, 2086]. identity-augmented [842]. Identity-Based [726, 1922]. IDS [625, 1161, 1866]. IEEE [135, 329, 380, 528, 604, 894, 1054, 1086, 1187, 1640, 1657, 2010]. IFIP [1649]. IFMIF [1716]. IFMIF-DONES [1716]. II [2083]. IIoT [590, 591, 634, 1703]. ILETC [1489]. illegal [2054]. ILP [412]. Image [120, 1865]. Image-based [120]. imagery [654]. images [67]. imbalance [625, 1962]. imbalanced [226]. IMCFN [120]. imitation [1333]. Impact [34, 133, 212, 365, 382, 508, 512, 696, 922, 1080, 1088, 1487, 1755, 1918, 1960, 2136, 2138]. impairment [721]. impairment-aware [721]. impairments [2097]. imperfect [515, 1130, 1480, 1761]. Implementability [1713]. Implementation [7, 29, 403, 410, 519, 701, 736, 1453, 1464, 1514, 1699, 1791, 1847, 1961, 1996, 2051]. implementation-agnostic [1961]. Implementing [1308, 1313, 1987]. implications [796, 1357, 2136]. importance [254, 1199, 1262, 1864, 1972]. Improve [99, 394, 853, 1129, 1511, 1806, 2125]. Improved [45, 79, 158, 254, 433, 625, 687, 788, 1048, 1137, 1305, 1325, 1837, 2062, 2083]. improvement [783, 1290, 1713]. improvements [1518]. Improving [255, 280, 375, 539, 753, 820, 888, 904, 1153, 1162, 1355, 1721, 1731, 1892, 2114]. Impulse [121]. IMSI [677]. In-band [495, 1229, 1232, 1470, 1828, 1973, 2098]. In-depth [1357]. in-kernel [1967]. in-kernel/user-space [1967]. In-network [1358, 1562, 1877]. Incentive [114, 170, 1160, 1671, 1756, 1858, 1999, 2056]. incentive-based [1671]. Incentive-driven [1756]. Incentivizing [461]. incident [1406]. incomplete [1201, 1395]. inconsistencies [2136]. incorporating [1916]. Increased [144, 1683]. increasing [1610]. Incremental [1489, 1517, 2121]. Independent [125, 378, 505, 522, 2112]. Index [1803]. Indicators [1666]. indistinguishability [1279]. Indoor [270, 449, 1134, 1854, 1885, 2024]. Indoor/Outdoor [1134]. induced [961, 1962]. induction [153, 358]. Industrial [68, 371, 401, 499, 513, 522, 589, 592, 607, 608, 635, 820, 821, 873, 1135, 1428, 1640, 1838, 1862, 2133]. industries [1320]. industry [1885, 2034, 2096]. inertial [804]. iNet [1464]. infeasibility [2007]. infection [1598]. inference [303, 369, 578, 1182, 1262, 1634, 1658, 1682, 1702, 1704, 1899]. inference-Sector [578]. Influence [1740, 1771]. Information [21, 142, 158, 214, 304, 529, 533, 643, 656, 662, 785, 915, 937, 955, 1023, 1080, 1098, 1129, 1143, 1201, 1225, 1247, 1287, 1353, 1395, 1457, 1782, 1883, 1961, 2102, 2113, 2114]. Information-Centric [533, 937]. information-theoretic [2114]. Informer [1808]. Informer-based [1808]. Infrastructure [85, 205, 464, 519, 794, 851, 879, 1302, 1600, 1640, 2120]. infrastructures [278, 1056]. initial [261, 1197]. initialization [687]. initiation [190]. Initiatives [810]. injection [1754]. innovation [204, 1375, 2033]. innovations [388, 507]. input [662, 733]. inputs [804]. insider [505]. inspired [273, 907, 983, 1761, 1802]. installation [682]. instance [159, 397, 1108, 1155]. Instant [2104]. institution [1963]. INT [1736]. INTaaS [1973]. integral [192]. Integrated [80, 172, 286, 497, 529, 679, 897, 1015, 1018, 1034, 1047, 1252, 1444, 1575, 1654, 1668, 2066]. Integrating [1924]. integration [195, 2063]. integrity [351, 848, 924, 1275, 1830]. integrity-preserving [1275]. Intelligence [389, 416, 448, 568, 613, 700, 705, 805, 983, 1271, 1343, 1373, 1378, 1496, 1511, 1587, 1604, 1739, 1942, 2032, 2045]. intelligence-based [389]. Intelligent [93, 318, 357, 389, 449, 475, 479, 480, 488, 520, 561, 732, 859, 872, 936, 1021, 1049, 1288, 1303, 1358, 1409, 1464, 1528, 1534, 1575, 1610, 1629, 1659, 1700, 1703, 1783, 1849, 1939, 1959, 1975, 1993, 1997, 2008, 2022, 2046]. Intent [1120, 1363, 1411, 1535]. Intent-based [1120, 1535]. Intent-driven [1411]. intents [1203]. inter [360, 452, 465, 732, 939, 1015, 1201, 1785, 1975, 2131]. inter-BBU [360]. inter-cluster [465]. inter-constellation [1015]. inter-datacenter [939, 1201, 1785]. inter-domain [1975]. inter-flow [2131]. inter-numerology [732]. inter-server [452]. interaction [1581]. interactions [862]. interactive [1118]. interactivity [1118]. interdomain [86, 750]. interest [5, 1163]. interests [411, 514]. interface [166, 904, 1524, 1661]. interfaces [993]. Interference [34, 414, 725, 732, 773, 1678, 1943, 1954, 2029]. interference-based [2029]. interfering [135]. Interlayer [932]. interleaved [1826]. intermittent [2110]. Internal [43, 1314]. Internet [21, 154, 161– 163, 167, 182, 185, 188, 219, 247, 258, 262, 274, 288, 295, 302, 382, 418, 425, 431, 432, 444, 458, 470, 480, 491, 496, 513, 515, 522, 568, 576, 592, 596, 635, 652, 756, 776, 807, 825, 852, 863, 873, 882, 884, 905, 933, 937, 940, 941, 945, 957, 967, 1012, 1017, 1036, 1135, 1159, 1160, 1168, 1262, 1293, 1314, 1319, 1331, 1342, 1394, 1424, 1427, 1471, 1511, 1545, 1582, 1587, 1672, 1683, 1710, 1719, 1760, 1782, 1783, 1799, 1802, 1821, 1837, 1839, 1857, 1862, 1895, 1910, 1917, 1925, 1940, 1962, 1965, 1993, 2048, 2083, 2096, 2117, 2133]. Internet-of-Things [247, 496, 1910]. Internet-scale [933]. Internet-wide [940]. interoperability [875]. interoperable [1482, 1763]. interpretability [727]. interrupt [997]. interval [343]. interworking [873]. IntOpt [1229]. Intra [465, 1577, 2131]. Intra- [465, 2131]. Intra-body [1577]. Introducing [1938]. Introduction [1412, 1587, 1649]. Intrusion [55, 137, 146, 176, 226, 283, 312, 441, 446, 469, 491, 540, 614, 625, 675, 715, 754, 772, 901, 916, 941, 983, 1084, 1172, 1199, 1204, 1287, 1529, 1531, 1594, 1689, 1783, 1812, 1821, 1839, 1853, 1866, 1938, 1939, 2090]. intrusions [1673]. invariant [146]. inverse [1867]. Inversion [1960]. investigation [794]. INVISIBLE [2029]. invocation [1919]. IoD [596, 710, 1917]. IoMT [1165]. IoT [48, 69, 112, 147, 183, 184, 211, 215, 238, 240, 251, 259, 281, 292, 425, 430, 449, 458, 485, 499, 505, 530, 545, 581, 607, 610, 629, 636, 682, 753, 802, 821, 822, 828, 837, 898, 903, 928, 935, 987, 996, 1019, 1021, 1028, 1042, 1067, 1134, 1153, 1154, 1166, 1189, 1268, 1277, 1283, 1299, 1374, 1376, 1398, 1429, 1455, 1476, 1477, 1482, 1520, 1524, 1525, 1539, 1549, 1571, 1585, 1613, 1614, 1633, 1673, 1679, 1689, 1726, 1763, 1779, 1787, 1821, 1832, 1838, 1841, 1852, 1870, 1924, 1944, 1948, 1979, 2010, 2057, 2060, 2093, 2109]. IoT [18, 71, 261, 312, 351, 352, 371, 385, 487, 497, 538, 589, 608, 660, 661, 668, 767, 820, 829, 836, 859, 897, 909, 968, 1008, 1089, 1123, 1143, 1184, 1194, 1204, 1237, 1260, 1322, 1329, 1343, 1369, 1416, 1493, 1531, 1538, 1594, 1719, 1731, 1738, 1796, 1834, 1866, 1875, 1908, 1934, 1956, 2041, 2087]. IoT-based [48, 281, 449, 581, 829, 987, 996, 1299]. IoT-blockchain [1277]. IoT-ecosystem [897]. IoT-enabled [71]. IoT-FoG [1633, 1908]. IoT-Gov [1719]. IoT-networks [487]. IoT-oriented [589]. IOTA [821, 2105]. IoTs [719, 920, 1038]. IoV [639, 1163, 1308]. IoVs [597]. IP [364, 457, 614, 938, 1259, 1770, 1953]. IP-network [1259]. IP/MPLS [457]. IPFS [1985]. IPro [93]. IPsec [1723]. IPv4 [1427]. IPv4/IPv6 [1427]. IPv6 [940, 1427, 1789]. iRED [2098]. IRIS [1540]. Irregular [820, 1744]. irrigation [49]. IRS [1814]. IRS-assisted [1814]. IRSA [1624]. ISCC [1187]. Island [1533]. isolated [1036]. isolation [126, 892, 1314, 2127]. ISP [122, 1479]. issue [658, 1587, 1649]. issues [81, 282, 585, 1022, 1463, 1513]. item [354]. Iterative [1626, 1735]. IXPs [369].

Jailhouse [2127]. Jain [1785]. jammer [729]. Jamming [427, 505, 1054, 1181, 1269, 1743, 1955, 2016]. January [1, 423, 913, 1381, 1906]. JANUS [2100]. JANUS-based [2100]. job [522, 588]. job-shop [588]. jobs [429, 1682]. joining [2057]. Joint [13, 15, 63, 193, 244, 353, 366, 374, 394, 413, 426, 573, 601, 619, 664, 724, 747, 757, 800, 826, 857, 888, 917, 1034, 1038, 1070, 1073, 1100, 1114, 1118, 1159, 1175, 1231, 1272, 1328, 1339, 1393, 1456, 1490, 1526, 1538, 1583, 1622, 1631, 1654, 1669, 1729, 1738, 1753, 1833, 1853, 1871, 1881, 1888, 1904, 1913, 1914, 1935, 1958, 2008, 2024, 2047]. Judging [989]. July [186, 202, 637, 666, 1093, 1111, 1636, 1651]. June [168, 615, 1077, 1589, 1615]. JurisNN [989].

Ka-band [1831]. kangaroo [446]. kangaroo-based [446]. Keeping [49]. Kernel [476, 612]. Kernel-level [612]. Kernel/eBPF [476]. kernel/user [1967]. Key [188, 291, 292, 328, 373, 405, 726, 765, 786, 824, 1018, 1144, 1236, 1238, 1325, 1492, 1666, 1818, 2027, 2028]. key-agreement [405]. keypads [2112]. keys [2027]. KeySFC [37]. keystroke [2112]. keyword [1851]. KFTO [691]. kicking [41]. Knacks [1556]. knowing [94]. knowledge [212, 865, 1820, 1834, 1836]. known [140]. Krill [845]. Kuhn [691].

L [431, 1107]. L-band [1107]. L-RPL [431]. L2 [2048]. L2-BiTCN-CNN [2048]. LA-Trickle [722]. label [737, 1142, 1687]. labeled [1043, 1565]. labelled [994]. Lagrange [31]. LAKA [1492]. LAKA-UAV [1492]. LAN [734]. language [809]. LANs [35]. LAPCHS [517]. Laplacian [431]. Large [12, 60, 68, 276, 382, 440, 489, 753, 832, 887, 1010, 1336, 1432, 1453, 1506, 1562, 1878, 2008, 2120]. large-scale [68, 276, 382, 753, 1010, 1336, 1432, 1506, 1562, 1878, 2008]. laser [1406]. lasting [1402]. latencies [347]. Latency [36, 85, 278, 344, 353, 382, 428, 453, 551, 567, 647, 766, 933, 971, 997, 1004, 1013, 1029, 1056, 1169, 1228, 1249, 1265, 1445, 1487, 1570, 1745, 1806, 1824, 1964, 1966, 2008]. latency-aware [278, 933, 1056]. latency-critical [1169]. Latency-energy [353]. Latency-Guaranteed [551]. latency-sensitive [1265]. Latin [773]. lattice [1173, 1283]. lattice-based [1173]. Layer [16, 99, 110, 368, 458, 600, 601, 617, 670, 781, 883, 1031, 1066, 1120, 1388, 1455, 1540, 1841, 1854, 1891, 2032, 2070, 2102]. layered [283, 795, 1420, 1748, 1952]. layering [603]. layers [904]. LDA [1476]. LDA-2IoT [1476]. Leaderless [1356, 1451]. leak [1099]. Leakage [158, 726, 1234]. Leakage-Resilient [726]. learned [809, 918]. Learning [55, 163, 290, 292, 330, 357, 400, 439, 444, 445, 447, 491, 537, 550, 569, 579, 609, 635, 636, 659, 688, 700, 707, 744, 772, 774, 790, 827, 830, 835, 854, 878, 883, 912, 916, 920, 929, 941, 967, 969, 1037, 1058, 1073, 1124, 1136, 1200, 1219, 1235, 1253, 1331, 1335, 1338, 1341, 1376, 1377, 1420, 1447, 1464, 1477, 1482, 1491, 1495, 1496, 1503, 1541, 1550, 1587, 1614, 1628, 1650, 1668, 1702, 1730, 1759, 1766, 1786, 1793, 1821, 1839, 1874, 1882, 1887, 1930, 1944, 1964, 1981, 2013, 2014, 2018, 2056, 2065, 2072, 2074, 2077, 2090, 2098, 2101, 2109, 2116, 2121, 2133]. learning [6, 8, 231, 280, 317, 329, 346, 368, 379, 387, 394, 406, 470, 492, 526, 540, 628, 654, 715, 757, 775, 776, 798, 819, 839, 872, 896, 908, 924, 968, 998, 1004, 1020, 1029, 1041, 1050, 1062, 1064, 1076, 1089, 1126, 1154, 1172, 1193, 1217, 1226, 1233, 1273, 1303, 1332, 1333, 1339, 1342, 1346, 1367, 1370, 1372, 1375, 1388, 1401, 1414, 1460, 1471, 1493, 1500, 1501, 1547, 1548, 1558, 1562, 1583, 1598, 1627, 1629, 1660, 1664, 1677, 1703, 1713, 1718, 1720, 1733, 1751, 1754, 1756, 1774, 1780, 1784, 1789, 1835, 1867, 1877, 1891, 1902, 1949, 1973, 2011, 2021, 2049, 2063, 2091, 2135]. learning [42, 63, 75, 156, 207, 209, 253, 478, 532, 588, 608, 660, 661, 723, 727, 737, 788, 844, 962, 966, 972, 978, 1024, 1033, 1081, 1135, 1202, 1213, 1265, 1281, 1399, 1408, 1475, 1489, 1527, 1578, 1580, 1602, 1622, 1659, 1707, 1732, 1790, 1796, 1812, 1823, 1862, 1957, 1980, 1999, 2006, 2012, 2037, 2064, 2126, 2132, 2136]. Learning-aided [1124]. Learning-based [357, 368, 387, 609, 636, 661, 790, 1341, 1375, 1388, 1471, 1527, 1548, 1627– 1629, 1751, 1789, 1796, 1930, 2126, 2133]. learning-driven [8, 207, 491]. learning-powered [2072]. learnt [1839]. leasing [237]. least [45]. least-squares [45]. LEDBAT [1119, 1365]. ledger [659]. legs [1935]. length [34]. LENS [1317]. LEO [151, 1110, 1186, 1533, 1842, 2017]. LEO-based [1842]. Less [94, 302, 1365, 1699, 1893]. less-than-best-effort [1699]. lessons [809, 1839]. Let [2054]. Level [191, 348, 350, 419, 421, 466, 590, 612, 791, 878, 929, 1270, 1309, 1377, 1428, 1465, 1476, 1575, 1618, 1689, 1777, 1818, 1909, 1932, 1938, 2014, 2049]. Leveraging [486, 522, 956, 1942]. Licensed [301]. LiDAR [1440]. Life [219, 1064, 1459]. lifecycle [1832]. lifetime [180, 1071]. Light [834, 1055, 1406, 1447, 1618, 1690, 1764, 2029]. light-weight [834, 1447]. lighting [449]. LightLog [944]. lightning [765, 2052]. Lightpath [1886]. lightpaths [1136]. LightPoW [1394]. Lightweight [29, 102, 240, 256, 285, 351, 485, 517, 657, 726, 828, 846, 848, 944, 949, 998, 1237, 1283, 1296, 1314, 1392, 1465, 1492, 1504, 1534, 1544, 1595, 1781, 1834, 1875, 1903, 1917, 1967, 2123]. like [277]. LiKey [2112]. LiM [219]. LiM-AHP-G-C [219]. Limitations [58, 1801]. limited [91, 726]. limits [508, 1617]. line [1609]. linear [46, 556, 1240, 1403, 1528]. linguistics [435]. linguistics-based [435]. Link [23, 32, 329, 431, 465, 652, 676, 679, 860, 932, 1339, 1350, 1678, 1815, 1864, 1898, 1966, 1982, 2006, 2114]. link-Configuration [329]. link-state [1350]. linkable [1850]. links [445, 961, 1621, 1974, 1989, 2081]. Linux [437, 612]. LIO [625]. LIO-IDS [625]. lion [761]. list [150]. literature [112, 983, 1204, 1411]. Live [255, 772, 867, 1074, 1415, 1626, 1806, 1905]. liveness [2030]. livestock [68]. Load [77, 105, 231, 249, 285, 360, 425, 427, 490, 545, 559, 599, 687, 791, 898, 981, 1039, 1080, 1180, 1184, 1209, 1403, 1526, 1579, 1619, 1700, 1732, 1882, 1950, 1962]. Load-aware [427, 1180, 1732]. Load-balanced [559, 1700]. Load-balancing [1039, 1526]. LOADER [436]. loads [1403, 1777]. Local [122, 308, 436, 664, 750, 1549, 1846]. localisation [322, 1994]. locality [884, 956]. Localization [45, 56, 117, 212, 289, 294, 817, 1082, 1134, 1355, 1539, 1559, 1794]. Locating [578, 852]. Location [65, 119, 350, 409, 604, 621, 652, 729, 806, 851, 1065, 1158, 1234, 1328, 1390, 1488, 1978, 2053, 2112]. location-authentication [119]. location-aware [65]. location-based [1158]. Location-independent [2112]. location-related [1065]. log [944]. Logchain [198]. Logging [614]. Logic [492, 731, 1593]. logic-based [1593]. logistic [1837]. logs [198, 419]. Long [645, 1402, 1478, 1540, 1687, 1756]. Long-lasting [1402]. long-range [1540]. long-term [1478, 1756]. Longest [1243]. Longitudinal [98]. lookup [354, 1926]. loop [1183, 1427]. loops [1427, 1455]. LoRa [1410, 1429, 1878, 1996]. LoRaWAN [383, 486, 925, 1224, 1243, 1432, 1526, 1681, 1804, 2045]. LoRaWAN-based [1681]. loss [467, 645]. Losslessly [1351]. lossy [285, 699]. Low [36, 123, 285, 321, 343, 453, 494, 511, 647, 716, 718, 719, 748, 759, 766, 879, 992, 1008, 1067, 1106, 1249, 1296, 1310, 1445, 1470, 1518, 1540, 1613, 1645, 1692, 1745, 1806, 1824, 1855, 1878, 1966, 1974, 1994, 2102]. low-altitude [1855]. Low-cost [716, 1296, 1613]. Low-delay [759, 1008, 1518, 1974, 2102]. low-infrastructure [879]. low-latency [766, 1249, 1445, 1806, 1966]. low-power [285, 321, 1878]. lowest [1557]. LPB [1898]. LPWAN [292, 378]. LPWANs [493]. LR [137, 674]. LR-PON [674]. LRS PKI [1850]. LRU [1976]. LRU-based [1976]. LSA [46]. LSTM [386, 625, 1157, 1777, 1959]. LSTM-based [1157, 1777]. LTE [47, 87, 155, 243, 296, 407, 559, 787, 904, 1058, 1606, 1988]. LTE-A [87]. LTE-U [787]. LTE-WiFi [47]. LTE/5G [1606]. LTE/WLAN [407]. LWSBFT [1356]. Lyapunov [975].

M [340]. M3F [1576]. MA [598]. MA-ABE [598]. MAC [110, 125, 780, 903, 994, 1086, 1324, 1351, 1683]. MAC-layer [110]. Machine [8, 163, 179, 330, 346, 357, 387, 394, 395, 444, 445, 447, 492, 526, 532, 540, 569, 609, 628, 688, 715, 730, 772, 835, 854, 916, 1033, 1041, 1058, 1064, 1073, 1074, 1172, 1226, 1299, 1303, 1335, 1447, 1503, 1587, 1598, 1622, 1650, 1660, 1702, 1703, 1718, 1887, 1902, 1944, 2011, 2072]. Machine-Learning [1335, 1598]. Machine-type [179]. machines [1933]. macro [580]. macro-femto [580]. MADDPG [2026]. made [282]. MADRLOM [2137]. MAESP [392]. MAGD [1822]. Magnetic [153, 358, 1273]. Maintaining [622, 624]. major [380]. Making [552, 598, 1561, 1597, 2033]. MalCon [1710]. Malicious [277, 751, 896, 929, 967, 1242, 1349, 1751, 1762, 1771, 1835, 2057]. malware [39, 74, 120, 570, 798, 968, 1012, 1286, 1576, 1710, 1798]. Management [9, 16, 18, 41, 88, 94, 118, 167, 220, 258, 290, 310, 318, 355, 361, 373, 378, 483, 510, 567, 591, 598, 630, 635, 640, 655, 656, 669, 731, 765, 808, 824, 832, 856, 862, 872, 873, 905, 945, 1064, 1113, 1144, 1186, 1236, 1302, 1357, 1363, 1411, 1530, 1546, 1638, 1746, 1776, 1818, 1832, 1857, 1862, 1869, 1888, 1924, 1943, 2011, 2024]. manager [61]. Managers [1302]. Managing [670]. MANET [1922]. manifolds [1266]. manipulators [1568]. MANO [500]. manufacturing [588, 589]. Many [189, 345, 1211]. many-field [345]. many-objective [1211]. Many-to-many [189]. map [1148, 1208]. map-aware [1148]. Mapping [917, 1209, 1261]. March [72, 502, 973, 1466, 1968]. Marconi [1464]. margin [1180]. maritime [564, 2093]. mark [760]. market [551]. market-oriented [551]. marketplace [977, 1366]. markets [1462]. marking [364, 760]. Markov [313, 364, 396, 1011, 2017]. Markovian [80, 1361]. Mars [1145]. MARVEL [231, 2109]. Massive [87, 218, 261, 391, 444, 898, 908, 1121, 1231, 1775]. Massive-Scale [391, 1121]. match [2033]. match-action [2033]. Matching [54, 189, 407, 1255, 1261, 1312, 1694]. matchings [1947]. Matchmaker [624]. MATEC [846]. materialized [936]. Mathematical [304, 396, 705]. MATILDA [670]. Matrix [887, 2019]. matter [2052]. max [1609]. maximal [2039]. Maximization [47, 215, 553, 702, 864, 1029, 1071, 1194, 1284, 1386, 1740, 1986, 2088]. maximize [1087, 1526, 1792]. Maximizing [74, 219, 587, 1190, 1787, 1861]. Maximum [129, 1441]. May [127, 148, 571, 593, 1045, 1059, 1551, 1566, 1764, 1978, 2079, 2106]. MAYA [1452]. MC [1029]. MC-NOMA-based [1029]. MCEN [1441]. MCS [1339]. MDCHD [798]. MDFD [1515]. MDP [1209]. measure [79, 159, 969].

Measurement [453, 634, 681, 683, 908, 1146, 1427, 1564, 1736, 1847, 2099, 2138]. measurement-based [2099]. Measurement-driven [681]. measurements [139, 225, 729, 1191, 1807, 2069]. Measuring [508, 1131, 1666, 1962]. MEC [8, 141, 147, 195, 306, 533, 844, 1070, 1129, 1217, 1310, 1362, 1490, 1497, 1601, 1677, 1703, 1741, 1958, 2094]. MEC-assisted [2094]. MEC-based [306, 1129]. MEC-enabled [8, 533, 1310]. MEC-IIoT [1703]. Mechanism [76, 114, 199, 224, 287, 301, 404, 475, 545, 566, 682, 692, 754, 802, 841, 915, 927, 1143, 1160, 1293, 1305, 1315, 1322, 1330, 1395, 1493, 1558, 1653, 1824, 1910, 1914, 1999, 2056, 2137]. mechanisms [170, 283, 1282, 1462, 1749, 1853]. Media [378, 1372]. Medical [351, 876, 970, 1105, 1113, 1874, 2117]. medium [833, 920]. Meet [1672, 1971]. meets [42, 134, 628, 872, 1235]. mega [1909]. MEH [1958]. membership [971]. MEMENTO [2121]. memetic [232]. Memory [645, 718, 755, 792, 1074]. MemOry-based [755]. Mesh [692, 1096, 1164, 1989]. message [83, 101, 845, 1156]. messages [284, 393, 1827]. messaging [1343]. meta [1602, 2012]. meta-learning [1602]. Metaheuristic [55, 697, 739, 749, 1011, 1367, 1823]. metaheuristic-based [697]. Method [152, 323, 372, 401, 463, 528, 592, 646, 662, 721, 727, 751, 785, 798, 821, 848, 891, 1050, 1227, 1232, 1299, 1320, 1458, 1495, 1597, 1608, 1625, 1669, 1705, 1731, 1770, 1781, 1786, 1808, 1816, 1818, 1852, 1903, 1923, 1959, 1963, 2021, 2023, 2090, 2116]. Methodology [681, 768, 1248, 1987]. Methods [369, 389, 447, 540, 556, 1612, 1826, 1940]. metric [1251, 1720]. metrics [564, 618, 2077]. metropolitan [2113]. metropolitan-wide [2113]. MFFusion [929]. MGKM [1236].

MGM [1193]. MGM-4-FL [1193]. MICN [514]. MicroCM [1919]. microservice [1873, 1919, 1932]. microservice-based [1873]. microservices [521, 2025]. microwave [32, 1378]. middlebox [649]. middleboxes [484]. MIF [785]. Migration [333, 360, 406, 671, 981, 1050, 1074, 1184, 1458, 1478, 1655, 2000, 2037, 2047]. migrations [942, 1220]. military [1534]. Millimeter [15, 780, 782, 961, 982, 1069, 1742, 2031]. millimeter-wave [15, 780, 782, 961, 1742]. MIMO [153, 313, 370, 515, 1231, 1260, 1601, 1775]. MIMO-NOMA [1260]. Min [1285, 1609]. min-max [1609]. mined [563]. mini [696]. mini-slots [696]. Minimal [960, 1822]. minimisation [1000]. minimization [289, 639, 732, 960, 1870, 1981, 2037, 2119]. minimize [445, 541]. minimized [189, 339]. Minimizing [285, 335, 860, 958, 1035, 1569, 1670, 1729]. Minimum [536, 651, 1179, 1210, 1350, 1661, 1735]. Minimum-cost [1735]. Mining [39, 225, 940, 1711, 2019]. mismatch [1983]. MISO [1814]. mistrusting [287]. mitigate [1678]. Mitigating [997, 1918]. mitigation [24, 76, 512, 611, 773, 818, 922, 1295, 1298, 1314, 1539, 1793]. mix [1843]. mix-zones [1843]. mixed [47, 1871, 2050]. mixed-timescale [1871]. mixture [226]. MJOA [1658]. MJOA-MU [1658]. ML [1090, 1097, 1407, 1520, 1813, 1900, 1998]. ML-based [1520, 1998]. MMT [462]. MMT-based [462]. mMTC [379, 1004]. mmWave [201, 413, 723, 804, 990, 1523, 1659]. mmWaves [145]. MobilE [9, 13, 94, 104, 114, 197, 206, 225, 267, 353, 359, 386, 387, 397, 402, 459, 485, 488, 516, 553, 577, 581, 586, 587, 639, 642, 699–701, 708, 724, 757, 764, 807, 817, 823, 826, 878, 882, 900, 903, 908, 923, 958, 961, 963, 977, 978, 984, 1013, 1016, 1035, 1042, 1062, 1129, 1141, 1190, 1211, 1213, 1248, 1270, 1328, 1357, 1359, 1377, 1384, 1386, 1420, 1423, 1430, 1434, 1446, 1454, 1473, 1474, 1481, 1485, 1497, 1498, 1512, 1543, 1544, 1558, 1560, 1561, 1596, 1606, 1625, 1628, 1709, 1722, 1737, 1747, 1753, 1807, 1858, 1860, 1878, 1911, 1949, 2054, 2067, 2070, 2110, 2116]. mobile [129, 143, 170, 207, 260, 366, 392, 439, 440, 602, 800, 864, 1088, 1108, 1337, 1425, 1435, 1472, 1609, 1655, 1712, 1732, 1765, 2012, 2037]. mobile-app [878, 1270, 1596]. Mobile-edge [639]. Mobility [88, 180, 227, 244, 281, 361, 365, 378, 392, 493, 533, 607, 886, 937, 1067, 1103, 1387, 1429, 1560, 1655, 1745, 1776, 1797, 1829]. Mobility-aware [281, 1655, 1829]. mobility-compliant [937]. Mode [310, 906, 1090, 1164, 1490, 1776, 2010]. Model [23, 41, 75, 80, 135, 155, 182, 226, 227, 252, 296, 340, 365, 467, 469, 487, 569, 576, 659, 694, 725, 827, 828, 854, 868, 916, 929, 1011, 1068, 1128, 1193, 1272, 1322, 1366, 1374, 1426, 1435, 1447, 1453, 1503, 1543, 1549, 1556, 1580, 1581, 1621, 1640, 1658, 1691, 1748, 1750, 1765, 1775, 1788, 1797, 1819, 1823, 1854, 1886, 1898, 1905, 1956, 1957, 1961, 1976, 2002, 2048, 2051, 2118, 2122, 2131, 2136]. model-free [1503, 1580]. Modeling [14, 166, 338, 364, 676, 764, 901, 1047, 1353, 1579, 1679, 1684, 1853, 1963, 1998, 2015, 2017, 2042]. Modelling [123, 331, 684, 957, 1280]. Models [101, 220, 235, 357, 509, 539, 564, 908, 1051, 1136, 1385, 1452, 1496, 1887, 2081]. modes [857]. Modified [717, 1011, 1295]. modular [67, 1791]. Modulation [77, 839, 1309, 1436, 1510, 1555]. Modulation-Level [1309]. module [1313, 2082]. molecular [1684]. Monero [980]. Monero-based [980]. monetary [255]. monitor [1229]. Monitoring [33, 93, 215, 228, 321, 412, 476, 606, 664, 679, 716, 795, 859, 923, 1072, 1221, 1268, 1546, 1681, 1919]. monitoring-cycle [412]. Monkey [1235]. Monte [1924]. months [1487]. morphological [1084]. MOVES [755]. moving [478, 498]. MP [641].

MP-CREDIT [641]. MPLS [457, 606].

MQTT [767, 1514, 1673, 1866].

MQTT-enabled [1866]. MR [1464]. MR-iNet [1464]. MRAI [1965]. MRAI-free [1965]. MSCCS [980]. MSYM [57]. MT [1309]. MTC [218]. MU [370, 1601, 1658]. MU-MIMO [370, 1601]. MuHoW [2035]. MUlti [7, 19, 31, 53, 65, 79, 103, 200, 229, 256, 276, 283, 340, 344, 347, 350, 365, 376, 407, 438, 463, 464, 636, 641, 707, 729, 781, 795, 803, 808, 816, 839, 929, 1019, 1031, 1034, 1047, 1052, 1125, 1130, 1169, 1224, 1241, 1255, 1268, 1278, 1328, 1351, 1405, 1420, 1421, 1428, 1432, 1440, 1454, 1464, 1485, 1491, 1505, 1507, 1515, 1524, 1534, 1537, 1576, 1582, 1584, 1591, 1597, 1621, 1660, 1661, 1668, 1671, 1685, 1686, 1689, 1702, 1722, 1777, 1778, 1786, 1787, 1791, 1818, 1829, 1870, 1872, 1874, 1875, 1885, 1889, 1896, 1937, 1947, 1956, 2002, 2013, 2014, 2022, 2048, 2082, 2114, 2120]. multi [18, 56, 70, 74, 115, 118, 192, 207, 224, 231, 233, 248, 260, 299, 318, 429, 456, 500, 552, 560, 621, 662, 685, 687, 698, 724, 732, 785, 805, 824, 844, 870, 899, 911, 964, 972, 1029, 1062, 1127, 1142, 1151, 1185, 1217, 1236, 1262, 1267, 1281, 1290, 1294, 1327, 1350, 1388, 1408, 1430, 1569, 1575, 1602, 1605, 1619, 1641, 1677, 1704, 1738, 1741, 1750, 1814, 1851, 1958, 1963, 1966, 1987, 2012, 2021, 2037, 2129]. Multi-access [560, 724, 972, 1029, 1584, 1591, 1987]. Multi-agent [231, 318, 707, 808, 1062, 1281, 1408, 1432, 1464, 1668, 2012, 2021]. Multi-ant [1872]. multi-antenna [805]. multi-AP [1937]. multi-armed [299, 1641, 2013]. Multi-authority [698, 1671, 1686]. multi-band [65]. multi-beam [1294, 1405]. multi-blockchain [1750]. multi-branch [732]. multi-cell [19, 207]. multi-centrality [795]. multi-channel [463, 552]. multi-channeled [1420]. multi-class [70]. multi-classification [2048]. multi-cloud [870, 1169]. Multi-criteria [464]. Multi-dimensional [1875, 2082].

Multi-domain [248, 256, 824, 1125, 1327, 1597, 2014]. multi-domains [500]. multi-edge [870]. multi-factor [1534]. Multi-failure [1052]. multi-frequency [1702]. multi-gNodeB [1885]. multi-group [1236]. Multi-hop [7, 65, 1019, 1031, 1034, 1151, 1569, 1660, 1661, 1685, 1947, 2013, 2114]. multi-information [785]. multi-input [662]. multi-institution [1963]. Multi-interface [1524, 1661]. multi-jammer [729]. multi-keyword [1851]. multi-label [1142]. multi-layer [781, 1388]. Multi-layered [283]. Multi-level [929, 1428, 1575, 1689, 1777, 1818]. multi-link [1966]. Multi-MEC [1741]. multi-node [2120]. Multi-objective [31, 79, 276, 350, 438, 456, 621, 687, 1047, 1241, 1267, 1787, 1896, 1956, 1958, 2129]. Multi-observable [376]. multi-operator [1224]. multi-optimality [636]. multi-pair [2021]. Multi-path [641, 964, 1127, 1537]. multi-player [1454]. multi-point [347]. multi-protocol [1576]. Multi-provider [1491]. multi-purpose [1268]. multi-quality [347]. multi-radio [103]. Multi-RAT [407, 803, 1619]. multi-region [233]. Multi-resource [1485, 1786, 1889]. multi-RIS [1290]. multi-router [192]. multi-scale [463, 1507]. multi-scanner [74]. Multi-Sensor [340]. MUlti-Service [118, 899, 1421]. multi-session [1576]. multi-smartphone [911]. Multi-source [1515, 2002]. multi-split [1791]. Multi-stage [429, 1704, 1722, 1778]. multi-step [785]. multi-task [839, 1262, 1602]. multi-tasks [1677]. Multi-tenancy [229]. multi-tenant [56, 344, 1255]. multi-tier [224, 365, 816, 1278]. Multi-timescale [795]. multi-topology [1350]. multi-traffic [18]. Multi-UAV [844, 1621]. multi-UAV-assisted [1870]. multi-UAV-enabled [1328]. Multi-User [200, 260, 685, 1130, 1290, 1351, 1430, 1505, 1605, 1738, 1814, 1829, 1851, 2022, 2037]. multi-users [1217]. multi-vendor [500]. multi-view [1185]. multibiometric [327]. Multicast [200, 713, 1100, 1198, 1294, 1309, 1583, 1896, 2002, 2049]. multicasting [1051, 1151]. multicell [1311]. multichannel [57]. multicolor [2024]. multidimensional [1293]. multiform [1452]. multigraph [303]. Multihop [1162, 1316, 1916]. multilayer [457, 1740]. Multimedia [159, 253, 359, 397, 708, 962, 1148, 1176, 1588]. multimetric [1148]. Multimodal [1377, 1549, 1611, 1625, 1712, 2131]. multiparameter [1391]. Multipath [26, 77, 165, 339, 425, 459, 462, 513, 541, 601, 673, 764, 796, 832, 1032, 1212, 1254, 1415, 1680, 1859, 1974, 2005, 2092, 2102]. Multiple [60, 86, 131, 275, 290, 335, 383, 395, 408, 462, 514, 558, 624, 744, 756, 791, 891, 932, 1036, 1117, 1240, 1395, 1426, 1500, 1506, 1816, 1878, 1954, 2022, 2065]. Multiple-level [791]. multiplexed [178, 339, 2092]. multiplexing [1057, 1315, 1390, 1701]. multiservice [434, 684]. Multisite [2101]. multisource [631]. multistory [822]. multitask [878, 1549]. Multiuser [943, 1744]. multivariate [2077]. Munkres [691]. MuSLi [1440]. MUTAA [1326]. Mutual [726, 834, 1327, 1504, 1646].

Nakagami [1904]. Nakagami- [1904]. name [471, 537, 547, 607]. name-based [607]. Named [25, 158, 173, 433, 736, 832, 1163, 1772, 1776, 1793, 1892]. named-data [1772]. names [435]. nano [181]. nano-networks [181]. Nanonetwork [1880]. Nanonetwork-based [1880]. nanonetworks [293]. narrow [458, 1319]. Narrowband [161]. Nash [381]. native [927, 1977]. Natural [809]. navigation [270, 488]. NB [458, 822, 837, 1679, 1934]. NB-IoT [458, 822, 1679, 1934]. NBA [607].

NC [174, 1018]. NC-enabled [1018]. NDN [25, 157, 174, 1815]. NDN-based [1815]. Near [207, 250, 358, 1121, 2046]. Near-optimal [207, 1121, 2046]. near-optimally [250]. needles [354]. neighbor [1971]. Neighborhood [254, 618, 1235, 1803]. Neighborhood-Understood [1803]. NemesisGuard [997]. nerve [116]. nervous [1577]. NESEPRIN [1189]. NEST [1792]. nested [39]. NetShield [1877]. NetTiSA [1953]. NetView [316]. Network [16, 37, 44, 55, 62, 98, 130, 174, 205, 252, 292, 307, 323, 333, 355, 483, 485, 495, 522, 529, 624, 627, 645, 647, 668, 670, 671, 675, 697, 716, 735, 762, 763, 769, 772, 828, 846, 861, 877, 892, 907, 942, 943, 965, 996, 999, 1037, 1043, 1079, 1081, 1085, 1145, 1194, 1220, 1229, 1232, 1235, 1241, 1274, 1287, 1288, 1338, 1384, 1385, 1396, 1400, 1420, 1491, 1501, 1544, 1656, 1674, 1689, 1693, 1706, 1746, 1749, 1766, 1807, 1812, 1887, 1898, 1913, 1916, 1930, 1932, 1964, 1973, 2010, 2023, 2038, 2067, 2068, 2077, 2078, 2088, 2094, 2097, 2098]. network [5, 15, 40, 64, 137, 151, 180, 192, 233, 236, 242, 265, 288, 309, 316, 317, 338, 346, 386, 420, 514, 545, 566, 570, 605, 620, 650, 713, 730, 821, 838, 912, 916, 933, 934, 956, 986, 998, 1064, 1066, 1092, 1095, 1110, 1117, 1133, 1162, 1172, 1175, 1188, 1239, 1249, 1250, 1292, 1312, 1334, 1349, 1350, 1352, 1358, 1367, 1379, 1389, 1406, 1411, 1471, 1497, 1532, 1562, 1579, 1581, 1591, 1596, 1600, 1618, 1630, 1669, 1713, 1736, 1759, 1770, 1792, 1797, 1816, 1820, 1835, 1848, 1896, 1901, 1923, 1939, 1953, 1959, 1967, 1975, 1994, 2000, 2001, 2007, 2024, 2057, 2066, 2115, 2124, 2137]. network [14, 56, 59, 75, 87, 120, 143, 155, 160, 201, 215, 226, 269, 296, 334, 411, 429, 431, 453, 456, 461, 512, 526, 621, 626, 643, 646, 683, 694, 738, 765, 780, 805, 844, 879, 894, 899, 900, 908, 909, 932, 944, 979, 989, 1010, 1014, 1041, 1076, 1089, 1101, 1114, 1115, 1143, 1157, 1163, 1184, 1185, 1197, 1203, 1214, 1247, 1259– 1261, 1265, 1266, 1271, 1281, 1346, 1355, 1362, 1363, 1386, 1390, 1431, 1470, 1474, 1500, 1543, 1564, 1570, 1578, 1634, 1643, 1738, 1757, 1758, 1778, 1866, 1877, 1890, 1920, 1936, 1938, 1950, 1952, 1993, 2002, 2036, 2052, 2078, 2103]. network-accelerated [1346]. network-assisted [1674]. network-aware [730]. network-based [143, 879, 916, 1866]. Network-coded [900, 1385, 2088]. networkcoding-based-multicast-wireless-sensor [713]. network-on-chip [934]. network-slicing [522]. network-wide [316]. networked [955, 1074, 1408, 2074]. Networking [25, 76, 158, 167, 173, 295, 433, 453, 455, 618, 630, 686, 736, 832, 951, 1106, 1363, 1364, 1428, 1456, 1607, 1649, 1664, 1755, 1772, 1776, 1793, 1892, 1966, 2033, 2051, 2071]. Networks [49, 79, 88, 97, 133, 139, 140, 211, 218, 263, 283, 289, 296, 300, 310, 327, 342, 361, 367, 390, 401, 438, 457, 462, 496, 497, 533, 535, 547, 553, 559, 580, 617, 686, 721, 722, 731, 738, 748, 764, 807, 858, 883, 887, 903, 967, 1034, 1051, 1058, 1069, 1096, 1098, 1141, 1166, 1176, 1189, 1278, 1292, 1309, 1357, 1378, 1385, 1404, 1421, 1447, 1455, 1464, 1468, 1498, 1502, 1504, 1509, 1588, 1601, 1603, 1620, 1621, 1661, 1668, 1673, 1764, 1828, 1853, 1866, 1870, 1903, 1908, 1911, 1955, 2015, 2035, 2062, 2065, 2070, 2088, 2089]. networks [7, 77, 103, 125, 138, 181, 232, 239, 240, 243, 264, 313, 335, 353, 370, 381, 384, 396, 408, 413, 427, 465, 479, 494, 508, 511, 515, 542, 550, 564, 636, 674, 701, 707, 749, 781, 806, 838, 868, 895, 917, 920, 925, 1031, 1047, 1085, 1097, 1125, 1180, 1219, 1268, 1274, 1275, 1291, 1310, 1311, 1320, 1323, 1432, 1440, 1461, 1507, 1517, 1518, 1525, 1526, 1532, 1535, 1540, 1560, 1568, 1571, 1606, 1614, 1626, 1631, 1650, 1670, 1674, 1685, 1730, 1759, 1813, 1859, 1861, 1874, 1882, 1905, 1918, 1948, 1971, 1989, 2013, 2050, 2061, 2068, 2076, 2092, 2096, 2104, 2114]. networks [10, 13, 23, 34, 38, 66, 83, 95, 101, 106, 124, 126, 171, 197, 217, 273, 329, 339, 349, 350, 360, 365, 383, 523, 536, 566, 582, 607, 655, 693, 699, 706, 713, 756, 770, 775, 782, 803, 857, 888, 898, 939, 963, 1019, 1057, 1082, 1120, 1121, 1129, 1137, 1144, 1148, 1154, 1211, 1215, 1246, 1251, 1270, 1286, 1301, 1315, 1373, 1391, 1393, 1411, 1453, 1460, 1473, 1523, 1529, 1583, 1599, 1610, 1646, 1657, 1660, 1682, 1704, 1709, 1720, 1740, 1745, 1747, 1780, 1784, 1801, 1805, 1818, 1824, 1840, 1846, 1871, 1876, 1888, 1897, 1949, 1961, 1998, 2026, 2029, 2047, 2060, 2095, 2120]. networks [6, 8, 9, 32, 47, 87, 105, 134, 164, 180, 189, 191, 213, 231, 237, 248, 275, 280, 285, 308, 319, 320, 352, 374, 375, 379, 385, 393, 442, 467, 473, 475, 487, 505–507, 510, 541, 562, 563, 583, 586, 623, 633, 642, 644, 665, 668, 673, 714, 761, 771, 790, 793, 800, 804, 833, 869, 906, 1004, 1007, 1036, 1062, 1127, 1147, 1183, 1218, 1226, 1238, 1256, 1303, 1305, 1316, 1329, 1342, 1375, 1397, 1412, 1478, 1500, 1531, 1543, 1569, 1584, 1641, 1733, 1761, 1769, 1833, 1886, 1891, 1943, 1972, 1982, 1984, 2011, 2039, 2058, 2091]. networks [17, 30, 45, 52, 56, 63, 115, 143, 146, 172, 178, 193, 199, 216, 253, 261, 366, 392, 426, 440, 443, 446, 478, 488, 520, 534, 538, 552, 558, 606, 613, 648, 651, 652, 657, 671, 679, 691, 723, 746, 754, 950, 960, 979, 982, 1015, 1024, 1027, 1032, 1054, 1088, 1102, 1152, 1153, 1181, 1210, 1269, 1353, 1390, 1408, 1410, 1435, 1444, 1445, 1457, 1475, 1492, 1506, 1508, 1531, 1537, 1557, 1574, 1580, 1594, 1653, 1654, 1663, 1680, 1681, 1683, 1693, 1698, 1701, 1705, 1732, 1735, 1765, 1790, 1794, 1823, 1827, 1904, 1909, 1912, 1914, 1934, 1986, 2008, 2012, 2073, 2102, 2118, 2140]. Neural [120, 143, 226, 269, 288, 338, 386, 496, 563, 570, 645, 646, 659, 675, 737, 846, 907, 912, 1188, 1249, 1270, 1274, 1286, 1288, 1292, 1396, 1453, 1531, 1634, 1779, 1866, 1932, 1967, 2023, 2115]. Neuro [739]. Neuro-Fuzzy [739]. neutrality [989]. neutrosophic [1780]. Never [825]. Next [409, 519, 606, 659, 661, 807, 821, 1372, 1509, 1614, 1619, 1975, 2068]. next-generation [606, 807, 1372, 1614, 2068]. NextG [1611]. NFMI [358]. NFStream [965]. NFV [44, 52, 141, 194, 195, 256, 283, 320, 344, 375, 406, 472, 543, 645, 770, 857, 917, 1056, 1176, 1302, 1373, 1422, 1462, 1491, 1588, 1607, 2061]. NFV-based [2061]. NFV-enabled [52, 917, 979, 1373, 1607, 1835]. NFV/SDN [256, 375]. NIC [2001]. NIDS [1812]. NLOS [1406, 1539]. NLOS-FSO [1406]. NNSplit [1453]. NNSplit-SØREN [1453]. no [140, 322]. No-GPS [322]. no-scheduling [140]. Node [243, 254, 267, 341, 372, 795, 864, 957, 967, 1068, 1082, 1452, 1532, 1782, 1977, 2120]. node-removal [372]. Nodes [7, 64, 377, 747, 1058, 1608, 2057]. noise [1847]. NOMA [217, 367, 515, 553, 565, 585, 724, 800, 822, 1029, 1069, 1130, 1210, 1233, 1260, 1290, 1385, 1456, 1474, 1480, 1601, 1605, 1691, 1761, 1871, 1912, 2024, 2088]. NOMA-based [724, 800, 1871]. NOMA-enabled [553]. NOMA-FDR [1480]. Non [46, 135, 142, 345, 380, 395, 408, 489, 509, 556, 1039, 1361, 1374, 1506, 1528, 1653, 1668, 1761, 1934, 1998]. non-anchor [1934]. non-backtracking [489]. non-collided [142]. Non-cooperative [1039]. non-deterministic [1653]. non-ID [1761]. non-IEEE [135]. non-IoT [1374]. non-linear [46, 556, 1528]. non-Markovian [1361]. Non-Orthogonal [395, 408, 1506]. Non-parametric [509]. non-SDN [605]. non-standalone [1998]. non-terrestrial [1668]. non-urban [380]. non-wildcard [345]. Nonlinear [759, 1403, 1954, 1986]. nonnegative [887]. nonorthogonal [756, 1954]. nonuniform [250]. NORD [1532]. Normal [1449]. north [2040]. NoSQL [1022]. notice [1766]. notification [808, 1714]. Novel [47, 75, 87, 273, 323, 401, 486, 551, 564, 591, 658, 662, 669, 688, 697, 701, 722, 792, 798, 975, 1066, 1097, 1133, 1158, 1163, 1251, 1293, 1299, 1309, 1397, 1405, 1560, 1563, 1576, 1603, 1641, 1802, 1850, 1853, 1867, 1891, 1966, 1988, 2121]. November [325, 814, 1263, 1767, 1809, 1903]. nowhere [2054]. NPN [1433]. NR [696, 1745, 1746, 1861]. NR-DC [1745].

NR-U [696]. NSGA [2083]. NSGA2 [137]. NSGA2-LR [137]. NTS4PTP [1144]. Nucleo [2041]. NUI [1803]. number [1501]. numeric [2112]. numerology [696, 732, 1861]. NUV [236]. Nyquist [961].

O [536, 892, 1303, 1407, 2089]. O-RAN [536, 1303, 1407]. O-RAN-based [2089]. OAUTH [2086]. Ob [1247]. Ob-EID [1247]. obfuscation [1208]. object [324, 498, 661, 1132]. objective [31, 79, 276, 350, 438, 456, 621, 687, 1047, 1211, 1241, 1267, 1787, 1896, 1956, 1958, 2129]. objectives [319]. observable [376]. observation [1692]. observer [2053]. Obstacle [1247]. obstructing [2050]. occupancy [1181, 2104]. OCDMA [556]. OCDMA-PON [556]. October [271, 297, 742, 778, 1205, 1222, 1727]. ODL [1347]. ODM [1561]. ODM-BCSA [1561]. ODNS [1879]. OFDM [565, 961, 1847]. OFDM-based [1847]. OFDM-NOMA [565]. OFDMA [310, 370, 1323, 1505]. Off [26, 556, 602, 1013, 1422, 1733]. off-path [26]. Off-site [1422]. offline [1359, 2007]. offload [1516, 2122]. offloaded [1572]. Offloading [170, 207, 228, 260, 353, 366, 375, 385, 387, 402, 459, 546, 553, 560, 609, 691, 705, 724, 764, 769, 781, 807, 830, 843, 844, 870, 921, 950, 958, 972, 978, 1035, 1070, 1087, 1089, 1123, 1165, 1178, 1183, 1190, 1194, 1233, 1340, 1388, 1399, 1430, 1468, 1471, 1472, 1477, 1490, 1538, 1561, 1669, 1677, 1709, 1730, 1747, 1752, 1753, 1796, 1826, 1908, 1911, 1936, 1987, 2026, 2066, 2110, 2124, 2129, 2137]. offloading-efficiency [553]. offloading/sharing [228]. offs [276, 982]. Offspeeding [415]. olive [1979]. on-chip [1989]. On-demand [26, 316, 339, 1736]. on-field [1847]. on-NIC [2001]. onboard [792]. one [625, 1819]. one-class [1819]. one-vs-one [625]. Online [78, 91, 237, 250, 290, 329, 346, 419, 648, 668, 671, 697, 727, 846, 915, 962, 979, 1041, 1080, 1084, 1233, 1326, 1353, 1432, 1475, 1641, 1722, 1890, 1914, 1930, 1970, 1973, 1981, 2007, 2049, 2084, 2085, 2094, 2129]. only [1667]. ONOS [1347]. Open [166, 300, 322, 390, 536, 546, 758, 847, 926, 991, 1024, 1055, 1146, 1161, 1167, 1221, 1370, 1374, 1463, 1634, 1742, 1799, 1811, 1839, 1923, 1942, 1996, 2033, 2042, 2067, 2076, 2138]. Open-access [847, 1742]. Open-RAN [1024]. open-set [1811]. open-source [166, 926, 1161, 1996, 2033, 2138]. OpenAirInterface [204]. openflow [43, 1897]. OpenRAN [1407]. operated [53]. operating [577, 813, 1623]. Operation [759, 1164]. Operations [853, 918, 1431, 1880, 2041, 2067]. operator [170, 899, 1224, 1278]. operator-biased [899]. operator-owned [170]. Opportunistic [83, 334, 341, 425, 542, 893, 923, 1123, 1153, 1383, 1409, 1506, 1545]. opportunities [197, 416, 584, 809, 1126, 1371, 1463, 1902, 2032]. opposite [1441]. oppositional [1162]. Opt [545]. Opt-ACM [545]. optic [961]. Optical [77, 124, 133, 178, 239, 339, 342, 457, 506, 541, 721, 748, 749, 771, 794, 869, 939, 1107, 1180, 1202, 1251, 1309, 1315, 1390, 1391, 1535, 1701, 1886, 1896, 2006, 2015, 2039, 2062, 2065, 2092]. optical-acoustic [124]. Optimal [100, 122, 151, 173, 207, 402, 415, 488, 552, 623, 674, 800, 816, 950, 971, 990, 1019, 1121, 1301, 1430, 1548, 1591, 1609, 1689, 1692, 1746, 1792, 1833, 1858, 1908, 2008, 2039, 2040, 2046]. optimal-latency [971]. optimality [636]. optimally [250]. optimisation [626, 1241, 1975]. Optimistic [2085]. Optimization [15, 83, 205, 232, 276, 289, 323, 338, 348, 353, 394, 422, 438, 457, 561, 619, 664, 681, 730, 761, 763, 766, 806, 821, 823, 835, 843, 845, 871, 885, 888, 894, 917, 972, 988, 1047, 1085, 1100, 1107, 1118, 1178, 1197, 1200, 1211, 1220, 1229, 1280, 1292, 1301, 1326, 1328, 1367, 1391, 1397, 1474, 1631, 1669, 1682, 1697, 1756, 1787, 1794, 1814, 1817, 1853, 1870, 1871, 1891, 1913, 1914, 1956, 1961, 2026, 2101, 2110, 2126, 2129, 2140]. Optimization-based [457, 763]. optimization-Machine [835]. Optimized [294, 295, 509, 545, 646, 928, 1013, 1064, 1066, 1166, 1226, 1691, 1738, 1775]. Optimizing [11, 114, 456, 1212, 1216, 1278, 1342, 1498]. Optimum [1403, 1559]. option [1413]. orbit [847, 1692]. orchestrate [1419]. Orchestrating [306]. orchestration [37, 248, 256, 309, 320, 504, 526, 990, 1056, 1114, 1302, 1485, 1975, 2025]. orchestrator [224]. order [1972]. ordering [749]. OrderSketch [902]. organizing [735, 838, 1706, 1712]. oriented [15, 167, 442, 551, 583, 589, 710, 949, 1372, 1817]. Orthogonal [395, 408, 1506, 1510]. OSDL [1202]. OSPF [46]. OSWireless [1876]. other [162]. Outage [66, 407, 1621, 1761]. Outage-aware [407]. Outdoor [267, 789, 1082, 1134]. outlook [1724]. output [662]. outsourced [698]. outsourcing [266, 922, 1548]. over-the-air [257, 1321]. overhead [1470, 2038]. overlapped [848]. overlapping [150]. overlaps [216]. overlay [933, 2074]. overlaying [426]. Overview [274, 493, 584, 672, 1106, 1330, 1612, 1799]. OVS [546]. OVS-CAB [546]. owned [170].

P2P [372, 1179, 1364]. P4 [801, 935, 1040, 1122, 1228, 1229, 1239, 1246, 1573, 1647, 2033, 2098]. P4-AQM [2098]. P4-based [1228, 1246]. P4-programmable [1122]. P4Resilience [1052]. P4TE [1218]. pack [30]. Packet [34, 106, 154, 345, 348, 364, 483, 538, 633, 675, 878, 919, 1019, 1115, 1117, 1195, 1384, 1405, 1530, 1667, 1745, 1754, 1800, 1884, 1921, 1938, 1967, 2123]. packet-based [633]. packet-chaining [1115]. packet-in-packet [106]. Packet-level [878, 1938]. packets [262, 567]. packing [1351]. PADAL [1994]. pair [2021]. pairing [1881, 1922]. pairing-free [1922]. pairwise [1569]. Pandemic [209, 382, 998, 1377, 1512]. papers [1187, 1649]. paradigm [1476]. paradigms [658]. parallel [532, 1444, 1867, 1909, 2125]. parallelism [891]. parallelized [2061]. Parameterized [1790, 1982]. parameters [126, 1384, 1448, 2045]. parametric [509]. PARFAIT [1008]. parking [987, 1804]. Partial [260, 827, 1085, 1128, 1190, 2000]. partially [1003, 1043]. participant [119, 411, 2090]. participant-selection [2090]. Participants [1416]. partition [92]. partitioning [239, 345, 1399, 1634, 1909, 2120]. parts [142]. party [29]. Passive [1617, 1623, 1994, 2065]. Past [734]. Path [26, 111, 154, 162, 181, 201, 303, 431, 467, 641, 664, 964, 1003, 1034, 1127, 1537, 1802, 1835, 1836, 1872, 1881, 1958, 2039, 2114]. paths [86, 488, 1121, 1212]. Patient [1985]. Patient-centric [1985]. pattern [940, 1702, 1712]. patterns [516, 886, 1711]. PAutoBotCatcher [863]. PAveMENT [488]. PAWR [1407, 1742]. payment [130, 1537, 1893, 2052]. PBCNN [675]. PBFT [1105]. PBFT-based [1105]. PC [1569]. PD [1210, 1663]. PD-CPS [1663]. PD-NOMA [1210]. peaks [1949]. PEDDA [1704]. peer [91, 729, 2052]. peer-to-peer [729, 2052]. Penalty [721]. Penalty-method [721]. penetration [1992]. PEP [1929]. PEP-based [1929]. per-flow [1868]. perception [1891]. perfect [515]. Performance [12, 56, 105, 143, 155, 220, 229, 261, 296, 313, 330, 331, 370, 380, 403, 410, 456, 467, 476, 506, 528, 543, 580, 610, 627, 672, 696, 718, 721, 725, 764, 767, 792, 794, 836, 888, 894, 919, 943, 979, 1088, 1129, 1145, 1151, 1290, 1305, 1310, 1319, 1385, 1401, 1412, 1474, 1511, 1518, 1541, 1548, 1600, 1605, 1618, 1656, 1721, 1731, 1749, 1828, 1842, 1863, 1962, 1998, 2022, 2041, 2125]. performance-aware [979]. performance-low [718]. performance-optimal [1548]. performances [1227]. period [1653]. periodic [1715]. permanent [1096]. permissionless [1364, 1608]. permutation [1189]. Persistent [354, 563, 1289, 1425, 1427, 1586, 1884, 2118]. person [1905]. Personalized [411, 1076, 1499, 1766]. personnel [1617]. perspective [85, 210, 303, 371, 387, 603, 962, 1438, 1633, 1992]. Perspectives [629, 1134]. perturbation [324]. pervasive [1271, 1942]. PETNet [1920]. PF [1922]. PF-IBDA [1922]. PGCE [1128]. pharmaceutical [351]. phase [242, 291, 470, 701, 2057]. Phishing [269, 890, 1020, 1335, 2130]. PhishNot [1335]. phone [1860]. PHP [2133]. Physical [418, 469, 505, 725, 927, 1031, 1042, 1332, 1336, 1546, 1933, 2032, 2070]. physically [1639]. Pi [741, 2041]. Pi-Radio [741]. PIE [113]. PIFO [1960]. piggybacking [1827]. piggybacking-based [1827]. pile [654]. PiNcH [59]. ping [1574]. ping-pong [1574]. Pinning [579]. pipe [2091]. pipeline [117]. PISA [1218]. PKI [1150, 1850]. placement [8, 17, 100, 122, 173, 241, 330, 335, 392, 397, 455, 456, 527, 536, 630, 639, 642, 687, 693, 730, 787, 835, 857, 895, 899, 900, 959, 979, 1037, 1108, 1147, 1215, 1268, 1358, 1367, 1457, 1503, 1708, 1732, 1816, 1829, 1849, 1935, 2094, 2140]. places [1055]. placing [671]. Plaintext [1920]. Plaintext-aware [1920]. plan [574]. Plane [249, 443, 476, 506, 695, 951, 1052, 1195, 1228, 1354, 1453, 1609, 1647, 1700, 1849, 1864]. planes [238, 1579, 1662, 1938]. planet [837]. Planning [111, 155, 201, 294, 296, 664, 674, 739, 805, 808, 928, 1429, 1431, 1722, 1747, 1802, 2007]. Planning-based [1722]. Platform [114, 286, 332, 421, 529, 530, 670, 758, 781, 836, 892, 987, 1091, 1140, 1638]. platform-centric [114]. platforms [112, 233, 420, 478, 498, 610, 629, 947, 991, 1330, 1407, 1548, 1991, 2135]. platoon [474]. play [1024, 2082]. player [1454]. PLC [468, 1847]. PLC/RF [468]. plenty [511]. plot [1788]. PLS [1238]. PLS-HECC-based [1238]. Plug [2055, 2082]. Plug-and-play [2082]. PMIPv6 [282]. Po-Fi [507]. Poa [1952]. PoiEvent [1289]. Point [35, 115, 299, 347, 806, 1351]. point-to-point [1351]. pointer [1471]. points [170, 501, 744, 1268]. POIs [812]. poisoning [2132]. Poisson [307]. PoLe [912]. policies [1929]. Policy [6, 125, 445, 649, 683, 750, 1166, 1363, 1423, 1856, 1871, 1976]. policy-based [1166]. Policy-rich [750]. political [371]. Pollard [1226]. polling [1249]. pollution [1793]. polynomial [1236]. PON [235, 556, 674, 2116]. pong [1574]. pool [985]. pool-hopping [985]. pooling [1662]. poor [279]. popular [433]. Popularity [94, 295]. Popularity-aware [94]. populations [12]. PoR [2125]. pornography [554]. port [305, 1663]. port-based [305]. portion [345]. portion-based [345]. position [747, 760, 1156, 1620]. position-based [1156, 1620]. Positioning [1068, 1132, 1688, 1854]. post [434, 1283, 1523]. post-disaster [434, 1523]. post-quantum [1283]. Potential [35, 82, 1801, 1924, 1996]. PoW [1174, 1394]. Powder [758]. Power [100, 196, 263, 275, 280, 285, 310, 321, 343, 349, 381, 426, 565, 650, 702, 713, 730, 860, 900, 961, 979, 982, 1107, 1164, 1210, 1231, 1281, 1311, 1339, 1386, 1397, 1456, 1526, 1540, 1691, 1733, 1825, 1860, 1878, 1904, 2099, 2137, 2138]. power-efficiency [280]. power-efficient [730, 979]. power-Splitting [275]. powered [4, 215, 431, 552, 702, 774, 1175, 1190, 1490, 1523, 1814, 2072]. pp [1588]. PPTA [1488]. PQCLP [1982]. Pr [643]. Pr-CAI [643]. PRACH [898]. Practicable [1099]. Practical [7, 115, 555, 602, 740, 1053, 1663, 1704, 1715, 1723]. Practicality [1967]. practice [865]. practices [1839]. pragmatic [1353]. pre [1539, 1714]. pre-congestion [1714]. pre-mitigation [1539]. preamble [1219]. Precise [453, 886]. precision [7, 68, 132]. Predicted [334]. predicting [409, 516, 1227]. Prediction [23, 143, 244, 330, 333, 357, 386, 389, 509, 599, 663, 878, 891, 932, 943, 981, 1010, 1011, 1103, 1297, 1421, 1441, 1458, 1463, 1527, 1528, 1579, 1581, 1698, 1777, 1797, 1808, 1830, 1837, 1898, 1949, 1959, 1964, 1982, 2056, 2078]. Prediction-based [981]. predictions [963, 1048]. Predictive [886, 1098, 1560, 1711, 1805, 1889]. predictor [645]. preemption [2051]. preference [1499, 1948]. preference-aware [1948]. prefetch [157]. prefix [922]. preliminary [672]. premise [865]. PrePass [249, 445]. PrePass-Flow [445]. preprocessing [1348]. prescription [1105]. presence [1664, 1761]. Present [734]. preservation [97, 478, 1113, 1910]. preserve [604]. preserved [1015]. Preserving [206, 404, 518, 597, 863, 952, 964, 966, 998, 1008, 1083, 1237, 1245, 1275, 1434, 1446, 1488, 1502, 1554, 1563, 1737, 1832, 1985, 1991, 2027, 2064, 2132]. prevent [1645]. prevention [283, 441, 449, 772, 1204]. Price [136]. Price-aware [136]. Pricing [404, 800, 870, 950, 1284, 1462]. Pricing-based [950, 1284]. PriDPM [404]. primal [1970]. primal-dual [1970]. primitive [291, 1798]. principle [956]. Principles [1258, 1303, 1369]. prioritization [1143, 1707]. prioritized [41]. prioritizing [891]. Priority [261, 643, 1050, 1088, 1315, 1950]. Priority-aware [1315]. Priority-awareness [1050]. Priority-based [261]. priority-queue [1088]. Privacy [97, 206, 256, 274, 404, 471, 478, 485, 518, 547, 581, 597, 604, 610, 629, 863, 915, 924, 952, 964, 966, 992, 998, 1008, 1015, 1022, 1083, 1113, 1158, 1208, 1213, 1237, 1245, 1353, 1414, 1438, 1446, 1488, 1499, 1502, 1513, 1554, 1563, 1737, 1750, 1760, 1826, 1832, 1843, 1882, 1910, 1978, 1985, 1991, 2027, 2056, 2064, 2132]. Privacy-aware [256, 1826, 1882, 2056]. privacy-preservation [1113]. privacy-preserved [1015]. Privacy-preserving [404, 518, 863, 964, 966, 1008, 1083, 1237, 1245, 1446, 1488, 1502, 1554, 1563, 1737, 1832, 1991, 2027, 2064, 2132]. Privacy-utility [1208]. Private [734, 915, 1065, 1149, 1337, 1387, 2028, 2113]. PRLAP [1917]. PRLAP-IoD [1917]. Proactive [32, 95, 344, 682, 748, 853, 936, 1254, 1431, 1826]. proactively [1700]. Probabilistic [364, 793, 1554]. Probability [181, 303, 334, 1069, 1180, 1181, 1349, 1523, 1621, 1665]. Probability-based [181, 303]. probe [793, 994]. ProbInfer [303]. probing [1006]. problem [243, 342, 350, 621, 787, 835, 1436]. problem-The [787]. Problems [394, 557, 1167, 1371]. procedure [528, 731]. procedures [417, 1106]. Process [105, 219, 496, 676, 1013, 1361, 2017]. processes [396]. Processing [42, 48, 348, 522, 620, 648, 708, 809, 1117, 1140, 1389, 1490, 1562, 1595, 1967, 2072, 2110, 2125]. processing/offloading [1490]. producting [1921]. production [1332]. Profile [182, 1133, 2019]. profiling [1656, 1860]. Profit [599, 1817]. profit-oriented [1817]. profitable [697]. programmability [624, 695, 1228]. Programmable [238, 390, 436, 623, 1040, 1052, 1122, 1195, 1453, 1516, 1530, 1579, 1647, 1662, 1936, 1938, 2123]. Programming [273, 334, 436, 556, 951]. programs [896]. Progress [1126, 1520]. progressive [1704]. prompt [1820]. prompt-based [1820]. Prompt-WNQA [1820]. Proof [842, 912, 998, 1174, 1277, 2030]. proof-of-concept [998]. Proof-of-Stake [842, 2030]. Proof-of-Work [1174]. ProofChain [1150]. Propagation [7, 101, 123, 1014, 1027, 1353, 1361, 1396, 1557, 1804, 1855, 1883, 2118]. properties [1113]. property [1245]. proportional [1048, 1553]. proportional-fair [1048]. proposed [334]. Prospects [955, 1873]. protect [825]. protected [239, 1104]. protection [190, 205, 314, 336, 484, 695, 1099, 1158, 1213, 1337, 1422, 1455, 1665, 2028]. PROTECTOR [336]. Protocol [181, 190, 219, 247, 268, 284, 285, 293, 328, 405, 452, 460, 514, 517, 533, 581, 597, 605, 652, 699, 704, 717, 726, 745, 761, 775, 780, 786, 788, 829, 834, 858, 875, 889, 910, 971, 976, 1017, 1018, 1028, 1086, 1137, 1148, 1162, 1168, 1173, 1181, 1237, 1277, 1325, 1327, 1333, 1356, 1361, 1404, 1405, 1434, 1460, 1461, 1465, 1494, 1502, 1540, 1544, 1563, 1571, 1576, 1592, 1606, 1610, 1624, 1639, 1643, 1660, 1673, 1680, 1805, 1874, 1917, 1922, 1952, 2027, 2030, 2035, 2058, 2100, 2111, 2128]. protocols [58, 69, 162, 184, 275, 351, 874, 903, 1044, 1080, 1545, 1683, 1823, 1842, 1940, 2113]. provable [460]. Provably [704, 1325, 1922]. Provenance [154]. Provenance-enabled [154]. provide [915, 1343, 1376]. provider [606, 1375, 1491]. providers [2086]. provision [349, 960]. Provisioning [42, 144, 541, 748, 939, 967, 984, 1118, 1202, 1282, 1320, 1462, 1497, 1886, 1935, 1973, 2025, 2062]. proximity [2113]. Proxy [1055, 1310, 2057, 2117]. PSM [1164]. PSM-DMO [1164]. pSMART [256]. PSO [55, 1474]. PTAP [1563]. PTP [1144]. Pub [1343, 1416, 1514, 1763]. Pub/sub [1343, 1416, 1514, 1763]. Public [188, 704, 824, 949, 1245, 1387, 1671, 1778]. publish [828, 937, 1380, 1674, 1694]. publish-subscribe [828, 937]. publish/subscribe [1694]. publishing [97, 1757]. PUF [418, 688, 834, 1327, 1717, 1917]. PUF-based [834, 1327, 1917]. pull [304]. pull-based [304]. pulsing [160]. purification [159]. purpose [1268]. purposes [1807]. push [1776]. Pushing [1617]. Pythia [1117].

Q [391, 738, 757, 775, 788, 871, 1004, 1460, 1500, 1503, 1793]. Q-FANET [788].

Q-ICAN [1793]. Q-learning [757, 775, 788, 1004, 1500, 1503]. Q-learning-based [1460]. Q-Network [738]. Q-SR [871]. QBDD [2128]. QoE [42, 136, 155, 296, 565, 1176, 1542, 1588, 1970, 2097]. QoE-aware [565, 1970]. QoE-driven [1176, 1588]. QoS [15, 20, 109, 139, 201, 455, 475, 508, 561, 617, 618, 650, 701, 853, 962, 1057, 1297, 1316, 1372, 1415, 1619, 1661, 1808, 1853, 1870, 1912, 1918, 1924, 2018, 2026]. QoS-Aware [20, 201, 853, 1057, 1415, 2026]. QoS-based [1912]. QoS-guaranteed [455, 475]. QoS-Oriented [15, 1372]. QoT [1136, 1180]. QRNG [2027]. QTAR [1460]. Quadtree [452]. Quadtree-based [452]. Quality [175, 215, 221, 224, 347, 430, 462, 492, 506, 652, 923, 1072, 1323, 1362, 1384, 1666, 1806, 1807, 1843, 1905, 2077]. Quality-driven [1323]. Quantifying [1972]. quantile [1136]. Quantum [1168, 1283, 1486, 1630, 1895, 1982, 2128]. Quantum-resistant [2128]. quasi [315]. quasi-random [315]. queries [797, 936, 1275]. Query [182, 825, 1177, 1893]. question [1820]. queue [113, 550, 872, 1088, 1530, 2104]. queue-aware [550]. Queueing [1960]. queues [1777]. Queuing [340, 2081]. QUIC [866, 1212, 1360, 1525]. Quick [511].

R [70]. R-CNN [70]. radar [1638]. Radio [18, 40, 103, 123, 155, 183, 213, 237, 262, 296, 318, 321, 323, 332, 335, 360, 374, 381, 391, 396, 427, 438, 552, 565, 655, 701, 728, 741, 853, 943, 960, 1032, 1051, 1079, 1129, 1175, 1181, 1271, 1298, 1329, 1370, 1384, 1500, 1543, 1638, 1670, 1678, 1729, 1746, 1804, 1807, 1915, 2042, 2073, 2076]. radio- [853]. radio-based [1329]. radios [1066, 1742, 1996]. Raft [1952]. Raft-Poa [1952]. railway [578]. rainfall [1384]. RAN [16, 126, 234, 315, 318, 536, 803, 851, 1024, 1209, 1219, 1220, 1280, 1303, 1310, 1407, 1718, 1729, 1791, 2067, 2089]. RAN-aware [1310]. Random [101, 218, 313, 315, 802, 1004, 1197, 1319, 1350, 1624]. random-based [101]. randomization [604, 994, 1324]. Randomized [1769]. randomly [1291]. range [729, 1082, 1275, 1540, 1893]. range-based [1082]. Rank [1314, 1530]. rank-aware [1530]. ranking [901, 1532]. Ranking-based [1532]. RANs [144, 2075]. RansomCare [611]. Ransomware [611, 1941]. Rapid [1310, 1464]. raptor [1591, 1624]. Raptor-IRSA [1624]. Raspberry [2041]. RAT [407, 803, 1619]. rate [110, 193, 343, 365, 511, 573, 759, 1029, 1067, 1087, 1190, 1311, 1526, 1684, 1771, 1966, 1990, 2011, 2022, 2088]. rate-splitting [2022]. Rateless802.11 [279]. rates [697]. ratio [864]. rational [2028]. raw [1286, 2010]. Rayleigh [134]. RB [1861]. RBT [2010]. RCAP [780]. RCAP/CTAP [780]. Re [71, 292, 309, 748, 1261, 2117]. Re-defining [71]. Re-Encryption [2117]. re-mapping [1261]. re-orchestration [309]. re-provisioning [748]. re-synchronization [292]. reachability [236]. Reaching [521]. reaction [730]. reactive [1826]. readable [1379]. readers [805]. reading [817]. ready [1739]. Real [393, 620, 677, 717, 860, 969, 1140, 1200, 1299, 1348, 1387, 1508, 1546, 1574, 1598, 1698, 1734, 1762, 1983]. Real-time [393, 620, 717, 860, 969, 1140, 1200, 1299, 1348, 1508, 1546, 1598, 1698, 1734]. Realistic [160, 213, 886, 1791, 1842]. reality [700, 867, 1183, 1808]. Realizing [1635]. Reallocation [459]. Realtime [386, 963]. reasoning [1643]. reasoning-based [1643]. rebalancing [86]. Receiver [134, 1341, 1699, 1815, 2111]. Receiver-based [1815]. receiver-driven [1699, 2111]. receivers [183]. recently [1884]. receptive [1182]. rechargeable [888, 1211]. recognition [288, 993, 2082, 2112, 2139]. Recommendation [576, 859, 1625]. reconfigurable [669, 956, 1152, 1997, 2033].

Reconfiguration [647, 990, 1215, 1216, 1501, 1517, 1813, 1849, 2000]. reconstruction [198, 785, 1898]. record [876]. Records [409]. Recovery [17, 623, 1052, 1246, 1665, 1864]. recruitment [119, 1279]. rectangles [773]. rectangles-based [773]. recurrence [1788]. Recurrent [659, 1062, 1531, 1556]. Redact4Trace [2108]. redactable [2108]. reduce [722, 1441]. reduced [130, 757]. reducer [640]. Reducing [692, 1180, 1405, 2038, 2052]. reduction [377, 454, 492, 1424]. reduction-of-quality [492]. redundancy [189, 454, 1272]. redundant [1600]. reference [230, 1082]. reflecting [2022]. reflective [2046]. regenerative [2017]. region [233, 1984]. regions [5, 1990]. Registered [2010]. Registered-backoff-time [2010]. registration [190]. regression [409, 1240]. regression-based [1240]. regular [54]. Regularized [156]. regulation [989]. Regulatory [1807]. reinforced [1581]. Reinforcement [231, 253, 280, 290, 379, 406, 550, 579, 588, 608, 636, 661, 700, 707, 723, 774, 776, 830, 844, 872, 920, 972, 1029, 1037, 1050, 1062, 1081, 1089, 1200, 1217, 1219, 1235, 1265, 1281, 1338, 1339, 1342, 1367, 1372, 1388, 1408, 1464, 1471, 1477, 1491, 1493, 1501, 1580, 1583, 1614, 1627, 1628, 1659, 1668, 1677, 1707, 1713, 1730, 1732, 1733, 1759, 1784, 1786, 1789, 1790, 1796, 1823, 1835, 1839, 1862, 1867, 1874, 1882, 1891, 1980, 1981, 2012, 2021, 2049, 2074, 2091, 2098, 2101, 2126]. reject [1413]. rekey [1723]. relabelling [1596]. related [493, 1065, 2041]. Relational [1643]. relationship [1259, 1270]. Relay [4, 134, 267, 461, 831, 1069, 1088, 1141, 1269, 1290, 1385, 1544, 1599, 1691, 1780, 1881, 1977, 1979, 2013, 2073]. Relay-Assisted [1069]. Relay-empowered [2073]. relaying [196, 1506]. relays [275, 1409, 1423, 1498, 2088]. Reliability [278, 349, 746, 1591, 1683, 1916, 2002]. reliability-guaranteed [1591]. Reliable [377, 454, 618, 766, 1068, 1127, 1156, 1220, 1244, 1327, 1469, 1558, 1620, 1745, 1825, 1874, 1885, 2039]. relocation [1535]. remote [53, 67, 1030, 1129]. remote-controlled [1129]. removal [372]. renewal [1697]. Rényi [2103]. repair [1773]. Repeated [640]. repetition [820, 1744]. replacement [1644]. replay [1489, 2040]. Replicated [436]. replication [61, 1128, 1435, 1765]. replications [942]. representation [1602, 1806]. representations [447]. Reputation [287, 376, 538, 1322, 1952, 1999, 2105]. Reputation-based [287, 1952]. request [1147, 1708]. requesters [1395]. requests [307, 439, 994, 2085]. required [541]. Requirements [84, 1509]. Rerouting [1647]. Res [1783]. Res-TranBiLSTM [1783]. rescue [1523, 1880]. Research [69, 319, 332, 529, 544, 584, 758, 946, 1134, 1257, 1438, 1520, 1799, 1901, 1979, 2135]. reservation [308, 1115]. Resilience [377, 405, 505, 703, 1052, 1457]. Resilient [287, 669, 673, 726, 784]. resistance [1879]. resistant [277, 477, 2128]. resisting [688]. resolution [1317, 1363, 1865, 2084]. resolvers [1146]. Resource [7, 15, 18, 40– 42, 91, 138, 144, 217, 237, 244, 249, 263, 290, 310, 318, 367, 381, 413, 429, 472, 479, 484, 526, 527, 536, 551, 553, 573, 577, 580, 592, 619, 645, 655, 706, 719, 723, 724, 726, 738, 800, 819, 823, 939, 950, 975, 978, 1024, 1034, 1151, 1157, 1159, 1217, 1233, 1240, 1255, 1272, 1284, 1305, 1315, 1328, 1331, 1340, 1371, 1432, 1444, 1456, 1472, 1477, 1485, 1490, 1505, 1528, 1536, 1541, 1546, 1583, 1584, 1631, 1638, 1654, 1668, 1692, 1706, 1718, 1729, 1746, 1775, 1786, 1829, 1888, 1889, 1908, 1912, 1975, 2024, 2026, 2035, 2056, 2084]. resource-allocation [1490]. Resource-Aware [1331]. resource-constrained [7, 819]. Resource-efficient [1151, 1536]. Resource-limited [91, 726]. resources [541, 1430, 1572, 1933]. response [121, 998]. restoration [457, 748, 2007]. restricted [341]. results [1612, 2038]. Retina [926]. Retraction [1766]. retrieval [1065, 1153]. retrievals [1711]. Reuse [1108, 1914, 1937]. Reuse-based [1914]. Reusing [196]. Reveal [1708]. Revelation [597]. Revenue [210, 1738]. reverse [1643]. Review [112, 180, 184, 282, 472, 783, 905, 941, 983, 1134, 1204, 1221, 1302, 1411, 1510, 1520, 1646, 1939, 1993, 2134, 2135]. revision [291]. revisited [1623]. Revisiting [113, 921, 1115, 1436]. revocable [698]. revocation [1894]. reward [1491]. rewriting [997]. RF [468, 552, 1175, 1370, 1611, 1898]. RF-powered [552, 1175]. RFFI [1915]. RFID [22, 351, 555, 805, 817, 1296, 1304, 1639]. RFID-based [817]. RFTRUST [802]. RICERCANDO [225]. rich [750]. right [598]. ring [1486, 1850]. RIS [1100, 1290, 1520, 1583, 1605, 1629, 2126]. RIS-aided [1583, 1605]. RIS-assisted [1100, 1629, 2126]. RIS-enhanced [1520]. risk [1593, 1852, 2118]. RL [1779]. RL-Adapt [1779]. RLS [1627]. RNN [2070]. RNN-based [2070]. RNTI [1357]. Road [141, 390, 474, 661, 1010, 1156, 1176, 1588]. road-network [1010]. roads [380]. roadside [782]. roaming [1434]. robot [270, 519, 2139]. robotic [70, 1925]. robotics [195]. robots [1129]. Robust [14, 22, 38, 59, 174, 217, 317, 404, 468, 623, 644, 650, 702, 720, 806, 868, 948, 971, 1214, 1276, 1312, 1329, 1338, 1352, 1364, 1534, 1617, 1665, 1811, 1917, 1932, 2063, 2064, 2132]. Robustness [1342, 1347, 1708, 1918]. Robustness-aware [1708]. Role [1303, 1585]. roles [736]. RoNS [1214]. root [1396]. Roots [620]. Roots-tracing [620]. Rosenblatt [1464]. Route [239, 285, 739, 1099, 1226]. RouteChain [1300]. router [192]. Routers [157, 410, 1389]. routes [884]. Routing [30, 37, 77, 95, 133, 151, 191, 216, 253, 278, 285, 293, 339, 341, 342, 412, 425, 457, 466, 475, 520, 527, 531, 542, 558, 618, 636, 652, 671, 673, 679, 745, 749, 750, 754, 761, 775, 788, 790, 857, 871, 1003, 1032, 1057, 1104, 1110, 1121, 1127, 1137, 1148, 1156, 1162, 1189, 1200, 1254, 1276, 1289, 1300, 1301, 1309, 1315, 1333, 1350, 1383, 1392, 1400, 1405, 1427, 1460, 1469, 1487, 1533, 1537, 1545, 1571, 1610, 1620, 1646, 1654, 1680, 1697, 1748, 1782, 1784, 1823, 1874, 1891, 1914, 1950, 1977, 1980, 1990, 2091, 2092]. RPL [154, 431, 802, 1314]. RPL-based [154, 802]. RRH [374, 1209]. RRH-Sector [1209]. RRH-Sector-BBU [1209]. RSSI [1539, 2045]. RSSI-based [1539]. RT [1825]. RTC [926]. RTIFed [1999]. RTS [893]. RTS/CTS [893]. rule [546, 1921]. rule-caching [546]. Rumors [651]. running [1956]. runtime [681, 1517]. rural [1049].

S [1107, 1419, 1721]. S-FoT [1721]. S-HIDRA [1419]. S2H [892]. S2X [1831]. SA [646]. safe [488]. safety [141, 1156]. SAIDE [414]. Salience [1816]. Salience-based [1816]. sampling [1771, 2038]. Sanitization [569]. SAPWSN [1404]. Satellite [80, 151, 163, 417, 731, 1110, 1186, 1197, 1257, 1294, 1533, 1538, 1654, 1692, 1831, 1862, 1909, 1929, 2008]. Satellite-assisted [1862]. Satellite-based [1257]. satellite-terrestrial [1538, 1654]. Satisfaction [129, 459, 2024]. Satisfaction-based [459]. satisfying [328]. saturation [76, 961, 1155]. Save [1164]. saving [1979]. SBERT [1398]. SBPG [836]. SBS [105]. SCADA [469]. scalability [753, 790, 1243]. Scalable [31, 68, 99, 156, 227, 287, 405, 539, 579, 601, 858, 920, 924, 967, 1007, 1052, 1125, 1140, 1277, 1451, 1482, 1608, 1618, 1639, 1717, 1921, 1985, 1997]. Scale [68, 276, 382, 391, 440, 463, 753, 868, 933, 1010, 1121, 1191, 1336, 1432, 1506– 1508, 1562, 1635, 1878, 1962, 2008, 2120]. scale-free [868, 1508]. ScaleDRL [579]. Scaling [8, 52, 344, 942, 2113]. scanner [74]. scanning [940]. scans [1663]. scenario [785, 1384, 1544, 1731, 1851, 1929]. scenarios [47, 294, 434, 604, 898, 1173, 1366, 1514, 1721, 1964]. schedule [140, 771]. Scheduled [218]. scheduler [155, 296, 2102]. Scheduling [15, 78, 92, 99, 140, 193, 194, 262, 281, 337, 403, 415, 429, 434, 439, 465, 470, 486, 550, 588, 643, 664, 773, 784, 885, 891, 1013, 1019, 1038, 1089, 1104, 1117, 1147, 1159, 1175, 1201, 1210, 1316, 1326, 1371, 1403, 1426, 1428, 1468, 1471, 1493, 1524, 1530, 1553, 1627, 1697, 1705, 1708, 1777, 1784, 1786, 1849, 1861, 1896, 1904, 1911, 1914, 1956, 1966, 1974, 1988, 2005, 2008, 2049, 2065, 2104, 2114, 2123]. Scheme [33, 65, 87, 102, 119, 150, 191, 194, 206, 240, 242, 277, 295, 344, 373, 384, 397, 408, 425, 477, 485, 486, 498, 520, 534, 541, 591, 596, 623, 636, 669, 698, 701, 710, 730, 828, 872, 876, 882, 886, 893, 938, 957, 959, 972, 981, 1015, 1068, 1101, 1133, 1156, 1158, 1183, 1188, 1189, 1198, 1211, 1236, 1238, 1245, 1249, 1269, 1271, 1283, 1296, 1304, 1424, 1425, 1457, 1486, 1492, 1504, 1620, 1628, 1638, 1663, 1688, 1717, 1773, 1780, 1815, 1817, 1825, 1834, 1835, 1851, 1864, 1893, 1908, 2020, 2024, 2028, 2059, 2062, 2120, 2124]. Schemes [7, 261, 688]. Schnorr [277]. Schnorr-like [277]. science [2011]. scientific [25, 290, 337]. SCM [961]. scoping [1646]. scores [2097]. scoring [376, 1379, 2063]. scripting [24]. SD [501, 976, 1392, 2074]. SD-CDN [976]. SD-WAN [501, 2074]. SD-WANs [1392]. SDBGPChain [1645]. SDFog [1096]. SDFog-Mesh [1096]. SDHW [545]. SDHW-IoT [545]. SDM [178, 1436]. SDM-EON [178]. SDM-EONs [1436]. SDN [16, 31, 33, 44, 86, 93, 102, 111, 118, 249, 256, 259, 283, 345, 355, 375, 414, 445, 456, 476, 499, 500, 507, 529, 579, 586, 599, 605, 614, 632, 639, 643, 647, 649, 657, 682, 709, 731, 763, 787, 791, 857, 877, 890, 979, 981, 1023, 1052, 1056, 1089, 1155, 1166, 1176, 1188, 1198, 1239, 1247, 1295, 1347, 1419, 1424, 1527, 1588, 1633, 1645, 1700, 1723, 1732, 1754, 1758, 1835, 1864, 1897, 1911, 1926, 1931, 1932, 1950, 1980, 1992]. SDN-aware [1166]. SDN-based [614, 643, 657, 682, 731, 1424, 1723, 1911]. SDN-enabled [111, 1089, 1732]. SDN-IoV [639]. SDN/NFV [283, 857, 979, 1835]. SDN/NFV-enabled [979, 1835]. SDN/non [605]. SDN/non-SDN [605]. SDNs [561, 795]. SDNSandbox [1375]. SDONs [1479]. SDR [16, 332, 1996]. SDR-based [332]. SDR-LoRa [1996]. SDTMA [25]. SDTMA-NDN [25]. SDWSN [640]. Seamless [500, 1151, 1197]. Search [263, 679, 1235, 1561, 1880]. searchable [876, 1851]. searching [804]. SecFHome [1030]. SECHO [1628]. second [1459]. Secrecy [1480]. secret [277, 2028]. section [1412]. Sector [578, 1209]. sectorization [578]. Secure [68, 119, 240, 247, 257, 259, 263, 455, 477, 505, 515, 519, 597, 598, 657, 673, 704, 710, 719, 756, 774, 776, 826, 828, 831, 856, 876, 882, 909, 970, 1003, 1008, 1018, 1030, 1038, 1083, 1102, 1110, 1113, 1127, 1173, 1196, 1269, 1296, 1300, 1304, 1321, 1325, 1343, 1404, 1435, 1486, 1555, 1563, 1628, 1680, 1685, 1686, 1717, 1748, 1765, 1875, 1881, 1893, 1922, 1951, 2034, 2105, 2117]. Secure-aware [455]. secured [518, 834]. Securing [401, 649, 753, 801, 1106, 1743, 2050]. Security [81, 82, 84, 158, 266, 274, 291, 351, 359, 418, 442, 458, 471, 474, 493, 540, 547, 564, 581, 610, 617, 627, 629, 630, 656, 658, 767, 768, 770, 802, 824, 836, 848, 904, 935, 980, 988, 992, 1007, 1021– 1023, 1040, 1047, 1106, 1114, 1133, 1146, 1166, 1203, 1255, 1257, 1313, 1332, 1425, 1437, 1438, 1508, 1513, 1529, 1603, 1633, 1646, 1723, 1757, 1822, 1939, 1987, 1992, 2032, 2041, 2059, 2135]. security-aware [1255]. security-enhanced [980]. security-related [2041]. SED [1717].

SED-SGC [1717]. segment [412, 871, 1121]. segmentation [270, 498, 1182]. segmented [1242]. select [311]. selected [1187, 1649]. selecting [1423, 1462]. Selection [13, 53, 114, 137, 159, 176, 267, 299, 310, 341, 383, 407, 431, 463, 497, 528, 566, 640, 749, 793, 795, 821, 831, 854, 890, 905, 977, 1034, 1082, 1090, 1175, 1180, 1209, 1224, 1226, 1231, 1269, 1299, 1339, 1436, 1448, 1490, 1531, 1549, 1608, 1691, 1700, 1712, 1775, 1782, 1835, 1887, 1888, 1958, 1977, 1979, 2013, 2074, 2090]. Selective [538, 614, 1736]. selective-edge [538]. Self [727, 735, 838, 951, 1024, 1056, 1064, 1298, 1301, 1359, 1455, 1546, 1642, 1706, 1712, 1769, 1775, 1841, 2025]. Self-adaptive [1056, 1546]. self-adjusting [1769]. self-attention [1775]. Self-attentive [727]. Self-backhauling [1298]. self-coordination [735]. self-driven [951]. Self-healing [1301]. Self-organizing [735, 838, 1706, 1712]. Self-play [1024]. self-protection [1455]. self-sovereign [1359]. self-stabilizing [2025]. self-sustainable [1841]. Self-verifiable [1642]. selfish [393]. Semantic [583, 694]. semi [6, 537, 877, 1043, 1096, 1586, 2060, 2125]. semi-parallel [2125]. semi-permanent [1096]. semi-supervised [6, 537, 877, 1043, 2060]. semi-synthetic [1586]. sensing [63, 67, 193, 311, 356, 430, 602, 733, 923, 1454, 1737, 2119]. sensitive [336, 499, 1265, 1428, 1517, 1553, 1653, 1705, 1755, 1784, 1851, 2036, 2051, 2094]. Sensor [7, 30, 34, 38, 40, 45, 83, 106, 124, 180, 215, 232, 253, 264, 275, 289, 319, 340, 350, 497, 523, 621, 679, 713, 716, 722, 761, 775, 806, 868, 888, 920, 967, 1027, 1032, 1036, 1082, 1085, 1101, 1127, 1137, 1162, 1211, 1226, 1268, 1269, 1273, 1275, 1301, 1336, 1348, 1355, 1404, 1435, 1440, 1460, 1540, 1680, 1681, 1765, 1794, 1818, 1823, 1833, 1874, 2013, 2058]. sensors [321, 631, 804, 1609]. sentence [590]. sentence-level [590]. September [245, 711, 1170, 1588, 1695]. sequences [505, 1402]. sequential [97, 228, 1711]. serendipity [371]. serial [1244]. serialized [1899]. series [156, 1336, 1528, 1953, 2077]. serve [1759]. served [1792]. server [234, 327, 419, 452, 484, 816, 1073, 1118, 1806, 2133]. server-side [1806]. Serverless [147, 1091]. servers [1644]. Service [8, 14, 20, 56, 118, 160, 164, 167, 173, 194, 251, 252, 256, 314, 317, 368, 375, 392, 397, 504, 560, 592, 606, 663, 678, 687, 693, 816, 899, 917, 939, 967, 977, 991, 1008, 1017, 1056, 1084, 1108, 1120, 1125, 1147, 1154, 1159, 1169, 1228, 1230, 1252, 1255, 1272, 1340, 1362, 1373, 1398, 1411, 1421, 1422, 1426, 1437, 1455, 1458, 1462, 1463, 1478, 1485, 1488, 1491, 1501, 1523, 1535, 1538, 1572, 1655, 1701, 1749, 1758, 1786, 1804, 1807, 1829, 1932, 1935, 1973, 1975, 2000, 2002, 2007, 2036, 2037, 2053, 2061, 2062, 2094]. Service-aware [173]. Service-Based [991, 1017]. Service-centric [592]. service-oriented [167]. Services [20, 88, 185, 255, 361, 380, 477, 734, 820, 918, 1083, 1114, 1158, 1176, 1214, 1266, 1413, 1434, 1542, 1588, 1666, 1924, 1935, 2002, 2081]. serving [267]. session [190, 1576]. sessions [376, 2049]. Set [1068, 1350, 1780, 1811]. SETCAP [1017]. setter [892]. setting [277]. SFCs [1792]. SGC [1717]. SGD [731]. SHA [753]. SHA-256 [753]. Shadowsocks [948]. Shannon [583]. SHAPARAK [405]. shape [1611]. shapelet [156]. shaper [1784]. sharding [2087]. share [96]. Shared [301, 484, 563, 833, 1268, 1830, 1859]. Sharing [4, 21, 122, 210, 228, 381, 396, 422, 776, 876, 955, 1051, 1098, 1102, 1140, 1496, 1635, 1846, 1851, 2028, 2035, 2056, 2117]. SHARP [597]. shelf [602]. Shell [1740]. Shining [1055]. shipping [519]. shop [588]. Short [645, 1800]. shot [799, 1368, 1602, 1632]. should [760]. shuffling [848]. SIC [1130, 1761]. Side [740, 811, 812, 997, 1042, 1073, 1234, 1806]. Side-channel [740, 811, 812, 1042, 1234]. sided [684]. sidelink [58]. sidelink-based [58]. SIDS [1821]. SIEM [656]. sightings [1604]. sigmoid [1861]. sigmoid-based [1861]. signal [839, 925, 1208, 2112]. signaling [757, 1529]. signals [946, 961, 1617]. signature [1173, 1486]. signatures [268, 1850]. signcryption [1101, 1795]. silent [1195]. SIM [1961]. similarity [895, 1794, 1976]. SIMO [1690, 1764]. SIMO-Underwater [1690, 1764]. simple [36]. Simulation [529, 1332, 1772, 1909]. simulations [539]. simulator [166, 230, 1885]. simulators [1759, 1901]. simultaneous [159, 1426]. sine [1162]. single [662, 684, 839, 1532]. single-output [662]. single-sided [684]. sink [1435, 1765]. sinkhole [802]. SIP [190, 935, 1957]. SIP-based [1957]. SIP-enabled [935]. site [24, 1422]. situation [909]. six [1854]. six-layer [1854]. Size [403, 1780]. Size-based [403]. skeleton [2091]. Sketch [683, 902, 1285, 1564, 1868]. skew [7]. Skipping [208]. Skipping-based [208]. SlaaS [1817]. Slardar [1201]. slate [1380]. sleep [310, 906, 1627]. sleeping [234, 1393]. slice [1202, 1320, 1497, 1501, 1981]. sliceable [869]. sliced [41]. slices [126, 526, 732, 1229, 1816, 1843]. Slicing [16, 44, 233, 522, 670, 851, 899, 1219, 1265, 1271, 1543, 1718, 1746, 1758, 1813, 1843, 1887, 2065]. sliding [1084, 1564]. slope [467]. slot [1881]. slot-pairing [1881]. slots [696]. Slotted [820, 1243, 1744]. slotted-ALOHA [820]. Slow [164]. small [138, 408, 562, 900, 1018, 1635, 1846]. small-cell [1846]. Smart [48, 49, 69, 71, 82, 95, 131, 145, 211, 322, 496, 518, 519, 568, 584, 588, 590, 609, 618, 655, 656, 708, 719, 774, 807, 826, 949, 987, 1030, 1071, 1096, 1105, 1113, 1140, 1238, 1267, 1387, 1592, 1624, 1645, 1752, 1801, 1804, 1859, 1925, 2020, 2023, 2071, 2098]. SmartDefense [1061]. SmartFCT [280]. SmartNIC [1079]. smartphone [611, 911, 1915]. smartphones [221]. SmartSBD [1859]. SmartTRO [1342].

SMOTE [226]. SNARKs [2027]. sniffing [12]. Snort [1161]. Social [10, 13, 97, 101, 182, 189, 258, 334, 393, 411, 422, 430, 558, 651, 755, 865, 905, 987, 1103, 1118, 1284, 1353, 1387, 1416, 1557, 1821, 1869, 1883, 2095]. Social-aware [10, 189, 422, 1869]. Society [1092]. SOFI [1259]. SoftCaching [1782]. Software [25, 76, 84, 95, 157, 164, 216, 231, 257, 349, 355, 410, 443, 446, 449, 455, 475, 512, 520, 523, 539, 543, 545, 606, 613, 624, 630, 658, 673, 706, 746, 790, 858, 978, 1037, 1066, 1096, 1102, 1120, 1163, 1184, 1246, 1313, 1321, 1535, 1607, 1641, 1693, 1742, 1782, 1796, 1876, 1897, 1918, 1926, 1977, 1996, 2103, 2135, 2137]. Software-Defined [76, 84, 95, 231, 446, 455, 475, 520, 523, 606, 613, 624, 658, 706, 858, 1096, 1120, 1184, 1246, 1535, 1641, 1742, 1782, 1796, 1876, 1918, 1926, 2137]. Software-Defined-Radios [1996]. softwarization [1993]. Softwarized [248, 300, 533, 1358, 2099]. solar [789]. solar-aware [789]. solution [59, 66, 218, 285, 378, 468, 585, 586, 987, 1095, 1144, 1806, 2083, 2108]. Solutions [82, 88, 361, 418, 705, 709, 1134, 1257, 1371, 1416, 1464, 1512, 1763, 1900, 2095]. Solving [282, 342, 621]. SOREC [1706]. SØREN [1453]. soundscape [716]. source [37, 135, 166, 193, 657, 926, 1161, 1198, 1515, 1634, 1996, 2002, 2033, 2138]. sourced [1225, 1739]. sources [131, 911]. sovereign [1359]. SP [2125]. SP-PoR [2125]. SP2F [518]. Space [172, 178, 339, 529, 765, 1015, 1057, 1234, 1315, 1390, 1457, 1510, 1701, 1790, 1967, 2066, 2069, 2092]. space-aerial-terrestrial [172]. space-air-ground [2066]. space-division [1315, 1390, 1701]. space-efficient [765]. Space-Ground [529, 1015]. spaces [1702]. Spam [420, 845, 1020]. sparse [1043, 1191]. sparsification [1622]. Spatial [178, 454, 600, 808, 1287, 1488, 1555, 1937, 1991, 2038, 2078, 2127]. Spatial-Temporal [2078]. spatially [1886, 2015]. Spatio [117, 359, 1797, 2048]. Spatio-temporal [117, 359, 1797, 2048]. spatiotemporal [555, 591, 1248]. SPC [943]. SPEAR [656]. Special [658, 1412, 1587, 1649]. Specific [745, 1972]. specification [252, 1876]. Spectral [323, 556, 1311, 1638, 1913]. Spectral-Energy [323]. spectrally [869, 1886, 2015]. spectrally-spatially [1886, 2015]. Spectrum [4, 109, 133, 178, 183, 229, 242, 243, 311, 332, 339, 342, 396, 422, 426, 489, 497, 541, 655, 701, 721, 732, 749, 943, 1051, 1057, 1309, 1391, 1479, 1500, 1661, 1702, 1733, 1904, 2119]. SpectrumChain [1479]. speculative [490]. speech [2097]. speed [377, 415, 498, 641, 699, 1663, 1868, 1884, 1953]. Speeding [1014]. speedtest [440]. Sphinx [699]. Split [1280, 1331, 1791]. SplitPath [1032]. Splittable [1194]. Splitting [136, 275, 702, 1311, 1682, 1974, 2005, 2022]. spoke [350]. Spoofed [369]. spoofing [110, 1225]. spray [558]. spray-and-wait [558]. spreader [718]. spreaders [1868]. spreading [383]. Spring [1068]. SQHCP [455]. squares [45]. Squatting [41]. SR [871]. SRE [836]. SRLG [674]. SRLG-aware [674]. SRRS [1361]. SRv6 [476, 1736]. SSAPPIDENTIFY [948]. SSDI [907]. Stability [1003, 1815]. Stability-aware [1815]. stabilizing [2025]. Stable [431, 1805]. stack [1168, 1618]. stacked [1043]. Stackelberg [1737, 2122]. Stackelberg-evolutionary [2122]. stacking [435]. stage [194, 429, 449, 868, 1029, 1316, 1704, 1722, 1778]. staged [1581]. Stake [842, 2030]. stakeholder [442]. stakeholder-oriented [442]. stale [1080]. staleness [2136]. stand [331]. stand-alone [331]. standalone [1998]. standard [1611, 1644]. standardisation [417]. standards [1359, 2134]. Start [1960]. Start-Time [1960]. State [390, 416, 452, 474, 543, 632, 653, 810, 1021, 1230, 1266, 1350, 1426, 1510, 1550, 2032].

State-of-art [1510]. State-of-the-Art [390, 474, 543, 632, 653, 1021, 1550, 2032]. stateful [17, 443]. States [436]. static [903, 997, 1012]. Station [19, 234, 626, 820, 1688, 1774, 1951, 2075]. stationary [307]. stations [111, 123]. statistical [1762]. Status [955]. Steerable [990, 1069]. Steerable-Beam [1069]. steering [37, 1974]. stegomalware [612]. Step [165, 785, 808, 1299, 1999]. STM32 [2041]. Stochastic [364, 380, 573, 2017, 2075]. stock [659]. storage [663, 959, 1128, 1245, 1429, 1670, 1773, 1830]. store [419]. strategies [171, 220, 832, 1024, 1332, 1357, 1550, 1612, 1648, 1726, 2077]. strategy [26, 92, 118, 260, 262, 360, 509, 640, 650, 720, 789, 811, 825– 827, 1035, 1087, 1127, 1132, 1261, 1338, 1393, 1526, 1538, 1644, 1700, 1709, 1715, 1752, 1858, 1881, 1891, 1898, 1948, 2066, 2089, 2129]. strawberry [49]. Stream [354, 555, 797, 1595]. stream-based [555]. Streaming [174, 462, 600, 601, 867, 1130, 1323, 1348, 1380, 1415, 1626, 1779, 1806, 1848, 2001, 2081, 2101]. Streams [354, 902, 1041, 1275, 1868, 1884]. strength [925]. strict [37]. stricter [508]. stricter-than-ICNIRP [508]. STRIDE [1040]. STRIDE-based [1040]. Strike [1920]. strong [2030]. strongly [882]. structural [932]. structure [36, 254, 582, 680, 938, 1757, 1794, 1890]. Structured [887, 959, 1411, 1719]. structures [684, 1972]. Study [10, 49, 161, 162, 382, 540, 634, 804, 884, 1041, 1056, 1105, 1172, 1357, 1387, 1427, 1547, 1681, 1804, 1855, 2042, 2045, 2113]. Studying [1721, 1807]. sub [332, 395, 804, 1265, 1343, 1416, 1514, 1763, 2049]. sub-6 [332, 804]. sub-action [1265]. sub-terahertz [395]. sub-tree [2049]. Subchannel [40]. subject [34]. subjective [2097]. subliminal [277]. submarine [2009]. subscribe [828, 937, 1380, 1674, 1694]. Subscriber [678]. subscription [1543]. substations [349, 441]. substrate [1532]. success [1349]. successful [134]. suitability [610]. suitable [311, 1553]. Sum [1190, 1814, 2088]. sum-rate [2088]. super [869, 1317, 1865, 1868]. super-channels [869]. super-spreaders [1868]. superlattice [1717]. supervised [6, 537, 877, 1043, 2060]. supplies [970]. supply [351, 865]. Support [224, 767, 1067, 1140, 1202, 1294, 1299]. Supporting [238, 315, 389, 533, 927, 1079, 1453, 1739, 1772]. surface [1145, 1367, 2022]. surfaces [1997, 2046]. surgery [1925]. Suricata [1161]. surrogate [1348]. surrogate-assisted [1348]. surveillance [573, 907, 1474, 1534, 1548]. Survey [21, 24, 44, 69, 81, 123, 172, 190, 200, 257, 302, 319, 356, 387, 388, 417, 418, 441, 443, 471, 473, 495, 547, 568, 585, 603, 608, 628– 630, 632, 653, 686, 705, 706, 734–736, 752, 769– 771, 837, 838, 874, 903, 904, 906, 945, 992, 1000, 1016, 1021, 1023, 1033, 1040, 1051, 1066, 1072, 1122, 1126, 1135, 1167, 1168, 1178, 1257, 1258, 1370, 1437, 1438, 1511–1513, 1545, 1564, 1585, 1611, 1633, 1647, 1648, 1724–1726, 1759, 1798, 1799, 1838, 1840, 1841, 1854, 1901, 1902, 1941– 1943, 2031, 2096]. Survivability [1110, 1137, 1162, 1412, 2016]. survivable [674]. suspicious [376]. sustainability [512]. sustainable [248, 371, 837, 1196, 1841]. Sustaining [890]. SUTSEC [259]. SVC [600]. SVM [152, 614, 1207]. Swarm [669, 728, 805, 883, 983, 1301, 1664, 1669]. swarms [48]. SWIPT [367, 1274, 2088]. SWIPT-empowered [1274]. SWIPT-enabled [367, 2088]. switch [981, 1184, 1218, 1700, 1733]. switch-migration [1184]. switch-off [1733]. switch-to-controller [1700]. Switches [43, 249, 543, 794, 1037, 1040, 1516, 1647, 2033]. switching [65, 275, 626, 684, 771, 792, 860, 1806, 1897, 1974]. Switzerland [1241]. Sybil [842, 1515, 1570]. Symmetric [328, 786, 1325]. symmetric-key [786, 1325]. symposium [1187]. Symptom [1259]. Symptom-fault [1259]. SYN [1662, 2103]. synchronisation [1630]. synchronization [7, 160, 292, 452, 1461, 1580, 1778]. synchronized [1613]. Synchronizing [1239]. synchronous [1318, 1356]. SYNCOP [456]. synthetic [1009, 1586]. SYNTROPY [2103]. System [57, 70, 86, 146, 176, 376, 446, 449, 469, 471, 547, 577, 580, 595, 613, 614, 625, 631, 656, 685, 703, 792, 856, 859, 907, 910, 923, 924, 937, 948, 952, 980, 1049, 1105, 1113, 1213, 1217, 1359, 1474, 1531, 1577, 1598, 1623, 1690, 1716, 1764, 1812, 1814, 1831, 1836, 1866, 1909, 1925, 1932, 2022, 2093, 2105, 2126, 2128]. system-level [1909]. Systematic [282, 472, 768, 905, 983, 1204, 1726, 1939]. Systematizing [1992]. Systems [71, 74, 78, 114, 117, 131, 147, 210, 265, 266, 321, 347, 357, 389, 396, 441, 453, 460, 466, 468, 517, 565, 576, 589, 658, 724, 766, 772, 774, 808, 813, 824, 831, 947, 949, 959, 983, 1042, 1067, 1072, 1074, 1107, 1130, 1196, 1231, 1258, 1290, 1297, 1332, 1336, 1343, 1376, 1409, 1416, 1448, 1472, 1480, 1494, 1520, 1546, 1555, 1575, 1593, 1605, 1625, 1634, 1653, 1659, 1678, 1689, 1694, 1724, 1725, 1763, 1773, 1800, 1847, 1854, 1873, 1900, 1957, 1958, 2017, 2031, 2037, 2090].

TABDeep [2049]. table [157, 2033]. table-based [157]. tables [60, 157, 1926]. Tabu [263]. Tactile [1159, 1672, 1925]. tag [817]. tags [1296]. tail [760]. tailored [522]. Tailoring [1731]. TAMBUS [401]. tangle [821]. Target [45, 806, 1557]. targets [400]. Task [129, 207, 281, 324, 347, 366, 400, 402, 589, 609, 648, 664, 691, 705, 707, 784, 830, 839, 843, 870, 1013, 1038, 1070, 1087, 1089, 1123, 1178, 1194, 1262, 1326, 1468, 1472, 1477, 1488, 1490, 1493, 1538, 1582, 1602, 1625, 1722, 1737, 1747, 1752, 1753, 1833, 1908, 1987, 1991, 2026, 2110, 2129]. task-processing [648]. tasks [78, 434, 1358, 1388, 1471, 1477, 1572, 1677, 1730, 1911, 1956, 2065]. tat [1179]. Taxonomy [44, 252, 472, 557, 688, 770, 1221]. TbDd [2087]. TCP [162, 192, 403, 437, 513, 796, 1122, 1518, 1731, 1929, 2103]. TCP/AQM [192]. TD [1392]. TDD [1678]. TDMA [125, 728]. TE [31, 1049]. TEA [1915]. TEA-RFFI [1915]. team [430]. tech [622]. Technical [230, 371, 1585]. technique [153, 310, 445, 625, 755, 824, 1199, 1275, 1299, 2029]. Techniques [71, 163, 526, 585, 741, 873, 906, 1000, 1021, 1041, 1178, 1226, 1647, 1798, 2007, 2101, 2135]. technologies [195, 378, 584, 585, 634, 1072, 1577, 1801, 1902, 2134]. Technology [653, 753, 1505, 1536, 1758, 1775, 1900]. telecom [670]. telecommunication [986, 1574]. Telemetry [316, 483, 495, 1229, 1232, 1389, 1470, 1573, 1973, 2038, 2098]. Temperature [1915]. Temporal [117, 359, 454, 944, 1017, 1324, 1473, 1797, 2048, 2078]. temporal- [1324]. tenancy [229]. tenant [56, 344, 1255, 1271]. tenant-driven [1271]. tensor [662, 861]. tensor-information [662]. Terahertz [286, 293, 395, 946, 1183]. TeraNova [286]. Term [645, 1478, 1756]. termination [1714]. terrestrial [172, 1538, 1654, 1668]. test [1808]. testbed [286, 665, 1742]. testbeds [847, 1613, 1901]. testing [332, 453, 1407, 2113]. text [590, 1242]. their [126, 1012]. theoretic [210, 800, 940, 1095, 2114]. theoretical [617, 1148]. Theory [260, 640, 705, 845, 1261, 1329, 1338, 1403, 1499, 1858, 2038, 2066]. theory-based [640]. thermal [1503]. thermal-aware [1503]. Things [21, 154, 161, 167, 182, 185, 188, 219, 247, 258, 262, 274, 302, 418, 425, 431, 444, 458, 470, 480, 491, 496, 513, 515, 522, 568, 576, 592, 635, 652, 753, 756, 776, 783, 807, 863, 873, 882, 905, 937, 941, 957, 967, 1012, 1036, 1075, 1135, 1160, 1262, 1293, 1314, 1319, 1331, 1342, 1394, 1424, 1511, 1545, 1587, 1672, 1683, 1710, 1719, 1782, 1783, 1799, 1821, 1839, 1857, 1862, 1910, 1993, 2096, 2117]. third [29]. Thomas [105]. Threat [563, 613, 770, 909, 1449, 1496, 1604, 1703, 2118]. Threats [82, 104, 988, 1106, 1257, 1425, 1437, 1586, 1646]. three [242, 770, 1284, 1571, 1794, 2124]. three-dimensional [770, 1794]. three-phase [242]. three-tier [1284, 2124]. threshold [791, 1677]. Throughput [11, 83, 494, 746, 780, 960, 1004, 1032, 1310, 1360, 1445, 1688, 1814, 2091]. tie [1883]. tie-driven [1883]. tier [217, 224, 365, 816, 1275, 1278, 1284, 2124]. tile [600]. Time [7, 125, 156, 178, 212, 219, 275, 393, 499, 620, 717, 722, 860, 891, 969, 1140, 1175, 1179, 1200, 1220, 1234, 1299, 1336, 1348, 1352, 1354, 1394, 1428, 1508, 1510, 1517, 1528, 1546, 1598, 1613, 1630, 1653, 1665, 1698, 1705, 1734, 1755, 1762, 1784, 1833, 1953, 1960, 2010, 2051, 2069, 2077, 2094]. time-aware [1784]. time-constrained [1394]. time-delay [1755]. time-frequency [1510]. time-of-arrival [212]. time-of-departure [212]. time-sensitive [1428, 1517, 1653, 1755, 1784, 2051, 2094]. Time-series [1336, 1528, 1953]. time-space [1234]. time-switching [275]. time-synchronized [1613]. Timely [1416]. times [130, 544, 1687]. timescale [795, 1871, 2094]. Timing [539, 1163, 1693, 1716]. TIoT [1672]. Tip [821]. tit [1179]. tit-for-tat-based [1179]. TLS [1413, 2055]. TLS-handshake [2055]. TMAGRL [2094]. TNN [1866]. TNN-IDS [1866]. Tolerance [541, 782, 961, 1908]. tolerant [17, 78, 191, 558, 840, 1092, 1137, 1272, 1301, 1608, 1777]. TONTA [668]. tool [155, 166, 296, 432, 453, 926]. toolkit [225]. tools [71, 265, 681, 2138]. top [718, 959, 1641]. top- [718, 1641]. top-down [959]. topological [1794]. topologies [331, 605, 1152, 1241, 1350, 1989]. Topology [102, 125, 211, 232, 252, 350, 739, 858, 868, 986, 1071, 1342, 1350, 1460, 1599, 1600, 1789, 1909, 1947, 2016]. topology-aware [1460]. TOPSIS [254]. Tor [1983]. touch [300, 1120, 1840, 1876, 2068]. Tower [1288]. TPM [291]. trace [798, 1357]. traceability [1836, 1894]. traceable [1563]. Traceback [364, 614]. Tracer [501]. traceroute [1789]. traces [995]. tracing [154, 612, 620, 1613, 1836, 2108]. Tracking [29, 45, 150, 192, 501, 747, 805, 1334, 1565, 1872, 1994, 2009]. TrackSign [1565]. TrackSign-labeled [1565]. tractable [250]. trade [276, 556, 982, 1013]. trade-off [556, 1013]. trade-offs [276, 982]. tradeoff [323, 617, 1311]. tradeoff-based [1311]. tradeoffs [136]. trades [1208]. trading [595, 1499]. Traditional [437, 1370]. Traffic [12, 18, 36, 47, 54, 59, 77, 118, 136, 163, 209, 214, 228, 288, 315, 338, 340, 346, 352, 357, 369, 389, 466, 489, 527, 554, 561, 574, 579, 591, 630, 668, 731, 751, 795, 846, 852, 878, 879, 887, 898, 911, 926, 929, 942, 969, 976, 995, 1005, 1009, 1011, 1041, 1049, 1076, 1116, 1191, 1218, 1250, 1276, 1286, 1295, 1312, 1315, 1377, 1421, 1433, 1507, 1535, 1547, 1549, 1553, 1576, 1581, 1596, 1607, 1669, 1670, 1762, 1766, 1771, 1774, 1797, 1811, 1819, 1890, 1899, 1923, 1929, 1949, 1953, 1958, 1959, 1974, 1980, 2005, 2054, 2072, 2078, 2104, 2115, 2121, 2124]. traffic [37, 42, 159, 261, 432, 509, 646, 727, 866, 877, 909, 948, 962, 989, 999, 1142, 1489, 1602, 1781, 1903, 1920, 2008, 2131]. Traffic-aware [1535, 1607]. traffic-processing [42]. traffics [1920]. training [509, 912, 1475, 1867]. trajectories [1010, 1088]. Trajectory [288, 294, 555, 843, 928, 1326, 1337, 1401, 1583, 1631, 1747, 1870]. trajectory-resource [1631]. Tram [851]. TranBiLSTM [1783]. transaction [130, 1175]. transactions [581, 659, 897, 2125]. transceivers [1971]. Transfer [100, 196, 661, 875, 1578]. Transferable [1421]. transferring [25]. transfers [832, 1785]. Transformer [1866]. transient [433]. transients [1216]. transition [171]. transitions [1151].

Transmission [115, 335, 343, 506, 713, 756, 860, 959, 1090, 1107, 1212, 1441, 1445, 1448, 1555, 1599, 1601, 1741, 1825, 1878, 1881, 2027, 2091, 2105]. transmission-aware [506]. transmissions [1351, 1591]. transmit [1775]. transmitting [1406]. TransMUSE [1421]. Transparent [96, 1776]. transponder [869]. transport [162, 278, 699, 883, 1699, 1842, 2111]. Transportation [357, 389, 949, 1409, 2093]. transports [641]. traversal [162]. Tree [64, 542, 679, 959, 1152, 1218, 1645, 1687, 1703, 1769, 1967, 2049]. Tree-structured [959]. Trend [668]. Trend-based [668]. trends [104, 1550, 1725, 1758]. trials [1066]. triangle [704]. Trickle [722]. trigger [2010]. trilateration [1082]. Triple [1999]. Triple-step [1999]. troubleshooting [198, 1379, 1431]. Truck [474]. true [286]. TruSD [251]. Trust [251, 258, 259, 355, 430, 448, 576, 802, 831, 862, 945, 951, 1133, 1150, 1162, 1293, 1303, 1322, 1394, 1748, 1857, 1897, 1987, 2087, 2118]. Trust-Aware [430, 831, 1897, 1987]. trust-based [2087]. Trusted [102, 377, 421, 660, 1343, 1712, 1750, 1851, 2053]. TrustedBaaS [421]. trustworthiness [1895]. trustworthy [876, 2095]. truth [432]. Truthful [1282, 1548]. TSBFT [1451]. TSCH [773]. TSCL [1234]. TSCRNN [591]. TSN [1640, 1966]. TSN-based [1640]. TSNMRA [1559]. TTL [1976]. tune [812, 1537]. tuned [120]. tuning [854]. tunnel [752]. tuple [919]. Tutorial [230, 1256]. TVWS [109]. twin [1292, 1655, 1840, 1984, 2042, 2096]. twin-assisted [1655]. twin-enabled [1840]. TwinPeaks [188]. twins [777]. Two [134, 217, 240, 291, 470, 701, 829, 1029, 1179, 1251, 1275, 1299, 1316, 1476, 1800, 1881, 1935, 2014, 2049, 2094]. two-class [1179]. two-dimensional [1251]. two-factor [240, 829]. two-level [2014, 2049]. two-phase [291, 470, 701]. two-stage [1029]. two-step [1299]. two-tier [217, 1275]. two-timescale [2094]. Two-way [134, 1800, 1881]. type [179, 395, 839, 1273]. types [126, 1665, 1972].

U [696, 787]. UAP [530]. UASNs

[988, 1748]. UAV [48, 132, 201, 294, 415, 480, 664, 738, 820, 843, 844, 853, 883, 888, 928, 982, 1035, 1036, 1038, 1088, 1141, 1151, 1326, 1328, 1383, 1410, 1472, 1492, 1502, 1523, 1583, 1621, 1629, 1631, 1664, 1669, 1688, 1738, 1747, 1753, 1787, 1800, 1840, 1855, 1870, 1881, 1939, 1952, 1955, 1986, 2021, 2042, 2126]. UAV-aided [1038, 1326, 1800]. UAV-assisted [415, 738, 1035, 1036, 1410, 1523, 1753, 1787, 1955, 2021]. UAV-based [1088, 1939]. UAV-CFmMIMOs [1629]. UAV-enabled [1141, 1472, 1631, 1747]. UAV-IoT [928]. UAV-to-ground [1855]. UAV-to-UAV [1621]. UAVs [228, 388, 394, 434, 479, 715, 728, 777, 826, 907, 975, 1280, 1327, 1513, 1753]. UAVs-assisted [479, 975]. ubiquitous [1529]. UDNs [900]. UDP [1825]. UDP-based [1825]. UDP-RT [1825]. UE [562]. UE-VBS [562]. Ukraine [1487]. Ultra [208, 286, 321, 453, 562, 580, 738, 766, 906, 1296, 1323, 1393, 1733, 1824, 1885]. ultra-broadband [286]. Ultra-Dense [208, 562, 580, 738, 906, 1323, 1733]. ultra-lightweight [1296]. ultra-low [453, 1824]. ultra-reliable [766, 1885]. ultrabroadband [665]. ultrasonic [121]. ultrasound [602]. unattended [1435, 1765]. Unavailability [1933]. Unavailability-aware [1933]. Unbalanced [1250]. Unbiased [902]. Unboxing [1895]. uncertain [1497, 1665, 2110]. uncertainty [764, 806, 1136]. unclonable [418, 1639]. Uncoded [1014]. Uncoordinated [147]. underground [1681]. underlaid [1278]. underlay [11, 1274]. underlaying [263]. Understanding [85, 230, 255, 554, 796, 922, 1012]. Understood [1803]. Underwater [83, 124, 153, 775, 1127, 1269, 1400, 1460, 1545, 1690, 1764, 1994, 1995, 2006, 2013, 2040, 2070, 2139]. unified [530, 803, 1644, 1705]. unit [1556]. Units [1505]. Universal [324, 1595, 1953]. Universally [177]. university [2069]. unknown [284, 563, 1664, 1923]. Unlicensed [229, 243, 1058, 1988]. unlinkability [328]. Unlocking [1924]. Unmanned [321, 488, 518, 631, 739, 998, 1095, 1109, 1492]. Unobtrusive [22]. Unraveled [1586]. unsaturated [780]. unseen [1136]. unstable [293]. unstructured [198]. Unsupervised [305, 633, 1751, 2130]. untrusted [1881]. Unveiling [680]. unwittingly [915]. up-offloaded [1572]. updatable [1662]. update [1062, 1071, 1321, 1426, 1607]. Updates [236, 257, 916, 1285, 1402]. updating [1923]. UPF [1215, 1450]. upgraded [1555]. uplink [367, 1167, 1210, 1385, 1448, 1481, 1631, 1912, 2088, 2126]. uplink/downlink [1481]. uploading [10, 1658]. Urban [380, 716, 822, 1582, 1721, 1940]. URL [269, 1335, 2130]. URLLC [179, 229, 261, 265, 820, 1448, 1964]. URLs [890]. usability [162]. usage [516, 554, 1459]. usages [1248]. Use [141, 230, 578, 851, 1378, 1525]. used [1901, 2104, 2135]. User [19, 104, 105, 129, 141, 200, 244, 247, 260, 327, 346, 374, 376, 409, 411, 428, 559, 565, 619, 685, 733, 744, 757, 833, 1130, 1231, 1276, 1290, 1328, 1351, 1430, 1434, 1505, 1543, 1605, 1618, 1668, 1669, 1711, 1712, 1738, 1760, 1814, 1829, 1849, 1851, 1883, 1948, 1951, 2022, 2024, 2037, 2069, 2124]. user-cell [2124]. user-centric [105]. user-level [1618]. user-space [1967]. user-to-multiple [744]. users [112, 267, 311, 516, 587, 701, 825, 1048, 1133, 1217, 1557, 2021, 2108]. uses [487]. Using [25, 37, 44, 55, 116, 141, 142, 152, 153, 183, 232, 266, 270, 283, 290, 292, 299, 310, 322, 346, 379, 386, 396, 409, 439, 531, 537, 542, 562, 614, 625, 719, 753, 801, 802, 808, 818, 826, 854, 878, 908, 970, 998, 1013, 1037, 1043, 1047, 1048, 1065– 1067, 1117, 1122, 1153, 1154, 1158, 1163, 1182, 1228, 1229, 1244, 1270, 1274, 1292, 1299, 1301, 1333, 1372, 1384, 1391, 1403, 1416, 1447, 1476, 1550, 1556, 1579, 1617, 1635, 1642, 1645, 1673, 1740, 1754, 1762, 1803, 1813, 1821, 1850, 1858, 1925, 2016, 2045, 2086, 2105, 2112]. using [22, 35, 70, 120, 206, 249, 467, 492, 497, 512, 519, 526, 563, 588, 642, 654, 761, 798, 805, 845, 890, 962, 1032, 1062, 1089, 1142, 1143, 1226, 1288, 1295, 1298, 1329, 1383, 1401, 1449, 1474, 1489, 1492, 1531, 1559, 1570, 1594, 1598, 1648, 1731, 1808, 1890, 1893, 2017, 2027, 2101, 2111, 2130]. Utility [47, 129, 587, 1048, 1194, 1208, 1266, 1843]. utility-maximizing [587]. Utilization [144, 360, 1505]. Utilization-aware [360]. Utilized [259]. utilizing [737, 1984]. UTL DGA22 [1439]. UTXO [461]. UTXO-based [461]. UVLC [1690, 1764]. UW [1400]. UW-GRE [1400]. UWB [894].

v1 [741]. V2G [1459]. V2I [921]. V2V [109]. V2X [81, 1016, 1297, 1440, 1808, 1898]. vague [841]. vagus [116]. Validation [165, 964, 1699, 1791, 1847, 1997]. validations [1723]. validator [971]. VALNET [964]. value [841, 1154]. value-added [1154]. VANET [208]. VANETs [84, 368, 1156, 1173, 1469, 1515, 1563, 1795, 1815, 2105]. Variable [869, 1501, 1657]. variation [151]. Variational [818]. VariBan [1657]. various [1241, 2048]. VB [360]. VBS [562]. VCMaker [867]. VEC [1468]. vector [36, 1299]. vector-coded [36]. Vehicle [206, 464, 808, 817, 851, 1095, 1123, 1273, 1409, 1492, 1572, 1795, 1898]. Vehicle-Based [1123]. Vehicle-cluster-based [1409]. vehicle-to-everything [206]. vehicle-to-infrastructure [464]. vehicle-to-vehicle [1795]. Vehicles [53, 257, 306, 321, 488, 509, 518, 631, 739, 826, 945, 998, 1109, 1211, 1225, 1321, 1471, 1582, 2050, 2083]. VEhicular [119, 145, 146, 381, 384, 389, 393, 433, 520, 534, 617, 643, 755, 769, 781, 830, 858, 949, 1048, 1098, 1148, 1247, 1304, 1362, 1529, 1620, 1730, 1772, 1780, 1805, 1827, 1871, 1888, 1904, 1913, 2018, 2026, 2029, 2050, 2055, 2128]. vendor [500]. ventilation [449]. Verifiable [177, 862, 1166, 1642, 1875, 1893]. verification [60, 102, 184, 236, 460, 957, 1608, 1808]. verifier [1245]. VeriTable [60]. versatile [665]. version [1803]. versus [1365]. vertex [679]. vertical [1320]. very [12, 58]. vessel [564, 1994]. via [59, 145, 179, 733, 888, 967, 1180, 1185, 1211, 1225, 1277, 1334, 1342, 1491, 1548, 1602, 1693, 1700, 1704, 1707, 1878, 1881, 1899, 1973, 2014, 2026, 2040, 2054]. vibration [685]. vibration-based [685]. Video [99, 208, 255, 380, 462, 573, 600, 601, 867, 1009, 1317, 1323, 1380, 1626, 1734, 1741, 1806, 1846, 1905, 1948, 2005, 2084]. Video-on-Demand [380, 2005]. videogame [627]. VideoTrain [1009]. view [577, 855, 1185, 1905]. views [936]. violation [445, 1180]. violations [989]. VIOS [1975]. Virtual [5, 68, 234, 273, 309, 333, 479, 562, 633, 697, 730, 734, 942, 979, 1074, 1183, 1235, 1261, 1338, 1351, 1367, 1390, 1485, 1503, 1532, 1603, 1656, 1735, 1926, 1930, 1933, 2000, 2075]. Virtualization [14, 592, 1157, 1536]. Virtualized [390, 442, 658, 671, 693, 694, 892, 1079, 1302, 1813, 1861, 1889, 2101]. Visible [1690, 1764, 2029]. vision [70]. visit [85]. visualization [896]. VLC [1898, 2024]. VLC-NOMA [2024]. VLC/RF [1898]. VM [885, 892]. VNE [1817]. VNF [52, 91, 241, 330, 331, 472, 527, 835, 939, 1050, 1282, 1708, 1816, 2047]. VNF-AAPC [241]. VNFs [17]. void [775, 1571]. VoIP [772]. Volume [361, 547, 1588, 1765, 1903]. vote [841]. Voting [177, 308]. VPLS [734]. VPN [1819]. VPNs [1414]. VR [1741]. vs [508, 625, 982, 1853]. VSI [1412]. VSI-networks [1412]. vSwitch [546]. vulnerabilities [1438]. vulnerability [986, 1023, 1040, 1267, 1606, 1852, 2023]. vulnerable [141].

wait [558]. waiting [1354, 1687]. wake [321]. wake-up [321]. Walking [1935]. WAN [501, 1200, 2074]. WANs [624, 1392, 1977]. WAPS [1825]. war [1487]. warning [314, 393]. WatchEDGE [2071]. Water [117, 215, 584]. watermarking [1352]. Wave [15, 437, 780, 782, 961, 982, 1069, 1742, 2031]. way [134, 1800, 1881]. WBAN [859, 1494]. WBAN-IoT [859]. WBANs [528, 1627]. weak [64]. weakly [1356]. Wearable [653]. weather [925, 2045]. web [221, 554, 680, 809, 1075, 1131, 1414, 1433, 1565, 2133]. Web3 [1570]. webpage [1242]. WebRTC [2097]. WebRTC-QoE [2097]. webshell [2133]. website [537, 799, 866, 1242, 1368, 1632, 1667, 1983]. websites [85]. weight [834, 981, 1447]. weighted [542, 737]. welfare [1284]. WGAN [1250]. whale [761]. Which [301, 1161]. White [1702]. Who [1568, 1843]. wholly [1792]. Wi [98, 696, 787, 994, 1445, 1617, 1937, 1988]. Wi-Fi [98, 696, 787, 994, 1445, 1617, 1937, 1988]. Wi-Gait [1617]. wide [316, 940, 1341, 1646, 1674, 1918, 2113]. wide-band [1341]. width [1955]. Wifi [47, 170, 279, 322, 353, 391, 507, 764, 804, 921, 1324, 1438, 1447, 2112]. Wifi-based [322]. wild [677, 921, 2009]. wildcard [345]. will [1611]. Window [295, 437, 1084, 1564, 1731]. wired [605]. wired/wireless [605]. Wireless [4, 7, 34, 35, 38, 40, 45, 100, 106, 125, 134, 179, 180, 200, 215, 232, 233, 240, 253, 264, 286, 289, 319, 329, 365, 379, 462, 473, 523, 542, 544, 545, 550, 573, 605, 618, 626, 674, 702, 713, 716, 722, 747, 758, 761, 831, 847, 868, 888, 903, 920, 947, 1019, 1082, 1096, 1101– 1103, 1137, 1159, 1162, 1183, 1190, 1211, 1226, 1253, 1268, 1301, 1346, 1355, 1404, 1412, 1460, 1464, 1490, 1504, 1523, 1540, 1555, 1560, 1569, 1631, 1660, 1685, 1724, 1759, 1794, 1801, 1818, 1820, 1823, 1828, 1833, 1865, 1874, 1876, 1916, 1984, 1994, 2006, 2058, 2114]. wireless [30, 63, 196, 679, 1032, 1073, 1210, 1408, 1580, 1680, 1681, 1698, 1743, 1814]. wireless-powered [4, 1490]. wise [1354]. within [715, 892]. without [1687, 1743, 2051]. WLAN [407, 725, 759, 768, 1086]. WLANs [65, 299, 720, 1124, 1249, 1372, 1505]. WLANs-based [1505]. WNQA [1820]. wolf [30]. work [497, 622, 1174, 1277, 1291]. worker [114, 420, 1279]. worker-selection [114]. workers [532]. workflow [198]. workflows [290, 337, 891, 933]. Workforce [1431]. workgroup [487]. workload [1528, 1665, 1709, 1777, 1889]. workload-dependent [1665]. workloads [175]. world [677, 1983]. worldwide [884]. WoTemu [1075]. WPCNs [11, 196]. wrapper [137]. WSN [454, 539, 717, 789, 834, 1071, 1441]. WSNs [117, 276, 456, 848, 1193, 1559, 1818].

X [1648]. X.509 [1150]. X.509-compatible [1150]. XG [2116]. XG-PON [2116]. XSS [24].

YANG [1452]. YB [1898]. Yinker [1445].

Zeek [1161]. Zero [300, 305, 951, 1120, 1133, 1303, 1596, 1834, 1840, 1876, 1987, 2068, 2118]. zero-day [305]. zero-knowledge [1834]. Zero-relabelling [1596]. Zero-touch [300, 1120, 1840, 1876, 2068]. zero-trust [951, 1133]. zk [2027]. zk-SNARKs [2027]. zone [384, 1534]. zones [1843]. zooming [1393].

References

Anonymous:2020:Ja

[1] Anonymous. 15 January 2020. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):??, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBa

[2] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):Article 107055, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                              sciencedirect.com/science/article/
                                              pii/S1389128619316652.

Melodia:2020:ED

[3] Tommaso Melodia and Antonio Iera. Editorial — December 2019. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):Article 107051, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                              URL http://www.sciencedirect.
                                              com/science/article/pii/S1389128619316615.

Tang:2020:BAC

[4] Kun Tang and Shaowei Liao. Bufferaided cooperative spectrum sharing with full-duplex wireless-powered relay. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): Article 106974, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                              //www.sciencedirect.com/science/
                                              article/pii/S1389128619306966.

He:2020:DEE

[5] Mengyang He, Lei Zhuang, Shuaikui Tian, Guoqing Wang, and Kunli Zhang. DROI: Energy-efficient virtual network embedding algorithm based on dynamic regions of interest. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): Article 106952, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304438.

Noferesti:2020:AAS

[6] Morteza Noferesti and Rasool Jalili. ACoPE: an adaptive semi-supervised learning approach for complex-policy enforcement in high-bandwidth networks. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): Article 106943, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304074.

Huan:2020:PIP

[7] Xintao Huan and Kyeong Soo Kim. On the practical implementation of propagation delay and clock skew compensated high-precision time synchronization schemes with resource-constrained sensor nodes in multi-hop wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):Article 106959, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300398.

Subramanya:2020:MLD

[8] Tejas Subramanya, Davit Harutyunyan, and Roberto Riggio. Machine learning-driven service function chain placement and scaling in MEC-enabled 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):Article 106980, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619310254.

Sharifi:2020:CAE

[9] Seyed Ali Sharifi and Seyed Morteza Babamir. The clustering algorithm for efficient energy management in mobile ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):Article 106983, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619302865.

Liu:2020:SAD

[10] Xiaolan Liu, Bin Yang, Xiaohong Jiang, Lisheng Ma, and Shikai Shen. On social-aware data uploading study of D2D-enabled cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): Article 106955, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619307145.

Yuan:2020:OTF

[11] Lina Yuan, Suzhi Bi, Xiaohui Lin, and Hui Wang. Optimizing throughput fairness of cluster-based cooperation in underlay cognitive WPCNs.

Computer Networks (Amsterdam, Netherlands: 1999), 166(??):Article 106853, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311514.

Roquero:2020:PEC

[12] P. Roquero, E. Magaña, R. Leira, and J. Aracil. Performance evaluation of client-based traffic sniffing for very large populations. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):Article 106985, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302336.

Mao:2020:JHS

[13] Zhifei Mao, Yuming Jiang, Xiaoqiang Di, and Yordanos Woldeyohannes. Joint head selection and airtime allocation for data dissemination in mobile social networks. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):Article 106990, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619303780.

Pham:2020:MAR

[14] Tuan-Minh Pham, Serge Fdida, Thi-Thuy-Lien Nguyen, and Hoai-Nam Chu. Modeling and analysis of robust service composition for network functions virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): Article 106989, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619305080.

Ma:2020:QOJ

[15] Zhongyu Ma, Bo Li, Zhongjiang Yan, and Mao Yang. QoS-Oriented joint optimization of resource allocation and concurrent scheduling in 5G millimeter-wave network. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): Article 106979, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619306668.

Barmpounakis:2020:NSE

[16] Sokratis Barmpounakis, Nikolaos Maroulis, Michael Papadakis, George Tsiatsios, Dimitrios Soukaras, and Nancy Alonistioti. Network slicing — enabled RAN management for 5G: Cross layer control based on SDN and SDR. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): Article 106987, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309776.

Yuan:2020:FTP

[17] Guochang Yuan, Zichuan Xu, Binxu Yang, Weifa Liang, Wei Koong Chai, Daphné Tuncer, Alex Galis, George Pavlou, and Guowei Wu. Fault tolerant placement of stateful VNFs and dynamic fault recovery in cloud networks. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): Article 106953, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300970.

Saddoud:2020:RRM

[18] Ahlem Saddoud, Wael Doghri, Emna Charfi, and Lamia Chaari Fourati. 5G radio resource management approach for multi-traffic IoT communications. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): Article 106936, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303876.

Kazi:2020:CMC

[19] Baha Uddin Kazi and Gabriel Wainer. Coordinated multi-cell cooperation with user centric dynamic coordination station. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): Article 106948, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304335.

Khanouche:2020:FQA

[20] Mohamed Essaid Khanouche, Hania Gadouche, Zoubeyr Farah, and Abdelkamel Tari. Flexible QoS-aware services composition for service computing environments. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):Article 106982, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307169.

deMatos:2020:CIS

[21] Everton de Matos, Ramão Tiago Tiburski, Carlos Roberto Moratelli, Sergio Johann Filho, Leonardo Albernaz Amaral, Gowri Ramachandran, Bhaskar Krishnamachari, and Fabiano Hessel. Context information sharing for the Internet of Things: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): Article 106988, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619310400.

Zhang:2020:URH

[22] Qian Zhang, Run Zhao, Dong Li, and Dong Wang. Unobtrusive and robust human identification using COTS RFID. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): Article 106818, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861831079X.

Li:2020:EMB

[23] Kuanyang Li, Lilan Tu, and Lang Chai. Ensemble-model-based link prediction of complex networks. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):Article 106978, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308710.

Rodriguez:2020:CSS

[24] Germán E. Rodríguez, Jenny G. Torres, Pamela Flores, and Diego E. Benavides. Cross-site scripting (XSS) attacks and mitigation: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): Article 106960, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311247.

Alhowaidi:2020:ESN

[25] Mohammad Alhowaidi, Byrav Ramamurthy, Brian Bockelman, and David Swanson. Enhancing the SDTMA-NDN architecture for transferring the scientific data software using named data networking. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):Article 106954, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302087.

Hu:2020:DPC

[26] Xiaoyan Hu, Shaoqi Zheng, Guoqiang Zhang, Lixia Zhao, Guang Cheng, Jian Gong, and Ruidong Li. An on-demand off-path cache exploration based multipath forwarding strategy. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): Article 107032, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304979.

Anonymous:2020:Fa

[27] Anonymous. 11 February 2020. Computer Networks (Amsterdam, Netherlands: 1999), 167(??):??, February 11, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBb

[28] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 167(??):Article 107075, February 11, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1389128619317529.

Cozza:2020:HLD

[29] Federico Cozza, Alfonso Guarino, Francesco Isernia, Delfina Malandrino, Antonio Rapuano, Raffaele Schiavone, and Rocco Zaccagnino. Hybrid and lightweight detection of third party tracking: Design, implementation, and evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 106993, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619311685.

Xiu-wu:2020:CRA

[30] Y. U. Xiu-wu, Y. U. Hao, Liu Yong, and Xiao Ren-rong. A clustering routing algorithm based on wolf pack algorithm for heterogeneous wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 106994, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308679.

Jaglarz:2020:ELD

[31] Piotr Jaglarz, Piotr Boryło, Andrzej Szymański, and Piotr Chołda. Enhanced Lagrange Decomposition for multi-objective scalable TE in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 106992, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308260.

Pan:2020:PML

[32] Lujia Pan, Jianfeng Zhang, Patrick P. C. Lee, Marcus Kalander, Junjian Ye, and Pinghui Wang. Proactive microwave link anomaly detection in cellular data networks. Computer Networks (Amsterdam, Netherlands: 1999), 167(??):Article 106969, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307662.

Yang:2020:FMS

[33] Ze Yang and Kwan L. Yeung. Flow monitoring scheme design in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 107007, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618306431.

Lattanzi:2020:EEI

[34] Emanuele Lattanzi, Paolo Capellacci, and Valerio Freschi. Experimental evaluation of the impact of packet length on wireless sensor networks subject to interference. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 106986, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619303755.

Raschella:2020:DAP

[35] Alessandro Raschellà, Faycal Bouhafs, Michael Mackay, Qi Shi, Jorge Ortín, José Ramón Gállego, and Maria Canales. A dynamic access point allocation algorithm for dense wireless LANs using potential game. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 106991, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302439.

Soylu:2020:BVC

[36] Tuncay Soylu, Oguzhan Erdem, and Aydin Carus. Bit vector-coded simple CART structure for low latency traffic classification on FP-GAs. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 106977, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304323.

Dominicini:2020:KTS

[37] Cristina K. Dominicini, Gilmar L. Vassoler, Rodolfo Valentim, Rodolfo S. Villaca, Moisés R. N. Ribeiro, Magnos Martinello, and Eduardo Zambon. KeySFC: Traffic steering using strict source routing for dynamic and efficient network orchestration. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 106975, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930194X.

Liu:2020:RDC

[38] Ren-Shiou Liu and Yen-Chen Chen. Robust data collection for energyharvesting wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 167(??):Article 107025, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618310594.

Alam:2020:MNF

[39] Shahid Alam, Soltan Abed Alharbi, and Serdar Yildirim. Mining nested flow of dominant APIs for detecting Android malware. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):Article 107026, February 11, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309855.

Liu:2020:SRA

[40] Zhixin Liu, Mingye Zhao, Yazhou Yuan, and Xinping Guan. Subchannel and resource allocation in cognitive radio sensor network with wireless energy harvesting. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):Article 107028, February 11, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309934.

El-mekkawi:2020:SKM

[41] Ahmed El-mekkawi, Xavier Hesselbach, and Jose Ramon Piney. Squatting and kicking model evaluation for prioritized sliced resource management. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 107006, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619305766.

Yu:2020:WQM

[42] Li Yu, Zongpeng Li, Yucun Zhong, Zhenzhou Ji, and Jiangchuan Liu. When QoE meets learning: a distributed traffic-processing framework for elastic resource provisioning in Het-Nets. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 106904, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618312829.

Singh:2020:FEI

[43] Deepak Singh, Bryan Ng, Yuan-Cheng Lai, Ying-Dar Lin, and Winston K. G. Seah. Full encapsulation or internal buffering in OpenFlow based hardware switches? Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 107033, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308424.

Barakabitze:2020:NSU

[44] Alcardo Alex Barakabitze, Arslan Ahmad, Rashid Mijumbi, and Andrew Hines. 5G network slicing using SDN and NFV: a survey of taxonomy, architectures and future challenges. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 106984, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304773.

Wang:2020:TLT

[45] Tao Wang, Xiang Wang, Wei Shi, Zongmin Zhao, Zhenxue He, and Tongsheng Xia. Target localization and tracking based on improved Bayesian enhanced leastsquares algorithm in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 106968, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618306406.

Al-Musawi:2020:IOL

[46] Bahaa Al-Musawi, Philip Branch, Mohammed Falih Hassan, and Shiva Raj Pokhrel. Identifying OSPF LSA falsification attacks through non-linear analysis. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 107031, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310333.

Swain:2020:NEA

[47] Siba Narayan Swain and C. Siva Ram Murthy. A novel energy-aware utility maximization for efficient deviceto-device communication in LTE-WiFi networks under mixed traffic scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 106995, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308035.

Mukherjee:2020:DAP

[48] Anandarup Mukherjee, Sudip Misra, Anumandala Sukrutha, and Narendra Singh Raghuwanshi. Distributed aerial processing for IoTbased edge UAV swarms in smart farming. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 107038, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619305328.

Angelopoulos:2020:KDE

[49] Constantinos Marios Angelopoulos, Gabriel Filios, Sotiris Nikoletseas, and Theofanis P. Raptis. Keeping data at the edge of smart irrigation networks: a case study in strawberry greenhouses. Computer Networks (Amsterdam, Netherlands: 1999), 167(??): Article 107039, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619305195.

Anonymous:2020:Fb

[50] Anonymous. 26 February 2020. Computer Networks (Amsterdam, Netherlands: 1999), 168(??):??, February 26, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBc

[51] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 168(??):Article 107113, February 26, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1389128620300505.

Yao:2020:FAV

[52] Yifu Yao, Songtao Guo, Pan Li, Guiyan Liu, and Yue Zeng. Forecasting assisted VNF scaling in NFV-enabled networks. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107040, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305705.

Gohar:2020:CEM

[53] Ali Gohar and Sanghwan Lee. A cost efficient multi remote driver selection for remote operated vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107029, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307728.

Sun:2020:ERE

[54] Xiuwen Sun, Hao Li, Dan Zhao, Xingxing Lu, Zheng Peng, and Chengchen Hu. Efficient regular expression matching over compressed traffic. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 106996, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311939.

Elmasry:2020:EDL

[55] Wisam Elmasry, Akhan Akbulut, and Abdul Halim Zaim. Evolving deep learning architectures for network intrusion detection using a double PSO metaheuristic. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):Article 107042, February 26, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930800X.

Vuletic:2020:LNS

[56] Pavle Vuletić, Bartosz Bosak, Marinos Dimolianis, Pascal Mérindol, David Schmitz, and Henrik Wessing. Localization of network service performance degradation in multi-tenant networks. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107050, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619303548.

Wang:2020:MMC

[57] Wenjie Wang, Donghai Tian, Weizhi Meng, Xiaoqi Jia, Runze Zhao, and Rui Ma. MSYM: a multichannel communication system for Android devices. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107024, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309983.

Castagno:2020:LSB

[58] Paolo Castagno, Vincenzo Mancuso, Matteo Sereno, and Marco Ajmone Marsan. Limitations and sidelinkbased extensions of 3GPP cellular access protocols for very crowded environments. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107046, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305377.

Sciancalepore:2020:PEE

[59] Savio Sciancalepore, Omar Adel Ibrahim, Gabriele Oligeri, and Roberto Di Pietro. PiNcH: an effective, efficient, and robust solution to drone detection via network traffic analysis. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107044, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311764.

Liu:2020:VFE

[60] Yaoqing Liu, Garegin Grigoryan, Jun Li, Guchuan Sun, and Tony Tauber. VeriTable: Fast equivalence verification of multiple large forwarding tables. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 106981, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618314105.

Heimovski:2020:FAH

[61] Gustavo B. Heimovski, Rogério C. Turchetti, Juliano A. Wickboldt, Lisandro Z. Granville, and Elias P. Duarte, Jr. FT-Aurora: a highly available IaaS cloud manager based on replication. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107041, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861831346X.

Cerda-Alabern:2020:GNC

[62] Llorenç Cerdà-Alabern, Roger Baig, and Leandro Navarro. On the Guifi.net community network economics. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107067, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309077.

Zhang:2020:LBJ

[63] Ping Zhang, Jianxin Wang, and Wenjun Li. A learning based joint compressive sensing for wireless sensing networks. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107030, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310217.

Madhja:2020:EAT

[64] Adelina Madhja, Sotiris Nikoletseas, and Alexandros A. Voudouris. Energy-aware tree network formation among computationally weak nodes. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107068, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309533.

Ali:2020:CLA

[65] Asad Ali and Faisal Ahmed Khan. Condition and location-aware channel switching scheme for multi-hop multiband WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):Article 107048, February 26, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306863.

Jie:2020:CBO

[66] Yu Jie, Ahmad Alsharoa, Ahmed E. Kamal, and Mohammad Alnuem. A CoMP-Based outage compensation solution for heterogeneous femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107061, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930060X.

Konstantinidis:2020:MCB

[67] Dimitrios Konstantinidis, Vasileios Argyriou, Tania Stathaki, and Nikolaos Grammalidis. A modular CNN-based building detector for remote sensing images. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107034, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304475.

Perakis:2020:CFP

[68] Konstantinos Perakis, Fenareti Lampathaki, Konstantinos Nikas, Yiannis Georgiou, Oskar Marko, and Jarissa Maselyne. CYBELE — Fostering precision agriculture and livestock farming through secure access to large-scale HPC enabled virtual industrial experimentation environments fostering scalable big data analytics. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107035, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305353.

Glaroudis:2020:SCR

[69] Dimitrios Glaroudis, Athanasios Iossifides, and Periklis Chatzimisios. Survey, comparison and research challenges of IoT application protocols for smart farming. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):Article 107037, February 26, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306942.

Wan:2020:FRC

[70] Shaohua Wan and Sotirios Goudos. Faster R-CNN for multi-class fruit detection using a robotic vision system. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107036, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306978.

Sarigiannidis:2020:BDE

[71] Panagiotis Sarigiannidis, Thomas Lagkas, Konstantinos Rantos, and Paolo Bellavista. The Big Data era in IoT-enabled smart farming: Redefining systems, tools, and techniques. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): Article 107043, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316251.

Anonymous:2020:Ma

[72] Anonymous. 14 March 2020. Computer Networks (Amsterdam, Netherlands: 1999), 169(??):??, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBd

[73] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 169(??):Article 107130, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1389128620301341.

Sakib:2020:MAM

[74] Muhammad N. Sakib, Chin-Tser Huang, and Ying-Dar Lin. Maximizing accuracy in multi-scanner malware detection systems. Computer Networks (Amsterdam, Netherlands: 1999), 169 (??):Article 107027, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302518.

Zhong:2020:HNN

[75] Ying Zhong, Wenqi Chen, Zhiliang Wang, Yifan Chen, Kai Wang, Yahui Li, Xia Yin, Xingang Shi, Jiahai Yang, and Keqin Li. HELAD: a novel network anomaly detection model based on heterogeneous ensemble learning. Computer Networks (Amsterdam, Netherlands: 1999), 169(??): Article 107049, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304086.

Xu:2020:ESA

[76] Jianfeng Xu, Liming Wang, and Zhen Xu. An enhanced saturation attack and its mitigation mechanism in software-defined networking. Computer Networks (Amsterdam, Netherlands: 1999), 169(??): Article 107092, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307625.

Ghazvini:2020:LBM

[77] Seyedeh Mina Hosseini Ghazvini, Akbar Ghaffarpour Rahbar, and Behrooz Alizadeh. Load balancing, multipath routing and adaptive modulation with traffic grooming in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 169(??): Article 107081, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309223.

Zhou:2020:OAS

[78] Chongyu Zhou, Chen-Khong Tham, and Mehul Motani. Online auction for scheduling concurrent delay tolerant tasks in crowdsourcing systems. Computer Networks (Amsterdam, Netherlands: 1999), 169(??): Article 107045, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930163X.

Abduljabbar:2020:IMO

[79] Dhuha Abdulhadi Abduljabbar, Siti Zaiton Mohd Hashim, and Roselina Sallehuddin. An improved multiobjective evolutionary algorithm for detecting communities in complex networks with graphlet measure. Computer Networks (Amsterdam, Netherlands: 1999), 169(??):Article 107070, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861931182X.

Kafiloglu:2020:MMS

[80] S. Sinem Kafiloglu, Gürkan Gür, and Fatih Alagöz. A Markovian model for satellite integrated cognitive and D2D HetNets. Computer Networks (Amsterdam, Netherlands: 1999), 169(??): Article 107083, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309235.

Ghosal:2020:SIC

[81] Amrita Ghosal and Mauro Conti. Security issues and challenges in V2X: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 169(??): Article 107093, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619305857.

Gunduz:2020:CSS

[82] Muhammed Zekeriya Gunduz and Resul Das. Cyber-security on smart grid: Threats and potential solutions. Computer Networks (Amsterdam, Netherlands: 1999), 169(??): Article 107094, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619311235.

Liu:2020:TOM

[83] Linfeng Liu, Ran Wang, Gaoxi Xiao, and Dongyue Guo. On the throughput optimization for message dissemination in opportunistic underwater sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 169(??): Article 107097, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618305711.

BenJaballah:2020:SDR

[84] Wafa Ben Jaballah, Mauro Conti, and Chhagan Lal. Security and design requirements for software-defined VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 169(??): Article 107099, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306553.

Zhuang:2020:ULV

[85] Shuying Zhuang, Jessie Hui Wang, Pei Zhang, and Jilong Wang. Understanding the latency to visit websites in China: an infrastructure perspective. Computer Networks (Amsterdam, Netherlands: 1999), 169(??): Article 107102, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308722.

Torres:2020:BSR

[86] Pedro R. Torres, Jr., Alberto García-Martínez, Marcelo Bagnulo, and Eduardo Parente Ribeiro. Bartolomeu: an SDN rebalancing system across multiple interdomain paths. Computer Networks (Amsterdam, Netherlands: 1999), 169(??): Article 107117, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310084.

Swain:2020:NCA

[87] Siba Narayan Swain and C. Siva Ram Murthy. A novel collision aware network assisted device discovery scheme empowering massive D2D communications in 3GPP LTE-A networks. Computer Networks (Amsterdam, Netherlands: 1999), 169(??): Article 107071, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619302130.

Akkari:2020:MMS

[88] Nadine Akkari and Nikos Dimitriou. Mobility management solutions for 5G networks: Architecture and services. Computer Networks (Amsterdam, Netherlands: 1999), 169(??): Article 107082, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619306346. See

corrigendum [361].

Anonymous:2020:Aa

[89] Anonymous. 7 April 2020. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):??, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBe

[90] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):Article 107203, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1389128620303364.

Xie:2020:OVC

[91] An Xie, Huawei Huang, Xiaoliang Wang, Zhuzhong Qian, and Sanglu Lu. Online VNF chain deployment on resource-limited edges by exploiting peer edge devices. Computer Networks (Amsterdam, Netherlands: 1999), 170 (??):Article 107069, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311454.

Li:2020:ESS

[92] Chunlin Li, Yihan Zhang, Zhiqiang Hao, and Youlong Luo. An effective scheduling strategy based on hypergraph partition in geographically distributed datacenters. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):Article 107096, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930684X.

Castillo:2020:IAI

[93] Edwin Ferney Castillo, Oscar Mauricio Caicedo Rendon, Armando Ordonez, and Lisandro Zambenedetti Granville. IPro: an approach for intelligent SDN monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):Article 107108, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128619304608.

Siris:2020:PAM

[94] Vasilios A. Siris. Popularity-aware mobile content management: When is knowing less better? Computer Networks (Amsterdam, Netherlands: 1999), 170(??):Article 107080, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128619309120.

Malik:2020:SRT

[95] Ali Malik, Benjamin Aziz, Mo Adda, and Chih-Heng Ke. Smart routing: Towards proactive fault handling of software-defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):Article 107104, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300271.

Chen:2020:BEB

[96] Kang Chen and Ning Yang. Bw-Share: Efficient bandwidth guarantee in cloud with transparent share adaptation. Computer Networks (Amsterdam, Netherlands: 1999), 170(??): Article 107095, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861931120X.

Bourahla:2020:PPS

[97] Safia Bourahla, Maryline Laurent, and Yacine Challal. Privacy preservation for social networks sequential publishing. Computer Networks (Amsterdam, Netherlands: 1999), 170(??): Article 107106, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305547.

Camacho:2020:LAC

[98] José Camacho, Chris McDonald, Ron Peterson, Xia Zhou, and David Kotz. Longitudinal analysis of a campus Wi-Fi network. Computer Networks (Amsterdam, Netherlands: 1999), 170(??): Article 107103, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308187.

Zhang:2020:ELC

[99] Guoqiang Zhang, Yue Wu, Xu Han, Qian Gao, and Jiasi Chen. Exploiting the layer correlation to improve DASH scheduling with scalable video coding. Computer Networks (Amsterdam, Netherlands: 1999), 170(??): Article 107116, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307777.

Ding:2020:OCP

[100] Xingjian Ding, Yongcai Wang, Guodong Sun, Chuanwen Luo, Deying Li, Wenping Chen, and Qian Hu. Optimal charger placement for wireless power transfer. Computer Networks (Amsterdam, Netherlands: 1999), 170(??): Article 107123, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307947.

Lai:2020:AER

[101] Wei Kuang Lai, Yi Uan Chen, and Tin-Yu Wu. Analysis and evaluation of random-based message propagation models on the social networks. Computer Networks (Amsterdam, Netherlands: 1999), 170(??): Article 107047, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619305602.

Huang:2020:TTE

[102] Xinli Huang, Peng Shi, Yufei Liu, and Fei Xu. Towards trusted and efficient SDN topology discovery: a lightweight topology verification scheme. Computer Networks (Amsterdam, Netherlands: 1999), 170(??): Article 107119, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619307558.

Chao:2020:ACH

[103] Chih-Min Chao, Chia-Tsun Chen, and Hsin-Chung Huang. An adjustable channel hopping algorithm for multi-radio cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 170(??): Article 107107, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619307704.

Wang:2020:UAM

[104] Chen Wang, Yan Wang, Yingying Chen, Hongbo Liu, and Jian Liu. User authentication on mobile devices: Approaches, threats and trends. Computer Networks (Amsterdam, Netherlands: 1999), 170(??): Article 107118, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618312799.

Ullah:2020:PAU

[105] Arif Ullah, Ziaul Haq Abbas, Ghulam Abbas, Fazal Muhammad, and Lei Jiao. Performance analysis of usercentric SBS deployment with load balancing in heterogeneous cellular networks: a Thomas cluster process approach. Computer Networks (Amsterdam, Netherlands: 1999), 170(??): Article 107120, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308874.

Ma:2020:GDP

[106] Xiaoyuan Ma, Peilin Zhang, Oliver Theel, and Jianming Wei. Gathering data with packet-in-packet in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 170 (??):Article 107124, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306255.

Anonymous:2020:Ab

[107] Anonymous. 22 April 2020. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):??, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBf

[108] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):Article 107217, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S138912862030387X.

Midya:2020:QAD

[109] Sadip Midya, Asmita Roy, Koushik Majumder, and Santanu Phadikar. QoS aware distributed dynamic channel allocation for V2V communication in TVWS spectrum. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):Article 107126, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618309228.

Yin:2020:ASM

[110] Wei Yin, Peizhao Hu, Jiahui Wen, and Hongjian Zhou. ACK spoofing on MAC-layer rate control: Attacks and defenses. Computer Networks (Amsterdam, Netherlands: 1999), 171(??): Article 107133, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619311387.

Bozkaya:2020:SED

[111] Elif Bozkaya and Berk Canberk. SDN-enabled deployment and path planning of aerial base stations. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):Article 107125, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313179.

Martin-Lopo:2020:LRI

[112] Miguel M. Martín-Lopo, Jaime Boal, and Álvaro Sánchez-Miralles. A literature review of IoT energy platforms aimed at end users. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):Article 107101, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912861931271X.

Imputato:2020:RDC

[113] Pasquale Imputato, Stefano Avallone, Mohit P. Tahiliani, and Gautam Ramakrishnan. Revisiting design choices in queue disciplines: the PIE case. Computer Networks (Amsterdam, Netherlands: 1999), 171(??): Article 107136, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313775.

Wang:2020:WSI

[114] Yingjie Wang, Yang Gao, Yingshu Li, and Xiangrong Tong. A workerselection incentive mechanism for optimizing platform-centric mobile crowdsourcing systems. Computer Networks (Amsterdam, Netherlands: 1999), 171 (??):Article 107144, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311557.

Zubow:2020:PCM

[115] Anatolij Zubow, Ahmad Rostami, and Suzan Bayhan. On practical cooperative multi point transmission for 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 171(??): Article 107105, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830803X.

Donohoe:2020:DBD

[116] Michael Donohoe, Brendan Jennings, and Sasitharan Balasubramaniam. Deep brain drug-delivery control using vagus nerve communications. Computer Networks (Amsterdam, Netherlands: 1999), 171(??): Article 107137, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861930982X.

Ayadi:2020:STC

[117] Aya Ayadi, Oussama Ghorbel, M. S. BenSalah, and Mohamed Abid. Spatiotemporal correlations for damages identification and localization in water pipeline systems based on WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 171(??): Article 107134, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309946.

Wang:2020:MSD

[118] Yaomin Wang, Xia Wang, Haiyan Li, Yi Dong, Qing Liu, and Xinling Shi.

A multi-service differentiation traffic management strategy in SDN cloud data center. Computer Networks (Amsterdam, Netherlands: 1999), 171(??): Article 107143, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309740.

Wang:2020:GLA

[119] Danxin Wang, Chuanhe Huang, Xieyang Shen, and Naixue Xiong. A general location-authentication based secure participant recruitment scheme for vehicular crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 171(??): Article 107152, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310618.

Vasan:2020:IIB

[120] Danish Vasan, Mamoun Alazab, Sobia Wassan, Hamad Naeem, Babak Safaei, and Qin Zheng. IMCFN: Imagebased malware classification using finetuned convolutional neural network architecture. Computer Networks (Amsterdam, Netherlands: 1999), 171(??): Article 107138, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304736.

Sciacca:2020:IRA

[121] Elisabetta C. Sciacca and Laura Galluccio. Impulse response analysis of an ultrasonic human body channel. Computer Networks (Amsterdam, Netherlands: 1999), 171(??): Article 107149, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619314938.

Mokryn:2020:OCP

[122] Osnat Mokryn, Adi Akavia, and Yossi Kanizo. Optimal cache placement with local sharing: an ISP guide to the benefits of the sharing economy. Computer Networks (Amsterdam, Netherlands: 1999), 171(??): Article 107153, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619301562.

Ahmad:2020:SRP

[123] Abrar Ahmad, Adnan Ahmad Cheema, and Dewar Finlay. A survey of radio propagation channel modelling for low altitude flying base stations. Computer Networks (Amsterdam, Netherlands: 1999), 171(??): Article 107122, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619310692.

Junior:2020:CDC

[124] Eduardo P. M. Câmara Júnior, Luiz F. M. Vieira, and Marcos A. M. Vieira. CAPTAIN: a data collection algorithm for underwater optical-acoustic sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 171(??): Article 107145, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309016.

Dragonas:2020:FAT

[125] Vasileios Dragonas, Georgios Tsoumanis, George Koufoudakis, Asterios Papamichail, Konstantinos Oikonomou, and Ioannis Stavrakakis. A Fairness-Aware topology independent TDMA MAC policy in time constrained wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):Article 107157, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128619316093.

Kotulski:2020:NTI

[126] Zbigniew Kotulski, Tomasz W. Nowak, Mariusz Sepczuk, and Marcin A. Tunia. 5G networks: Types of isolation and their parameters in RAN and CN slices. Computer Networks (Amsterdam, Netherlands: 1999), 171(??): Article 107135, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304797.

Anonymous:2020:Mb

[127] Anonymous. 8 May 2020. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBg

[128] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):Article 107241, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128620304722.

Yucel:2020:USA

[129] Fatih Yucel and Eyuphan Bulut. User satisfaction aware maximum utility task assignment in mobile crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 172(??): Article 107156, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311697.

Erdin:2020:BPN

[130] Enes Erdin, Mumin Cebe, Kemal Akkaya, Senay Solak, Eyuphan Bulut, and Selcuk Uluagac. A Bitcoin payment network with reduced transaction fees and confirmation times. Computer Networks (Amsterdam, Netherlands: 1999), 172(??): Article 107098, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308850.

Lytos:2020:TSF

[131] Anastasios Lytos, Thomas Lagkas, Panagiotis Sarigiannidis, Michalis Zervakis, and George Livanos. Towards smart farming: Systems, frameworks and exploitation of multiple sources. Computer Networks (Amsterdam, Netherlands: 1999), 172(??): Article 107147, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301201.

Radoglou-Grammatikis:2020:CUA

[132] Panagiotis Radoglou-Grammatikis, Panagiotis Sarigiannidis, Thomas Lagkas, and Ioannis Moscholios. A compilation of UAV applications for precision agriculture. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):Article 107148, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912862030116X.

Adhikari:2020:IBF

[133] Devlina Adhikari, Debasish Datta, and Raja Datta. Impact of BER in fragmentation-aware routing and spectrum assignment in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 172(??): Article 107167, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311776.

Nguyen:2020:WEH

[134] Tan N. Nguyen, Phuong T. Tran, and Miroslav Voznak. Wireless energy harvesting meets receiver diversity: a successful approach for two-way half-duplex relay networks over block Rayleigh fading channel. Computer Networks (Amsterdam, Netherlands: 1999), 172(??): Article 107176, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314689.

Bosch:2020:AMI

[135] Patrick Bosch, Steven Latré, and Chris Blondia. An analytical model for IEEE 40802.11 with non-IEEE 40802.11 interfering source. Computer Networks (Amsterdam, Netherlands: 1999), 172 (??):Article 107154, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861931148X.

Abbas:2020:PAT

[136] Nadine Abbas, Sanaa Sharafeddine, Hazem Hajj, and Zaher Dawy. Priceaware traffic splitting in D2D Het-Nets with cost-energy-QoE tradeoffs. Computer Networks (Amsterdam, Netherlands: 1999), 172(??): Article 107169, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619312307.

Khammassi:2020:NLW

[137] Chaouki Khammassi and Saoussen Krichen. A NSGA2-LR wrapper approach for feature selection in network intrusion detection. Computer Networks (Amsterdam, Netherlands: 1999), 172(??): Article 107183, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619315270.

Hou:2020:CRA

[138] Ronghui Hou, Kaiwen Huang, Huilin Xie, King-Shan Lui, and Hongyan Li. Caching and resource allocation in small cell networks.

Computer Networks (Amsterdam, Netherlands: 1999), 172(??):Article 107100, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309351.

Boz:2020:HAQ

[139] Eren Boz and Jukka Manner. A hybrid approach to QoS measurements in cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 172 (??):Article 107158, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314240.

Abbasloo:2020:SSW

[140] Soheil Abbasloo, Yang Xu, and H. Jonathan Chao. To schedule or not to schedule: When no-scheduling can beat the best-known flow scheduling algorithm in datacenter networks. Computer Networks (Amsterdam, Netherlands: 1999), 172(??): Article 107177, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309788.

Barmpounakis:2020:CAU

[141] Sokratis Barmpounakis, George Tsiatsios, Michael Papadakis, Evangelos Mitsianis, Nikolaos Koursioumpas, and Nancy Alonistioti. Collision avoidance in 5G using MEC and NFV: the vulnerable road user safety use case. Computer Networks (Amsterdam, Netherlands: 1999), 172(??): Article 107150, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619315816.

Abbasian:2020:CNA

[142] Amir Abbasian and Masoumeh Safkhani. CNCAA: a new anti-collision algorithm using both collided and non-collided parts of information. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):Article 107159, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619312241.

Xu:2020:GBN

[143] Lingwei Xu, Tianqi Quan, Jingjing Wang, T. Aaron Gulliver, and Khoa N. Le. GR and BP neural networkbased performance prediction of dualantenna mobile communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 172(??): Article 107172, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309545.

Hajisami:2020:ERP

[144] Abolfazl Hajisami, Tuyen X. Tran, Ayman Younis, and Dario Pompili. Elastic resource provisioning for increased energy efficiency and resource utilization in cloud-RANs. Computer Networks (Amsterdam, Netherlands: 1999), 172(??): Article 107170, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316019.

Li:2020:SVC

[145] Xiaotong Li, Ruiting Zhou, Ying-Jun Angela Zhang, Lei Jiao, and Zongpeng Li. Smart vehicular communication via 5G mmWaves. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):Article 107173, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313672.

Zhou:2020:DCI

[146] Man Zhou, Lansheng Han, Hongwei Lu, and Cai Fu. Distributed collaborative intrusion detection system for vehicular ad hoc networks based on invariant. Computer Networks (Amsterdam, Netherlands: 1999), 172(??): Article 107174, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311879.

Cicconetti:2020:UAS

[147] Claudio Cicconetti, Marco Conti, and Andrea Passarella. Uncoordinated access to serverless computing in MEC systems for IoT. Computer Networks (Amsterdam, Netherlands: 1999), 172 (??):Article 107184, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313684.

Anonymous:2020:Mc

[148] Anonymous. 22 May 2020. Computer Networks (Amsterdam, Netherlands: 1999), 173(??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBh

[149] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 173(??):Article 107263, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1389128620305363.

Tu:2020:TAL

[150] Shanshan Tu, Muhammad Waqas, Qiangqiang Lin, Sadaqat Ur Rehman, Muhammad Hanif, Chuangbai Xiao, M. Majid Butt, and Chin-Chen Chang. Tracking area list allocation scheme based on overlapping community algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107182, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309326.

Geng:2020:ODR

[151] Sunyue Geng, Sifeng Liu, Zhigeng Fang, and Su Gao. An optimal delay routing algorithm considering delay variation in the LEO satellite communication network. Computer Networks (Amsterdam, Netherlands: 1999), 173 (??):Article 107166, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313040.

Hosseini:2020:NHM

[152] Soodeh Hosseini and Behnam Mohammad Hasani Zade. New hybrid method for attack detection using combination of evolutionary algorithms, SVM, and ANN. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107168, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302191.

Li:2020:UCM

[153] Zhangyu Li, Soham Desai, Vaishnendr D. Sudev, Pu Wang, Jinsong Han, and Zhi Sun. Underwater cooperative MIMO communications using hybrid acoustic and magnetic induction technique. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107191, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300165.

Suhail:2020:PEP

[154] Sabah Suhail, Rasheed Hussain, Mohammad Abdellatif, Shashi Raj Pandey, Abid Khan, and Choong Seon Hong. Provenance-enabled packet path tracing in the RPL-based Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107189, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316123.

Sanchez:2020:DDS

[155] P. A. Sánchez, S. Luna-Ramírez, M. Toril, C. Gijón, and J. L. Bejarano-Luque. A data-driven scheduler performance model for QoE assessment in a LTE radio network planning tool. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107186, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309338. See

corrigendum [296].

Zhao:2020:RSL

[156] Huiyun Zhao, Zhisong Pan, and Wei Tao. Regularized shapelet learning for scalable time series classification. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107171, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309181.

Takemasa:2020:DPF

[157] Junji Takemasa, Yuki Koizumi, and Toru Hasegawa. Data prefetch for fast NDN software routers based on hash table-based forwarding tables. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107188, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861931196X.

Kondo:2020:NDN

[158] Daishi Kondo, Vassilis Vassiliades, Thomas Silverston, Hideki Tode, and Tohru Asami. The named data networking flow filter: Towards improved security over information leakage attacks. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107187, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316524.

Wu:2020:CMB

[159] Zheng Wu, Yu ning Dong, Hua-Liang Wei, and Wei Tian. Consistency measure based simultaneous feature selection and instance purification for multimedia traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107190, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311624.

Park:2020:AEP

[160] Jeman Park, Manar Mohaisen, Dae-Hun Nyang, and Aziz Mohaisen. Assessing the effectiveness of pulsing denial of service attacks under realistic network synchronization assumptions. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107146, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313118.

Rastogi:2020:NIT

[161] Eshita Rastogi, Navrati Saxena, Abhishek Roy, and Dong Ryeol Shin. Narrowband Internet of Things: a comprehensive study. Computer Networks (Amsterdam, Netherlands: 1999), 173 (??):Article 107209, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619313593.

Barik:2020:UTP

[162] Runa Barik, Michael Welzl, Gorry Fairhurst, Ahmed Elmokashfi, Thomas Dreibholz, and Stein Gjessing. On the usability of transport protocols other than TCP: a home gateway and Internet path traversal study. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107211, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861931446X.

Pacheco:2020:FCH

[163] Fannia Pacheco, Ernesto Exposito, and Mathieu Gineste. A framework to classify heterogeneous Internet traffic with machine learning and deep learning techniques for satellite communications. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107213, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619313544.

Pascoal:2020:SDS

[164] Túlio A. Pascoal, Iguatemi E. Fonseca, and Vivek Nigam. Slow denialof-service attacks on software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 173(??):

Article 107223, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619312927.

Ma:2020:AFS

[165] Lin Ma, Kai Bu, Ningchao Wu, Tianxiang Luo, and Kui Ren. Atlas: a first step toward multipath validation. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107224, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310667.

Martiradonna:2020:ASO

[166] Sergio Martiradonna, Alessandro Grassi, Giuseppe Piro, and Gennaro Boggia. 5G-air-simulator: an opensource tool modeling the 5G air interface. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107151, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619317359.

Huang:2020:ESO

[167] Mingfeng Huang, Anfeng Liu, Neal N. Xiong, Tian Wang, and Athanasios V. Vasilakos. An effective serviceoriented networking management architecture for 5G-enabled Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 173(??): Article 107208, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619311442.

Anonymous:2020:Jb

[168] Anonymous. 19 June 2020. Computer Networks (Amsterdam, Netherlands: 1999), 174(??):??, June 19, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBi

[169] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 174(??):Article 107306, June 19, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1389128620307386.

Zhao:2020:IMM

[170] Yi Zhao, Ke Xu, Yifeng Zhong, Xiang-Yang Li, Ning Wang, Hui Su, Meng Shen, and Ziwei Li. Incentive mechanisms for mobile data offloading through operator-owned WiFi access points. Computer Networks (Amsterdam, Netherlands: 1999), 174(??): Article 107226, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619316858.

Mazzenga:2020:ESG

[171] Franco Mazzenga, Romeo Giuliano, and Francesco Vatalaro. Effective strategies for gradual copperto-fiber transition in access networks. Computer Networks (Amsterdam, Netherlands: 1999), 174(??): Article 107225, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301183.

Zhang:2020:SSA

[172] Shunliang Zhang, Dali Zhu, and Yongming Wang. A survey on spaceaerial-terrestrial integrated 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 174(??): Article 107212, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314045.

Zhang:2020:SAO

[173] Ran Zhang, Jiang Liu, Renchao Xie, Tao Huang, F. Richard Yu, and Yunjie Liu. Service-aware optimal caching placement for named data networking. Computer Networks (Amsterdam, Netherlands: 1999), 174(??): Article 107193, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311193.

Hassan:2020:NBN

[174] Muhammad Hassan, Mauro Conti, and Chhagan Lal. NC based DAS– NDN: Network coding for robust dynamic adaptive streaming over NDN. Computer Networks (Amsterdam, Netherlands: 1999), 174(??): Article 107222, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619312551.

Carrega:2020:CEE

[175] Alessandro Carrega and Matteo Repetto. Coupling energy efficiency and quality for consolidation of cloud workloads. Computer Networks (Amsterdam, Netherlands: 1999), 174(??): Article 107210, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619316226.

Zhou:2020:BEI

[176] Yuyang Zhou, Guang Cheng, Shanqing Jiang, and Mian Dai. Building an efficient intrusion detection system based on feature selection and ensemble classifier. Computer Networks (Amsterdam, Netherlands: 1999), 174(??): Article 107247, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619314203.

Dimitriou:2020:ECF

[177] Tassos Dimitriou. Efficient, coercionfree and universally verifiable Blockchainbased voting. Computer Networks (Amsterdam, Netherlands: 1999), 174 (??):Article 107234, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619317414.

Yousefi:2020:FTA

[178] Faezeh Yousefi, Akbar Ghaffarpour Rahbar, and Amin Ghadesi. Fragmentation and time aware algorithms in spectrum and spatial assignment for space division multiplexed elastic optical networks (SDM-EON). Computer Networks (Amsterdam, Netherlands: 1999), 174(??): Article 107232, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861931597X.

Nouri:2020:MTW

[179] Parisa Nouri, Hirley Alves, Mikko A. Uusitalo, Onel Alcaraz López, and Matti Latva-aho. Machine-type wireless communications enablers for beyond 5G: Enabling URLLC via diversity under hard deadlines. Computer Networks (Amsterdam, Netherlands: 1999), 174(??): Article 107227, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619315361.

Nguyen:2020:MBN

[180] Linh Nguyen and Hoc T. Nguyen. Mobility based network lifetime in wireless sensor networks: a review. Computer Networks (Amsterdam, Netherlands: 1999), 174(??): Article 107236, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619315865.

Galal:2020:PBP

[181] Akram Galal and Xavier Hesselbach. Probability-based path discovery protocol for electromagnetic nanonetworks. Computer Networks (Amsterdam, Netherlands: 1999), 174(??): Article 107246, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308801.

Marche:2020:HES

[182] Claudio Marche, Luigi Atzori, Virginia Pilloni, and Michele Nitti. How to exploit the Social Internet of Things: Query generation model and device profiles’ dataset. Computer Networks (Amsterdam, Netherlands: 1999), 174 (??):Article 107248, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861931730X.

Calvo-Palomino:2020:ECR

[183] Roberto Calvo-Palomino, Héctor Cordobés, Markus Engel, Markus Fuchs, Pratiksha Jain, Marc Liechti, Sreeraj Rajendran, Matthias Schäfer, Bertold Van den Bergh, Sofie Pollin, Domenico Giustiniano, and Vincent Lenders. Electrosense+: Crowdsourcing radio spectrum decoding using IoT receivers. Computer Networks (Amsterdam, Netherlands: 1999), 174(??): Article 107231, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619314471.

Hofer-Schmitz:2020:TFV

[184] Katharina Hofer-Schmitz and Branka Stojanović. Towards formal verification of IoT protocols: a review. Computer Networks (Amsterdam, Netherlands: 1999), 174(??): Article 107233, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619317116.

Curado:2020:EEI

[185] Marilia Curado, Giacomo Tanganelli, Antonio A. F. Loureiro, and Eirini Eleni Tsiropoulou. Editorial: 5G-Enabled Internet of Things, applications and services. Computer Networks (Amsterdam, Netherlands: 1999), 174(??): Article 107229, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620304333.

Anonymous:2020:Jc

[186] Anonymous. 5 July 2020. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):??, July 5, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBj

[187] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):Article 107322, July 5, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128620307957.

Cho:2020:TAC

[188] Eunsang Cho, Jeongnyeo Kim, Minkyung Park, Hyeonmin Lee, Chorom Hamm, Soobin Park, Sungmin Sohn, Minhyeok Kang, and Ted Taekyoung Kwon. TwinPeaks: an approach for certificateless public key distribution for the Internet and Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):

Article 107268, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618314051.

Qian:2020:MMM

[189] Shenshen Qian, Bowen Wang, Song Li, Yanjing Sun, Yi Yu, and Jingjing Wang. Many-to-many matching for social-aware minimized redundancy caching in D2D-enabled cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 175(??): Article 107249, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316299.

Naeem:2020:SRH

[190] Makhdoom Muhammad Naeem, Intesab Hussain, and Malik Muhammad Saad Missen. A survey on registration hijacking attack consequences and protection for session initiation protocol (SIP). Computer Networks (Amsterdam, Netherlands: 1999), 175(??): Article 107250, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619312332.

Sharma:2020:AFT

[191] Kapil Sharma and Ram Narayan Yadav. An adaptive, fault tolerant, flowlevel routing scheme for data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 175(??): Article 107235, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309107.

Ma:2020:CTC

[192] Lujuan Ma, Xiaoping Liu, Huanqing Wang, and Xiaoping Deng. Congestion tracking control for multi-router TCP/ AQM network based on integral backstepping. Computer Networks (Amsterdam, Netherlands: 1999), 175(??): Article 107278, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300013.

Yang:2020:JSC

[193] Yang Yang, Songtao Guo, Guiyan Liu, and Quyuan Wang. Joint source coding rate allocation and flow scheduling for data aggregation in collaborative sensing networks. Computer Networks (Amsterdam, Netherlands: 1999), 175 (??):Article 107269, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314707.

Shen:2020:FSA

[194] Gengbiao Shen, Qing Li, Yong Jiang, Yu Wu, and Jianhui Lv. A four-stage adaptive scheduling scheme for service function chain in NFV. Computer Networks (Amsterdam, Netherlands: 1999), 175(??): Article 107259, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314276.

Antevski:2020:INM

[195] Kiril Antevski, Carlos J. Bernardos, Luca Cominardi, Antonio de la Oliva, and Alain Mourad. On the integration of NFV and MEC technologies: architecture analysis and benefits for edge robotics. Computer Networks (Amsterdam, Netherlands: 1999), 175(??): Article 107274, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620300797.

Zheng:2020:RWP

[196] Yuan Zheng, Suzhi Bi, Xiaohui Lin, and Hui Wang. Reusing wireless power transfer for backscatter-assisted relaying in WPCNs. Computer Networks (Amsterdam, Netherlands: 1999), 175 (??):Article 107277, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309399.

Mollahasani:2020:DAM

[197] Shahram Mollahasani, Alperen Eroglu, Ilker Demirkol, and Ertan Onur. Density-aware mobile networks: Opportunities and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):Article 107271, July 5, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128619309910.

Zhou:2020:LCW

[198] Pengpeng Zhou, Yang Wang, Zhenyu Li, Gareth Tyson, Hongtao Guan, and Gaogang Xie. Logchain: Cloud workflow reconstruction and troubleshooting with unstructured logs. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):Article 107279, July 5, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316731.

Yin:2020:FFA

[199] Wei Yin, Peizhao Hu, Wenbo Wang, Jiahui Wen, and Hongjian Zhou. FA-SUS: a fast association mechanism for 802.11ah networks. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):Article 107287, July 5, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128619312083.

BenHassouna:2020:MUD

[200] Asma Ben Hassouna, Hend Koubaa, and Leila Azouz Saidane. Multi-user diversity wireless multicast: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 175(??): Article 107282, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308564.

Shi:2020:QAU

[201] Lin Shi, Shoukun Xu, Haoyu Liu, and Zhongxu Zhan. QoS-Aware UAV Coverage path planning in 5G mmWave network. Computer Networks (Amsterdam, Netherlands: 1999), 175(??): Article 107207, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619313015.

Anonymous:2020:Jd

[202] Anonymous. 20 July 2020. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBk

[203] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):Article 107354, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1389128620308987.

Kaltenberger:2020:ODI

[204] Florian Kaltenberger, Aloizio P. Silva, Abhimanyu Gosain, Luhan Wang, and Tien-Thinh Nguyen. OpenAirInterface: Democratizing innovation in the 5G Era. Computer Networks (Amsterdam, Netherlands: 1999), 176(??): Article 107284, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619314410.

Dwivedi:2020:ECG

[205] Shubhra Dwivedi, Manu Vardhan, and Sarsij Tripathi. An effect of chaos grasshopper optimization algorithm for protection of network infrastructure. Computer Networks (Amsterdam, Netherlands: 1999), 176(??): Article 107251, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310175.

Rasheed:2020:PPS

[206] Iftikhar Rasheed, Lin Zhang, and Fei Hu. A privacy preserving scheme for vehicle-to-everything communications using 5G mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 176(??): Article 107283, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619315658.

Xia:2020:NOL

[207] Qiufen Xia, Zheng Lou, Wenzheng Xu, and Zichuan Xu. Near-optimal and learning-driven task offloading in a 5G multi-cell mobile edge cloud. Computer Networks (Amsterdam, Netherlands: 1999), 176(??): Article 107276, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301699.

Costa:2020:SBH

[208] Allan Costa, Lucas Pacheco, Denis Rosário, Leandro Villas, Antonio A. F. Loureiro, Susana Sargento, and Eduardo Cerqueira. Skippingbased handover algorithm for video distribution over ultra-dense VANET. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):Article 107252, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620302188.

Favale:2020:CTL

[209] Thomas Favale, Francesca Soro, Martino Trevisan, Idilio Drago, and Marco Mellia. Campus traffic and e-Learning during COVID-19 pandemic. Computer Networks (Amsterdam, Netherlands: 1999), 176(??): Article 107290, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620306046.

Cao:2020:RSE

[210] Zhi Cao, Honggang Zhang, Benyuan Liu, and Bo Sheng. Revenue sharing in edge-cloud systems: a game-theoretic perspective. Computer Networks (Amsterdam, Netherlands: 1999), 176(??): Article 107286, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861931028X.

Desikan:2020:TCF

[211] K. E. Srinivasa Desikan, Vijeth J. Kotagi, and C. Siva Ram Murthy. Topology control in fog computing enabled IoT networks for smart cities. Computer Networks (Amsterdam, Netherlands: 1999), 176(??): Article 107270, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620301468.

Rezar:2020:ITD

[212] Matija Rezar and Fabio Ricciato. On the impact of time-of-departure knowledge on the accuracy of time-of-arrival localization. Computer Networks (Amsterdam, Netherlands: 1999), 176(??): Article 107285, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313532.

Trevisan:2020:ERE

[213] Martino Trevisan, Ali Safari Khatouni, and Danilo Giordano. ERRANT: Realistic emulation of radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 176(??): Article 107289, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301420.

Dong:2020:CCE

[214] Cong Dong, Chen Zhang, Zhigang Lu, Baoxu Liu, and Bo Jiang. CETAnalytics: Comprehensive effective traffic information analytics for encrypted traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 176 (??):Article 107258, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309466.

Olatinwo:2020:EEM

[215] Segun O. Olatinwo and Trudi-H. Joubert. Energy efficiency maximization in a wireless powered IoT sensor network for water quality monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 176(??): Article 107237, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308709.

Zhao:2020:FAT

[216] Zhipeng Zhao, Weidong Yang, and Bin Wu. Flow aggregation through dynamic routing overlaps in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 176 (??):Article 107293, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311818.

Liu:2020:RRA

[217] Zhixin Liu, Guochen Hou, Yazhou Yuan, Kit Yan Chan, Kai Ma, and Xinping Guan. Robust resource allocation in two-tier NOMA heterogeneous networks toward 5G. Computer Networks (Amsterdam, Netherlands: 1999), 176 (??):Article 107299, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314422.

AlKhansa:2020:HSG

[218] Rasha Al Khansa, Hassan A. Artail, Mohamad Assaad, and Karim Kabalan. A Hybrid Scheduled and group-based random access solution for massive MTC networks. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):Article 107253, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310679.

Darabkh:2020:LAG

[219] Khalid A. Darabkh, Wafa’a K. Kassab, and Ala’ F. Khalifeh. LiM-AHP-G-C: Life time maximizing based on analytical hierarchal process and genetic clustering protocol for the Internet of Things environment. Computer Networks (Amsterdam, Netherlands: 1999), 176(??): Article 107257, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316111.

Stamou:2020:CAH

[220] Adamantia Stamou, Nikos Dimitriou, Kimon Kontovasilis, and Symeon Papavassiliou. Context-aware handover management for HetNets: Performance evaluation models and comparative assessment of alternative context acquisition strategies. Computer Networks (Amsterdam, Netherlands: 1999), 176 (??):Article 107272, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619315075.

Alami:2020:EQE

[221] Abdul Jawad Alami and Hassan Artail. Energy and quality of experience aware web browsing for smartphones. Computer Networks (Amsterdam, Netherlands: 1999), 176(??): Article 107256, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619317013.

Anonymous:2020:Ac

[222] Anonymous. 4 August 2020. Computer Networks (Amsterdam, Netherlands: 1999), 177(??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBl

[223] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 177(??):Article 107406, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1389128620310811.

Santos:2020:MTF

[224] Hugo Santos, Derian Alencar, Rodolfo Meneguette, Denis Rosário, Jéferson Nobre, Cristiano Both, Eduardo Cerqueira, and Torsten Braun. A multi-tier fog content orchestrator mechanism with quality of experience support. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107288, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620304448.

Pejovic:2020:RDM

[225] Veljko Pejović, Ivan Majhen, Miha Janez, and Blaz Zupan. RICER-CANDO: Data mining toolkit for mobile broadband measurements. Computer Networks (Amsterdam, Netherlands: 1999), 177(??):Article 107294, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301535.

Zhang:2020:ECN

[226] Hongpo Zhang, Lulu Huang, Chase Q. Wu, and Zhanbo Li. An effective convolutional neural network based on SMOTE and Gaussian mixture model for intrusion detection in imbalanced dataset. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107315, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300712.

Gharib:2020:ASH

[227] Mohammed Gharib, Ahmad Foroozani, Shahbaz Rezaei, Ali Mohammad Afshin Hemmatyar, and Ali Movaghar. An area-scalable human-based mobility model. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107300, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308667.

Alioua:2020:UTM

[228] Ahmed Alioua, Houssem eddine Djeghri, Mohammed Elyazid Tayeb Cherif, Sidi-Mohammed Senouci, and Hichem Sedjelmaci. UAVs for traffic monitoring: a sequential gamebased computation offloading/sharing approach. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107273, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619315798.

Zaki-Hindi:2020:MTU

[229] Ayat Zaki-Hindi, Salah-Eddine Elayoubi, and Tijani Chahed. Multitenancy and URLLC on unlicensed spectrum: Performance and design. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107311, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620300220.

Martiradonna:2020:UAS

[230] Sergio Martiradonna, Alessandro Grassi, Giuseppe Piro, and Gennaro Boggia. Understanding the 5G-air-simulator: a tutorial on design criteria, technical components, and reference use cases. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107314, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619317347.

Sun:2020:MEC

[231] Penghao Sun, Zehua Guo, Gang Wang, Julong Lan, and Yuxiang Hu. MAR-VEL: Enabling controller load balancing in software-defined networks with multi-agent reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107230, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861931566X.

Fu:2020:TOA

[232] Xiuwen Fu, Pasquale Pace, Gianluca Aloi, Lin Yang, and Giancarlo Fortino. Topology optimization against cascading failures on wireless sensor networks using a memetic algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107327, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619317888.

Marinova:2020:EEN

[233] Simona Marinova, Thomas Lin, Hadi Bannazadeh, and Alberto Leon-Garcia. End-to-end network slicing for future wireless in multi-region cloud platforms. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107298, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316081.

Sigwele:2020:EEC

[234] Tshiamo Sigwele, Yim Fun Hu, and Misfa Susanto. Energy-efficient 5G cloud RAN with virtual BBU server consolidation and base station sleeping. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107302, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301742.

Wu:2020:BAA

[235] Lijuan Wu, Chaoqin Gan, Hubao Qiao, Jianqiang Hui, and Zhongsen Xu. Bandwidth allocation algorithm based on differential BRP models in Ethernet PON. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107291, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619315993.

Li:2020:ARV

[236] Yahui Li, Zhiliang Wang, Xia Yin, Xingang Shi, Jianping Wu, Fangdan Ye, Jiangyuan Yao, and Han Zhang. Assisting reachability verification of network configurations updates with NUV. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107326, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619314458.

Sai:2020:EOA

[237] Yinghui Sai, Xiaotong Li, Ruiting Zhou, and Zongpeng Li. An efficient online auction for resource leasing in cloud radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107316, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861930790X.

Madureira:2020:SID

[238] André Luiz R. Madureira, Francisco Renato C. Araújo, and Leobino N.

Sampaio. On supporting IoT data aggregation through programmable data planes. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107330, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301195.

Chatterjee:2020:DBR

[239] Bijoy Chand Chatterjee and Eiji Oki. Defragmentation based on route partitioning in 1+1 protected elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107317, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310138.

Fotouhi:2020:LST

[240] Mahdi Fotouhi, Majid Bayat, Ashok Kumar Das, Hossein Abdi Nasib Far, S. Morteza Pournaghi, and M. A. Doostari. A lightweight and secure two-factor authentication scheme for wireless body area networks in healthcare IoT. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107333, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316457.

Sharma:2020:VAA

[241] Gourav Prateek Sharma, Wouter Tavernier, Didier Colle, and Mario Pickavet. VNF-AAPC: Acceleratoraware VNF placement and chaining. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107329, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862030058X.

Kim:2020:HNS

[242] Sungwook Kim. Heterogeneous network spectrum allocation scheme based on three-phase bargaining game. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107301, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309727.

Campos:2020:AHN

[243] Pablo Campos, Ángela Hernández-Solana, and Antonio Valdovinos-Bardají. Analysis of hidden node problem in LTE networks deployed in unlicensed spectrum. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):Article 107280, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620300396.

Cheng:2020:JUA

[244] Zhipeng Cheng, Ning Chen, Bang Liu, Zhibin Gao, Lianfen Huang, Xiaojiang Du, and Mohsen Guizani. Joint user association and resource allocation in HetNets based on user mobility prediction. Computer Networks (Amsterdam, Netherlands: 1999), 177(??): Article 107312, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862030298X.

Anonymous:2020:S

[245] Anonymous. 4 September 2020. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):??, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBm

[246] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):Article 107458, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311385.

Trivedi:2020:DSA

[247] Hiral S. Trivedi and Sankita J. Patel. Design of secure authentication protocol for dynamic user addition in distributed Internet-of-Things. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107335, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316949.

Uniyal:2020:ETS

[248] Navdeep Uniyal, Abubakar Siddique Muqaddas, Dimitrios Gkounis, Anderson Bravalheri, Shadi Moazzeni, Fragkiskos Sardis, Mischa Dohler, Reza Nejabati, and Dimitra Simeonidou. 5GUK exchange: Towards sustainable end-to-end multi-domain orchestration of softwarized 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107297, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316287.

Wang:2020:PLB

[249] Haibo Wang, Hongli Xu, Chen Qian, Juncheng Ge, Jianchun Liu, and He Huang. PrePass: Load balancing with data plane resource constraints using commodity SDN switches. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107339, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620305272.

Banditwattanawong:2020:FOA

[250] Thepparit Banditwattanawong and Masawee Masdisornchote. On formulation of online algorithm and framework of near-optimally tractable eviction for nonuniform caches. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):Article 107332, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301870.

Kalkan:2020:TTF

[251] Kübra Kalkan and Kasper Rasmussen. TruSD: Trust framework for service discovery among IoT devices. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107318, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311636.

Fulber-Garcia:2020:NST

[252] Vinicius Fulber-Garcia, Elias P. Duarte, Alexandre Huff, and Carlos R. P. dos Santos. Network service topology: Formalization, taxonomy and the CUSTOM specification model. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107337, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619312460.

Wang:2020:EED

[253] Denghui Wang, Jian Liu, and Dezhong Yao. An energy-efficient distributed adaptive cooperative routing based on reinforcement learning in wireless multimedia sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107313, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313568.

Lu:2020:NIE

[254] Mengke Lu. Node importance evaluation based on neighborhood structure hole and improved TOP-SIS. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107336, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861931031X.

Hu:2020:CCU

[255] Yuhan Hu, Ye Tian, Wen Yang, Xiaodong Wang, and Xinming Zhang. Content to cash: Understanding and improving crowdsourced live video broadcasting services with monetary donations. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107281, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619311545.

Joshi:2020:PLP

[256] Kalpana D. Joshi and Kotaro Kataoka. pSMART: a lightweight, privacyaware service function chain orchestration in multi-domain NFV/ SDN. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107295, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619311181.

Halder:2020:SAS

[257] Subir Halder, Amrita Ghosal, and Mauro Conti. Secure over-the-air software updates in connected vehicles: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107343, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619314963.

Jafarian:2020:DAT

[258] Besat Jafarian, Nasser Yazdani, and Mohammad Sayad Haghighi.

Discrimination-aware trust management for social Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):Article 107254, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316743.

Kalkan:2020:SSU

[259] Kübra Kalkan. SUTSEC: SDN utilized trust based secure clustering in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107328, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310783.

Zhou:2020:PCO

[260] Shuchen Zhou and Waqas Jadoon. The partial computation offloading strategy based on game theory for multiuser in mobile edge computing environment. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107334, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316482.

Weerasinghe:2020:PBI

[261] Thilina N. Weerasinghe, Indika A. M. Balapuwaduge, and Frank Y. Li. Priority-based initial access for URLLC traffic in massive IoT networks: Schemes and performance analysis. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):Article 107360, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619317967.

Tarek:2020:NSP

[262] Dina Tarek, Abderrahim Benslimane, M. Darwish, and Amira M. Kotb. A new strategy for packets scheduling in cognitive radio Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107292, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619313994.

Ashtiani:2020:PAR

[263] Amirhossein Feizi Ashtiani and Samuel Pierre. Power allocation and resource assignment for secure D2D communication underlaying cellular networks: a tabu search approach. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107350, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619314392.

Cui:2020:CDA

[264] Jin Cui, Khaled Boussetta, and Fabrice Valois. Classification of data aggregation functions in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107342, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620307143.

Karamyshev:2020:FAA

[265] A. Karamyshev, E. Khorov, A. Krasilov, and I. F. Akyildiz. Fast and accurate analytical tools to estimate network capacity for URLLC in 5G systems. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107331, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316445.

Benil:2020:CBS

[266] T. Benil and J. Jasper. Cloud based security on outsourcing using blockchain in e-health systems. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):Article 107344, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316986.

Sarieddine:2020:FMR

[267] Khaled Sarieddine, Malak Charaf, Mohammad Ayad, and Hassan Artail. A framework for mobile relay node selection for serving outdoor cell edge users. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107359, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620302164.

Lopes:2020:GAP

[268] Ana Paula Golembiouski Lopes and Paulo R. L. Gondim. Group authentication protocol based on aggregated signatures for D2D communication. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107192, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313076.

Wei:2020:AFU

[269] Wei Wei, Qiao Ke, Jakub Nowak, Marcin Korytkowski, Rafał Scherer, and Marcin Woźniak. Accurate and fast URL phishing detector: a convolutional neural network approach. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107275, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301109.

Gupta:2020:CSA

[270] Surbhi Gupta, Sangeeta R, Ravi Shankar Mishra, Gaurav Singal, Tapas Badal, and Deepak Garg. Corridor segmentation for automatic robot navigation in indoor environment using edge devices. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): Article 107374, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620306782.

Anonymous:2020:Oa

[271] Anonymous. 9 October 2020. Computer Networks (Amsterdam, Netherlands: 1999), 179(??):??, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBn

[272] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 179(??):Article 107500, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1389128620311671.

Kibalya:2020:NDP

[273] Godfrey Kibalya, Joan Serrat, Juan-Luis Gorricho, Haipeng Yao, and Peiying Zhang. A novel dynamic programming inspired algorithm for embedding of virtual networks in future networks. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107349, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619317980.

Sicari:2020:ITE

[274] Sabrina Sicari, Alessandra Rizzardi, and Alberto Coen-Porisini. 5G in the Internet of Things era: an overview on security and privacy challenges. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107345, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300827.

Singh:2020:ATS

[275] Vikash Singh, Roshan Kumar, and Zhao Wei. Adaptive time-switching and power-splitting protocols for energy harvesting sensor networks with multiple relays. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):Article 107341, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930862X.

Jeske:2020:DTO

[276] Marlon Jeske, Valério Rosset, and Mariá C. V. Nascimento. Determining the trade-offs between data delivery and energy consumption in large-scale WSNs by multi-objective evolutionary optimization. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):Article 107347, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620303741.

Krzywiecki:2020:SLI

[277] Lukasz Krzywiecki, Adam Bobowski, Marta Słowik, Marcin Słowik, and Patryk Kozieł. Schnorr-like identification scheme resistant to malicious subliminal setting of ephemeral secret. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107346, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301560.

Palmieri:2020:RLA

[278] Francesco Palmieri. A reliability and latency-aware routing framework for 5G transport infrastructures. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107365, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620305132.

Huang:2020:REW

[279] Tao Huang, Bin Tang, Baoliu Ye, Zhihao Qu, and Sanglu Lu. Rateless802.11: Extending WiFi applicability in extremely poor channels. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107361, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316202.

Sun:2020:SIP

[280] Penghao Sun, Zehua Guo, Sen Liu, Julong Lan, Junchao Wang, and Yuxiang Hu. SmartFCT: Improving power-efficiency for data center networks with deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107255, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300384.

Abdelmoneem:2020:MAT

[281] Randa M. Abdelmoneem, Abderrahim Benslimane, and Eman Shaaban. Mobility-aware task scheduling in cloud-fog IoT-based healthcare architectures. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107348, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619313106.

Hussain:2020:APE

[282] Anwar Hussain, Shah Nazir, Sulaiman Khan, and Ayaz Ullah. Analysis of PMIPv6 extensions for identifying and assessing the efforts made for solving the issues in the PMIPv6 domain: a systematic review. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):Article 107366, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619315889.

Abdulqadder:2020:MLI

[283] Ihsan H. Abdulqadder, Shijie Zhou, Deqing Zou, Israa T. Aziz, and Syed Muhammad Abrar Akber. Multilayered intrusion detection and prevention in the SDN/NFV enabled cloud of 5G networks using AI-based defense mechanisms. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107364, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619310205.

Sun:2020:CUP

[284] Fanghui Sun, Shen Wang, Chunrui Zhang, and Hongli Zhang. Clustering of unknown protocol messages based on format comparison. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107296, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862030445X.

Seyfollahi:2020:LLB

[285] Ali Seyfollahi and Ali Ghaffari. A lightweight load balancing and route minimizing solution for routing protocol for low-power and lossy networks. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107368, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310357.

Sen:2020:TPI

[286] Priyangshu Sen, Dimitris A. Pados, Stella N. Batalama, Erik Einarsson, Jonathan P. Bird, and Josep M. Jornet. The TeraNova platform: an integrated testbed for ultra-broadband wireless communications at true Terahertz frequencies. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):Article 107370, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620304473.

deOliveira:2020:BRB

[287] Marcela T. de Oliveira, Lúcio H. A. Reis, Dianne S. V. Medeiros, Ricardo C. Carrano, Sílvia D. Olabarriaga, and Diogo M. F. Mattos. Blockchain reputation-based consensus: a scalable and resilient mechanism for distributed mistrusting applications. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107367, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620300360.

Li:2020:ATR

[288] Ding Li, Wenzhong Li, Xiaoliang Wang, Cam-Tu Nguyen, and Sanglu Lu. App trajectory recognition over encrypted Internet traffic based on deep neural network. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):Article 107372, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620302140.

Bhat:2020:OBL

[289] Soumya J. Bhat and Santhosh K. Venkata. An optimization based localization with area minimization for heterogeneous wireless sensor networks in anisotropic fields. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):Article 107371, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620306083.

Asghari:2020:CRM

[290] Ali Asghari, Mohammad Karim Sohrabi, and Farzin Yaghmaee. A cloud resource management framework for multiple online scientific workflows using cooperative reinforcement learning agents. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107340, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619305791.

Zhang:2020:CFS

[291] Qianying Zhang and Shijun Zhao. A comprehensive formal security analysis and revision of the two-phase key exchange primitive of TPM 2.0. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107369, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619317918.

Alizadeh:2020:CCK

[292] Faezeh Alizadeh and Amir Jalaly Bidgoly. Cipher chaining key resynchronization in LPWAN IoT network using a deep learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107373, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301006.

Aliouat:2020:ERP

[293] Lina Aliouat, Hakim Mabed, and Julien Bourgeois. Efficient routing protocol for concave unstable terahertz nanonetworks. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):Article 107375, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620304205.

Demiane:2020:OUT

[294] Freddy Demiane, Sanaa Sharafeddine, and Omar Farhat. An optimized UAV trajectory planning for localization in disaster scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):Article 107378, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313234.

Kumar:2020:OCC

[295] Sumit Kumar and Rajeev Tiwari. Optimized content centric networking for future Internet: Dynamic popularity window based caching scheme. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107434, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311233.

Sanchez:2020:CDD

[296] P. A. Sánchez, S. Luna-Ramírez, M. Toril, C. Gijón, and J. L. Bejarano-Luque. Corrigendum to “A data-driven scheduler performance model for QoE assessment in a LTE radio network planning tool” Computer Networks 173 (2020) 107186. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): Article 107228, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620304321. See

[155].

Anonymous:2020:Ob

[297] Anonymous. 24 October 2020. Computer Networks (Amsterdam, Netherlands: 1999), 180(??):??, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBo

[298] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 180(??):Article 107540, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1389128620311956.

Lopez-Raventos:2020:CDC

[299] Álvaro López-Raventós and Boris Bellalta. Concurrent decentralized channel allocation and access point selection using multi-armed bandits in multi BSS WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107381, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314823.

Bonati:2020:CZT

[300] Leonardo Bonati, Salvatore D’Oro, Lorenzo Bertizzolo, Emrecan Demirors, Zhangyu Guan, Stefano Basagni, and Tommaso Melodia. CellOS: Zerotouch softwarized open cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107380, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862030503X.

Chouayakh:2020:LSA

[301] Ayman Chouayakh, Aurélien Bechler, Isabel Amigo, Loutfi Nuaymi, and Patrick Maillé. Licensed shared access for 5G: Which auction mechanism to choose? Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107358, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311491.

Muratkar:2020:BLI

[302] Tushar S. Muratkar, Ankit Bhurane, and Ashwin Kothari. Batteryless Internet of Things — a survey. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107385, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620303984.

Li:2020:PPB

[303] Xionglve Li, Zhiping Cai, Bingnan Hou, Ning Liu, Fang Liu, and Jieren Cheng. ProbInfer: Probability-based AS path inference from multigraph perspective. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107377, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620302036.

Sasabe:2020:MEA

[304] Masahiro Sasabe. Mathematical epidemiological analysis of dynamics of delay attacks on pullbased competitive information diffusion. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107383, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310904.

Blaise:2020:DZD

[305] Agathe Blaise, Mathieu Bouet, Vania Conan, and Stefano Secci. Detection of zero-day attacks: an unsupervised port-based approach. Computer Networks (Amsterdam, Netherlands: 1999), 180(??):Article 107391, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300761.

Giannone:2020:OHM

[306] Francesco Giannone, Pantelis A. Frangoudis, Adlen Ksentini, and Luca Valcarenghi. Orchestrating heterogeneous MEC-based applications for connected vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107402, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301997.

Panigrahy:2020:NCD

[307] Nitish K. Panigrahy, Jian Li, Don Towsley, and C. V. Hollot. Network cache design under stationary requests: Exact analysis and Poisson approximation. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107379, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313271.

Vergados:2020:LVN

[308] Dimitrios J. Vergados, Katina Kralevska, Yuming Jiang, and Angelos Michalas. Local voting: a new distributed bandwidth reservation algorithm for 6TiSCH networks. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):Article 107384, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619314628.

Garcia-Aviles:2020:AFF

[309] Gines Garcia-Aviles, Carlos Donato, Marco Gramaglia, Pablo Serrano, and Albert Banchs. ACHO: a framework for flexible re-orchestration of virtual network functions. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):Article 107382, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619317165.

Alostad:2020:DPR

[310] Jasem M. Alostad. Design of power and resource management in OFDMA networks using sleep mode selection technique. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107411, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309600.

Rajaguru:2020:HSS

[311] R. Rajaguru, K. Vimala Devi, and P. Marichamy. A hybrid spectrum sensing approach to select suitable spectrum band for cognitive users. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107387, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620304850.

deSouza:2020:HAI

[312] Cristiano Antonio de Souza, Carlos Becker Westphall, Renato Bobsin Machado, João Bosco Mangueira Sobral, and Gustavo dos Santos Vieira. Hybrid approach to intrusion detection in fog-based IoT environments. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107417, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619315439.

Hu:2020:PAB

[313] Zhiqun Hu, Hang Qi, Xiangming Wen, Zhaoming Lu, and Wenpeng Jing. Performance analysis based Markov chain in random access heterogeneous MIMO networks. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107415, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305341.

Ma:2020:EWD

[314] Lisheng Ma, Wei Su, Bin Wu, Bin Yang, and Xiaohong Jiang. Early warning disaster-aware service protection in geo-distributed data centers. Computer Networks (Amsterdam, Netherlands: 1999), 180(??):

Article 107419, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620300074.

Chousainov:2020:AFC

[315] Iskanter-Alexandros Chousainov, Ioannis Moscholios, Panagiotis Sarigiannidis, Alexandros Kaloxylos, and Michael Logothetis. An analytical framework of a C-RAN supporting random, quasi-random and bursty traffic. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107410, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620306484.

Lin:2020:NTD

[316] Yunsenxiao Lin, Yu Zhou, Zhengzheng Liu, Ke Liu, Yangyang Wang, Mingwei Xu, Jun Bi, Ying Liu, and Jianping Wu. NetView: Towards ondemand network-wide telemetry in the data center. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107386, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620302449.

Kasim:2020:ERD

[317] Ömer Kasim. An efficient and robust deep learning based network anomaly detection against distributed denial of service attacks. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):Article 107390, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620304114.

Xu:2020:IMA

[318] Jin Xu, Zbigniew Dziong, Yan Luxin, Zhe Huang, Ping Xu, and Adnane Cabani. Intelligent multi-agent based C-RAN architecture for 5G radio resource management. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):Article 107418, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314872.

Shahraki:2020:COW

[319] Amin Shahraki, Amir Taherkordi, Øystein Haugen, and Frank Eliassen. Clustering objectives in wireless sensor networks: a survey and research direction analysis. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107376, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620303121.

Spinnewyn:2020:DCN

[320] Bart Spinnewyn, Steven Latré, and Juan Felipe Botero. Delay-constrained NFV orchestration for heterogeneous cloud networks. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):Article 107420, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311099.

Trotta:2020:BDU

[321] Angelo Trotta, Marco Di Felice, Luca Perilli, Eleonora Franchi Scarselli, and Tullio Salmon Cinotti. BEE-DRONES: Ultra low-power monitoring systems based on unmanned aerial vehicles and wake-up radio ground sensors. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107425, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311142.

Nassar:2020:WBL

[322] Mohamed A. Nassar, Mahmud Hasan, Md Khan, Mirza Sultana, Md Hasan, Len Luxford, Peter Cole, Giles Oatley, and Polychronis Koutsakis. Wifi-based localisation datasets for No-GPS open areas using smart bins. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107422, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
acompnetamsterdam2020rticle/pii/
S1389128620311117.

Sasikumar:2020:NMO

[323] Syama Sasikumar and J. Jayakumari. A novel method for the optimization of spectral-energy efficiency tradeoff in 5G heterogeneous cognitive radio network. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107389, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620302115.

Zhang:2020:TCT

[324] Quanxin Zhang, Yuhang Zhao, Yajie Wang, Thar Baker, Jian Zhang, and Jingjing Hu. Towards cross-task universal perturbation against blackbox object detectors in autonomous driving. Computer Networks (Amsterdam, Netherlands: 1999), 180(??): Article 107388, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862030606X.

Anonymous:2020:N

[325] Anonymous. 9 November 2020. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):??, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBp

[326] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):Article 107612, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1389128620312433.

Babamir:2020:MCU

[327] Faezeh Sadat Babamir and Mürvet Kirci. A multibiometric cryptosystem for user authentication in client-server networks. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107427, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311166.

Braeken:2020:SKB

[328] An Braeken. Symmetric key based 5G AKA authentication protocol satisfying anonymity and unlinkability. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107424, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311130.

Karmakar:2020:OLA

[329] Raja Karmakar, Samiran Chattopadhyay, and Sandip Chakraborty. An online learning approach for auto linkconfiguration in IEEE 802.11ac wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107426, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311154.

Bunyakitanon:2020:AAV

[330] Monchai Bunyakitanon, Aloizio Pereira da Silva, Xenofon Vasilakos, Reza Nejabati, and Dimitra Simeonidou. Auto-3P: an autonomous VNF performance prediction and placement framework based on machine learning. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107433, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311221.

VanRossem:2020:VPM

[331] Steven Van Rossem, Wouter Tavernier, Didier Colle, Mario Pickavet, and Piet Demeester. VNF performance modelling: From stand-alone to chained topologies. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107428, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311178.

Bertizzolo:2020:AAS

[332] Lorenzo Bertizzolo, Leonardo Bonati, Emrecan Demirors, Amani Alshawabka, Salvatore D’Oro, Francesco Restuccia, and Tommaso Melodia. Arena: a 64-antenna SDR-based ceiling grid testing platform for sub-6 GHz 5G-and-beyond radio spectrum research. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107436, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311257.

Martins:2020:VNF

[333] Rafael de Jesus Martins, Cristiano Bonato Both, Juliano Araújo Wickboldt, and Lisandro Zambenedetti Granville. Virtual network functions migration cost: from identification to prediction. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107429, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031118X.

Yu:2020:PEP

[334] Genghua Yu, Zhigang Chen, Jia Wu, and Jian Wu. Predicted encounter probability based on dynamic programming proposed probability algorithm in opportunistic social network. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107465, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311440.

Hu:2020:CPM

[335] Chia-Cheng Hu, Jeng-Shyang Pan, Chin-Feng Lai, and Yueh-Min Huang. Content placement for minimizing transmission cost in multiple cloud radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):Article 107430, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311191.

Celdran:2020:PTP

[336] Alberto Huertas Celdrán, Manuel Gil Pérez, Izidor Mlakar, Jose M. Alcaraz Calero, Félix J. García Clemente, Gregorio Martínez Pérez, and Zakirul A. Bhuiyan. PROTECTOR: Towards the protection of sensitive data in Europe and the US. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107448, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311300.

Pasdar:2020:HSS

[337] Amirmohammad Pasdar, Young Choon Lee, and Khaled Almi’ani. Hybrid scheduling for scientific workflows on hybrid clouds. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):Article 107438, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311269.

Li:2020:TMO

[338] Junfei Li, Penghao Sun, and Yuxiang Hu. Traffic modeling and optimization in datacenters with graph neural network. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107528, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311865.

Jafari-Beyrami:2020:DFA

[339] Mohammad Jafari-Beyrami, Akbar Ghaffarpour Rahbar, and Soheil Hosseini. On-demand fragmentation-aware spectrum allocation in space division multiplexed elastic optical networks with minimized crosstalk and multipath routing. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107531, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311889.

Ali:2020:MSB

[340] Hamida Qumber Ali and Sayeed Ghani. Multi-sensor based M_{K} /Hyper_{K} /1/ M queuing model for heterogeneous traffic. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107512, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311774.

Anees:2020:DAE

[341] Junaid Anees, Hao-Chun Zhang, Bachirou Guene Lougou, Sobia Baig, and Yabibal Getahun Dessie. Delay aware energy-efficient opportunistic node selection in restricted routing. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107536, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311919.

Araujo:2020:SCR

[342] Carlos M. Araújo, João Marcos P. Silva, Anand Subramanian, and Iguatemi E. Fonseca. On solving the capacitated routing and spectrum allocation problem for flexgrid optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107535, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311907.

Park:2020:AAC

[343] Eunjeong Park, Myung-Sup Lee, Hyung-Sin Kim, and Saewoong Bahk. AdaptaBLE: Adaptive control of data rate, transmission power, and connection interval in Bluetooth low energy. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107520, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311816.

Hu:2020:PAS

[344] Guangwu Hu, Qing Li, Shuo Ai, Tan Chen, Jingpu Duan, and Yu Wu. A proactive auto-scaling scheme with latency guarantees for multi-tenant NFV cloud. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107552, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312032.

Alimohammadi:2020:CNW

[345] Hamed Alimohammadi and Mahmood Ahmadi. Common non-wildcard portion-based partitioning approach to SDN many-field packet classification. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107534, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311890.

Labayen:2020:OCU

[346] Víctor Labayen, Eduardo Magaña, Daniel Morató, and Mikel Izal. Online classification of user activities using machine learning on network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):Article 107557, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312081.

Gorlatova:2020:CTC

[347] Maria Gorlatova, Hazer Inaltekin, and Mung Chiang. Characterizing task completion latencies in multipoint multi-quality fog computing systems. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107526, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311853.

Tanaka:2020:EAL

[348] Kyosuke Tanaka, Hayato Yamaki, Shinobu Miwa, and Hiroki Honda. Evaluating architecture-level optimization in packet processing caches. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):Article 107550, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312020.

Leal:2020:RPS

[349] Alexánder Leal, Mario Durán, and Juan Felipe Botero. Reliability provision in software defined power substations communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):Article 107560, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312093.

Karatas:2020:MOB

[350] Mumtaz Karatas. A multi-objective bi-level location problem for heterogeneous sensor networks with hub-spoke topology. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107551, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312019.

Safkhani:2020:IMP

[351] Masoumeh Safkhani, Samad Rostampour, Ygal Bendavid, and Nasour Bagheri. IoT in medical and pharmaceutical: Designing lightweight RFID security protocols for ensuring supply chain integrity. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):Article 107558, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312068.

Vikhrova:2020:GBD

[352] Olga Vikhrova, Sara Pizzi, Antonella Molinaro, Antonio Iera, Konstantin Samouylov, and Giuseppe Araniti. Group-based delivery of critical traffic in cellular IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107563, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312111.

Fan:2020:LEO

[353] Wenhao Fan, Junting Han, Le Yao, Fan Wu, and Yuan’an Liu. Latency-energy optimization for joint WiFi and cellular offloading in mobile edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):

Article 107570, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312159.

Chen:2020:FNH

[354] Lin Chen, Haipeng Dai, Lei Meng, and Jihong Yu. Finding needles in a hay stream: On persistent item lookup in data streams. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):Article 107518, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311804.

Aliyu:2020:TMF

[355] Aliyu Lawal Aliyu, Adel Aneiba, Mohammad Patwary, and Peter Bull. A trust management framework for Software Defined Network (SDN) controller and network applications. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107421, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311105.

Bai:2020:ABS

[356] Yang Bai, Li Lu, Jerry Cheng, Jian Liu, Yingying Chen, and Jiadi Yu. Acoustic-based sensing and applications: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):Article 107447, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311282.

Boukerche:2020:MLB

[357] Azzedine Boukerche and Jiahao Wang. Machine learning-based traffic prediction models for Intelligent Transportation Systems. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):Article 107530, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311877.

Pal:2020:NNF

[358] Amitangshu Pal and Krishna Kant. NFMI: Near Field Magnetic Induction based communication. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):Article 107548, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128620312007.

Chen:2020:SMM

[359] Yuanyuan Chen, Jing Qiu, Xiaojiang Du, Lihua Yin, and Zhihong Tian. Security of mobile multimedia data: The adversarial examples for spatiotemporal data. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):Article 107432, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031121X.

Mahapatra:2020:UAV

[360] Byomakesh Mahapatra, Ashok Kumar Turuk, Sanket Kumar Panda, and Sarat Kumar Patra. Utilization-aware VB migration strategy for inter-BBU load balancing in 5G cloud radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107507, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311725.

Akkari:2020:CMM

[361] Nadine Akkari and Nikos Dimitriou. Corrigendum to “Mobility Management Solutions for 5G: Architecture and Services” Computer Networks Volume 169 (2020) 107082. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): Article 107561, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862031210X. See

[88].

Anonymous:2020:Da

[362] Anonymous. 9 December 2020. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):??, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBq

[363] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):Article 107664, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1389128620312779.

Fazio:2020:PMM

[364] Peppino Fazio, Mauro Tropea, Miroslav Voznak, and Floriano De Rango. On packet marking and Markov modeling for IP traceback: a deep probabilistic and stochastic analysis. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107464, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311439.

Khaki:2020:IMM

[365] Mahmood Khaki and Abdorasoul Ghasemi. The impact of mobility model on handover rate in heterogeneous multi-tier wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107454, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311348.

Zhang:2020:JTO

[366] Ni Zhang, Songtao Guo, Yifan Dong, and Defang Liu. Joint task offloading and data caching in mobile edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):Article 107446, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311312.

Baidas:2020:RAS

[367] Mohammed W. Baidas, Emad Alsusa, and Yao Shi. Resource allocation for SWIPT-enabled energy-harvesting downlink/uplink clustered NOMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):

Article 107471, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311476.

Mchergui:2020:BBS

[368] Abir Mchergui, Tarek Moulahi, and Salem Nasri. BaaS: Broadcast as a service cross-layer learningbased approach in cloud assisted VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107468, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311464.

Muller:2020:STI

[369] Lucas Müller, Matthew Luckie, Bradley Huffaker, kc claffy, and Marinho Barcellos. Spoofed traffic inference at IXPs: Challenges, methods and analysis. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107452, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311336.

Daldoul:2020:PEO

[370] Yousri Daldoul, Djamal-Eddine Meddour, and Adlen Ksentini. Performance evaluation of OFDMA and MU-MIMO in 802.11ax networks. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107477, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311531.

Tuysuz:2020:SSG

[371] Mehmet Fatih Tuysuz and Ramona Trestian. From serendipity to sustainable green IoT: Technical, industrial and political perspective. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107469, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311452.

Dehkordi:2020:ENR

[372] Mohammad Jafari Dehkordi and Babak Sadeghiyan. An effective noderemoval method against P2P botnets. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107488, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311592.

Iqbal:2020:DGK

[373] Salman Iqbal, Miss Laiha Mat Kiah, Aqeel ur Rehman, Zahid Abbas, and Babak Daghighi. DM-GKM: a key management scheme for dynamic group based applications. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107476, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031152X.

Taleb:2020:FDA

[374] Hussein Taleb, Kinda Khawam, Samer Lahoud, Melhem El Helou, and Steven Martin. A fully distributed approach for joint user association and RRH clustering in cloud radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107445, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311294.

Polverini:2020:IDS

[375] Marco Polverini, Jaime Galán-Jiménez, Francesco G. Lavacca, Antonio Cianfrani, and Vincenzo Eramo. Improving dynamic service function chaining classification in NFV/SDN networks through the offloading concept. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107480, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311543.

Lalouani:2020:MOR

[376] Wassila Lalouani and Mohamed Younis. Multi-observable reputation scoring system for flagging suspicious user sessions. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107474, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311506.

Tseng:2020:RBT

[377] Lewis Tseng, Yingjian Wu, Haochen Pan, Moayad Aloqaily, and Azzedine Boukerche. Reliable broadcast with trusted nodes: Energy reduction, resilience, and speed. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):Article 107486, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311579.

Ayoub:2020:MIS

[378] Wael Ayoub, Abed Ellatif Samhat, Fabienne Nouvel, Mohamad Mroue, Hassan Jradi, and Jean-Christophe Prévotet. Media independent solution for mobility management in heterogeneous LPWAN technologies. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):Article 107423, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311129.

Pacheco-Paramo:2020:DAD

[379] Diego Pacheco-Paramo and Luis Tello-Oquendo. Delay-aware dynamic access control for mMTC in wireless networks using deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107493, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311610.

Begin:2020:DVD

[380] Thomas Begin, Anthony Busson, Isabelle Guérin Lassous, and Azzedine Boukerche. Delivering video-on-demand services with IEEE 802.11p to major non-urban roads: a stochastic performance analysis. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):Article 107440, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128620311270.

Chouikhi:2020:GNE

[381] Samira Chouikhi, Lyes Khoukhi, Moez Esseghir, and Leila Merghem-Boulahia. Generalized Nash equilibrium approach for radio resource sharing and power allocation in vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107490, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311580.

Candela:2020:ICP

[382] Massimo Candela, Valerio Luconi, and Alessio Vecchio. Impact of the COVID-19 pandemic on the Internet latency: a large-scale study. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):Article 107495, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311622.

Loubany:2020:AAS

[383] Ali Loubany, Samer Lahoud, and Rida El Chall. Adaptive algorithm for spreading factor selection in Lo-RaWAN networks with multiple gateways. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107491, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311609.

Gallego-Madrid:2020:EZE

[384] Jorge Gallego-Madrid, Ramon Sanchez-Iborra, Jose Santa, and Antonio Skarmeta. Evaluation of a zone encryption scheme for vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107523, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031183X.

Shahryari:2020:EED

[385] Om-Kolsoom Shahryari, Hossein Pedram, Vahid Khajehvand, and Mehdi Dehghan TakhtFooladi. Energy-efficient and delay-guaranteed computation offloading for fog-based IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107511, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311762.

Mei:2020:RMB

[386] Lifan Mei, Runchen Hu, Houwei Cao, Yong Liu, Zifan Han, Feng Li, and Jin Li. Realtime mobile bandwidth prediction using LSTM neural network and Bayesian fusion. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):Article 107515, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311798.

Shakarami:2020:SCO

[387] Ali Shakarami, Mostafa Ghobaei-Arani, and Ali Shahidinejad. A survey on the computation offloading approaches in mobile edge computing: a machine learning-based perspective. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107496, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311634.

Mishra:2020:SCC

[388] Debashisha Mishra and Enrico Natalizio. A survey on cellular-connected UAVs: Design challenges, enabling 5G/ B5G innovations, and experimental advancements. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107451, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311324.

Boukerche:2020:AIB

[389] Azzedine Boukerche, Yanjie Tao, and Peng Sun. Artificial intelligence-based vehicular traffic flow prediction methods for supporting intelligent transportation systems. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):Article 107484, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311567.

Bonati:2020:OPV

[390] Leonardo Bonati, Michele Polese, Salvatore D’Oro, Stefano Basagni, and Tommaso Melodia. Open, programmable, and virtualized 5G networks: State-of-the-art and the road ahead. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107516, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311786.

Al-shawabka:2020:MSQ

[391] Amani Al-shawabka, Francesco Restuccia, Salvatore D’Oro, and Tommaso Melodia. Massive-scale I/Q datasets for WiFi radio fingerprinting. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107566, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312123.

Zhao:2020:MMA

[392] Xuhui Zhao, Yan Shi, and Shanzhi Chen. MAESP: Mobility aware edge service placement in mobile edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107435, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311245.

Ullah:2020:RTD

[393] Noor Ullah, Xiangjie Kong, Limei Lin, Mubarak Alrashoud, Amr Tolba, and Feng Xia. Real-time dissemination of emergency warning messages in 5G enabled selfish vehicular social networks. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107482, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311555.

Ullah:2020:UJO

[394] Zaib Ullah, Fadi Al-Turjman, Uzair Moatasim, Leonardo Mostarda, and Roberto Gagliardi. UAVs joint optimization problems and machine learning to improve the 5G and beyond communication. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107478, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311518.

Sabuj:2020:ANO

[395] Saifur Rahman Sabuj, A. M. Musa Shakib Khan, and Masanori Hamamura. Application of non-orthogonal multiple access for machine type communication in sub-terahertz band. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):Article 107508, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311737.

Briones-Reyes:2020:MES

[396] Arturo Briones-Reyes, Luis Alberto Vásquez-Toledo, Alfonso Prieto-Guerrero, and Rafael Aguilar-Gonzalez. Mathematical evaluation of spectrum sharing in cognitive radio networks for 5G systems using Markov processes. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107521, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311828.

Roy:2020:AEM

[397] Palash Roy, Sujan Sarker, Md. Abdur Razzaque, Mohammad Mehedi Hassan, Salman A. AlQahtani, Gianluca Aloi, and Giancarlo Fortino. AI-enabled mobile multimedia service instance placement scheme in mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): Article 107573, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312160.

Anonymous:2020:Db

[398] Anonymous. 24 December 2020. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2020:EBr

[399] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):Article 107716, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313116.

Jiang:2020:CFL

[400] Hui Jiang, Min Liu, Bo Yang, Qingxiang Liu, Jizhong Li, and Xiaobing Guo. Customized federated learning for accelerated edge computing with heterogeneous task targets. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): Article 107569, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312135.

Bernieri:2020:TNA

[401] Giuseppe Bernieri, Stefano Cecconello, Mauro Conti, and Gianluca Lain. TAMBUS: a novel authentication method through covert channels for securing industrial networks. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): Article 107583, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312214.

Mashhadi:2020:OAD

[402] Farshad Mashhadi, Sergio A. Salinas Monroy, Arash Bozorgchenani, and Daniele Tarchi. Optimal auction for delay and energy constrained task offloading in mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): Article 107527, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311841.

Marin:2020:SBS

[403] Andrea Marin, Sabina Rossi, and Carlo Zen. Size-based scheduling for TCP flows: Implementation and performance evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):Article 107574, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312172.

Liu:2020:PPP

[404] Yuxian Liu, Fagui Liu, Hao-Tian Wu, Xinglin Zhang, Bowen Zhao, and Xingfu Yan. PriDPM: Privacypreserving dynamic pricing mechanism for robust crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):Article 107582, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312226.

Hajian:2020:SSH

[405] R. Hajian, S. ZakeriKia, S. H. Erfani, and M. Mirabi. SHAPARAK: Scalable healthcare authentication protocol with attack-resilience and anonymous key-agreement. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):Article 107567, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312147.

Sun:2020:EFM

[406] Penghao Sun, Julong Lan, Junfei Li, Zehua Guo, Yuxiang Hu, and Tao Hu. Efficient flow migration for NFV with graph-aware deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): Article 107575, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312184.

Abdulshakoor:2020:OAM

[407] Ahmed I. Abdulshakoor, M. G. Anany, and Mahmoud M. Elmesalawy. Outage-aware matching game approach for cell selection in LTE/ WLAN multi-RAT HetNets. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):Article 107596, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312317.

Calhan:2020:HSS

[408] Ali Çalhan and Murtaza Cicioglu. Handover scheme for 5G small cell networks with non-orthogonal multiple access. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): Article 107601, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312342.

Mahdizadeh:2020:RFP

[409] Mohammad Saleh Mahdizadeh and Behnam Bahrak. A regression framework for predicting user’s next location using Call Detail Records. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):Article 107618, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312470.

Redzovic:2020:IPC

[410] Hasan Redzović, Aleksandra Smiljanić, and Mihailo Vesović. Implementation and performance comparison of high-capacity software routers. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): Article 107585, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312238.

Sun:2020:PEP

[411] Heli Sun, Shan Zhang, Jianbin Huang, Liang He, Xiaolong Jiang, and Zhangtian Duan. A personalized eventparticipant arrangement framework based on user interests in social network. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): Article 107607, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031238X.

Li:2020:IFM

[412] Xiaoqian Li and Kwan L. Yeung. ILP formulations for monitoringcycle design based on segment routing. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):Article 107605, December 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312378.

Dai:2020:JAB

[413] Xiangwen Dai and Jinsong Gui. Joint access and backhaul resource allocation for D2D-assisted dense mmWave cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): Article 107602, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312354.

Hu:2020:SEA

[414] Tao Hu, Peng Yi, Yuxiang Hu, Julong Lan, Zhen Zhang, and Ziyong Li. SAIDE: Efficient application interference detection and elimination in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): Article 107619, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312494.

Ye:2020:OOE

[415] Weidu Ye, Junzhou Luo, Feng Shan, Wenjia Wu, and Ming Yang. Offspeeding: Optimal energy-efficient flight speed scheduling for UAV-assisted edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): Article 107577, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312196.

Zhang:2020:TAI

[416] Shunliang Zhang and Dali Zhu. Towards artificial intelligence enabled 6G: State of the art, challenges, and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): Article 107556, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031207X.

Guidotti:2020:ASP

[417] A. Guidotti, S. Cioni, G. Colavolpe, M. Conti, T. Foggi, A. Mengali, G. Montorsi, A. Piemontese, and A. Vanelli-Coralli. Architectures, standardisation, and procedures for 5G satellite communications: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):Article 107588, December 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031224X.

Shamsoshoara:2020:SPU

[418] Alireza Shamsoshoara, Ashwija Korenda, Fatemeh Afghah, and Sherali Zeadally. A survey on physical unclonable function (PUF)-based security solutions for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): Article 107593, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312275.

Chodak:2020:HLC

[419] Grzegorz Chodak, Grazyna Suchacka, and Yash Chawla. HTTP-level ecommerce data based on server access logs for an online store. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):Article 107589, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312263.

Kuang:2020:SWD

[420] Li Kuang, Huan Zhang, Ruyi Shi, Zhifang Liao, and Xiaoxian Yang. A spam worker detection approach based on heterogeneous network embedding in crowdsourcing platforms. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): Article 107587, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312251.

Ma:2020:TBE

[421] Zhaofeng Ma, Weizhe Zhao, Shoushan Luo, and Lingyun Wang. TrustedBaaS: Blockchain-enabled distributed and higher-level trusted platform. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):Article 107600, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312330.

Llerena:2020:SAS

[422] Yarisley Peña Llerena and Paulo R. L. Gondim. Social-aware spectrum sharing for D2D communication by artificial bee colony optimization. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): Article 107581, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312202.

Anonymous:2021:Ja

[423] Anonymous. 15 January 2021. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):??, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBa

[424] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):Article 107775, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1389128620313517.

Adil:2021:CFO

[425] Muhammad Adil. Congestion free opportunistic multipath routing load balancing scheme for Internet of Things (IoT). Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107707, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313049.

Zeb:2021:JPS

[426] Jehan Zeb, Aamir Hassan, and Muhammad Danish Nisar. Joint power and spectrum allocation for D2D communication overlaying cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107683, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312913.

Chao:2021:LAA

[427] Chih-Min Chao and Wei-Che Lee. Load-aware anti-jamming channel hopping design for cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107681, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312901.

Palumbo:2021:CAC

[428] Fabio Palumbo, Giuseppe Aceto, Alessio Botta, Domenico Ciuonzo, Valerio Persico, and Antonio Pescapé. Characterization and analysis of cloudto-user latency: the case of Azure and AWS. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107693, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312962.

Zeng:2021:SCM

[429] Yue Zeng, Baoliu Ye, Bin Tang, Songtao Guo, and Zhihao Qu. Scheduling coflows of multi-stage jobs under network resource constraints. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):Article 107686, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312937.

Liu:2021:TAS

[430] Xiuwen Liu, Jianming Fu, Yanjiao Chen, Weichen Luo, and Zihan Tang. Trust-aware sensing quality estimation for team crowdsourcing in social IoT. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107695, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312986.

Pushpalatha:2021:RRP

[431] M. Pushpalatha, T. Anusha, T. Rama Rao, and Revathi Venkataraman. L-RPL: RPL powered by Laplacian energy for stable path selection during link failures in an Internet of Things network. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107697, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312998.

Zaki:2021:GGG

[432] Faiz Zaki, Abdullah Gani, Hamid Tahaei, Steven Furnell, and Nor Badrul Anuar. Grano-GT: a granular ground truth collection tool for encrypted browser-based Internet traffic. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107617, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312482.

Amadeo:2021:DIC

[433] Marica Amadeo, Giuseppe Ruggeri, Claudia Campolo, and Antonella Molinaro. Diversity-improved caching of popular transient contents in Vehicular Named Data Networking. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107625, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312548.

Rottondi:2021:SET

[434] Cristina Rottondi, Francesco Malandrino, Andrea Bianco, Carla Fabiana Chiasserini, and Ioannis Stavrakakis. Scheduling of emergency tasks for multiservice UAVs in post-disaster scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107644, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312639.

Zeng:2021:FDD

[435] Yuwei Zeng, Xiaochun Yun, Xunxun Chen, Boquan Li, Haiwei Tsang, Yipeng Wang, Tianning Zang, and Yongzheng Zhang. Finding disposable domain names: a linguistics-based stacking approach. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):Article 107642, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312627.

Sviridov:2021:LDR

[436] German Sviridov, Marco Bonola, Angelo Tulumello, Paolo Giaccone, Andrea Bianco, and Giuseppe Bianchi. LOcAl DEcisions on Replicated States (LOADER) in programmable dataplanes: Programming abstraction and experimental evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):Article 107637, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312597.

Abdelsalam:2021:TWL

[437] A. Abdelsalam, M. Luglio, N. Patriciello, C. Roseti, and F. Zampognaro. TCP Wave over Linux: a disruptive alternative to the traditional TCP window approach. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107633, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312585.

Alonso:2021:MOO

[438] Rodney Martinez Alonso, David Plets, Margot Deruyck, Luc Martens, Glauco Guillen Nieto, and Wout Joseph. Multi-objective optimization of cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):Article 107651, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862031269X.

Chen:2021:LSB

[439] Ning Chen, Sheng Zhang, Jie Wu, Zhuzhong Qian, and Sanglu Lu. Learning scheduling bursty requests in Mobile Edge Computing using DeepLoad. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107655, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312706.

Midoglu:2021:LSS

[440] Cise Midoglu, Konstantinos Kousias, Özgü Alay, Andra Lutu, Antonios Argyriou, Michael Riegler, and Carsten Griwodz. Large scale “speedtest” experimentation in Mobile Broadband Networks. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107629, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312561.

Quincozes:2021:SID

[441] Silvio E. Quincozes, Célio Albuquerque, Diego Passos, and Daniel Mossé. A survey on intrusion detection and prevention systems in digital substations. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107679, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312895.

Suraci:2021:SOS

[442] Chiara Suraci, Giuseppe Araniti, Andrea Abrardo, Giuseppe Bianchi, and Antonio Iera. A stakeholder-oriented security analysis in virtualized 5G cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):Article 107604, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312366.

Zhang:2021:SSD

[443] Xiaoquan Zhang, Lin Cui, Kaimin Wei, Fung Po Tso, Yangyang Ji, and Weijia Jia. A survey on stateful data plane in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107597, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312305.

Balci:2021:MCM

[444] Abdullah Balci and Radosveta Sokullu. Massive connectivity with machine learning for the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):Article 107646, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312652.

Ibrar:2021:PFM

[445] Muhammad Ibrar, Lei Wang, Gabriel-Miro Muntean, Aamir Akbar, Nadir Shah, and Kaleem Razzaq Malik. PrePass-Flow: a machine learning based technique to minimize ACL policy violation due to links failure in hybrid SDN. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107706, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313025.

Yazdinejadna:2021:KBI

[446] Abbas Yazdinejadna, Reza M. Parizi, Ali Dehghantanha, and Mohammad S. Khan. A kangaroo-based intrusion detection system on software-defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107688, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312949.

Bai:2021:CAF

[447] Yude Bai, Zhenchang Xing, Duoyuan Ma, Xiaohong Li, and Zhiyong Feng. Comparative analysis of feature representations and machine learning methods in Android family classification. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107639, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312603.

Qureshi:2021:ADT

[448] Kashif Naseer Qureshi, Gwanggil Jeon, and Francesco Piccialli. Anomaly detection and trust authority in artificial intelligence and cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107647, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312664.

Mahbub:2021:CEI

[449] Mobasshir Mahbub, M. Mofazzal Hossain, and Md. Shamrat Apu Gazi. Cloud-enabled IoT-based embedded system and software for intelligent indoor lighting, ventilation, early stage fire detection and prevention. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): Article 107673, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312858.

Anonymous:2021:Fa

[450] Anonymous. 11 February 2021. Computer Networks (Amsterdam, Netherlands: 1999), 185(??):??, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBb

[451] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 185(??):Article 107797, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1389128620313682.

Moll:2021:QBS

[452] Philipp Moll, Selina Isak, Hermann Hellwagner, and Jeff Burke. A quadtree-based synchronization protocol for inter-server game state synchronization. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107723, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313177.

Park:2021:FFB

[453] Taejune Park, Seungwon Shin, Insik Shin, and Kilho Lee. Formullar: an FPGA-based network testing tool for flexible and precise measurement of ultra-low latency networking systems. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107689, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312950.

Yemeni:2021:RST

[454] Zaid Yemeni, Haibin Wang, Waleed M. Ismael, Yanan Wang, and Zhengming Chen. Reliable spatial and temporal data redundancy reduction approach for WSN. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107701, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313013.

Yi:2021:SSA

[455] Peng Yi, Tao Hu, Yuxiang Hu, Julong Lan, Zhen Zhang, and Ziyong Li. SQHCP: Secure-aware and QoSguaranteed heterogeneous controller placement for software-defined networking. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107740, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313268.

Tahmasebi:2021:SEM

[456] Shirin Tahmasebi, Nayereh Rasouli, Amir Hosein Kashefi, Elmira Rezabeyk, and Hamid Reza Faragardi. SYNCOP: an evolutionary multiobjective placement of SDN controllers for optimizing cost and network performance in WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):Article 107727, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313190.

Amorim:2021:ACO

[457] Kelvin Santos Amorim and Gustavo Sousa Pavani. Ant colony optimization-based distributed multilayer routing and restoration in IP/MPLS over optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 185(??):Article 107747, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031330X.

Jha:2021:LBS

[458] Rakesh Kumar Jha, Puja, Haneet Kour, Manoj Kumar, and Shubha Jain. Layer based security in Narrow Band Internet of Things (NB-IoT). Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107592, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312299.

Bhooanusas:2021:SBD

[459] Nuntanut Bhooanusas, Sok-Ian Sou, and Kai-Chun Cheng. Satisfactionbased dynamic bandwidth reallocation for multipath mobile data offloading. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107594, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312287.

Asari:2021:NPH

[460] Amirhossein Asari, Mahdi R. Alagheband, Majid Bayat, and Maryam Rajabzadeh Asaar. A new provable hierarchical anonymous certificateless authentication protocol with aggregate verification in ADS-B systems. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107599, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312329.

Wang:2021:ICR

[461] Xu Wang, Yanjiao Chen, and Qian Zhang. Incentivizing cooperative relay in UTXO-based blockchain network. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107631, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312573.

Afzal:2021:MMB

[462] Samira Afzal, Christian Esteve Rothenberg, Vanessa Testoni, Prakash Kolan, and Imed Bouazizi. Multipath MMTbased approach for streaming high quality video over multiple wireless access networks. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):Article 107638, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312615.

Huang:2021:MCA

[463] Lisheng Huang, Jinye Ran, Wenyong Wang, Tan Yang, and Yu Xiang. A multi-channel anomaly detection method with feature selection and multi-scale analysis. Computer Networks (Amsterdam, Netherlands: 1999), 185(??):Article 107645, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312640.

Ndashimye:2021:MCB

[464] Emmanuel Ndashimye, Nurul I. Sarkar, and Sayan Kumar Ray. A multicriteria based handover algorithm for vehicle-to-infrastructure communications. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107652, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312688.

Eksert:2021:IIC

[465] M. Levent Eksert, Hamdullah Yücel, and Ertan Onur. Intra- and intercluster link scheduling in CUPS-based ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):Article 107659, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312718.

Geng:2021:FLE

[466] Nan Geng, Yuan Yang, and Mingwei Xu. Flow-level and efficient traffic engineering in conventional routing systems. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107671, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312834.

Shehzad:2021:PAC

[467] Khurram Shehzad, Noor M. Khan, and Junaid Ahmed. Performance analysis of coverage-centric heterogeneous cellular networks using dual-slope path loss model. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107672, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312846.

Noura:2021:ERD

[468] Hassan N. Noura, Reem Melki, Ali Chehab, and Javier Hernandez Fernandez. Efficient and robust data availability solution for hybrid PLC/RF systems. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107675, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031286X.

Sheng:2021:CPM

[469] Chuan Sheng, Yu Yao, Qiang Fu, and Wei Yang. A cyber-physical model for SCADA system and its intrusion detection. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107677, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312883.

Shadroo:2021:TPS

[470] Shabnam Shadroo, Amir Masoud Rahmani, and Ali Rezaee. The two-phase scheduling based on deep learning in the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 185(??):Article 107684, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312925.

Khormali:2021:DNS

[471] Aminollah Khormali, Jeman Park, Hisham Alasmary, Afsah Anwar, Muhammad Saad, and David Mohaisen. Domain name system security and privacy: a contemporary survey. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107699, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313001. See

corrigendum [547].

Schardong:2021:NRA

[472] Frederico Schardong, Ingrid Nunes, and Alberto Schaeffer-Filho. NFV resource allocation: a systematic review and taxonomy of VNF forwarding graph embedding. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):Article 107726, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313189.

Terzi:2021:GWD

[473] Caglar Terzi and Ibrahim Korpeoglu. 60 GHz wireless data center networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107730, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313207.

Ghosal:2021:TPS

[474] Amrita Ghosal, Sang Uk Sagong, Subir Halder, Kalana Sahabandu, Mauro Conti, Radha Poovendran, and Linda Bushnell. Truck platoon security: State-of-the-art and road ahead. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107658, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031272X.

Sun:2021:QGI

[475] Weifeng Sun, Zun Wang, and Guanghao Zhang. A QoS-guaranteed intelligent routing mechanism in softwaredefined networks. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):Article 107709, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313050.

Mayer:2021:PMH

[476] Andrea Mayer, Pierpaolo Loreti, Lorenzo Bracciale, Paolo Lungaroni, Stefano Salsano, and Clarence Filsfils. Performance monitoring with H^{2}: Hybrid Kernel/eBPF data plane for SRv6 based hybrid SDN. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):Article 107705, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313037.

Ali:2021:CRS

[477] Zeeshan Ali, Shehzad Ashraf Chaudhry, Khalid Mahmood, Sahil Garg, Zhihan Lv, and Yousaf Bin Zikria. A clogging resistant secure authentication scheme for fog computing services. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107731, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313219.

Wu:2021:DLP

[478] Yulei Wu, Yuxiang Ma, Hong-Ning Dai, and Hao Wang. Deep learning for privacy preservation in autonomous moving platforms enhanced 5G heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):Article 107743, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862031327X.

Cao:2021:TIV

[479] Haotong Cao, Yue Hu, and Longxiang Yang. Towards intelligent virtual resource allocation in UAVs-assisted 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107660, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312731.

Wang:2021:IUB

[480] Xiaoding Wang, Sahil Garg, Hui Lin, Georges Kaddoum, Jia Hu, and Mohammed F. Alhamid. An intelligent UAV based data aggregation algorithm for 5G-enabled Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 185(??): Article 107628, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031255X.

Anonymous:2021:Fb

[481] Anonymous. 26 February 2021. Computer Networks (Amsterdam, Netherlands: 1999), 186(??):??, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBc

[482] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 186(??):Article 107850, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1389128621000293.

AlGhadhban:2021:FFD

[483] Amer AlGhadhban. F4Tele: FSO for data center network management and packet telemetry. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):Article 107711, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313062.

Fujita:2021:SBR

[484] Risa Fujita, Fujun He, and Eiji Oki. Shared backup resource assignment for middleboxes considering server protection capabilities. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):Article 107734, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313220.

Choudhury:2021:HLS

[485] Hiten Choudhury. HashXor: a lightweight scheme for identity privacy of IoT devices in 5G mobile network. Computer Networks (Amsterdam, Netherlands: 1999), 186(??): Article 107753, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313335.

Triantafyllou:2021:LFL

[486] Anna Triantafyllou, Panagiotis Sarigiannidis, Thomas Lagkas, Ioannis D. Moscholios, and Antonios Sarigiannidis. Leveraging fairness in LoRaWAN: a novel scheduling scheme for collision avoidance. Computer Networks (Amsterdam, Netherlands: 1999), 186(??): Article 107735, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313232.

Ramirez:2021:ING

[487] Pedro Luis González Ramírez, Jaime Lloret, Jesús Tomás, and Mikel Hurtado. IoT-networks group-based model that uses AI for workgroup allocation. Computer Networks (Amsterdam, Netherlands: 1999), 186(??): Article 107745, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313293.

Malboubi:2021:PFI

[488] Mehdi Malboubi, Abhijeet Bhorkar, and Frank Jiang. PAveMENT: a framework for the intelligent and safe navigation of unmanned aerial vehicles over optimal PAths in MobilE NeTworks. Computer Networks (Amsterdam, Netherlands: 1999), 186(??): Article 107785, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313591.

Majumdar:2021:NBS

[489] Prasanta Majumdar and Tanmay De. A non-backtracking spectrum allocation algorithm in a large EON under dynamic traffic grooming. Computer Networks (Amsterdam, Netherlands: 1999), 186(??): Article 107757, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313359.

Li:2021:CLB

[490] Chunlin Li, Mingyang Song, Qingchuan Zhang, and Youlong Luo. Cluster load based content distribution and speculative execution for geographically distributed cloud environment. Computer Networks (Amsterdam, Netherlands: 1999), 186(??): Article 107807, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000025.

Ge:2021:TDL

[491] Mengmeng Ge, Naeem Firdous Syed, Xiping Fu, Zubair Baig, and Antonio Robles-Kelly. Towards a deep learningdriven intrusion detection approach for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):Article 107784, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862031358X.

Rios:2021:DRQ

[492] Vinícius de Miranda Rios, Pedro R. M. Inácio, Damien Magoni, and Mário M. Freire. Detection of reduction-of-quality DDoS attacks using fuzzy logic and machine learning algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 186(??): Article 107792, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313633.

Jradi:2021:OMR

[493] Hassan Jradi, Abed Ellatif Samhat, Fabienne Nouvel, Mohamad Mroue, and Jean-Christophe Prévotet. Overview of the mobility related security challenges in LPWANs. Computer Networks (Amsterdam, Netherlands: 1999), 186(??): Article 107761, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313384.

Haile:2021:EEC

[494] Habtegebreil Haile, Karl-Johan Grinnemo, Simone Ferlin, Per Hurtig, and Anna Brunstrom. End-to-end congestion control approaches for high throughput and low delay in 4G/5G cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):Article 107692, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312974.

Tan:2021:BNT

[495] Lizhuang Tan, Wei Su, Wei Zhang, Jianhui Lv, Zhenyi Zhang, Jingying Miao, Xiaoxi Liu, and Na Li. In-band network telemetry: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 186(??): Article 107763, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313396.

Chen:2021:CNN

[496] Chen Chen, Qiang Hui, Wenxuan Xie, Shaohua Wan, Yang Zhou, and Qingqi Pei. Convolutional neural networks for forecasting flood process in Internet-of-Things enabled smart city. Computer Networks (Amsterdam, Netherlands: 1999), 186(??): Article 107744, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313281.

Sekaran:2021:ISS

[497] Ramesh Sekaran, Surya Narayana Goddumarri, Suresh Kallam, Manikandan Ramachandran, Rizwan Patan, and Deepak Gupta. 5G integrated spectrum selection and spectrum access using AI-based frame work for IoT based sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 186(??):

Article 107649, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312676.

Talati:2021:ADO

[498] Shreya Talati, Darshan Vekaria, Aparna Kumari, and Sudeep Tanwar. An AI-driven object segmentation and speed control scheme for autonomous moving platforms. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):Article 107783, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313566.

Balasubramanian:2021:SAT

[499] Venkatraman Balasubramanian, Moayad Aloqaily, and Martin Reisslein. An SDN architecture for time sensitive industrial IoT. Computer Networks (Amsterdam, Netherlands: 1999), 186(??): Article 107739, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313256.

Neto:2021:SMM

[500] Emidio P. Neto, Felipe S. Dantas Silva, Lucas M. Schneider, Augusto V. Neto, and Roger Immich. Seamless MANO of multi-vendor SDN controllers across federated multidomains. Computer Networks (Amsterdam, Netherlands: 1999), 186(??): Article 107752, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313311.

Dayal:2021:SWF

[501] Neelam Dayal and Shashank Srivastava. SD-WAN flood tracer: Tracking the entry points of DDoS attack flows in WAN. Computer Networks (Amsterdam, Netherlands: 1999), 186(??): Article 107813, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000050.

Anonymous:2021:Ma

[502] Anonymous. 14 March 2021. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):??, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBd

[503] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):Article 107882, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1389128621000505.

Toumi:2021:CDS

[504] Nassima Toumi, Olivier Bernier, Djamal-Eddine Meddour, and Adlen Ksentini. On cross-domain Service Function Chain orchestration: an architectural framework. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):Article 107806, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621000013.

Navas:2021:PRI

[505] Renzo E. Navas, Frédéric Cuppens, Nora Boulahia Cuppens, Laurent Toutain, and Georgios Z. Papadopoulos. Physical resilience to insider attacks in IoT networks: Independent cryptographically secure sequences for DSSS anti-jamming. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):Article 107751, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313323.

Rezaee:2021:QTA

[506] Arash Rezaee, Omid Akbari Sheikhabad, and Lotfollah Beygi. Quality of transmission-aware control plane performance analysis for elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107755, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313347.

Shi:2021:PFF

[507] Zhou Shi, Ye Tian, Xiaodong Wang, Jiangyu Pan, and Xinming Zhang. Po-fi: Facilitating innovations on WiFi networks with an SDN approach. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107781, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313578.

Galan-Jimenez:2021:MII

[508] Jaime Galán-Jiménez and Luca Chiaraviglio. Measuring the impact of IC-NIRP vs. stricter-than-ICNIRP exposure limits on QoS and EMF from cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107824, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000128.

Wang:2021:NPM

[509] Jiahao Wang and Azzedine Boukerche. Non-parametric models with optimized training strategy for vehicles traffic flow prediction. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107791, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313621.

Thomdapu:2021:DCM

[510] Srujan Teja Thomdapu, Palash Katiyar, and Ketan Rajawat. Dynamic cache management in content delivery networks. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107822, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000116.

Hassani:2021:QPA

[511] Hamid Hassani, Francesco Gringoli, and Douglas J. Leith. Quick and plenty: Achieving low delay and high rate in 802.11ac edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107820, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000098.

Shah:2021:IMI

[512] Sayed Qaiser Ali Shah, Farrukh Zeeshan Khan, and Muneer Ahmad. The impact and mitigation of ICMP based economic denial of sustainability attack in cloud computing environment using software defined network. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107825, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100013X.

Morawski:2021:GMT

[513] Michał Morawski and Przemysław Ignaciuk. A green multipath TCP framework for industrial Internet of Things applications. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107831, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000189.

Malik:2021:MNC

[514] Hirah Malik, Cédric Adjih, Claudio Weidmann, and Michel Kieffer. MICN: a network coding protocol for ICN with multiple distinct interests per generation. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107816, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000074.

Deng:2021:SBD

[515] Yanlin Deng, Quanzhong Li, Qi Zhang, Liang Yang, and Jiayin Qin. Secure beamforming design in MIMO NOMA networks for Internet of Things with perfect and imperfect CSI. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107839, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000220.

Walelgne:2021:CPD

[516] Ermias Andargie Walelgne, Alemnew Sheferaw Asrese, Jukka Manner, Vaibhav Bajpai, and Jörg Ott. Clustering and predicting the data usage patterns of geographically diverse mobile users. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107737, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313244.

Nikkhah:2021:LLA

[517] Fahimeh Nikkhah and Masoumeh Safkhani. LAPCHS: a lightweight authentication protocol for cloudbased health-care systems. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):Article 107833, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128621000190.

Kumar:2021:SSP

[518] Randhir Kumar, Prabhat Kumar, Rakesh Tripathi, Govind P. Gupta, Thippa Reddy Gadekallu, and Gautam Srivastava. SP2F: a secured privacy-preserving framework for smart agricultural unmanned aerial vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107819, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000086.

Yang:2021:DAI

[519] Jiapie Yang, Prosanta Gope, Yongqiang Cheng, and Li Sun. Design, analysis and implementation of a smart next generation secure shipping infrastructure using autonomous robot. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107779, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313554.

Zhao:2021:IFB

[520] Liang Zhao, Zhenguo Bi, Mingwei Lin, Ammar Hawbani, Junling Shi, and Yunchong Guan. An intelligent fuzzy-based routing scheme for software-defined vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107837, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000219.

Xue:2021:RCD

[521] Gang Xue, Shuiguang Deng, Di Liu, and Zeming Yan. Reaching consensus in decentralized coordination of distributed microservices. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):Article 107786, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313608.

Al-Makhadmeh:2021:ITN

[522] Zafer Al-Makhadmeh and Amr Tolba. Independent and tailored networkslicing architecture for leveraging industrial Internet of Things job processing. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107827, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000153.

Liu:2021:CBF

[523] Qingzhi Liu, Long Cheng, Renan Alves, Tanir Ozcelebi, Fernando Kuipers, Guixian Xu, Johan Lukkien, and Shanzhi Chen. Cluster-based flow control in hybrid software-defined wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 187(??): Article 107788, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862031361X.

Anonymous:2021:Aa

[524] Anonymous. 7 April 2021. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBe

[525] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):Article 107964, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1389128621000980.

Salhab:2021:NSR

[526] Nazih Salhab, Rami Langar, and Rana Rahim. 5G network slices resource orchestration using machine learning techniques. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107829, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000165.

Golkarifard:2021:DVP

[527] Morteza Golkarifard, Carla Fabiana Chiasserini, Francesco Malandrino, and Ali Movaghar. Dynamic VNF placement, resource allocation and traffic routing in 5G. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107830, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000177.

Touijer:2021:ICC

[528] Bethaina Touijer, Yann Ben Maissa, and Salma Mouline. IEEE 802.15.6 CSMA/CA access method for WBANs: Performance evaluation and new backoff counter selection procedure. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):Article 107759, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313372.

Zhang:2021:RIS

[529] Yu Zhang, Bo Wang, Bingli Guo, Yabo Yuan, Tao Dong, Jie Yin, Kexin Li, Xiang Guo, and Shanguo Huang. A research on integrated space-ground information network simulation platform based on SDN. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):Article 107821, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000104.

Hendaoui:2021:UUA

[530] Fatma Hendaoui, Hamdi Eltaief, and Habib Youssef. UAP: a unified authentication platform for IoT environment. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107811, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000049.

Moriano:2021:UBA

[531] Pablo Moriano, Raquel Hill, and L. Jean Camp. Using bursty announcements for detecting BGP routing anomalies. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):Article 107835, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000207.

Xu:2021:DBW

[532] Chuan Xu, Giovanni Neglia, and Nicola Sebastianelli. Dynamic backup workers for parallel machine learning. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107846, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000256.

Benedetti:2021:SME

[533] Paolo Benedetti, Giuseppe Piro, and Luigi Alfredo Grieco. A softwarized and MEC-enabled protocol architecture supporting consumer mobility in information-centric networks. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107867, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000396.

Wang:2021:DFG

[534] Tao Wang, Li Kang, and Jiang Duan. Dynamic fine-grained access control scheme for vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107872, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000402.

Amaizu:2021:CED

[535] G. C. Amaizu, C. I. Nwakanma, S. Bhardwaj, J. M. Lee, and D. S. Kim. Composite and efficient DDoS attack detection framework for B5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107871, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000438.

Kazemifard:2021:MDF

[536] Nasim Kazemifard and Vahid Shah-Mansouri. Minimum delay function placement and resource allocation for Open RAN (O-RAN) 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107809, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000037.

Faroughi:2021:TWD

[537] Azadeh Faroughi, Andrea Morichetta, Luca Vassio, Flavio Figueiredo, Marco Mellia, and Reza Javidan. Towards website domain name classification using graph based semi-supervised learning. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107865, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000384.

Yang:2021:DSE

[538] Lei Yang, Liang Liu, Zuchao Ma, and Youwei Ding. Detection of selective-edge packet attack based on edge reputation in IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107842, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000232.

Volnes:2021:IAT

[539] Espen Volnes, Stein Kristiansen, and Thomas Plagemann. Improving the accuracy of timing in scalable WSN simulations with communication software execution models. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):Article 107855, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000335.

Kilincer:2021:MLM

[540] Ilhan Firat Kilincer, Fatih Ertam, and Abdulkadir Sengur. Machine learning methods for cyber security intrusion detection: Datasets and comparative study. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107840, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000141.

Takeda:2021:MPS

[541] Kenta Takeda, Takehiro Sato, Ryoichi Shinkuma, and Eiji Oki. Multipath provisioning scheme for fault tolerance to minimize required spectrum resources in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107895, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000591.

Derakhshanfard:2021:ORW

[542] Nahideh Derakhshanfard and Reza Soltani. Opportunistic routing in wireless networks using bitmap-based weighted tree. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):Article 107892, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100058X.

Zhang:2021:PBS

[543] Tianzhu Zhang, Leonardo Linguaglossa, Paolo Giaccone, Luigi Iannone, and James Roberts. Performance benchmarking of state-of-the-art software switches for NFV. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):Article 107861, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000372.

Doost-Mohammady:2021:GTW

[544] Rahman Doost-Mohammady, Oscar Bejarano, and Ashutosh Sabharwal. Good times for wireless research. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107870, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000414.

Kumar:2021:OAO

[545] Rohit Kumar, Venkanna U., and Vivek Tiwari. Opt-ACM: an optimized load balancing based admission control mechanism for software defined hybrid wireless based IoT (SDHW-IoT) network. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107888, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000542.

Gao:2021:OCE

[546] Peixuan Gao, Yang Xu, and H. Jonathan Chao. OVS-CAB: Efficient rulecaching for Open vSwitch hardware offloading. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107844, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000244.

Khormali:2021:CDN

[547] Aminollah Khormali, Jeman Park, Hisham Alasmary, Afsah Anwar, Muhammad Saad, and David Mohaisen. Corrigendum to “Domain name system security and privacy: a contemporary survey” Computer Networks Volume 185 (2020)

107699. Computer Networks (Amsterdam, Netherlands: 1999), 188(??): Article 107814, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000062. See

[471].

Anonymous:2021:Ab

[548] Anonymous. 22 April 2021. Computer Networks (Amsterdam, Netherlands: 1999), 189(??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBf

[549] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 189(??):Article 108024, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1389128621001341.

Fawaz:2021:RLA

[550] Hassan Fawaz, Melhem El Helou, Samer Lahoud, and Kinda Khawam. A reinforcement learning approach to queue-aware scheduling in full-duplex wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):Article 107893, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000578.

Lin:2021:NLG

[551] Jie Lin, Lin Huang, Hanlin Zhang, Xinyu Yang, and Peng Zhao. A novel latency-guaranteed based resource double auction for market-oriented edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107873, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000426.

Zakariya:2021:ODM

[552] Ahmed Y. Zakariya, Sherif I. Rabia, and W. K. Zahra. Optimal decision making in multi-channel RF-powered cognitive radio networks with ambient backscatter capability. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107907, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000669.

Baidas:2021:RAO

[553] Mohammed W. Baidas. Resource allocation for offloading-efficiency maximization in clustered NOMAenabled mobile edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107919, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000712.

Morichetta:2021:UWP

[554] Andrea Morichetta, Martino Trevisan, Luca Vassio, and Julia Krickl. Understanding web pornography usage from traffic analysis. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):Article 107909, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000670.

Feng:2021:DRC

[555] Yue Feng, Weiqing Huang, Siye Wang, Yanfang Zhang, and Shang Jiang. Detection of RFID cloning attacks: a spatiotemporal trajectory data stream-based practical approach. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107922, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000736.

Martinez:2021:ESE

[556] Cristiane A. Pendeza Martinez, Taufik Abrão, and André Luís Machado Martinez. Energy and spectral efficiency trade-off in OCDMA-PON assisted by non-linear programming methods. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107920, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000724.

Chuprikov:2021:FTC

[557] Pavel Chuprikov, Alex Davydow, Kirill Kogan, Sergey I. Nikolenko, and Alexander Sirotkin. Formalization and taxonomy of compute-aggregate problems for cloud computing applications. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107915, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000694.

Wu:2021:AMS

[558] Libing Wu, Shuqin Cao, Yanjiao Chen, Jianqun Cui, and Yanan Chang. An adaptive multiple sprayand-wait routing algorithm based on social circles in delay tolerant networks. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107901, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000633.

Biswas:2021:LBU

[559] Soumadip Biswas, Arobinda Gupta, and Sandip Chakraborty. Loadbalanced user associations in dense LTE networks. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):Article 107928, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000761.

Zhong:2021:CSC

[560] Shijie Zhong, Songtao Guo, Hongyan Yu, and Quyuan Wang. Cooperative service caching and computation offloading in multi-access edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107916, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000566.

Huang:2021:ITC

[561] Xiaohong Huang, Man Zeng, and Kun Xie. Intelligent traffic control for QoS optimization in hybrid SDNs. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107877, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000451.

Venkateswararao:2021:UUV

[562] Kuna Venkateswararao, Pravati Swain, Christophoros Christophorou, and Andreas Pitsillides. Using UE-VBS for dynamic virtual small cells deployment and backhauling in 5G ultra-dense networks. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107926, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000748.

Shang:2021:DUA

[563] Longkang Shang, Dong Guo, Yuede Ji, and Qiang Li. Discovering unknown advanced persistent threat using shared features mined by neural networks. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107937, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000803.

Enoch:2021:NSM

[564] Simon Yusuf Enoch, Jang Se Lee, and Dong Seong Kim. Novel security models, metrics and security assessment for maritime vessel networks. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107934, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000797.

Rahdari:2021:QAP

[565] Farhad Rahdari, Naser Movahhedinia, Mohammad Reza Khayyambashi, and Shahrokh Valaee. QoEaware power control and user grouping in cognitive radio OFDM-NOMA systems. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107906, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000657.

Modeas:2021:ADN

[566] Ioannis Modeas, Alexandros Kaloxylos, Lazaros Merakos, and Dimitris Tsolkas. An adaptive and distributed network selection mechanism for 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107943, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000840.

Davydow:2021:CBM

[567] Alex Davydow, Pavel Chuprikov, Sergey I. Nikolenko, and Kirill Kogan. Competitive buffer management for packets with latency constraints. Computer Networks (Amsterdam, Netherlands: 1999), 189(??):

Article 107942, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000839.

Farrokhi:2021:AIT

[568] Alireza Farrokhi, Reza Farahbakhsh, Javad Rezazadeh, and Roberto Minerva. Application of Internet of Things and artificial intelligence for smart fitness: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):Article 107859, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000360.

Ahmed:2021:MLM

[569] Usman Ahmed, Gautam Srivastava, and Jerry Chun-Wei Lin. A machine learning model for data sanitization. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107914, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000700.

Lu:2021:ECD

[570] Ning Lu, Dan Li, Wenbo Shi, Pandi Vijayakumar, Francesco Piccialli, and Victor Chang. An efficient combined deep neural network based malware detection framework in 5G environment. Computer Networks (Amsterdam, Netherlands: 1999), 189(??): Article 107932, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000785.

Anonymous:2021:Mb

[571] Anonymous. 8 May 2021. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBg

[572] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001602.

Huang:2021:SJR

[573] Guanglun Huang, Baoxian Zhang, Zheng Yao, and Cheng Li. Stochastic joint rate control and resource allocation for wireless video surveillance. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000645.

Shaer:2021:EEP

[574] Ibrahim Shaer, Greg Sidebottom, Anwar Haque, and Abdallah Shami. Efficient execution plan for egress traffic engineering. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000815.

Femminella:2021:EDT

[575] Mauro Femminella, Matteo Pergolesi, and Gianluca Reali. 5G experiment design through Blueprint. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000864.

Chen:2021:ATM

[576] Guozhu Chen, Fanping Zeng, Jian Zhang, Tingting Lu, Jingfei Shen, and Wenjuan Shu. An adaptive trust model based on recommendation filtering algorithm for the Internet of Things systems. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100089X.

Chen:2021:MEC

[577] Xincheng Chen. Mobile edge computing resource allocation: an operating system view. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100075X.

Sundberg:2021:LET

[578] Simon Sundberg and Johan Garcia. Locating eNodeBs through sectorization inference-sector fitting evaluated on a railway use case. Computer Networks (Amsterdam, Netherlands:

1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000852.

Sun:2021:SSD

[579] Penghao Sun, Zehua Guo, Julong Lan, Junfei Li, Yuxiang Hu, and Thar Baker. ScaleDRL: a scalable deep reinforcement learning approach for traffic engineering in SDN with pinning control. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000554.

Adebayo:2021:PEU

[580] Adekunle Adebayo, Ibhadode Osagie, Adekunle Oluwaseun Tope, and Ibhadode Akhere Pauline. Performance evaluation of ultra-dense macrofemto cellular networks based on efficient channel resource allocation system. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000876.

Attarian:2021:ACP

[581] Reyhane Attarian and Sattar Hashemi. An anonymity communication protocol for security and privacy of clients in IoT-based mobile health transactions. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001079.

Li:2021:ADD

[582] Huichun Li, Chengli Zhao, Yangyang Liu, and Xue Zhang. Anomaly detection by discovering bipartite structure on complex networks. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000621.

Strinati:2021:NBS

[583] Emilio Calvanese Strinati and Sergio Barbarossa. 6G networks: Beyond Shannon towards semantic and goaloriented communications. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000773.

Lalle:2021:CTS

[584] Yandja Lalle, Mohamed Fourati, Lamia Chaari Fourati, and João Paulo Barraca. Communication technologies for Smart Water Grid applications: Overview, opportunities, and research directions. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000827.

Akbar:2021:NET

[585] Aamina Akbar, Sobia Jangsher, and Farrukh A. Bhatti. NOMA and 5G emerging technologies: a survey on issues and solution techniques. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000888.

Silva:2021:HSS

[586] Rui Silva, David Santos, Flávio Meneses, Daniel Corujo, and Rui L. Aguiar. A hybrid SDN solution for mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000931.

Li:2021:DAD

[587] Yu Li and Wenjian Xu. D-AdFeed: a diversity-aware utility-maximizing advertising framework for mobile users. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000906.

Wang:2021:DJS

[588] Libing Wang, Xin Hu, Yin Wang, Sujie Xu, Shijun Ma, Kexin Yang, Zhijun Liu, and Weidong Wang. Dynamic job-shop scheduling in smart manufacturing using deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001031.

Velusamy:2021:FTA

[589] Nandagopal Velusamy, Fadi Al-Turjman, Rajagopal Kumar, and Jothilakshmi Ramakrishnan. A framework for task allocation in IoToriented industrial manufacturing systems. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001043.

Dong:2021:SLT

[590] Jialiang Dong, Zhitao Guan, Longfei Wu, Xiaojiang Du, and Mohsen Guizani. A sentence-level text adversarial attack algorithm against IIoT based smart grid. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100092X.

Lin:2021:TNC

[591] Kunda Lin, Xiaolong Xu, and Honghao Gao. TSCRNN: a novel classification scheme of encrypted traffic based on flow spatiotemporal features for efficient management of IIoT. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001067.

Alqahtani:2021:ECR

[592] Fayez Alqahtani, Mohammed Al-Maitah, Khaldoun Besoul, and S. K. Elagan. Elastic computing resource virtualization method for a service-centric industrial Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000918.

Anonymous:2021:Mc

[593] Anonymous. 22 May 2021. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBh

[594] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001870.

Hu:2021:BBT

[595] Donghui Hu, Yifan Li, Lixuan Pan, Meng Li, and Shuli Zheng. A blockchain-based trading system for big data. Computer Networks (Amsterdam, Netherlands: 1999), 191 (??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100116X.

Chaudhry:2021:GIC

[596] Shehzad Ashraf Chaudhry, Khalid Yahya, Marimuthu Karuppiah, Rupak Kharel, Ali Kashif Bashir, and Yousaf Bin Zikria. GCACS-IoD: a certificate based generic access control scheme for Internet of drones. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001195.

Vasudev:2021:SPP

[597] Harsha Vasudev and Debasis Das. P^{2}-SHARP: Privacy preserving secure hash based authentication and revelation protocol in IoVs. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001146.

Hei:2021:MMA

[598] Yiming Hei, Jianwei Liu, Hanwen Feng, Dawei Li, Yizhong Liu, and Qianhong Wu. Making MA-ABE fully accountable: a blockchain-based approach for secure digital right management. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001389.

Zhong:2021:APP

[599] Hong Zhong, Jinpeng Fan, Jie Cui, Yan Xu, and Lu Liu. Assessing profit of prediction for SDN controllers load balancing. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001158.

Zhang:2021:ELS

[600] Guoqiang Zhang, Chenyi Wu, and Qian Gao. Exploiting layer and spatial correlations to enhance SVC and tile based 360-degree video streaming. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001122.

Ergiz:2021:JMF

[601] Yamac Ergiz, Ali Murat Demirtas, and Tolga Girici. Joint multipath flow and layer allocation for scalable video streaming. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001171.

Zhai:2021:TPU

[602] Shuangjiao Zhai, Guixin Ye, Zhanyong Tang, Jie Ren, Dingyi Fang, Baoying Liu, and Zheng Wang. Towards practical 3D ultrasound sensing on commercial-off-the-shelf mobile devices. Computer Networks (Amsterdam, Netherlands: 1999), 191 (??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001134.

Wen:2021:ACB

[603] Yujuan Wen, Fengyuan Lu, Yufei Liu, and Xinli Huang. Attacks and countermeasures on blockchains: a survey from layering perspective. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001080.

Cominelli:2021:ICR

[604] Marco Cominelli, Felix Kosterhon, Francesco Gringoli, Renato Lo Cigno, and Arash Asadi. IEEE 802.11 CSI randomization to preserve location privacy: an empirical evaluation in different scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100102X.

Martinez-Yelmo:2021:EEH

[605] Isaias Martinez-Yelmo, Joaquin Alvarez-Horcajo, Juan Antonio Carral, and Diego Lopez-Pajares. eHDDP: Enhanced Hybrid Domain Discovery Protocol for network topologies with both wired/wireless and SDN/non-SDN devices. Computer Networks (Amsterdam, Netherlands: 1999), 191 (??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001110.

Zaballa:2021:TMH

[606] Eder Ollora Zaballa, David Franco, Signe Erdman Thomsen, Marivi Higuero, Henrik Wessing, and Michael S. Berger. Towards monitoring hybrid next-generation software-defined and service provider MPLS networks. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000943.

Ming:2021:NNB

[607] Zhongxing Ming and Mingwei Xu. NBA: a name-based approach to device mobility in industrial IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 191 (??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001055.

Wu:2021:DRL

[608] Yulei Wu, Zehua Wang, Yuxiang Ma, and Victor C. M. Leung. Deep reinforcement learning for blockchain in industrial IoT: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001213.

Aazam:2021:TOE

[609] Mohammad Aazam, Sherali Zeadally, and Eduardo Feo Flushing. Task offloading in edge computing for machine learning-based smart healthcare. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001298.

Brotsis:2021:SBP

[610] Sotirios Brotsis, Konstantinos Limniotis, Gueltoum Bendiab, Nicholas Kolokotronis, and Stavros Shiaeles. On the suitability of blockchain platforms for IoT applications: Architectures, security, privacy, and performance. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001225.

Faghihi:2021:RDC

[611] Farnood Faghihi and Mohammad Zulkernine. RansomCare: Data-centric detection and mitigation against smartphone crypto-ransomware. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001250.

Caviglione:2021:KLT

[612] Luca Caviglione, Wojciech Mazurczyk, Matteo Repetto, Andreas Schaffhauser, and Marco Zuppelli. Kernel-level tracing for detecting stegomalware and covert channels in Linux environments. Computer Networks (Amsterdam, Netherlands: 1999), 191 (??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001249.

Yurekten:2021:CCT

[613] Ozgur Yurekten and Mehmet Demirci. Citadel: Cyber threat intelligence assisted defense system for softwaredefined networks. Computer Networks (Amsterdam, Netherlands: 1999), 191 (??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001262.

Hadem:2021:SBI

[614] Pynbianglut Hadem, Dilip Kumar Saikia, and Soumen Moulik. An SDN-based intrusion detection system using SVM with selective logging for IP traceback. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001274.

Anonymous:2021:Jb

[615] Anonymous. 19 June 2021. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBi

[616] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912862100222X.

Sun:2021:CLT

[617] Zemin Sun, Yanheng Liu, Jian Wang, Rundong Yu, and Dongpu Cao. Crosslayer tradeoff of QoS and security in vehicular ad hoc networks: a game theoretical approach. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001390.

Ashraf:2021:RQA

[618] Usman Ashraf, Adnan Ahmed, Mohammed Al-Naeem, and Unsa Masood. Reliable and QoS aware routing metrics for wireless neighborhood area networking in smart grids. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100150X.

Jain:2021:UAR

[619] Akshay Jain, Elena Lopez-Aguilera, and Ilker Demirkol. User association and resource allocation in 5G (AURA-5G): a joint optimization framework. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001547.

Liu:2021:RTC

[620] Meili Liu, Xiaogang Qi, Lifang Liu, and Hao Pan. Roots-tracing of communication network alarm: a realtime processing framework. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001432.

Yakici:2021:SMO

[621] Ertan Yakici and Mumtaz Karatas. Solving a multi-objective heterogeneous sensor network location problem with genetic algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001456.

deVos:2021:CMF

[622] Martijn de Vos and Johan Pouwelse. ConTrib: Maintaining fairness in decentralized big tech alternatives by accounting work. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001705.

Tomassilli:2021:DRP

[623] Andrea Tomassilli, Giuseppe Di Lena, Frédéric Giroire, Issam Tahiri, Damien Saucez, Stéphane Pérennes, Thierry Turletti, Ruslan Sadykov, François Vanderbeck, and Chidung Lac. Design of robust programmable networks with bandwidth-optimal failure recovery scheme. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001468.

Dou:2021:MMN

[624] Songshi Dou, Guochun Miao, Zehua Guo, Chao Yao, Weiran Wu, and Yuanqing Xia. Matchmaker: Maintaining network programmability for software-defined WANs under multiple controller failures. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100147X.

Gupta:2021:LIH

[625] Neha Gupta, Vinita Jindal, and Punam Bedi. LIO-IDS: Handling class imbalance using LSTM and improved one-vs-one technique in intrusion detection system. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001675.

Vallero:2021:BSS

[626] Greta Vallero, Margot Deruyck, Michela Meo, and Wout Joseph. Base station switching and edge caching optimisation in high energy-efficiency wireless access network. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100181X.

Bryant:2021:EVN

[627] Blake Bryant and Hossein Saiedian. An evaluation of videogame network architecture performance and security. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002000.

Jiang:2021:WML

[628] Huiling Jiang, Qing Li, Yong Jiang, GengBiao Shen, Richard Sinnott, Chen Tian, and Mingwei Xu. When machine learning meets congestion control: a survey and comparison. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001407.

Babun:2021:SIP

[629] Leonardo Babun, Kyle Denney, Z. Berkay Celik, Patrick McDaniel, and A. Selcuk Uluagac. A survey on IoT platforms: Communication, security, and privacy perspectives. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001444.

Priyadarsini:2021:SDN

[630] Madhukrishna Priyadarsini and Padmalochan Bera. Software defined networking architecture, traffic management, security, and placement: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001481.

Yang:2021:DAS

[631] Jiachen Yang, Shan Liu, Hansong Su, and Ying Tian. Driving assistance system based on data fusion of multisource sensors for autonomous unmanned ground vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001511.

Khorsandroo:2021:HSE

[632] Sajad Khorsandroo, Adrián Gallego Sánchez, Ali Saman Tosun, JM Arco, and Roberto Doriguzzi-Corin. Hybrid SDN evolution: a comprehensive survey of the state-of-the-art. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001109.

Spiekermann:2021:UPB

[633] Daniel Spiekermann and Jörg Keller. Unsupervised packet-based anomaly detection in virtual networks. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128621001286.

Li:2021:MSD

[634] Fuliang Li, Zhenbei Guo, Bocheng Liang, Xiushuang Yi, Xingwei Wang, Weichao Li, and Yi Wang. A measurement study on device-to-device communication technologies for IIoT. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100164X.

Guo:2021:EFM

[635] Yinghao Guo, Zichao Zhao, Ke He, Shiwei Lai, Junjuan Xia, and Lisheng Fan. Efficient and flexible management for industrial Internet of Things: a federated learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001961.

Cong:2021:DRL

[636] Peizhuang Cong, Yuchao Zhang, Zheli Liu, Thar Baker, Hissam Tawfik, Wendong Wang, Ke Xu, Ruidong Li, and Fuliang Li. A deep reinforcement learning-based multi-optimality routing scheme for dynamic IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001535.

Anonymous:2021:Jc

[637] Anonymous. 5 July 2021. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBj

[638] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002371.

Li:2021:MEC

[639] Bo Li, Xiaoheng Deng, and Yiqin Deng. Mobile-edge computing-based delay minimization controller placement in SDN-IoV. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001493.

Shiny:2021:RGT

[640] S. Suja Golden Shiny, S. Sathya Priya, and K. Murugan. Repeated game theory-based reducer selection strategy for energy management in SD-WSN. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100178X.

Huang:2021:MCM

[641] Shan Huang, Dezun Dong, Zejia Zhou, and Xiangke Liao. MP-CREDIT: Multi-path credit for highspeed data center transports. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001560.

Somesula:2021:CDA

[642] Manoj Kumar Somesula, Rashmi Ranjan Rout, and D. V. L. N. Somayajulu. Contact duration-aware cooperative cache placement using genetic algorithm for mobile edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001559.

Sharma:2021:PCP

[643] Abhilasha Sharma and Lalit Kumar Awasthi. Pr-CAI: Priority basedcontext aware information scheduling for SDN-based vehicular network. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001791.

Vanderhallen:2021:RAA

[644] Stien Vanderhallen, Jo Van Bulck, Frank Piessens, and Jan Tobias Mühlberg. Robust authentication for automotive control networks through covert channels. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001699.

Eramo:2021:ALS

[645] Vincenzo Eramo, Francesco Giacinto Lavacca, Tiziana Catena, and Paul Jaime Perez Salazar. Application of a long short term memory neural predictor with asymmetric loss function for the resource allocation in NFV network architectures. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001833.

Yang:2021:NTF

[646] Hanyu Yang, Xutao Li, Wenhao Qiang, Yuhan Zhao, Wei Zhang, and Chang Tang. A network traffic forecasting method based on SA optimized ARIMA-BP neural network. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001821.

Song:2021:FFC

[647] Huangfei Song, Songtao Guo, Pan Li, and Guiyan Liu. FCNR: Fast and consistent network reconfiguration with low latency for SDN. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001912.

Yu:2021:OEE

[648] Li Yu, Zongpeng Li, Jiangchuan Liu, and Ruiting Zhou. Online and energy-efficient task-processing for distributed edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100044X.

Bu:2021:SMP

[649] Kai Bu, Yutian Yang, Zixuan Guo, Yuanyuan Yang, Xing Li, and Shigeng Zhang. Securing middlebox policy enforcement in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001808.

Liu:2021:GBR

[650] Zhixin Liu, Xiaopin Li, Yazhou Yuan, Yi Yang, and Xinping Guan. Game based robust power allocation strategy with QoS guarantee in D2D communication network. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002012.

Yao:2021:RCM

[651] Xiaopeng Yao, Guangxian Liang, Chonglin Gu, and Hejiao Huang. Rumors clarification with minimum credibility in social networks. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001973.

Yarinezhad:2021:EER

[652] Ramin Yarinezhad and Sadoon Azizi. An energy-efficient routing protocol for the Internet of Things networks based on geographical location and link quality. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001936.

Ometov:2021:SWT

[653] Aleksandr Ometov, Viktoriia Shubina, Lucie Klus, Justyna Skibińska, Salwa Saafi, Pavel Pascacio, Laura Flueratoru, Darwin Quezada Gaibor, Nadezhda Chukhno, Olga Chukhno, Asad Ali, Asma Channa, Ekaterina Svertoka, Waleed Bin Qaim, Raúl Casanova-Marqués, Sylvia Holcer, Joaquín Torres-Sospedra, Sven Casteleyn, Giuseppe Ruggeri, Giuseppe Araniti, Radim Burget, Jiri Hosek, and Elena Simona Lohan. A survey on wearable technology: History, state-of-the-art and current challenges. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001651.

Shamsoshoara:2021:AIP

[654] Alireza Shamsoshoara, Fatemeh Afghah, Abolfazl Razi, Liming Zheng, Peter Z. Fulé, and Erik Blasch. Aerial imagery pile burn detection using deep learning: the FLAME dataset. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001201.

Lopez-Benitez:2021:SSR

[655] Miguel López-Benítez, Alessandro Raschellà, Sara Pizzi, Li Wang, Marco Di Felice, and Kaushik Roy Chowdhury. Smart spectrum and radio resource management for future 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001754.

Radoglou-Grammatikis:2021:SSS

[656] Panagiotis Radoglou-Grammatikis, Panagiotis Sarigiannidis, Eider Iturbe, Erkuden Rios, Saturnino Martinez, Antonios Sarigiannidis, Georgios Eftathopoulos, Yannis Spyridis, Achilleas Sesis, Nikolaos Vakakis, Dimitrios Tzovaras, Emmanouil Kafetzakis, Ioannis Giannoulakis, Michalis Tzifas, Alkiviadis Giannakoulias, Michail Angelopoulos, and Francisco Ramos. SPEAR SIEM: a security information and event management system for the smart grid. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001237.

Zhou:2021:DLF

[657] Qizhao Zhou, Junqing Yu, and Dong Li. A dynamic and lightweight framework to secure source addresses in the SDN-based networks. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001663.

Valenza:2021:GES

[658] Fulvio Valenza, Matteo Repetto, and Stavros Shiaeles. Guest editorial: Special issue on novel cyber-security paradigms for software-defined and virtualized systems. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001997.

Bansal:2021:NGS

[659] Gaurang Bansal, Vinay Chamola, Georges Kaddoum, Md. Jalil Piran, and Mubarak Alrashoud. Next generation stock exchange: Recurrent neural learning model for distributed ledger transactions. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001183.

Zhang:2021:TCF

[660] Qingyang Zhang, Hong Zhong, Weisong Shi, and Lu Liu. A trusted and collaborative framework for deep learning in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001523.

Wang:2021:TRL

[661] Ke Wang, Chien-Ming Chen, M. Shamim Hossain, Ghulam Muhammad, Sachin Kumar, and Saru Kumari. Transfer reinforcement learning-based road object detection in next generation IoT domain. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001687.

Lu:2021:NTI

[662] Liangfu Lu, Xiaohan Ren, Chenwei Cui, Zhiyuan Tan, Yulei Wu, and Zhizhen Qin. A novel tensorinformation bottleneck method for multi-input single-output applications. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001742.

Wang:2021:DEC

[663] Jiansi Wang, Haopeng Chen, Fuxiao Zhou, Meng Sun, Ziang Huang, and Zhengtong Zhang. A-DECS: Enhanced collaborative edge-edge data storage service for edge computing with adaptive prediction. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001730.

Tang:2021:JGL

[664] Yuan Tang, Yiming Miao, Ahmed Barnawi, Bander Alzahrani, Reem Alotaibi, and Kai Hwang. A joint global and local path planning optimization for UAV task scheduling towards crowd air monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000682.

Sen:2021:VET

[665] Priyangshu Sen, Viduneth Ariyarathna, Arjuna Madanayake, and Josep M. Jornet. A versatile experimental testbed for ultrabroadband communication networks above 100 GHz. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001778.

Anonymous:2021:Jd

[666] Anonymous. 20 July 2021. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBk

[667] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002577.

Shahraki:2021:TTB

[668] Amin Shahraki, Amir Taherkordi, and Øystein Haugen. TONTA: Trendbased online network traffic analysis in ad-hoc IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001985.

Wubben:2021:NRR

[669] Jamie Wubben, Francisco Fabra, Carlos T. Calafate, Juan-Carlos Cano, and Pietro Manzoni. A novel resilient and reconfigurable swarm management scheme. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100195X.

Bruschi:2021:MNS

[670] Roberto Bruschi, Jane Frances Pajo, Franco Davoli, and Chiara Lombardo. Managing 5G network slicing and edge computing with the MATILDA telecom layer platform. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001766.

Xie:2021:OAM

[671] Yanghao Xie, Sheng Wang, Binbin Wang, Shizhong Xu, Xiong Wang, and Jing Ren. Online algorithm for migration aware virtualized network function placing and routing in dynamic 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001924.

Panicker:2021:AEO

[672] Ashwin Panicker, Ozgur Ozdemir, Mihail L. Sichitiu, Ismail Guvenc, Rudra Dutta, Vuk Marojevic, and Brian Floyd. AERPAW emulation overview and preliminary performance evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001717.

Ren:2021:MRR

[673] Quan Ren, Tao Hu, Jiangxing Wu, Yuxiang Hu, Lei He, and Julong Lan. Multipath resilient routing for endogenous secure software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002048.

Gupta:2021:OPD

[674] Jitendra Gupta, Md Shahbaz Akhtar, Aneek Adhya, and Sudhan Majhi. Optimal planning and design of SRLGaware survivable LR-PON for wireless and FTTx networks. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002097.

Yu:2021:PPB

[675] Lian Yu, Jingtao Dong, Lihao Chen, Mengyuan Li, Bingfeng Xu, Zhao Li, Lin Qiao, Lijun Liu, Bei Zhao, and Chen Zhang. PBCNN: Packet bytesbased convolutional neural network for network intrusion detection. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001948.

Polverini:2021:EDL

[676] Marco Polverini, Juan Luis Herrera, Pierpaolo Salvo, and Jaime Galán-Jiménez. Early detection of link failures through the modeling of the hardware deterioration process. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002115.

Palama:2021:ICW

[677] Ivan Palamà, Francesco Gringoli, Giuseppe Bianchi, and Nicola Blefari-Melazzi. IMSI catchers in the wild: a real world 4G/5G assessment. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002061.

Ozdem:2021:SAD

[678] Mehmet Özdem and Mustafa Alkan. Subscriber aware dynamic service function chaining. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002073.

Yigit:2021:BFS

[679] Yasin Yigit, Vahid Khalilpour Akram, and Orhan Dagdeviren. Breadth-first search tree integrated vertex cover algorithms for link monitoring and routing in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002103.

Magan-Carrion:2021:UIW

[680] Roberto Magán-Carrión, Alberto Abellán-Galera, Gabriel Maciá-Fernández, and Pedro García-Teodoro. Unveiling the I2P web structure: a connectivity analysis. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002188.

Caiazza:2021:MDD

[681] Chiara Caiazza, Claudio Cicconetti, Valerio Luconi, and Alessio Vecchio. Measurement-driven design and runtime optimization in edge computing: Methodology and tools. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002085.

Cai:2021:DAP

[682] Yun-Zhan Cai, Yu-Ting Wang, and Meng-Hsun Tsai. Dynamic adjustment for proactive flow installation mechanism in SDN-based IoT. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002279.

Wang:2021:DSA

[683] Rui Wang, Hongchao Du, Zhaoyan Shen, and Zhiping Jia. DAP-Sketch: an accurate and effective network measurement sketch with Deterministic Admission Policy. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002164.

Glabowski:2021:MSS

[684] Mariusz Głabowski, Maciej Sobieraj, and Maciej Stasiak. Modelling of single-sided multiservice switching structures. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100205X.

Xing:2021:VBM

[685] Tianzhang Xing, Chase Wu, Jie Wang, Fei Shang, Ruilin Li, and Xiaojiang Chen. A vibration-based multiuser concurrent communication system with commercial devices. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002280.

Navidan:2021:GAN

[686] Hojjat Navidan, Parisa Fard Moshiri, Mohammad Nabati, Reza Shahbazian, Seyed Ali Ghorashi, Vahid Shah-Mansouri, and David Windridge. Generative adversarial networks (GANs) in networking: a comprehensive survey and evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002139.

Maia:2021:IMO

[687] Adyson M. Maia, Yacine Ghamri-Doudane, Dario Vieira, and Miguel Franklin de Castro. An improved multiobjective genetic algorithm with heuristic initialization for service placement and load distribution in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002127.

El-Hajj:2021:TPS

[688] Mohammed El-Hajj, Ahmad Fadlallah, Maroun Chamoun, and Ahmed Serhrouchni. A taxonomy of PUF schemes with a novel arbiter-based PUF resisting machine learning attacks. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002036.

Anonymous:2021:A

[689] Anonymous. 4 August 2021. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBl

[690] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621003169.

Yao:2021:KKM

[691] Yingbiao Yao, Yuancheng Qin, Wei Feng, Pei Li, Xiaorong Xu, Xin Xu, and Xuesong Liang. KFTO: Kuhn–Munkres based fair task offloading in fog networks. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002024.

Hortelano:2021:REC

[692] Diego Hortelano, Teresa Olivares, and M. Carmen Ruiz. Reducing the energy consumption of the friendship mechanism in Bluetooth mesh. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002309.

Magoula:2021:GAA

[693] Lina Magoula, Sokratis Barmpounakis, Ioannis Stavrakakis, and Nancy Alonistioti. A genetic algorithm approach for service function chain placement in 5G and beyond, virtualized edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002176.

elhoudaNouar:2021:SVN

[694] Nour el houda Nouar, Sami Yangui, Noura Faci, Khalil Drira, and Saïd Tazi. A semantic virtualized network functions description and discovery model. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002140.

Steadman:2021:DED

[695] Jacob Steadman and Sandra Scott-Hayward. DNSxP: Enhancing data exfiltration protection through data plane programmability. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002310.

Kosek-Szott:2021:DCA

[696] Katarzyna Kosek-Szott, Alice Lo Valvo, Szymon Szott, Pierluigi Gallo, and Ilenia Tinnirello. Downlink channel access performance of NR-U: Impact of numerology and mini-slots on coexistence with Wi-Fi in the 5 GHz band. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002437.

Aguilar-Fuster:2021:NEF

[697] Christian Aguilar-Fuster and Javier Rubio-Loyola. A novel evaluation function for higher acceptance rates and more profitable metaheuristicbased online virtual network embedding. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002449.

Tu:2021:ROM

[698] Shanshan Tu, Muhammad Waqas, Fengming Huang, Ghulam Abbas, and Ziaul Haq Abbas. A revocable and outsourced multi-authority attributebased encryption scheme in fog computing. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002474.

Li:2021:STP

[699] Junfeng Li, Dan Li, Wenfei Wu, K. K. Ramakrishnan, Jinkun Geng, Fanzhao Wang, and Kai Zheng. Sphinx: a transport protocol for high-speed and lossy mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002462.

Chen:2021:EIC

[700] Miaojiang Chen, Wei Liu, Tian Wang, Anfeng Liu, and Zhiwen Zeng. Edge intelligence computing for mobile augmented reality with deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002425.

Chakraborty:2021:DIN

[701] Tamal Chakraborty and Iti Saha Misra. Design and implementation of a novel two-phase spectrum handoff scheme for QoS aware mobile users in cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002450.

Liu:2021:REE

[702] Zhixin Liu, Meihua Zhou, Yanyan Shen, Yazhou Yuan, Kit Yan Chan, and Yi Yang. Robust energy efficient maximization in wireless powered CRNs based on power splitting. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002528.

Junior:2021:DRS

[703] Franklin Magalhães Ribeiro Junior and Carlos Alberto Kamienski. Data resilience system for fog computing. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002656.

Tian:2021:PSP

[704] Junfeng Tian and Haoning Wang. A provably secure and public auditing protocol based on the bell triangle for cloud data. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100267X.

Saeik:2021:TOE

[705] Firdose Saeik, Marios Avgeris, Dimitrios Spatharakis, Nina Santi, Dimitrios Dechouniotis, John Violos, Aris Leivadeas, Nikolaos Athanasopoulos, Nathalie Mitton, and Symeon Papavassiliou. Task offloading in edge and cloud computing: a survey on mathematical, artificial intelligence and control theory solutions. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002322.

Mohamed:2021:SDN

[706] Arwa Mohamed, Mosab Hamdan, Suleman Khan, Ahmed Abdelaziz, Sharief F. Babiker, Muhammad Imran, and M. N. Marsono. Software-defined networks for resource allocation in cloud computing: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002152.

Huang:2021:MAR

[707] Binbin Huang, Xiao Liu, Shangguang Wang, Linxuan Pan, and Victor Chang. Multi-agent reinforcement learning for cost-aware collaborative task execution in energyharvesting D2D networks. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002334.

Gao:2021:EAB

[708] Honghao Gao, Walayat Hussain, Yuyu Yin, Wenbing Zhao, and Muddesar Iqbal. Editorial: AI-based mobile multimedia computing for datasmart processing. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002486.

Rojas:2021:CSH

[709] Elisa Rojas, Rashid Amin, Carmen Guerrero, Marco Savi, and Adib Rastegarnia. Challenges and solutions for hybrid SDN. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002498.

Irshad:2021:SBO

[710] Azeem Irshad, Shehzad Ashraf Chaudhry, Anwar Ghani, and Muhammad Bilal. A secure blockchain-oriented data delivery and collection scheme for 5Genabled IoD environment. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002668.

Anonymous:2021:S

[711] Anonymous. 4 September 2021. Computer Networks (Amsterdam, Netherlands: 1999), 196(??):??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBm

[712] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 196(??):??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621003650.

Khalily-Dermany:2021:TPA

[713] M. Khalily-Dermany. Transmission power assignment in network-coding-based-multicast-wireless-sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100253X.

Sun:2021:CCD

[714] Rong Sun, Huihui Zheng, and Jingwei Liu. Coded caching design for fogaided networks. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002760.

Khan:2021:BBD

[715] Ammar Ahmed Khan, Muhammad Mubashir Khan, Kashif Mehboob Khan, Junaid Arshad, and Farhan Ahmad. A blockchain-based decentralized machine learning framework for collaborative intrusion detection within UAVs. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002644.

Arce:2021:FBL

[716] Pau Arce, David Salvo, Gema Piñero, and Alberto Gonzalez. FIWARE based low-cost wireless acoustic sensor network for monitoring and classification of urban soundscape. Computer Networks (Amsterdam, Netherlands: 1999), 196(??):??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002504.

Achroufene:2021:MCC

[717] Achour Achroufene, Mourad Chelik, and Nassima Bouadem. Modified CSMA/CA protocol for realtime data fusion applications based on clustered WSN. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002802.

Bruschi:2021:FHP

[718] Valerio Bruschi, Salvatore Pontarelli, Jerome Tollet, Dave Barach, and Giuseppe Bianchi. FlowFight: High performance-low memory top-k spreader detection. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002772.

Paul:2021:BBS

[719] Rourab Paul, Nimisha Ghosh, Suman Sau, Amlan Chakrabarti, and Prasant Mohapatra. Blockchain based secure smart city architecture using low resource IoTs. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002759.

Chadda:2021:ACW

[720] Amel Chadda, Marija Stojanova, Thomas Begin, Anthony Busson, and Isabelle Guérin Lassous. Assigning channels in WLANs with channel bonding: a fair and robust strategy. Computer Networks (Amsterdam, Netherlands: 1999), 196(??): ??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002516.

Behera:2021:PMB

[721] Sadananda Behera and Goutam Das. Penalty-method based impairmentaware performance of elastic optical networks with spectrum conversion. Computer Networks (Amsterdam, Netherlands: 1999), 196(??): ??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002784.

Aghaei:2021:TNA

[722] Abdollah Aghaei, Javad Akbari Torkestani, Hamidreza Kermajani, and Abbas Karimi. LA-Trickle: a novel algorithm to reduce the convergence time of the wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002796.

Zhang:2021:RAM

[723] Bibo Zhang, Francesco Devoti, Ilario Filippini, and Danilo De Donno. Resource allocation in mmWave 5G IAB networks: a reinforcement learning approach based on column generation. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002838.

Wan:2021:JCO

[724] Zhilan Wan, Ding Xu, Dahu Xu, and Ishtiaq Ahmad. Joint computation offloading and resource allocation for NOMA-based multi-access mobile edge computing systems. Computer Networks (Amsterdam, Netherlands: 1999), 196(??):??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002863.

Uddin:2021:DPA

[725] Md. Forkan Uddin. Downlink performance analysis of a CSMA based WLAN under physical interference model. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002887.

Tseng:2021:LLR

[726] Yuh-Min Tseng, Jian-Lun Chen, and Sen-Shan Huang. A lightweight leakage-resilient identity-based mutual authentication and key exchange protocol for resource-limited devices. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002826.

Xie:2021:SAD

[727] Guorui Xie, Qing Li, and Yong Jiang. Self-attentive deep learning method for online traffic classification and its interpretability. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002930.

Touati:2021:CRD

[728] Haifa Touati, Amira Chriki, Hichem Snoussi, and Farouk Kamoun. Cognitive radio and dynamic TDMA for efficient UAVs swarm communications. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002929.

Krishnamurthy:2021:AFA

[729] Prashanth Krishnamurthy, Farshad Khorrami, and Rahul Kumar. An approximate factorization approach to multi-jammer location and range estimation from peer-to-peer connectivity measurements. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002942.

Kiani:2021:NAP

[730] Mohsen Kiani and Mohammad Reza Khayyambashi. A network-aware and power-efficient virtual machine placement scheme in cloud datacenters based on chemical reaction optimization. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002954.

Aurizzi:2021:SBT

[731] Matteo Maria Aurizzi, Tommaso Rossi, Emanuele Raso, Ludovico Funari, and Ernestina Cianca. An SDN-based traffic handover control procedure and SGD management logic for EHF satellite networks. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002905.

Zambianco:2021:IMB

[732] Marco Zambianco and Giacomo Verticale. Intelligent multi-branch allocation of spectrum slices for inter-numerology interference minimization. Computer Networks (Amsterdam, Netherlands: 1999), 196(??):??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002875.

Tang:2021:UIE

[733] Yutao Tang, Yue Li, Qun Li, Kun Sun, Haining Wang, and Zhengrui Qin. User input enrichment via sensing devices. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002917.

Gaur:2021:SVP

[734] Kuntal Gaur, Anshuman Kalla, Jyoti Grover, Mohammad Borhani, Andrei Gurtov, and Madhusanka Liyanage. A survey of virtual private LAN services (VPLS): Past, present and future. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002814.

Bayazeed:2021:SSC

[735] Adnan Bayazeed, Khaldoun Khorzom, and Mohamad Aljnidi. A survey of self-coordination in self-organizing network. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002681.

Nayak:2021:SRB

[736] Sabuzima Nayak, Ripon Patgiri, and Angana Borah. A survey on the roles of Bloom filter in implementation of the named data networking. Computer Networks (Amsterdam, Netherlands: 1999), 196(??):??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002747.

Wang:2021:DNL

[737] Dongsheng Wang, Prayag Tiwari, Mohammad Shorfuzzaman, and Ingo Schmitt. Deep neural learning on weighted datasets utilizing label disagreement from crowdsourcing. Computer Networks (Amsterdam, Netherlands: 1999), 196(??): ??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002711.

Chen:2021:DQN

[738] Xin Chen, Xu Liu, Ying Chen, Libo Jiao, and Geyong Min. Deep Q-Network based resource allocation for UAV-assisted ultra-dense networks. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100284X.

Pustokhina:2021:EEN

[739] Irina V. Pustokhina, Denis A. Pustokhin, E. Laxmi Lydia, Mohamed Elhoseny, and K. Shankar. Energy efficient neuro-fuzzy cluster based topology construction with metaheuristic route planning algorithm for unmanned aerial vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002632.

Paguada:2021:TPA

[740] Servio Paguada, Lejla Batina, and Igor Armendariz. Toward practical autoencoder-based side-channel analysis evaluations. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002735.

Dhananjay:2021:PRV

[741] Aditya Dhananjay, Kai Zheng, Marco Mezzavilla, Lorenzo Iotti, Dennis Shasha, and Sundeep Rangan. Pi-Radio v1: Calibration techniques to enable fully-digital beamforming at 60 GHz. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002620.

Anonymous:2021:Oa

[742] Anonymous. 9 October 2021. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBn

[743] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621004060.

Dinh:2021:DUM

[744] Thi Ha Ly Dinh, Megumi Kaneko, Keisuke Wakao, Kenichi Kawamura, Takatsune Moriyama, Hirantha Abeysekera, and Yasushi Takatori. Distributed user-to-multiple access points association through deep learning for beyond 5G. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002899.

Rahman:2021:AAS

[745] Md. Rashedur Rahman, Md. Motaharul Islam, Ahmed Iqbal Pritom, and Yazed Alsaawy. ASRPH: Application specific routing protocol for health care. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002978.

Yang:2021:AHR

[746] Xuwei Yang, Hongli Xu, Jianchun Liu, Chen Qian, Xingpeng Fan, He Huang, and Haibo Wang. Achieving high reliability and throughput in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002966.

Campeon:2021:JPC

[747] Juan Pablo Grisales Campeón and Pablo I. Fierens. Joint position and clock tracking of wireless nodes. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003091.

Shakouri:2021:PRL

[748] Saba Shakouri and Akbar Ghaffarpour Rahbar. Proactive restoration with low re-provisioning in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003042.

Dinarte:2021:RSA

[749] Henrique A. Dinarte, Bruno V. A. Correia, Daniel A. R. Chaves, and Raul C. Almeida. Routing and spectrum assignment: a metaheuristic for hybrid ordering selection in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003054.

Shao:2021:PRI

[750] Xiaozhe Shao and Lixin Gao. Policyrich interdomain routing with local coordination. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100308X.

Fang:2021:CCB

[751] Yong Fang, Kai Li, Rongfeng Zheng, Shan Liao, and Yue Wang. A communication-channel-based method for detecting deeply camouflaged malicious traffic. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100311X.

Wang:2021:CSD

[752] Yue Wang, Anmin Zhou, Shan Liao, Rongfeng Zheng, Rong Hu, and Lei Zhang. A comprehensive survey on DNS tunnel detection. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003248.

Fotohi:2021:SCB

[753] Reza Fotohi and Fereidoon Shams Aliee. Securing communication between things using blockchain technology based on authentication and SHA-256 to improving scalability in large-scale IoT. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003303.

Yin:2021:IDC

[754] Sheng lin Yin, Xing lan Zhang, and Shuo Liu. Intrusion detection for capsule networks based on dual routing mechanism. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003327.

Vegni:2021:MMB

[755] Anna Maria Vegni, Carlos Borrego Iglesias, and Valeria Loscrí. MOVES: a MemOry-based VEhicular Social forwarding technique. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003236.

Li:2021:SBN

[756] Quanzhong Li and Liang Yang. Secure beamforming with nonorthogonal multiple access transmission in cooperative CR networks for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003340.

Thang:2021:JBU

[757] Ha Duc Thang, Lila Boukhatem, Megumi Kaneko, and Nhan Nguyen-Thanh. Joint beamforming and user association with reduced CSI signaling in mobile environments: a Deep Q-learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003078.

Breen:2021:PPO

[758] Joe Breen, Andrew Buffmire, Jonathon Duerig, Kevin Dutt, Eric Eide, Anneswa Ghosh, Mike Hibler, David Johnson, Sneha Kumar Kasera, Earl Lewis, Dustin Maas, Caleb Martin, Alex Orange, Neal Patwari, Daniel Reading, Robert Ricci, David Schurig, Leigh B. Stoller, Allison Todd, Jacobus Van der Merwe, Naren Viswanathan, Kirk Webb, and Gary Wong. Powder: Platform for open wireless data-driven experimental research. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003017.

Gringoli:2021:LDH

[759] Francesco Gringoli and Douglas J. Leith. Low-delay high-rate operation of 802.11ac WLAN downlink: Nonlinear controller analysis and design. Computer Networks (Amsterdam, Netherlands: 1999), 197(??): ??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003261.

Lu:2021:CCM

[760] Yifei Lu, Xu Ma, and Zhengzhi Xu. Choose a correct marking position: ECN should be freed from tail mark. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003297.

SureshKumar:2021:EER

[761] K. SureshKumar and P. Vimala. Energy efficient routing protocol using exponentially-ant lion whale optimization algorithm in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002851.

Belhamra:2021:ECN

[762] Mohamed Amine Belhamra and El Mamoun Souidi. Error correcting network codes. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002991.

Chai:2021:NCO

[763] Rong Chai, Xizheng Yang, Chunling Du, and Qianbin Chen. Network cost optimization-based capacitated controller deployment for SDN. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003273.

Bhooanusas:2021:PMM

[764] Nuntanut Bhooanusas and Sok-Ian Sou. Performance modeling of multipath mobile data offloading in cellular/WiFi networks with bandwidth uncertainty. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003431.

Wei:2021:ASE

[765] Guiyi Wei, Xiaohang Mao, Rongxing Lu, Jun Shao, Yunguo Guan, and Genhua Lu. Achieve spaceefficient key management in lightning network. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003406.

Lezzar:2021:OUR

[766] Mohamed Yacine Lezzar and Mustafa Mehmet-Ali. Optimization of ultrareliable low-latency communication systems. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003339.

Seoane:2021:PEC

[767] Victor Seoane, Carlos Garcia-Rubio, Florina Almenares, and Celeste Campo. Performance evaluation of CoAP and MQTT with security support for IoT environments. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003364.

Lindroos:2021:SMC

[768] Saku Lindroos, Antti Hakkala, and Seppo Virtanen. A systematic methodology for continuous WLAN abundance and security analysis. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003479.

Dziyauddin:2021:COC

[769] Rudzidatul Akmam Dziyauddin, Dusit Niyato, Nguyen Cong Luong, Ahmad Ariff Aizuddin Mohd Atan, Mohd Azri Mohd Izhar, Marwan Hadri Azmi, and Salwani Mohd Daud. Computation offloading and content caching and delivery in vehicular edge network: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002723.

Madi:2021:NSS

[770] Taous Madi, Hyame Assem Alameddine, Makan Pourzandi, and Amine Boukhtouta. NFV security survey in 5G networks: a three-dimensional threat taxonomy. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003066.

Tang:2021:EFS

[771] Yinan Tang, Tongtong Yuan, Bo Liu, and Chuangbai Xiao. Effective *- flow schedule for optical circuit switching based data center networks: a comprehensive survey. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003224.

Alvares:2021:DAL

[772] Christabelle Alvares, Dristi Dinesh, Syed Alvi, Tannish Gautam, Maheen Hasib, and Ali Raza. Dataset of attacks on a live enterprise VoIP network for machine learning based intrusion detection and prevention systems. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003030.

Boucetta:2021:LRB

[773] Chérifa Boucetta, Boubakr Nour, Michel Sortais, and Hassine Moungla. A Latin rectangles-based TSCH scheduling and interference mitigation design. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100298X.

Al-Marridi:2021:RLA

[774] Abeer Z. Al-Marridi, Amr Mohamed, and Aiman Erbad. Reinforcement learning approaches for efficient and secure blockchain-powered smart health systems. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003005.

Khan:2021:QLB

[775] Zahoor Ali Khan, Obaida Abdul Karim, Shahid Abbas, Nadeem Javaid, Yousaf Bin Zikria, and Usman Tariq. Q-learning based energyefficient and void avoidance routing protocol for underwater acoustic sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003212.

Miao:2021:FDR

[776] Qinyang Miao, Hui Lin, Xiaoding Wang, and Mohammad Mehedi Hassan. Federated deep reinforcement learning based secure data sharing for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003285.

Lv:2021:BDT

[777] Zhihan Lv, Dongliang Chen, Hailing Feng, Ranran Lou, and Huihui Wang. Beyond 5G for digital twins of UAVs. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003534.

Anonymous:2021:Ob

[778] Anonymous. 24 October 2021. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBo

[779] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128621004278.

Parsa:2021:TAM

[780] Mahshad Parsa and Seyed Ahmad Motamedi. Throughput analysis in a millimeter-wave network for a hybrid unsaturated RCAP/CTAP MAC protocol. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003443.

Busacca:2021:DML

[781] Fabio Busacca, Giuseppe Faraci, Christian Grasso, Sergio Palazzo, and Giovanni Schembra. Designing a multi-layer edge-computing platform for energy-efficient and delayaware offloading in vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003315.

Liu:2021:BTR

[782] Yuchen Liu and Douglas M. Blough. Blockage tolerance in roadside millimeterwave backhaul networks. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003571.

Xu:2021:GAC

[783] Rui Xu, Xu Wang, and Kirill Morozov. Group authentication for cloudto-things computing: Review and improvement. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100356X.

Cai:2021:FRD

[784] Lingfeng Cai, Xianglin Wei, Changyou Xing, Xia Zou, Guomin Zhang, and Xiulei Wang. Failure-resilient DAG task scheduling in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003480.

Mao:2021:MMS

[785] Beifeng Mao, Jing Liu, Yingxu Lai, and Motong Sun. MIF: a multi-step attack scenario reconstruction and attack chains extraction method based on multi-information fusion. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003376.

Munilla:2021:ESK

[786] Jorge Munilla, Mike Burmester, and Raquel Barco. An enhanced symmetric-key based 5G-AKA protocol. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003546.

Zilberman:2021:HSC

[787] Aviram Zilberman, Yoram Haddad, Sefi Erlich, Yossi (Joseph) Peretz, and Amit Dvir. Heterogeneous SDN controller placement problem-the Wi-Fi and 4G LTE-U case. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003558.

daCosta:2021:QFI

[788] Luis Antonio L. F. da Costa, Rafael Kunst, and Edison Pignaton de Freitas. Q-FANET: Improved Qlearning based routing protocol for FANETs. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003595.

Ma:2021:ASA

[789] Dongchao Ma, Xiaofu Huang, Yuekun Hu, Pengyu Wang, Mingwei Xu, and Li Ma. An adaptive solar-aware framework and strategy for outdoor deployment of WSN. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003583.

Nayyer:2021:LBH

[790] Amit Nayyer, Aman Kumar Sharma, and Lalit Kumar Awasthi. Learningbased hybrid routing for scalability in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003492.

Mokhtar:2021:MLT

[791] Hamza Mokhtar, Xiaoqiang Di, Ying Zhou, Alzubair Hassan, Ziyi Ma, and Shafiu Musa. Multiple-level threshold load balancing in distributed SDN controllers. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003510.

Zheng:2021:ADP

[792] Ling Zheng, Weitao Pan, Ya Gao, Huan Liu, and Jing Jiang. Architecture design and performance analysis of a novel memory system for high-bandwidth onboard switching fabric. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003509.

Qi:2021:PPS

[793] Xiaogang Qi, Jiahui Li, Zhiping Wang, and Lifang Liu. Probabilistic probe selection algorithm for fault diagnosis in communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003522.

Yan:2021:PIF

[794] Fulong Yan, Hugo Meyer, Changshun Yuan, Xuwei Xue, Bitao Pan, Xiaotao Guo, Nicola Calabretta, and Chongjin Xie. On the performance investigation of a fast optical switches based optical high performance computing infrastructure. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003418.

Feng:2021:MTM

[795] Li Feng, Yiru Yao, Liangmin Wang, and Geyong Min. Multi-timescale and multi-centrality layered node selection for efficient traffic monitoring in SDNs. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003601.

Ding:2021:UCI

[796] Lan Ding, Ye Tian, Tong Liu, Zhongxiang Wei, and Xinming Zhang. Understanding commercial 5G and its implications to (multipath) TCP. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003765.

Hou:2021:DTF

[797] Changsheng Hou, Bingnan Hou, Tongqing Zhou, and Zhiping Cai. DMatrix: Toward fast and accurate queries in graph stream. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003777.

Tian:2021:MNM

[798] Donghai Tian, Qianjin Ying, Xiaoqi Jia, Rui Ma, Changzhen Hu, and Wenmao Liu. MDCHD: a novel malware detection method in cloud using hardware trace and deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003728.

Chen:2021:FSW

[799] Mantun Chen, Yongjun Wang, Hongzuo Xu, and Xiatian Zhu. Few-shot website fingerprinting attack. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003108.

Roostaei:2021:GTJ

[800] Razie Roostaei, Zahra Dabiri, and Zeinab Movahedi. A game-theoretic joint optimal pricing and resource allocation for Mobile Edge Computing in NOMA-based 5G networks and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100342X.

He:2021:TSD

[801] Lin He, Peng Kuang, Ying Liu, Gang Ren, and Jiahai Yang. Towards securing duplicate address detection using P4. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100325X.

Prathapchandran:2021:TAS

[802] K. Prathapchandran and T. Janani. A trust aware security mechanism to detect sinkhole attack in RPL-based IoT environment using random forest — RFTRUST. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003832.

Khaturia:2021:FUM

[803] Meghna Khaturia, Pranav Jha, and Abhay Karandikar. 5G-Flow: a unified multi-RAT RAN architecture for beyond 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003820.

Rea:2021:BSM

[804] Maurizio Rea, Domenico Giustiniano, Pablo Jiménez Mateo, Yago Lizarribar, and Joerg Widmer. Beam searching for mmWave networks with sub-6 GHz WiFi and inertial sensors inputs: an experimental study. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100339X.

Shokouhifar:2021:SIR

[805] Mohammad Shokouhifar. Swarm intelligence RFID network planning using multi-antenna readers for asset tracking in hospital environments. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003911.

Eriskin:2021:PCH

[806] Levent Eriskin. Point coverage with heterogeneous sensor networks: a robust optimization approach under target location uncertainty. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003844.

Ali:2021:SCO

[807] Zaiwar Ali, Ziaul Haq Abbas, Ghulam Abbas, Abdullah Numani, and Muhammad Bilal. Smart computational offloading for mobile edge computing in next-generation Internet of Things networks. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003467.

Gharrad:2021:FSD

[808] Hana Gharrad, Nafaa Jabeur, Ansar Ul-Haque Yasar, Stephane Galland, and Mohammed Mbarki. A fivestep drone collaborative planning approach for the management of distributed spatial events and vehicle notification using multi-agent systems and firefly algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003029.

Perdices:2021:NLP

[809] Daniel Perdices, Javier Ramos, José L. García-Dorado, Iván González, and Jorge E. López de Vergara. Natural language processing for web browsing analytics: Challenges, lessons learned, and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003455.

Alam:2021:BBI

[810] Shadab Alam, Mohammed Shuaib, Wazir Zada Khan, Sahil Garg, Georges Kaddoum, M. Shamim Hossain, and Yousaf Bin Zikria. Blockchain-based initiatives: Current state and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100373X.

Jin:2021:ESC

[811] Shan Jin and Riccardo Bettati. Efficient side-channel attacks beyond divide-and-conquer strategy. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003807.

Rioja:2021:ATP

[812] Unai Rioja, Lejla Batina, Jose Luis Flores, and Igor Armendariz. Autotune POIs: Estimation of distribution algorithms for efficient side-channel analysis. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003789.

Esfahani:2021:ECA

[813] Mahdi Esfahani, Hadi Soleimany, and Mohammad Reza Aref. Enhanced cache attack on AES applicable on ARM-based devices with new operating systems. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003790.

Anonymous:2021:N

[814] Anonymous. 9 November 2021. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBp

[815] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621004679.

Ahat:2021:OSS

[816] Betül Ahat, Ahmet Cihat Baktır, Necati Aras, İ. Kuban Altınel, Atay Özgövde, and Cem Ersoy. Optimal server and service deployment for multi-tier edge cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 199(??): ??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003716.

Qin:2021:CAM

[817] Hua Qin, Weihong Chen, Weimin Chen, Ni Li, Min Zeng, and Yang Peng. A collision-aware mobile tag reading algorithm for RFID-based vehicle localization. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003881.

Baarli:2021:DMU

[818] Eirik Molde Bårli, Anis Yazidi, Enrique Herrera Viedma, and Hårek Haugerud. DoS and DDoS mitigation using variational autoencoders. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003753.

Liu:2021:CEA

[819] Jianchun Liu, Hongli Xu, Yang Xu, Zhenguo Ma, Zhiyuan Wang, Chen Qian, and He Huang. Communicationefficient asynchronous federated learning in resource-constrained edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003923.

Salehi:2021:IUB

[820] Shavbo Salehi and Behdis Eslamnour. Improving UAV base station energy efficiency for industrial IoT URLLC services by irregular repetition slotted-ALOHA. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003856.

Halgamuge:2021:OFB

[821] Malka N. Halgamuge. Optimization framework for best approver selection method (BASM) and best tip selection method (BTSM) for IOTA tangle network: Blockchain-enabled next generation industrial IoT. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003868.

Popli:2021:GNA

[822] Sakshi Popli, Rakesh Kumar Jha, and Sanjeev Jain. Green NOMA assisted NB-IoT based urban farming in multistory buildings. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003819.

Xiang:2021:RCM

[823] Bin Xiang, Jocelyne Elias, Fabio Martignon, and Elisabetta Di Nitto. Resource calendaring for mobile edge computing: Centralized and decentralized optimization approaches. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912862100390X.

Nait-Hamoud:2021:CPK

[824] Othmane Nait-Hamoud, Tayeb Kenaza, and Yacine Challal. Certificateless public key systems aggregation: an enabling technique for 5G multidomain security management and delegation. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004011.

Arana:2021:NQA

[825] Oscar Arana, Hector Benítez-Pérez, Javier Gomez, and Miguel Lopez-Guerrero. Never query alone: a distributed strategy to protect Internet users from DNS fingerprinting attacks. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004023.

Deng:2021:SDC

[826] Qingyong Deng, Shaobo Huang, Zhetao Li, Bin Guo, Liyao Xiang, and Rong Ran. A secure data collection strategy using mobile vehicles joint UAVs in smart city. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003996.

Liu:2021:FAP

[827] Juncai Liu, Jessie Hui Wang, Chenghao Rong, Yuedong Xu, Tao Yu, and Jilong Wang. FedPA: an adaptively partial model aggregation strategy in federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004199.

Amanlou:2021:LSA

[828] Sanaz Amanlou, Mohammad Kamrul Hasan, and Khairul Azmi Abu Bakar. Lightweight and secure authentication scheme for IoT network based on publish-subscribe fog computing model. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004175.

Sadri:2021:ATF

[829] Mohammad Javad Sadri and Maryam Rajabzadeh Asaar. An anonymous two-factor authentication protocol for IoT-based applications. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004151.

Vemireddy:2021:FRL

[830] Satish Vemireddy and Rashmi Ranjan Rout. Fuzzy reinforcement learning for energy efficient task offloading in vehicular fog computing. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004163.

Liao:2021:TAB

[831] Xuening Liao, Zhenqiang Wu, Yuanyu Zhang, and Xiaohong Jiang. Trustaware buffer-aided relay selection for secure communications in cooperative wireless systems. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004126.

Alhowaidi:2021:CML

[832] Mohammad Alhowaidi, Deepak Nadig, Boyang Hu, Byrav Ramamurthy, and Brian Bockelman. Cache management for large data transfers and multipath forwarding strategies in named data networking. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003972.

Sheshjavani:2021:CCS

[833] Abdollah Ghaffari Sheshjavani, Ahmad Khonsari, Seyed Pooya Shariatpanahi, and Masoumeh Moradian. Content caching for shared medium networks under heterogeneous users’ behaviors. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004114.

Mall:2021:CPB

[834] Priyanka Mall, Ruhul Amin, Mohammad S. Obaidat, and Kuei-Fang Hsiao. CoMSeC++: PUF-based secured lightweight mutual authentication protocol for drone-enabled WSN. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004230.

Araujo:2021:HOM

[835] Samuel M. A. Araújo, Fernanda S. H. de Souza, and Geraldo R. Mateus. A hybrid optimization-machine learning approach for the VNF placement and chaining problem. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004229.

Samaila:2021:PES

[836] Musa G. Samaila, Carolina Lopes, Édi Aires, João B. F. Sequeiros, Tiago Simões, Mário M. Freire, and Pedro R. M. Inácio. Performance evaluation of the SRE and SBPG components of the IoT hardware platform security advisor framework. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004370.

Popli:2021:CSG

[837] Sakshi Popli, Rakesh Kumar Jha, and Sanjeev Jain. A comprehensive survey on green ICT with 5G-NB-IoT: Towards sustainable planet. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003959.

Fourati:2021:CSS

[838] Hasna Fourati, Rihab Maaloul, Lamia Chaari, and Mohamed Jmaiel. Comprehensive survey on self-organizing cellular network approaches applied to 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003960.

Jagannath:2021:DMC

[839] Anu Jagannath and Jithin Jagannath. Dataset for modulation classification and signal type classification for multitask and single task learning. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100400X.

Arnold:2021:CDB

[840] Rachel Arnold and Dave Longley. Continuity: a deterministic Byzantine fault tolerant asynchronous consensus algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003935.

Liu:2021:FDD

[841] Yong Liu and Guangxia Xu. Fixed degree of decentralization DPoS consensus mechanism in blockchain based on adjacency vote and the average fuzziness of vague value. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003947.

Platt:2021:SAI

[842] Moritz Platt and Peter McBurney. Sybil attacks on identity-augmented proof-of-stake. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003893.

Liu:2021:TOO

[843] Peng Liu, Han He, Tingting Fu, Huijuan Lu, Abdulhameed Alelaiwi, and Md Wasif Islam Wasi. Task offloading optimization of cruising UAV with fixed trajectory. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003741.

Wei:2021:COM

[844] Dawei Wei, Jianfeng Ma, Linbo Luo, Yunbo Wang, Lei He, and Xinghua Li. Computation offloading over multi-UAV MEC network: a distributed deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003984.

Sharaff:2021:SMD

[845] Aakanksha Sharaff, Chandramani Kamal, Siddhartha Porwal, Surbhi Bhatia, Kuljeet Kaur, and Mohammad Mehendi Hassan. Spam message detection using Danger theory and Krill herd optimization. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004102.

Cheng:2021:MLN

[846] Jin Cheng, Yulei Wu, Yuepeng E, Junling You, Tong Li, Hui Li, and Jingguo Ge. MATEC: a lightweight neural network for online encrypted traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004217.

Kohli:2021:OAF

[847] Manav Kohli, Tingjun Chen, Mahmood Baraani Dastjerdi, Jackson Welles, Ivan Seskar, Harish Krishnaswamy, and Gil Zussman. Openaccess full-duplex wireless in the OR-BIT and COSMOS testbeds. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100387X.

Velasco:2021:LMS

[848] Francisco Alcaraz Velasco, Jose Manuel Palomares, and Joaquin Olivares. Lightweight method of shuffling overlapped data-blocks for data integrity and security in WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004205.

Anonymous:2021:Da

[849] Anonymous. 9 December 2021. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBq

[850] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621005028.

Tamang:2021:ARS

[851] Dinesh Tamang, Sergio Martiradonna, Andrea Abrardo, Gianluca Mandó, Gabriele Roncella, and Gennaro Boggia. Architecting 5G RAN slicing for location aware vehicle to infrastructure communications: the Autonomous Tram use case. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004400.

Garrett:2021:HAL

[852] Thiago Garrett, Luis C. E. Bona, and Elias P. Duarte. A holistic approach for locating traffic differentiation in the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004333.

Montero:2021:PRQ

[853] Emanuel Montero, Carlos Rocha, Helder Oliveira, Eduardo Cerqueira, Paulo Mendes, Aldri Santos, and Denis Rosário. Proactive radio- and QoSaware UAV as BS deployment to improve cellular operations. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100431X.

Batchu:2021:GML

[854] Raj Kumar Batchu and Hari Seetha. A generalized machine learning model for DDoS attacks detection using hybrid feature selection and hyperparameter tuning. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004394.

Wang:2021:DGV

[855] Shengling Wang, Yu Pu, Hongwei Shi, Jianhui Huang, and Yinhao Xiao. A differential game view of antagonistic dynamics for cybersecurity. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004357.

Zaabar:2021:HSB

[856] Bessem Zaabar, Omar Cheikhrouhou, Faisal Jamil, Meryem Ammi, and Mohamed Abid. HealthBlock: a secure blockchain-based healthcare data management system. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004382.

Moosavi:2021:EET

[857] Reza Moosavi, Saeedeh Parsaeefard, Mohammad Ali Maddah-Ali, Vahid Shah-Mansouri, Babak Hossein Khalaj, and Mehdi Bennis. Energy efficiency through joint routing and function placement in different modes of SDN/NFV networks. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004345.

Aljeri:2021:EES

[858] Noura Aljeri and Azzedine Boukerche. EDiPSo: an efficient scalable topology discovery protocol for softwaredefined vehicular networks. Computer Networks (Amsterdam, Netherlands:

1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003388.

Sharmila:2021:EIA

[859] A. Helen Sharmila and N. Jaisankar. Edge intelligent agent assisted hybrid hierarchical blockchain for continuous healthcare monitoring and recommendation system in 5G WBAN-IoT. Computer Networks (Amsterdam, Netherlands: 1999), 200(??): ??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004448.

Jung:2021:RTD

[860] SungHoon Jung and SaeWoong Bahk. Real-time dual-link transmission control for minimizing power and link switching cost. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004187.

Streit:2021:NAD

[861] Ananda Streit, Gustavo H. A. Santos, Rosa M. M. Leão, Edmundo de Souza e Silva, Daniel Menasché, and Don Towsley. Network anomaly detection based on tensor decomposition. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004412.

Pal:2021:BBT

[862] Shantanu Pal, Ambrose Hill, Tahiry Rabehaja, and Michael Hitchens. A blockchain-based trust management framework with verifiable interactions. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004436.

Lekssays:2021:PBB

[863] Ahmed Lekssays, Luca Landa, Barbara Carminati, and Elena Ferrari. PAutoBotCatcher: a blockchain-based privacy-preserving botnet detector for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100445X.

Yuan:2021:CHR

[864] Peiyan Yuan, Saike Shao, Lijuan Geng, and Xiaoyan Zhao. Caching hit ratio maximization in mobile edge computing with node cooperation. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004424.

Powell:2021:PPS

[865] Warwick Powell, Shoufeng Cao, Thomas Miller, Marcus Foth, Xavier Boyen, Barry Earsman, Santiago del Valle, and Charles Turner-Morris. From premise to practice of social consensus: How to agree on common knowledge in blockchain-enabled supply chains. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004606.

Zhan:2021:WFE

[866] Pengwei Zhan, Liming Wang, and Yi Tang. Website fingerprinting on early QUIC traffic. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004618.

Chen:2021:VCA

[867] Ning Chen, Sheng Zhang, Siyi Quan, Zhi Ma, Zhuzhong Qian, and Sanglu Lu. VCMaker: Content-aware configuration adaptation for video streaming and analysis in live augmented reality. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004461.

Fu:2021:TRE

[868] Xiuwen Fu, Pasquale Pace, Gianluca Aloi, Wenfeng Li, and Giancarlo Fortino. Toward robust and energy-efficient clustering wireless sensor networks: a double-stage scale-free topology evolution model. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004527.

Ujjwal:2021:DAH

[869] Ujjwal, Neha Kumari, Jaisingh Thangaraj, and Chhandita Roy. Distance adaptive hybrid super-channels enabled by sliceable bandwidth variable transponder for spectrally efficient elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004588.

Su:2021:GBD

[870] Yi Su, Wenhao Fan, Yuan’an Liu, and Fan Wu. Game-based distributed pricing and task offloading in multi-cloud and multi-edge environments. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004539.

Zhang:2021:QSE

[871] Jianwei Zhang and Chenwei Zhao. Q-SR: an extensible optimization framework for segment routing. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004497.

Ma:2021:ISC

[872] Huihui Ma, Du Xu, Yueyue Dai, and Qing Dong. An intelligent scheme for congestion control: When active queue management meets deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004485.

Wu:2021:TEF

[873] Yulei Wu, Laizhong Cui, Victor C. M. Leung, Tarik Taleb, and Sangheon Pack. Towards efficient and flexible management and interworking techniques for industrial Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004631.

Robinson:2021:SCC

[874] Peter Robinson. Survey of crosschain communications protocols. Computer Networks (Amsterdam, Netherlands:

1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004321.

Pillai:2021:BCA

[875] Babu Pillai, Kamanashis Biswas, Zhé Hóu, and Vallipuram Muthukkumarasamy. Burn-to-claim: an asset transfer protocol for blockchain interoperability. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004369.

Tang:2021:STM

[876] Xinyu Tang, Cheng Guo, Kim-Kwang Raymond Choo, Yining Liu, and Long Li. A secure and trustworthy medical record sharing scheme based on searchable encryption and blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100462X.

Wang:2021:BSS

[877] Pan Wang, Zixuan Wang, Feng Ye, and Xuejiao Chen. ByteSGAN: a semisupervised generative adversarial network for encrypted traffic classification in SDN edge gateway. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100459X.

Montieri:2021:PLP

[878] Antonio Montieri, Giampaolo Bovenzi, Giuseppe Aceto, Domenico Ciuonzo, Valerio Persico, and Antonio Pescapè. Packet-level prediction of mobile-app traffic using multitask deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 200(??): ??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004564.

Shusterman:2021:CBC

[879] Anatoly Shusterman, Chen Finkelstein, Ofir Gruner, Yarin Shani, and Yossi Oren. Cache-based characterization: a low-infrastructure, distributed alternative to network-based traffic and application characterization. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004710.

Anonymous:2021:Db

[880] Anonymous. 24 December 2021. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2021:EBr

[881] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621005272.

Zhang:2021:DSS

[882] Zhishuo Zhang and Shijie Zhou. A decentralized strongly secure attributebased encryption and authentication scheme for distributed Internet of mobile things. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004722.

Mao:2021:DLD

[883] Qian Mao, Lin Zhang, Fei Hu, Elizabeth Serena Bentley, and Sunil Kumar. Deep learning (DL)-based adaptive transport layer control in UAV swarm networks. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004473.

Candela:2021:WSG

[884] Massimo Candela, Valerio Luconi, and Alessio Vecchio. A worldwide study on the geographic locality of Internet routes. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004734.

Wang:2021:EEC

[885] Bin Wang, Fagui Liu, Weiwei Lin, Zhenjiang Ma, and Dishi Xu. Energyefficient collaborative optimization for VM scheduling in cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004783.

Palma:2021:PPC

[886] Noelia Pérez Palma, Falko Dressler, and Vincenzo Mancuso. Precise: Predictive content dissemination scheme exploiting realistic mobility patterns. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004746.

Atif:2021:SNM

[887] Syed Muhammad Atif, Nicolas Gillis, Sameer Qazi, and Imran Naseem. Structured nonnegative matrix factorization for traffic flow estimation of large cloud networks. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004771.

Liang:2021:CUD

[888] Shuang Liang, Zhiyi Fang, Geng Sun, Chi Lin, Jiahui Li, Songyang Li, and Aimin Wang. Charging UAV deployment for improving charging performance of wireless rechargeable sensor networks via joint optimization approach. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004849.

Braeken:2021:DDG

[889] An Braeken. Device-to-device group authentication compatible with 5G AKA protocol. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004850.

Wazirali:2021:SAD

[890] Raniyah Wazirali, Rami Ahmad, and Ashraf Abdel-Karim Abu-Ein. Sustaining accurate detection of phishing URLs using SDN and feature selection approaches. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004941.

Davami:2021:DSM

[891] Fatemeh Davami, Sahar Adabi, Ali Rezaee, and Amir Masoud Rahmani. Distributed scheduling method for multiple workflows with parallelism prediction and DAG prioritizing for time constrained cloud applications. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100476X.

Yang:2021:SHS

[892] Ye Yang, Haiyang Jiang, Guangxing Zhang, Xin Wang, Yilong Lv, Xing Li, Serge Fdida, and Gaogang Xie. S2H: Hypervisor as a setter within Virtualized Network I/O for VM isolation on cloud platform. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004862.

Sari:2021:CRC

[893] T. Tolga Sari and Gökhan Seçinti. Chain RTS/CTS scheme with opportunistic channel allocation. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004874.

Sharma:2021:POU

[894] Abhishek Sharma, Amik Garg, Sanjay Kumar Sharma, Vibhav Kumar Sachan, and Parvin Kumar. Performance optimization for UWB communication network under IEEE 802.15.4a channel conditions. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621004904.

Garetto:2021:CPN

[895] Michele Garetto, Emilio Leonardi, and Giovanni Neglia. Content placement in networks of similarity caches. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004825.

Kumar:2021:DMP

[896] Sanjeev Kumar and B. Janet. Distinguishing malicious programs based on visualization and hybrid learning algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004953.

Rodrigues:2021:ABI

[897] Carlo Kleber da Silva Rodrigues. Analyzing blockchain integrated architectures for effective handling of IoTecosystem transactions. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005089.

Miuccio:2021:DBE

[898] Luciano Miuccio, Daniela Panno, and Salvatore Riolo. A DNN-based estimate of the PRACH traffic load for massive IoT scenarios in 5G networks and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005077.

Shinde:2021:NOB

[899] Swapnil Sadashiv Shinde, Dania Marabissi, and Daniele Tarchi. A network operator-biased approach for multi-service network function placement in a 5G network slicing architecture. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004989.

Torre:2021:PEM

[900] Roberto Torre, Muhammad Tayyab, George Koudouridis, Xavier Gelabert, Riccardo Bassoli, and Frank H. P. Fitzek. Power efficient mobile small cell placement for network-coded cooperation in UDNs. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004515.

Dhooge:2021:HFB

[901] Laurens D’hooge, Miel Verkerken, Tim Wauters, Bruno Volckaert, and Filip De Turck. Hierarchical feature block ranking for data-efficient intrusion detection modeling. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005107.

Jie:2021:OUF

[902] Lu Jie, Chen Hongchang, Sun Penghao, Hu Tao, and Zhang Zhen. OrderSketch: an unbiased and fast sketch for frequency estimation of data streams. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004795.

Abid:2021:SRC

[903] Khaled Abid, Hicham Lakhlef, and Abdelmadjid Bouabdallah. A survey on recent contention-free MAC protocols for static and mobile wireless decentralized networks in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004886.

Yu:2021:IAI

[904] Chuan Yu, Shuhui Chen, Fei Wang, and Ziling Wei. Improving 4G/ 5G air interface security: a survey of existing attacks on different LTE layers. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004576.

Farhadi:2021:FSM

[905] Babak Farhadi, Amir Masoud Rahmani, Parvaneh Asghari, and Mehdi Hosseinzadeh. Friendship selection and management in social internet of things: a systematic review. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004813.

Salahdine:2021:SSM

[906] Fatima Salahdine, Johnson Opadere, Qiang Liu, Tao Han, Ning Zhang, and Shaohua Wu. A survey on sleep mode techniques for ultra-dense networks in 5G and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004801.

Akter:2021:CSC

[907] Rubina Akter, Van-Sang Doan, Jae-Min Lee, and Dong-Seong Kim. CNN-SSDI: Convolution neural network inspired surveillance system for UAVs detection and identification. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004503.

Qian:2021:DMN

[908] Bing Qian and Shun Lu. Detection of mobile network abnormality using deep learning models on massive network measurement data. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004837.

Zhao:2021:SIE

[909] Yuyu Zhao, Guang Cheng, Yu Duan, Zhouchao Gu, Yuyang Zhou, and Lu Tang. Secure IoT edge: Threat situation awareness based on network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004540.

Ma:2021:GAP

[910] Baolin Ma, Chao Yang, Mingzhe Chen, and Jianfeng Ma. GramMatch: an automatic protocol feature extraction and identification system. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004552.

Ren:2021:AIB

[911] Qiuning Ren, Chao Yang, and Jianfeng Ma. App identification based on encrypted multi-smartphone sources traffic fingerprints. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100493X.

Liu:2021:PLP

[912] Yuan Liu, Yixiao Lan, Boyang Li, Chunyan Miao, and Zhihong Tian. Proof of learning (PoLe): Empowering neural network training with consensus building on blockchains. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004965.

Anonymous:2022:Ja

[913] Anonymous. 15 January 2022. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBa

[914] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128621005788.

Guarino:2022:AMP

[915] Alfonso Guarino, Delfina Malandrino, and Rocco Zaccagnino. An automatic mechanism to provide privacy awareness and control over unwittingly dissemination of online private information. Computer Networks (Amsterdam, Netherlands: 1999), 202 (??):??, January 15, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005090.

Horchulhack:2022:TFM

[916] Pedro Horchulhack, Eduardo K. Viegas, and Altair O. Santin. Toward feasible machine learning model updates in network-based intrusion detection. Computer Networks (Amsterdam, Netherlands: 1999), 202 (??):??, January 15, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005120.

He:2022:JOS

[917] Yexiao He, Xiaoning Zhang, Zixiang Xia, Yutao Liu, Keshav Sood, and Shui Yu. Joint optimization of service chain graph design and mapping in NFV-enabled networks. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005168.

Aureli:2022:ADO

[918] Davide Aureli, Antonio Cianfrani, Marco Listanti, Marco Polverini, and Stefano Secci. Augmenting DiffServ operations with dynamically learned classes of services. Computer Networks (Amsterdam, Netherlands: 1999), 202 (??):??, January 15, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005144.

Zhang:2022:DDA

[919] Chunyang Zhang, Gaogang Xie, and Xin Wang. DynamicTuple: the dynamic adaptive tuple for high-performance packet classification. Computer Networks (Amsterdam, Netherlands: 1999), 202(??): ??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005193.

Dutta:2022:DRL

[920] Hrishikesh Dutta and Subir Biswas. Distributed reinforcement learning for scalable wireless medium access in IoTs and sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005417.

Yang:2022:RWO

[921] Furong Yang, Andrea Ferlini, Davide Aguiari, Davide Pesavento, Rita Tse, Suman Banerjee, Gaogang Xie, and Giovanni Pau. Revisiting WiFi offloading in the wild for V2I applications. Computer Networks (Amsterdam, Netherlands: 1999), 202 (??):??, January 15, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005211.

Zeng:2022:UIO

[922] Man Zeng, Xiaohong Huang, Pei Zhang, and Dandan Li. Understanding the impact of outsourcing mitigation against BGP prefix hijacking. Computer Networks (Amsterdam, Netherlands: 1999), 202 (??):??, January 15, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005338.

Nguyen:2022:CAO

[923] Viet Dung Nguyen, Phi Le Nguyen, Kien Nguyen, and Phan Thuan Do. Constant approximation for opportunistic sensing in mobile air quality monitoring system. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005314.

Ruckel:2022:FIP

[924] Timon Rückel, Johannes Sedlmeir, and Peter Hofmann. Fairness, integrity, and privacy in a scalable blockchain-based federated learning system. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005132.

Goldoni:2022:CBW

[925] Emanuele Goldoni, Pietro Savazzi, Lorenzo Favalli, and Anna Vizziello. Correlation between weather and signal strength in LoRaWAN networks: an extensive dataset. Computer Networks (Amsterdam, Netherlands: 1999), 202 (??):??, January 15, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100517X.

Perna:2022:ROS

[926] Gianluca Perna, Dena Markudova, Martino Trevisan, Paolo Garza, Michela Meo, and Maurizio M. Munafò. Retina: an open-source tool for flexible analysis of RTC traffic. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005235.

Min:2022:BNM

[927] Xinping Min, Lanju Kong, Qingzhong Li, Yuan Liu, Baochen Zhang, Yongguang Zhao, Zongshui Xiao, and Bin Guo. Blockchain-native mechanism supporting the circulation of complex physical assets. Computer Networks (Amsterdam, Netherlands: 1999), 202 (??):??, January 15, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004928.

Banerjee:2022:EED

[928] Anuradha Banerjee, Abu Sufian, Krishna Keshob Paul, and Sachin Kumar Gupta. EDTP: Energy and delay optimized trajectory planning for UAV-IoT environment. Computer Networks (Amsterdam, Netherlands:

1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005156.

Lin:2022:MML

[929] Kunda Lin, Xiaolong Xu, and Fu Xiao. MFFusion: a multi-level features fusion model for malicious traffic detection based on deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005399.

Anonymous:2022:Fa

[930] Anonymous. 11 February 2022. Computer Networks (Amsterdam, Netherlands: 1999), 203(??):??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBb

[931] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 203(??):??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128622000123.

Tang:2022:ILP

[932] Rui Tang, Xingshu Chen, Chuancheng Wei, Qindong Li, Wenxian Wang, Haizhou Wang, and Wei Wang. Interlayer link prediction based on multiple network structural attributes. Computer Networks (Amsterdam, Netherlands: 1999), 203(??):??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912862100534X.

Kathiravelu:2022:TIS

[933] Pradeeban Kathiravelu, Zachary Zaiman, Judy Gichoya, Luís Veiga, and Imon Banerjee. Towards an Internetscale overlay network for latencyaware decentralized workflows at the edge. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005363.

Momeni:2022:EEN

[934] M. Momeni and H. S. Shahhoseini. Energy efficient 3D network-on-chip based on approximate communication. Computer Networks (Amsterdam, Netherlands: 1999), 203(??): ??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005351.

Febro:2022:ESS

[935] Aldo Febro, Hannan Xiao, Joseph Spring, and Bruce Christianson. Edge security for SIP-enabled IoT devices with P4. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005612.

Xia:2022:PIE

[936] Qiufen Xia, Lizhen Zhou, Wenhao Ren, and Yi Wang. Proactive and intelligent evaluation of big data queries in edge clouds with materialized views. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005429.

Gundogan:2022:MCP

[937] Cenk Gündogan, Peter Kietzmann, Thomas C. Schmidt, and Matthias Wählisch. A mobility-compliant publish-subscribe system for an informationcentric Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 203(??):??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005375.

Luo:2022:FIA

[938] Wanming Luo, Shihao Liu, Yihao Jia, Zhe Chen, and Sheng Jiang. Flexible IP: an adaptable IP address structure and its efficient addressing scheme. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005636.

Khatiri:2022:BRA

[939] Atefeh Khatiri, Ghasem Mirjalily, and Zhi-Quan Luo. Balanced resource allocation for VNF service chain provisioning in inter-datacenter elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005727.

Yang:2022:GTA

[940] Tao Yang, Bingnan Hou, Zhiping Cai, Kui Wu, Tongqing Zhou, and Chengyu Wang. 6Graph: a graphtheoretic approach to address pattern mining for Internet-wide IPv6 scanning. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005430.

Campos:2022:EFL

[941] Enrique Mármol Campos, Pablo Fernández Saura, Aurora González-Vidal, José L. Hernández-Ramos, Jorge Bernal Bernabé, Gianmarco Baldini, and Antonio Skarmeta. Evaluating federated learning for intrusion detection in Internet of Things: Review and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005405.

Carpio:2022:SMR

[942] Francisco Carpio, Wolfgang Bziuk, and Admela Jukan. Scaling migrations and replications of Virtual Network Functions based on network traffic forecasting. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004898.

Kumar:2022:HSH

[943] Alok Kumar, Prabhat Thakur, Shweta Pandit, and Ghanshyam Singh. HSA-SPC: Hybrid spectrum access with spectrum prediction and cooperation for performance enhancement of multiuser cognitive radio network. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004977.

Wang:2022:LLT

[944] Zumin Wang, Jiyu Tian, Hui Fang, Liming Chen, and Jing Qin. Light-Log: a lightweight temporal convolutional network for log anomaly detection on the edge. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005119.

Hbaieb:2022:STM

[945] Amal Hbaieb, Samiha Ayed, and Lamia Chaari. A survey of trust management in the Internet of vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004758.

Bodet:2022:DST

[946] Duschia Bodet, Jacob Hall, Priyangshu Sen, Rachel Johnson, Isabelle Brandicourt, Xavier Cantos Roman, Omar Shoura, and Josep Miquel Jornet. Data signals for terahertz communications research. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005181.

Bonati:2022:EAE

[947] Leonardo Bonati, Stefano Basagni, and Tommaso Melodia. Editorial: Advances in experimental wireless platforms and systems. Computer Networks (Amsterdam, Netherlands: 1999), 203(??):??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005624.

Wang:2022:SRS

[948] Suixing Wang, Chao Yang, Gang Guo, Mingzhe Chen, and Jianfeng Ma. SSAPPIDENTIFY: a robust system identifies application over Shadowsocks’s traffic. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005387.

Baker:2022:BBF

[949] Thar Baker, Muhammad Asim, Hezekiah Samwini, Nauman Shamim, Mohammed M. Alani, and Rajkumar Buyya. A blockchain-based fogoriented lightweight framework for smart public vehicular transportation systems. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100548X.

Zhang:2022:OPB

[950] Kaiyuan Zhang, Xiaolin Gui, Dewang Ren, Tianjiao Du, and Xin He. Optimal pricing-based computation offloading and resource allocation for blockchain-enabled beyond 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005478.

Hireche:2022:DDP

[951] Othmane Hireche, Chafika Benzaïd, and Tarik Taleb. Deep data plane programming and AI for zero-trust self-driven networking in beyond 5G. Computer Networks (Amsterdam, Netherlands: 1999), 203(??): ??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005442.

Zhang:2022:BBP

[952] Guipeng Zhang, Zhenguo Yang, and Wenyin Liu. Blockchain-based privacy preserving e-health system for healthcare data in cloud. Computer Networks (Amsterdam, Netherlands: 1999), 203(??):??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004916.

Anonymous:2022:Fb

[953] Anonymous. 26 February 2022. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBc

[954] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128622000366.

Aski:2022:ANE

[955] Vidyadhar Jinnappa Aski, Vijaypal Singh Dhaka, Sunil Kumar, Sahil Verma, and Danda B. Rawat. Advances on networked ehealth information access and sharing: Status, challenges and prospects. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005545.

Griner:2022:CLP

[956] Chen Griner, Stefan Schmid, and Chen Avin. CacheNet: Leveraging the principle of locality in reconfigurable network design. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005326.

Hameed:2022:TFM

[957] Khizar Hameed, Saurabh Garg, Muhammad Bilal Amin, and Byeong Kang. Towards a formal modelling, analysis and verification of a clone node attack detection scheme in the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005648.

Guo:2022:EHC

[958] Mian Guo, Qirui Li, Zhiping Peng, Xiushan Liu, and Delong Cui. Energy harvesting computation offloading game towards minimizing delay for mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005491.

Zhou:2022:TSD

[959] Anan Zhou, Benshun Yi, and Laigan Luo. Tree-structured data placement scheme with cluster-aided topdown transmission in erasure-coded distributed storage systems. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005703.

Zheng:2022:EPM

[960] Kechen Zheng, Haijiang Ge, Kaikai Chi, and Xiaoying Liu. Energy provision minimization of energy-harvesting cognitive radio networks with minimal throughput demands. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005740.

Dien:2022:TSN

[961] Nguyen V. Dien, Nguyen V. Tuan, Le T. P. Mai, Nguyen V. Hieu, Vuong Q. Phuoc, Nguyen Q. N. Quynh, and Nguyen T. Hung. Tolerance of SCM Nyquist and OFDM signals for heterogeneous fiber-optic and millimeter-wave mobile backhaul links under the effect of power amplifier saturation induced clipping. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005600.

Wu:2022:OMT

[962] Zheng Wu, Yu ning Dong, Xiaohui Qiu, and Jiong Jin. Online multimedia traffic classification from the QoS perspective using deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005715.

Mei:2022:RMB

[963] Lifan Mei, Jinrui Gou, Yujin Cai, Houwei Cao, and Yong Liu. Realtime mobile bandwidth and handoff predictions in 4G/5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005879.

Sengupta:2022:VPP

[964] Binanda Sengupta. VALNET: Privacy-preserving multi-path validation. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005594.

Aouini:2022:NFN

[965] Zied Aouini and Adrian Pekar. NF-Stream: a flexible network data analysis framework. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005739.

Wan:2022:PPB

[966] Yichen Wan, Youyang Qu, Longxiang Gao, and Yong Xiang. Privacypreserving blockchain-enabled federated learning for B5G-driven edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005454.

Abubaker:2022:BSP

[967] Zain Abubaker, Nadeem Javaid, Ahmad Almogren, Mariam Akbar, Mansour Zuair, and Jalel Ben-Othman. Blockchained service provisioning and malicious node detection via federated learning in scalable Internet of Sensor Things networks. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005570.

Rey:2022:FLM

[968] Valerian Rey, Pedro Miguel Sánchez Sánchez, Alberto Huertas Celdrán, and Gérôme Bovet. Federated learning for malware detection in IoT devices. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005582.

Salman:2022:TER

[969] Ola Salman, Imad H. Elhajj, Ali Chehab, and Ayman Kayssi. Towards efficient real-time traffic classifier: a confidence measure with ensemble deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005521.

Cheema:2022:BBS

[970] Muhammad Asaad Cheema, Rafay Iqbal Ansari, Nouman Ashraf, Syed Ali Hassan, Hassaan Khaliq Qureshi, Ali Kashif Bashir, and Christos Politis. Blockchain-based secure delivery of medical supplies using drones. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005661.

Saltini:2022:BRO

[971] Roberto Saltini. BigFooT: a robust optimal-latency BFT blockchain consensus protocol with dynamic validator membership. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100520X.

Wang:2022:OCO

[972] Jian Wang, Hongchang Ke, Xuejie Liu, and Hui Wang. Optimization for computational offloading in multiaccess edge computing: a deep reinforcement learning scheme. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005569.

Anonymous:2022:Ma

[973] Anonymous. 14 March 2022. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBd

[974] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128622000573.

Lin:2022:NLB

[975] Jie Lin, Lin Huang, Hanlin Zhang, Xinyu Yang, and Peng Zhao. A novel Lyapunov based dynamic resource allocation for UAVs-assisted edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005685.

Kuk:2022:APA

[976] Ekber Çetin Kük and Müge Erel-Özçevik. Access protocol aware controller design for eMBB traffic in SD-CDN. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005533.

Li:2022:CAS

[977] Wenhao Li, Hamid Faragardi, Mustafa Ozger, Cicek Cavdar, and Björn Skubic. Cost aware service selection in a mobile edge marketplace. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005508.

Wang:2022:COR

[978] Zhongyu Wang, Tiejun Lv, and Zheng Chang. Computation offloading and resource allocation based on distributed deep learning and software defined mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005843.

Zahedi:2022:PEP

[979] Seyed Reza Zahedi, Shahram Jamali, and Peyman Bayat. A power-efficient and performance-aware online virtual network function placement in SDN/ NFV-enabled networks. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005971.

Liu:2022:MMB

[980] Liang Liu, Lin Liu, Beibei Li, Yi Zhong, Shan Liao, and Lei Zhang. MSCCS: a monero-based securityenhanced covert communication system. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621006010.

Zhong:2022:PBD

[981] Hong Zhong, Jinshan Xu, Jie Cui, Xiuwen Sun, Chengjie Gu, and Lu Liu. Prediction-based dual-weight switch migration scheme for SDN load balancing. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005946.

Wang:2022:BDM

[982] Yang Wang, Marco Giordani, Xiangming Wen, and Michele Zorzi. On the beamforming design of millimeter wave UAV networks: Power vs. capacity trade-offs. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005922.

Nasir:2022:SII

[983] Muhammad Hassan Nasir, Salman A. Khan, Muhammad Mubashir Khan, and Mahawish Fatima. Swarm intelligence inspired intrusion detection systems — a systematic literature review. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005673.

Cao:2022:APM

[984] Tuo Cao, Yibo Jin, Xiongfeng Hu, Sheng Zhang, Zhuzhong Qian, Baoliu Ye, and Sanglu Lu. Adaptive provisioning for mobile cloud gaming at edges. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100565X.

Cortesi:2022:NAB

[985] Eugenio Cortesi, Francesco Bruschi, Stefano Secci, and Sami Taktak. A new approach for Bitcoin pool-hopping detection. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621006009.

Liu:2022:CNT

[986] Meili Liu, Xiaogang Qi, and Hao Pan. Construction of network topology and geographical vulnerability for telecommunication network. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000032.

Floris:2022:SIB

[987] Alessandro Floris, Simone Porcu, Luigi Atzori, and Roberto Girau. A social IoT-based platform for the deployment of a smart parking solution. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005995.

Liu:2022:EOD

[988] Linfeng Liu, Zhipeng Zhang, Jiagao Wu, and Jia Xu. Entropy optimization of degree distributions against security threats in UASNs. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005934.

deCarvalho:2022:JJT

[989] Márcio Barbosa de Carvalho and Lisandro Zambenedetti Granville. JurisNN: Judging traffic differentiations as network neutrality violations according to the regulation. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005880.

Santos:2022:TOO

[990] Ricardo Santos, Nina Skorin-Kapov, Hakim Ghazzai, Andreas Kassler, and Gia Khanh Tran. Towards the optimal orchestration of steerable mmWave backhaul reconfiguration. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005958.

Aumayr:2022:SBA

[991] Erik Aumayr, Giuseppe Caso, Anne-Marie Bosneag, Almudena Diaz Zayas, Özgü Alay, Bruno Garcia, Konstantinos Kousias, Anna Brünstrom, Pedro Merino Gomez, and Harilaos Koumaras. Service-based analytics for 5G open experimentation platforms. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005892.

Casar:2022:SBL

[992] Matthias Cäsar, Tobias Pawelke, Jan Steffan, and Gabriel Terhorst. A survey on Bluetooth low energy security and privacy. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005697.

Fronteddu:2022:DHG

[993] Graziano Fronteddu, Simone Porcu, Alessandro Floris, and Luigi Atzori. A dynamic hand gesture recognition dataset for human-computer interfaces. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000172.

Pintor:2022:DLD

[994] Lucia Pintor and Luigi Atzori. A dataset of labelled device Wi-Fi probe requests for MAC address derandomization. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000196.

Barcelo-Armada:2022:AAT

[995] Rubén Barceló-Armada, Ismael Castell-Uroz, and Pere Barlet-Ros. Amazon Alexa traffic traces. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000184.

Bagies:2022:CDN

[996] Enas Bagies, Ahmed Barnawi, Saoucene Mahfoudh, and Neeraj Kumar. Content delivery network for IoT-based fog computing environment. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005557.

Salehi:2022:NMI

[997] Majid Salehi, Gilles De Borger, Danny Hughes, and Bruno Crispo. Nemesis-Guard: Mitigating interrupt latency side channel attacks with static binary rewriting. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005910.

Nasser:2022:LFL

[998] Nidal Nasser, Zubair Md Fadlullah, Mostafa M. Fouda, Asmaa Ali, and Muhammad Imran. A lightweight federated learning based privacy preserving B5G pandemic response network using unmanned aerial vehicles: a proof-of-concept. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005466.

Wang:2022:NTA

[999] Chonghua Wang, Hao Zhou, Zhiqiang Hao, Shu Hu, Jun Li, Xueying Zhang, Bo Jiang, and Xuehong Chen. Network traffic analysis over clustering-based collective anomaly detection. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000019.

Mukta:2022:SDM

[1000] Rahma Mukta, Hye young Paik, Qinghua Lu, and Salil S. Kanhere. A survey of data minimisation techniques in blockchain-based healthcare. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000044.

Anonymous:2022:Aa

[1001] Anonymous. 7 April 2022. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBe

[1002] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200086X.

Yang:2022:PSP

[1003] Yan Yang, Xingang Shi, Qiang Ma, Yahui Li, Xia Yin, and Zhiliang Wang. Path stability in partially deployed secure BGP routing. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000020.

Silva:2022:TLD

[1004] Giovanni Maciel Ferreira Silva and Taufik Abrão. Throughput and latency in the distributed Q-learning random access mMTC networks. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000214.

Dahanayaka:2022:DTF

[1005] Thilini Dahanayaka, Guillaume Jourjon, and Suranga Seneviratne. Dissecting traffic fingerprinting CNNs with filter activations. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000068.

Yang:2022:BTB

[1006] Furong Yang, Qinghua Wu, Zhenyu Li, Yanmei Liu, Giovanni Pau, and Gaogang Xie. BBRv2+: Towards balancing aggressiveness and fairness with delay-based bandwidth probing. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000226.

Sabur:2022:TSG

[1007] Abdulhakim Sabur, Ankur Chowdhary, Dijiang Huang, and Adel Alshamrani. Toward scalable graph-based security analysis for cloud networks. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000251.

Sciancalepore:2022:PPP

[1008] Savio Sciancalepore. PARFAIT: Privacypreserving, secure, and low-delay service access in fog-enabled IoT ecosystems. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000275.

Madarasingha:2022:VGB

[1009] Chamara Madarasingha, Shashika R. Muramudalige, Guillaume Jourjon, Anura Jayasumana, and Kanchana Thilakarathna. VideoTrain++: GANbased adaptive framework for synthetic video traffic generation. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000202.

deSousa:2022:PLS

[1010] Roniel S. de Sousa, Azzedine Boukerche, and Antonio A. F. Loureiro. On the prediction of large-scale roadnetwork constrained trajectories. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003352.

Raskar:2022:MEM

[1011] Charushila Raskar and Shikha Nema.

Metaheuristic enabled modified hidden Markov model for traffic flow prediction. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000160.

Choi:2022:UIT

[1012] Jinchun Choi, Afsah Anwar, Abdulrahman Alabduljabbar, Hisham Alasmary, Jeffrey Spaulding, An Wang, Songqing Chen, DaeHun Nyang, Amro Awad, and David Mohaisen. Understanding Internet of Things malware by analyzing endpoints in their static artifacts. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000056.

Hosseini:2022:OTS

[1013] Entesar Hosseini, Mohsen Nickray, and Shamsollah Ghanbari. Optimized task scheduling for cost-latency trade-off in mobile fog computing using fuzzy analytical hierarchy process. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100596X.

Zhang:2022:SBP

[1014] Lihao Zhang, Taotao Wang, and Soung Chang Liew. Speeding up block propagation in Bitcoin network: Uncoded and coded designs. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000238.

Xiong:2022:BBP

[1015] Ting Xiong, Ran Zhang, Jiang Liu, Tao Huang, Yunjie Liu, and F. Richard Yu. A blockchain-based and privacy-preserved authentication scheme for inter-constellation collaboration in Space-Ground Integrated Networks. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200024X.

Brehon-Grataloup:2022:MEC

[1016] Lucas Bréhon-Grataloup, Rahim Kacimi, and André-Luc Beylot. Mobile edge computing for V2X architectures and applications: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000263.

El-Zawawy:2022:SSB

[1017] Mohamed A. El-Zawawy, Alessandro Brighente, and Mauro Conti. SETCAP: Service-based energy-efficient temporal credential authentication protocol for Internet of Drones. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000305.

Vasudevan:2022:IAE

[1018] Vipindev Adat Vasudevan, Muhammad Tayyab, George P. Koudouridis, Xavier Gelabert, and Ilias Politis. An integrated approach for energy efficient handover and key distribution protocol for secure NC-enabled small cells. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000317.

Kim:2022:CFO

[1019] Hak-Jin Kim, Marie S. Kim, and Seung-Jae Han. Collision-free optimal packet scheduling algorithm for multi-hop wireless IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000408.

Magdy:2022:ESP

[1020] Safaa Magdy, Yasmine Abouelseoud, and Mervat Mikhail. Efficient spam and phishing emails filtering based on deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000469.

Ahanger:2022:SAS

[1021] Tariq Ahamed Ahanger, Abdullah Aljumah, and Mohammed Atiquzzaman. State-of-the-art survey of artificial intelligent techniques for IoT security. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200007X.

Sicari:2022:SIC

[1022] Sabrina Sicari, Alessandra Rizzardi, and Alberto Coen-Porisini. Security & privacy issues and challenges in NoSQL databases. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000470.

Deb:2022:CSV

[1023] Raktim Deb and Sudipta Roy. A comprehensive survey of vulnerability and information security in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000299.

Wang:2022:SPL

[1024] Xiaoyang Wang, Jonathan D. Thomas, Robert J. Piechocki, Shipra Kapoor, Raúl Santos-Rodríguez, and Arjun Parekh. Self-play learning strategies for resource assignment in Open-RAN networks. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100551X.

Anonymous:2022:Ab

[1025] Anonymous. 22 April 2022. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBf

[1026] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128622001050.

Zhang:2022:CSN

[1027] Ke Zhang, Guang Zhang, Xiuwu Yu, Shaohua Hu, and Moxiao Li. Clustering the sensor networks based on energy-aware affinity propagation. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000652.

Makarem:2022:DEC

[1028] Nabil Makarem, Wafaa Bou Diab, Imad Mougharbel, and Naceur Malouch. On the design of efficient congestion control for the Constrained Application Protocol in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000457.

Nduwayezu:2022:LEA

[1029] Maurice Nduwayezu and Ji-Hoon Yun. Latency and energy aware rate maximization in MC-NOMA-based multi-access edge computing: a two-stage deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000500.

Guo:2022:SSR

[1030] Yimin Guo, Zhenfeng Zhang, and Yajun Guo. SecFHome: Secure remote authentication in fog-enabled smart home environment. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200041X.

Huang:2022:PLF

[1031] Cheng Huang, Aimin Tang, Bangzhao Zhai, and Xudong Wang. Physical layer forwarding for 5G multihop Backhaul networks. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000482.

Ribeiro:2022:SHT

[1032] Nildo dos Santos Ribeiro, Marcos A. M.

Vieira, Luiz F. M. Vieira, and Omprakash Gnawali. SplitPath: High throughput using multipath routing in dual-radio Wireless Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000494.

Nassef:2022:SDM

[1033] Omar Nassef, Wenting Sun, Hakimeh Purmehdi, Mallik Tatipamula, and Toktam Mahmoodi. A survey: Distributed Machine Learning for 5G and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000421.

Alghafari:2022:DJR

[1034] Hadeel Alghafari and Mohammad Sayad Haghighi. Decentralized joint resource allocation and path selection in multi-hop integrated access backhaul 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000524.

Tang:2022:UAM

[1035] Qiang Tang, Lixin Liu, Caiyan Jin, Jin Wang, Zhuofan Liao, and Yuansheng Luo. An UAV-assisted mobile edge computing offloading strategy for minimizing energy consumption. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000688.

Pei:2022:UAC

[1036] Jiahui Pei, Hongbin Chen, and Lei Shu. UAV-assisted connectivity enhancement algorithms for multiple isolated sensor networks in agricultural Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000664.

Bouzidi:2022:DCS

[1037] EL Hocine Bouzidi, Abdelkader Outtagarts, Rami Langar, and Raouf Boutaba. Dynamic clustering of software defined network switches and controller placement using deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000640.

Wang:2022:JFS

[1038] Zuyan Wang, Jun Tao, Yang Gao, Yifan Xu, Weice Sun, Yu Gao, and Wenqiang Li. Joint flight scheduling and task allocation for secure data collection in UAV-aided IoTs. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000627.

Mondal:2022:ENC

[1039] Sourav Mondal, Goutam Das, and Elaine Wong. An economic and noncooperative load-balancing framework among federated cloudlets. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000615.

AlSabeh:2022:SSA

[1040] Ali AlSabeh, Joseph Khoury, Elie Kfoury, Jorge Crichigno, and Elias Bou-Harb. A survey on security applications of P4 programmable switches and a STRIDE-based vulnerability assessment. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000287.

Shahraki:2022:CSO

[1041] Amin Shahraki, Mahmoud Abbasi, Amir Taherkordi, and Anca Delia Jurcut. A comparative study on online machine learning techniques for network traffic streams analysis. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000512.

Conti:2022:SCA

[1042] Mauro Conti, Eleonora Losiouk, Radha Poovendran, and Riccardo Spolaor. Side-channel attacks on mobile and IoT devices for cyber-physical systems. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200069X.

Aouedi:2022:HPL

[1043] Ons Aouedi, Kandaraj Piamrat, and Dhruvjyoti Bagadthey. Handling partially labeled network data: a semi-supervised approach using stacked sparse autoencoder. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005909.

Hyland-Wood:2022:GEB

[1044] David Hyland-Wood and Sandra Johnson. Guest editorial: Blockchain consensus protocols. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000718.

Anonymous:2022:Mb

[1045] Anonymous. 8 May 2022. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBg

[1046] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001281.

Enoch:2022:ISH

[1047] Simon Yusuf Enoch, Júlio Mendonça, Jin B. Hong, Mengmeng Ge, and Dong Seong Kim. An integrated security hardening optimization for dynamic networks using security and availability modeling with multi-objective algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000731.

Nguyen:2022:UCP

[1048] Thi Thuy Nga Nguyen, Olivier Brun, and Balakrishna J. Prabhu. Using channel predictions for improved proportional-fair utility for vehicular users. Computer Networks (Amsterdam, Netherlands: 1999), 208 (??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000780.

Islam:2022:ITE

[1049] Nazrul Islam and Chris Phillips. Intelligent traffic engineering (TE) system for rural broadband. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000901.

Qu:2022:PAV

[1050] Hua Qu, Ke Wang, and Jihong Zhao. Priority-awareness VNF migration method based on deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 208 (??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000755.

Bhattacharjee:2022:CRB

[1051] Sangeeta Bhattacharjee, Tamaghna Acharya, and Uma Bhattacharya. Cognitive radio based spectrum sharing models for multicasting in 5G cellular networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000779.

Li:2022:PSR

[1052] Ziyong Li, Yuxiang Hu, Jiangxing Wu, and Jie Lu. P4Resilience: Scalable resilience for multi-failure recovery in SDN with programmable data plane. Computer Networks (Amsterdam, Netherlands: 1999), 208 (??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000950.

Enoch:2022:PFC

[1053] Simon Yusuf Enoch, Chun Yong Moon, Donghwan Lee, Myung Kil Ahn, and Dong Seong Kim. A practical framework for cyber defense generation, enforcement and evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000810.

Yin:2022:JAD

[1054] Wei Yin, Peizhao Hu, Hongjian Zhou, Guoqiang Xing, and Jiahui Wen. Jamming attacks and defenses for fast association in IEEE 802.11ah networks. Computer Networks (Amsterdam, Netherlands: 1999), 208 (??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000913.

Bian:2022:SLD

[1055] Rui Bian, Shuai Hao, Haining Wang, and Chase Cotton. Shining a light on dark places: a comprehensive analysis of open proxy ecosystem. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000937.

Gharbaoui:2022:ESL

[1056] M. Gharbaoui, C. Contoli, G. Davoli, D. Borsatti, G. Cuffaro, F. Paganelli, W. Cerroni, P. Cappanera, and B. Martini. An experimental study on latencyaware and self-adaptive service chaining orchestration in distributed NFV and SDN infrastructures. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000822.

Mahmoudi:2022:QAR

[1057] Arezou Mahmoudi, Akbar Ghaffarpour Rahbar, and Mohammad Jafari-Beyrami. QoS-aware routing, space, and spectrum assignment in space division multiplexing networks. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001116.

Campos:2022:MLH

[1058] Pablo Campos, Ángela Hernández-Solana, and Antonio Valdovinos-Bardají. Machine learning for hidden nodes detection in unlicensed LTE networks. Computer Networks (Amsterdam, Netherlands: 1999), 208 (??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200072X.

Anonymous:2022:Mc

[1059] Anonymous. 22 May 2022. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBh

[1060] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200144X.

Myneni:2022:SDD

[1061] Sowmya Myneni, Ankur Chowdhary, Dijiang Huang, and Adel Alshamrani. SmartDefense: a distributed deep defense against DDoS attacks with edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000792.

Somesula:2022:CCU

[1062] Manoj Kumar Somesula, Rashmi Ranjan Rout, and D. V. L. N. Somayajulu. Cooperative cache update using multi-agent recurrent deep reinforcement learning for mobile edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 209 (??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000809.

Zhan:2022:DDH

[1063] Mengqi Zhan, Yang Li, Guangxi Yu, Bo Li, and Weiping Wang. Detecting DNS over HTTPS based data exfiltration. Computer Networks (Amsterdam, Netherlands: 1999), 209 (??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001104.

Nacef:2022:MLB

[1064] Abdelhakim Nacef, Abdellah Kaci, Youcef Aklouf, and Diego Leonel Cadette Dutra. Machine learning based fast self optimized and life cycle management network. Computer Networks (Amsterdam, Netherlands: 1999), 209 (??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000949.

Kita:2022:PRL

[1065] Kentaro Kita, Yuki Koizumi, and Toru Hasegawa. Private retrieval of locationrelated content using k-anonymity and application to ICN. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001013.

Jagannath:2022:DFC

[1066] Jithin Jagannath, Anu Jagannath, Justin Henney, Tyler Gwin, Zackary Kane, Noor Biswas, and Andrew Drozd. Design of fieldable crosslayer optimized network using embedded software defined radios: Survey and novel architecture with field trials. Computer Networks (Amsterdam, Netherlands: 1999), 209 (??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001098.

Katila:2022:AMS

[1067] Risto Katila, Tuan Nguyen Gia, and Tomi Westerlund. Analysis of mobility support approaches for edgebased IoT systems using high data rate Bluetooth Low Energy 5. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001141.

Chen:2022:EDH

[1068] Tianfei Chen, Shuaixin Hou, and Lijun Sun. An enhanced DV-hop positioning scheme based on spring model and reliable beacon node set. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001153.

Tang:2022:CPR

[1069] Kun Tang, Feiyu Jiao, Xuxun Liu, Wenquan Che, and Quan Xue. Coverage probability of relay-assisted NOMA millimeter wave networks with steerable-beam. Computer Networks (Amsterdam, Netherlands: 1999), 209 (??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001177.

Abbas:2022:JCC

[1070] Nadine Abbas, Sanaa Sharafeddine, Azzam Mourad, Chadi Abou-Rjeily, and Wissam Fawaz. Joint computing, communication and cost-aware task offloading in D2D-enabled Het-MEC. Computer Networks (Amsterdam, Netherlands: 1999), 209 (??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000974.

Serper:2022:CCB

[1071] Elif Zeynep Serper and Aysegül Altin-Kayhan. Coverage and connectivity based lifetime maximization with topology update for WSN in smart grid applications. Computer Networks (Amsterdam, Netherlands: 1999), 209 (??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001232.

Sassi:2022:CSA

[1072] Mohamed Saifeddine Hadj Sassi and Lamia Chaari Fourati. Comprehensive survey on air quality monitoring systems based on emerging computing and communication technologies. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000998.

Kaur:2022:FBJ

[1073] Jasneet Kaur and M. Arif Khan. Federation based joint client and server side machine learning for 5G and beyond Wireless Channel Estimation. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000986.

Haris:2022:LMV

[1074] Raseena M. Haris, Khaled M. Khan, and Armstrong Nhlabatsi. Live migration of virtual machine memory content in networked systems. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000962.

Mangas:2022:WEF

[1075] Andrés García Mangas, Francisco José Suárez Alonso, Daniel Fernando García Martínez, and Fidel Díez Díaz. WoTemu: an emulation framework for edge computing architectures based on the Web of Things. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000767.

Pei:2022:PFL

[1076] Jiaming Pei, Kaiyang Zhong, Mian Ahmad Jan, and Jinhai Li. Personalized federated learning framework for network traffic anomaly detection. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001001. See

retraction notice [1766].

Anonymous:2022:Jb

[1077] Anonymous. 19 June 2022. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBi

[1078] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128622001797.

Borromeo:2022:FAS

[1079] Justine Cris Borromeo, Koteswararao Kondepu, Nicola Andriolli, and Luca Valcarenghi. FPGA-accelerated Smart-NIC for supporting 5G virtualized radio access network. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001189.

Beraldi:2022:ISI

[1080] Roberto Beraldi, Claudia Canali, Riccardo Lancellotti, and Gabriele Proietti Mattia. On the impact of stale information on distributed online load balancing protocols for edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 210 (??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001207.

Wang:2022:TEE

[1081] Yang Wang, Yutong Li, Ting Wang, and Gang Liu. Towards an energyefficient data center network based on deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001220.

Luomala:2022:ARB

[1082] Jari Luomala and Ismo Hakala. Adaptive range-based localization algorithm based on trilateration and reference node selection for outdoor wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000743.

Ogundoyin:2022:SPP

[1083] Sunday Oyinlola Ogundoyin and Ismaila Adeniyi Kamil. Secure and privacy-preserving D2D communication in fog computing services. Computer Networks (Amsterdam, Netherlands: 1999), 210 (??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001244.

Baldini:2022:ODD

[1084] Gianmarco Baldini and Irene Amerini. Online Distributed Denial of Service (DDoS) intrusion detection based on adaptive sliding window and morphological fractal dimension. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200113X.

Charr:2022:PCO

[1085] Jean-Claude Charr, Karine Deschinkel, Rania Haj Mansour, and Mourad Hakem. Partial coverage optimization under network connectivity constraints in heterogeneous sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 210 (??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001165.

Gupta:2022:DAI

[1086] Ankit Kumar Gupta and T. G.

Venkatesh. Design and analysis of IEEE 802.11 based full duplex WLAN MAC protocol. Computer Networks (Amsterdam, Netherlands: 1999), 210 (??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001190.

Li:2022:TOS

[1087] Zhehao Li, Lei Shi, Yi Shi, Zhenchun Wei, and Yang Lu. Task offloading strategy to maximize task completion rate in heterogeneous edge computing environment. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001219.

Yang:2022:DPP

[1088] Hailiang Yang, Rukhsana Ruby, and Kaishun Wu. Delay performance of priority-queue equipped UAV-based mobile relay networks: Exploring the impact of trajectories. Computer Networks (Amsterdam, Netherlands:

1999), 210(??):??, June 19, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000676.

Sellami:2022:EAT

[1089] Bassem Sellami, Akram Hakiri, Sadok Ben Yahia, and Pascal Berthou. Energy-aware task scheduling and offloading using deep reinforcement learning in SDN-enabled IoT network. Computer Networks (Amsterdam, Netherlands: 1999), 210 (??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001359.

Ioannou:2022:DAM

[1090] Iacovos Ioannou, Christophoros Christophorou, Vasos Vassiliou, and Andreas Pitsillides. A distributed AI/ML framework for D2D transmission mode selection in 5G and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001396.

Sabbioni:2022:DDD

[1091] Andrea Sabbioni, Lorenzo Rosa, Armir Bujari, Luca Foschini, and Antonio Corradi. DIFFUSE: a DIstributed and decentralized platForm enabling Function composition in Serverless Environments. Computer Networks (Amsterdam, Netherlands: 1999), 210 (??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200161X.

Ghosh:2022:BED

[1092] Timam Ghosh, Arijit Roy, and Sudip Misra. B2H: Enabling delay-tolerant blockchain network in healthcare for Society 5.0. Computer Networks (Amsterdam, Netherlands: 1999), 210 (??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000706.

Anonymous:2022:Jc

[1093] Anonymous. 5 July 2022. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBj

[1094] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002018.

Mairaj:2022:GTS

[1095] Aakif Mairaj and Ahmad Y. Javaid.

Game theoretic solution for an unmanned aerial vehicle network host under DDoS attack. Computer Networks (Amsterdam, Netherlands: 1999), 211 (??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001384.

Ali:2022:SMS

[1096] Shabir Ali, Mayank Pandey, and Neeraj Tyagi. SDFog-Mesh: a software-defined fog computing architecture over wireless mesh networks for semi-permanent smart environments. Computer Networks (Amsterdam, Netherlands: 1999), 211 (??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001578.

Ioannou:2022:NDA

[1097] Iacovos Ioannou, Christophoros Christophorou, Vasos Vassiliou, and Andreas Pitsillides. A novel distributed AI framework with ML for D2D communication in 5G/6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200158X.

Gomides:2022:PCC

[1098] Thiago S. Gomides, Robson E. De Grande, Rodolfo I. Meneguette, Fernanda S. H. de Souza, and Daniel L. Guidoni. Predictive congestion control based on collaborative information sharing for vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 211 (??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001347.

Bagnulo:2022:PRL

[1099] Marcelo Bagnulo, Alberto García-Martínez, Stefano Angieri, Andra Lutu, and Jinze Yang. Practicable route leak detection and protection with ASIRIA. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001402.

Ji:2022:JOR

[1100] Pengshuo Ji, Jie Jia, Jian Chen, Yunhe Xie, and Xingwei Wang. Joint optimization for RIS-assisted multicast D2D communications. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001499.

Niu:2022:BSS

[1101] Shufen Niu, Siwei Zhou, Lizhi Fang, Ying Hu, and Caifen Wang. Broadcast signcryption scheme based on certificateless in wireless sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 211 (??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001621.

Hasan:2022:BBS

[1102] Khalid Hasan, Mohammad Jabed Morshed Chowdhury, Kamanashis Biswas, Khandakar Ahmed, Md. Saiful Islam, and Muhammad Usman. A blockchain-based secure data-sharing framework for Software Defined Wireless Body Area Networks. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001670.

Ojha:2022:BWC

[1103] Tamoghna Ojha, Theofanis P. Raptis, Marco Conti, and Andrea Passarella. Balanced wireless crowd charging with mobility prediction and social awareness. Computer Networks (Amsterdam, Netherlands: 1999), 211 (??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001591.

Sharma:2022:RSP

[1104] Gourav Prateek Sharma, Wouter Tavernier, Didier Colle, and Mario Pickavet. Routing and scheduling for 1 + 1 protected DetNet flows. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001372.

Garcia:2022:ESC

[1105] Rodrigo D. Garcia, Gowri Ramachandran, and Jó Ueyama. Exploiting smart contracts in PBFT-based blockchains: a case study in medical prescription system. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001669.

Lacava:2022:SBL

[1106] Andrea Lacava, Valerio Zottola, Alessio Bonaldo, Francesca Cuomo, and Stefano Basagni. Securing Bluetooth low energy networking: an overview of security procedures and threats. Computer Networks (Amsterdam, Netherlands: 1999), 211 (??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001335.

Uzunidis:2022:BPO

[1107] D. Uzunidis, C. Matrakidis, E. Kosmatos, A. Stavdas, and A. Lord. On the benefits of power optimization in the S, C and L-band optical transmission systems. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001360.

Zhang:2022:EIR

[1108] Songli Zhang, Weijia Jia, Zhiqing Tang, Jiong Lou, and Wei Zhao. Efficient instance reuse approach for service function chain placement in mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001700.

Vegni:2022:GER

[1109] Anna Maria Vegni, Chaker Abdelaziz Kerrache, Waleed Ejaz, Enrico Natalizio, Jiming Chen, and Houbing Song. Guest editorial: Recent advances in connected and autonomous unmanned aerial/ground vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001724.

Li:2022:SRL

[1110] Hui Li, Dongcong Shi, Weizheng Wang, Dan Liao, Thippa Reddy Gadekallu, and Keping Yu. Secure routing for LEO satellite network survivability. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001712.

Anonymous:2022:Jd

[1111] Anonymous. 20 July 2022. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBk

[1112] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200233X.

Chang:2022:SMD

[1113] Jinyong Chang, Qiaochuan Ren, Yanyan Ji, Maozhi Xu, and Rui Xue. Secure medical data management with privacy-preservation and authentication properties in smart healthcare system. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001736.

Qin:2022:JOS

[1114] Yudong Qin, Deke Guo, Lailong Luo, and Ming Xu. A joint orchestration of security and functionality services at network edge. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001323.

Wu:2022:RNC

[1115] Ke Wu, Dezun Dong, Cunlu Li, and Weixia Xu. Revisiting network congestion avoidance through adaptive packet-chaining reservation. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001694.

Li:2022:TCC

[1116] Ying Li, Yi Huang, Suranga Seneviratne, Kanchana Thilakarathna, Adriel Cheng, Guillaume Jourjon, Darren Webb, David B. Smith, and Richard Yi Da Xu. From traffic classes to content: a hierarchical approach for encrypted traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200175X.

Giakoumakis:2022:SMN

[1117] Giannis Giakoumakis, Eva Papadogiannaki, Giorgos Vasiliadis, and Sotiris Ioannidis. Scheduling of multiple network packet processing applications using Pythia. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001682.

Tsipis:2022:JOS

[1118] Athanasios Tsipis and Konstantinos Oikonomou. Joint optimization of social interactivity and server provisioning for interactive games in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001840.

Bagnulo:2022:EEL

[1119] Marcelo Bagnulo and Alberto García-Martínez. An experimental evaluation of LEDBAT++. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001888.

Martini:2022:IBZ

[1120] B. Martini, M. Gharbaoui, and P. Castoldi. Intent-based zero-touch service chaining layer for software-defined edge cloud networks. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001876.

Luttringer:2022:DNO

[1121] Jean-Romain Luttringer, Thomas Alfroy, Pascal Mérindol, Quentin Bramas, François Clad, and Cristel Pelsser. Deploying near-optimal delayconstrained paths with Segment Routing in massive-scale networks. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001748.

Gomez:2022:STE

[1122] Jose Gomez, Elie F. Kfoury, Jorge Crichigno, and Gautam Srivastava. A survey on TCP enhancements using P4-programmable devices. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001852.

Sarieddine:2022:OVB

[1123] Khaled Sarieddine, Hassan Artail, and Haidar Safa. An opportunistic vehiclebased task assignment for IoT offloading. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200189X.

Wu:2022:LAC

[1124] Wenjia Wu, Yujing Liu, Jiazhi Yao, Xiaolin Fang, Feng Shan, Ming Yang, Zhen Ling, and Junzhou Luo. Learning-aided client association control for high-density WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200192X.

Chen:2022:DFS

[1125] Chen Chen, Lars Nagel, Lin Cui, and Fung Po Tso. Distributed federated service chaining: a scalable and cost-aware approach for multidomain networks. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001931.

Macas:2022:SDL

[1126] Mayra Macas, Chunming Wu, and Walter Fuertes. A survey on deep learning for cybersecurity: Progress, challenges, and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001864.

Uyan:2022:RSM

[1127] Osman Gokhan Uyan, Ayhan Akbas, and Vehbi Cagri Gungor. A reliable and secure multi-path routing strategy for underwater acoustic sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002122.

Junfeng:2022:PDS

[1128] Tian Junfeng, Bai Wenqing, and Jia Haoyi. PGCE: a distributed storage causal consistency model based on partial geo-replication and cloudedge collaboration architecture. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002092.

Nakimuli:2022:ERA

[1129] Winnie Nakimuli, Jaime Garcia-Reinoso, J. Enrique Sierra-Garcia, and Pablo Serrano. Exploiting radio access information to improve performance of remote-controlled mobile robots in MEC-based 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002067.

Dao:2022:ABS

[1130] Nhu-Ngoc Dao, Duc-Nghia Vu, Woongsoo Na, Trong-Minh Hoang, Dinh-Thuan Do, and Sungrae Cho. Adaptive bitrate streaming in multi-user downlink NOMA edge caching systems with imperfect SIC. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002080.

Callejo:2022:MDW

[1131] Patricia Callejo, Marcelo Bagnulo, Jaime González Ruiz, Andra Lutu, Alberto García-Martínez, and Rubén Cuevas. Measuring DoH with web ads. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001943.

Li:2022:FHD

[1132] Chunlin Li, Yong Zhang, and Youlong Luo. Flexible heterogeneous data fusion strategy for object positioning applications in edge computing environment. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002201.

Garcia-Teodoro:2022:NZT

[1133] P. García-Teodoro, J. Camacho, G. Maciá-Fernández, J. A. Gómez-Hernández, and V. J. López-Marín. A novel zero-trust network access control scheme based on the security profile of devices and users. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002109.

Asaad:2022:CRI

[1134] Safar M. Asaad and Halgurd S.

Maghdid. A comprehensive review of indoor/outdoor localization solutions in IoT era: Research challenges and future perspectives. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001918.

Boobalan:2022:FFL

[1135] Parimala Boobalan, Swarna Priya Ramu, Quoc-Viet Pham, Kapal Dev, Sharnil Pandya, Praveen Kumar Reddy Maddikunta, Thippa Reddy Gadekallu, and Thien Huynh-The. Fusion of Federated Learning and Industrial Internet of Things: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001955.

Maryam:2022:LQQ

[1136] Hafsa Maryam, Tania Panayiotou, and Georgios Ellinas. Learning quantile QoT models to address uncertainty over unseen lightpaths. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001608.

Mansour:2022:EAF

[1137] Romany F. Mansour, Suliman A. Alsuhibany, Sayed Abdel-Khalek, Randa Alharbi, Thavavel Vaiyapuri, Ahmed J. Obaid, and Deepak Gupta. Energy aware fault tolerant clustering with routing protocol for improved survivability in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001967.

Anonymous:2022:Ac

[1138] Anonymous. 4 August 2022. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBl

[1139] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128622003024.

Vitor:2022:SAS

[1140] Gonçalo Vítor, Pedro Rito, Susana Sargento, and Filipe Pinto. A scalable approach for smart city data platform: Support of real-time processing and data sharing. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001839.

Basturk:2022:EEC

[1141] Ilhan Bastürk. Energy-efficient communication for UAV-enabled mobile relay networks. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002134.

Zaki:2022:GGM

[1142] Faiz Zaki, Firdaus Afifi, Shukor Abd Razak, Abdullah Gani, and Nor Badrul Anuar. GRAIN: Granular multi-label encrypted traffic classification using classifier chain. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002213.

Zahedinia:2022:FBC

[1143] Marzieh Sadat Zahedinia, Mohammad Reza Khayyambashi, and Ali Bohlooli. Fog-based caching mechanism for IoT data in information centric network using prioritization. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002195.

Langer:2022:NCK

[1144] Martin Langer and Rainer Bermbach. NTS4PTP — a comprehensive key management solution for PTP networks. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002158.

Bonafini:2022:EEP

[1145] Stefano Bonafini, Claudio Sacchi, Riccardo Bassoli, Koteswararao Kondepu, Fabrizio Granelli, and Frank H. P. Fitzek. End-to-end performance assessment of a 3D network for 6G connectivity on Mars surface. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002171.

Luo:2022:MEO

[1146] Meng Luo, Yepeng Yao, Liling Xin, Zhengwei Jiang, Qiuyun Wang, and Wenchang Shi. Measurement for encrypted open resolvers: Applications and security. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002183.

Su:2022:DSP

[1147] Lina Su, Ne Wang, Ruiting Zhou, and Zongpeng Li. Dynamic service placement and request scheduling for edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001633.

Mezher:2022:GGT

[1148] Ahmad Mohamad Mezher and Mónica Aguilar Igartua. G-3MRP: a game-theoretical multimedia multimetric map-aware routing protocol for vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002225.

Liang:2022:PDC

[1149] Yihuai Liang, Yan Li, and Byeong-Seok Shin. Private decentralized crowdsensing with asynchronous blockchain access. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002237.

Saleem:2022:PXC

[1150] Tania Saleem, Muhammad Umar Janjua, Muhammad Hassan, Talha Ahmad, Filza Tariq, Khadija Hafeez, Muhammad Ahsan Salal, and Muhammad Danish Bilal. ProofChain: an x.509-compatible blockchain-based PKI framework with decentralized trust. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002110.

Tu:2022:RES

[1151] Wanqing Tu. Resource-efficient seamless transitions for high-performance multi-hop UAV multicasting. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128622001979.

Souza:2022:CDA

[1152] Otavio A. de O. Souza, Olga Goussevskaia, and Stefan Schmid. CB-Net: Demand-aware tree topologies for Reconfigurable Datacenter Networks. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002249.

Cremonezi:2022:IAR

[1153] Bruno Cremonezi, Airton Ribeiro Gomes Filho, Edelberto Franco Silva, José Augusto Miranda Nacif, Alex Borges Vieira, and Michele Nogueira. Improving the attribute retrieval on ABAC using opportunistic caches for Fog-Based IoT Networks. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001657.

Mian:2022:VAI

[1154] Adnan Noor Mian, Syed Waqas Haider Shah, Sanaullah Manzoor, Anwar Said, Kurtis Heimerl, and Jon Crowcroft. A value-added IoT service for cellular networks using federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002262.

Ran:2022:DSA

[1155] Longyan Ran, Yunhe Cui, Chun Guo, Qing Qian, Guowei Shen, and Huanlai Xing. Defending saturation attacks on SDN controller: a confusable instance analysis-based algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002298.

Abbas:2022:PBR

[1156] Ghulam Abbas, Sami Ullah, Muhammad Waqas, Ziaul Haq Abbas, and Muhammad Bilal. A positionbased reliable emergency message routing scheme for road safety in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002286.

Catena:2022:DLB

[1157] Tiziana Catena, Vincenzo Eramo, Massimo Panella, and Antonello Rosato. Distributed LSTM-based cloud resource allocation in Network Function Virtualization Architectures. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002390.

Nisha:2022:NPP

[1158] Nisha Nisha, Iynkaran Natgunanathan, Shang Gao, and Yong Xiang. A novel privacy protection scheme for location-based services using collaborative caching. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002377.

Guo:2022:JWR

[1159] Mian Guo, Mithun Mukherjee, Jaime Lloret, Jiangtao Ou, and Chengyuan Fan. Joint wireless resource allocation and service function chaining scheduling for Tactile Internet. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002407.

Alioua:2022:IMC

[1160] Ahmed Alioua, Roumayssa Hamiroune, Oumayma Amiri, Manel Khelifi, Sidi-Mohammed Senouci, Mikael Gidlund, and Sarder Fakhrul Abedin. Incentive mechanism for competitive edge caching in 5G-enabled Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002274.

Waleed:2022:WOS

[1161] Abdul Waleed, Abdul Fareed Jamali, and Ammar Masood. Which open-source IDS? Snort, Suricata or Zeek. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002420.

Hilal:2022:TAO

[1162] Anwer Mustafa Hilal, Amani Abdulrahman Albraikan, Sami Dhahbi, Saud S. Alotaibi, Rana Alabdan, Mesfer Al Duhayyim, Abdelwahed Motwakel, and Ishfaq Yaseen. Trust aware oppositional sine cosine based multihop routing protocol for improving survivability of wireless sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002444.

Arsalan:2022:IBT

[1163] Ahmad Arsalan and Rana Asif Rehman. Interest Broadcasting and Timing Attack in IoV (IBTA-IoV): a novel architecture using Named Software Defined Network. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002456.

Hernandez-Solana:2022:PDP

[1164] Á. Hernández-Solana, D. Perez-Diaz de Cerio, M. García-Lozano, A. Valdovinos, and J. L. Valenzuela. PSM-DMO: Power save mode and discontinuous BLE mesh operation. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002419.

Singh:2022:EEF

[1165] Namrata Singh and Ayan Kumar Das.

Energy-efficient fuzzy data offloading for IoMT. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002493.

Bringhenti:2022:AVO

[1166] Daniele Bringhenti, Jalolliddin Yusupov, Alejandro Molina Zarca, Fulvio Valenza, Riccardo Sisto, Jorge Bernal Bernabe, and Antonio Skarmeta. Automatic, verifiable and optimized policybased security enforcement for SDNaware IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002468.

Shi:2022:SDU

[1167] Yao Shi, Emad Alsusa, and Mohammed W. Baidas. A survey on downlink-uplink decoupled access: Advances, challenges, and open problems. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001906.

Illiano:2022:QIP

[1168] Jessica Illiano, Marcello Caleffi, Antonio Manzalini, and Angela Sara Cacciapuoti. Quantum Internet protocol stack: a comprehensive survey. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002250.

Hegyi:2022:SDD

[1169] Péter Hegyi. Service deployment design in latency-critical multi-cloud environment. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001487.

Anonymous:2022:S

[1170] Anonymous. 4 September 2022. Computer Networks (Amsterdam, Netherlands: 1999), 214(??):??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBm

[1171] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 214(??):??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128622003474.

Jmila:2022:AML

[1172] Houda Jmila and Mohamed Ibn Khedher. Adversarial machine learning for network intrusion detection: a comparative study. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002146.

Cao:2022:FSE

[1173] Yibo Cao, Shiyuan Xu, Xue Chen, Yunhua He, and Shuo Jiang. A forward-secure and efficient authentication protocol through lattice-based group signature in VANETs scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 214(??): ??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002626.

Lasla:2022:GPE

[1174] Noureddine Lasla, Lina Al-Sahan, Mohamed Abdallah, and Mohamed Younis. Green-PoW: an energy-efficient blockchain Proof-of-Work consensus algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002432.

Luong:2022:JTS

[1175] Nguyen Cong Luong, Tran The Anh, Zehui Xiong, Dusit Niyato, and Dong In Kim. Joint time scheduling and transaction fee selection in blockchain-based RFpowered backscatter cognitive radio network. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002535.

Barakabitze:2022:SNQ

[1176] Alcardo Alex Barakabitze and Ray Walshe. SDN and NFV for QoEdriven multimedia services delivery: the road towards 6G and beyond networks. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002523. See

corrigendum [1588].

Yang:2022:DDD

[1177] Donghui Yang, Zhenyu Li, Haiyang Jiang, Gareth Tyson, Hongtao Li, Gaogang Xie, and Yu Zeng. A deep dive into DNS behavior and query failures. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002511.

Kumari:2022:TOF

[1178] Nidhi Kumari, Anirudh Yadav, and Prasanta K. Jana. Task offloading in fog computing: a survey of algorithms and optimization techniques. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002547.

Sasabe:2022:AMD

[1179] Masahiro Sasabe and Masaki Kiyomitsu. Analysis of minimum distribution time of two-class tit-for-tat-based P2P file distribution. Computer Networks (Amsterdam, Netherlands: 1999), 214(??):??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002572.

Habibi:2022:RBP

[1180] Mehdi Habibi and Hamzeh Beyranvand. Reducing blocking probability and QoT violation in dynamic elastic optical networks via load-aware margin selection. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002602.

Chen:2022:AJC

[1181] Kuan-Wei Chen, Chih-Min Chao, Chih-Yu Lin, and Chun-Chao Yeh. Anti-jamming channel hopping protocol design based on channel occupancy probability for Cognitive Radio Networks. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200247X.

Li:2022:CEC

[1182] Nan Li, Alexandros Iosifidis, and Qi Zhang. Collaborative edge computing for distributed CNN inference acceleration using receptive field-based segmentation. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002638.

Picano:2022:HLV

[1183] Benedetta Picano and Romano Fantacci. Human-in-the-loop virtual reality offloading scheme in wireless 6G terahertz networks. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200264X.

Zafar:2022:DDS

[1184] Samra Zafar, Zefeng Lv, Nizam Hussain Zaydi, Muhammad Ibrar, and Xiaopeng Hu. DSMLB: Dynamic switch-migration based load balancing for software-defined IoT network. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002596.

Peng:2022:MVF

[1185] Songtao Peng, Jiaqi Nie, Xincheng Shu, Zhongyuan Ruan, Lei Wang, Yunxuan Sheng, and Qi Xuan. A multi-view framework for BGP anomaly detection via graph attention network. Computer Networks (Amsterdam, Netherlands: 1999), 214(??):??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200250X.

Markovitz:2022:LSB

[1186] Oren Markovitz and Michael Segal.

LEO satellite beam management algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002687.

Tsiropoulou:2022:GEI

[1187] Eirini Eleni Tsiropoulou, Christos Douligeris, Luca Foschini, Gang Li, and Theofanis P. Raptis. Guest editorial: 26th IEEE symposium on computers and communications (ISCC 2021) selected papers. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002729.

Fouladi:2022:DAD

[1188] Ramin Fadaei Fouladi, Orhan Ermis, and Emin Anarim. A DDoS attack detection and countermeasure scheme based on DWT and auto-encoder neural network for SDN. Computer Networks (Amsterdam, Netherlands: 1999), 214(??):??, September 4, 2022.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002560.

Bomgni:2022:NNS

[1189] Alain Bertrand Bomgni, Miguel Landry Foko Sindjoung, Dhalil Kamdem Tchibonsou, Mthulisi Velempini, and Jean Frédéric Myoupo. NE-SEPRIN: a new scheme for energyefficient permutation routing in IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002699.

Chen:2022:DBP

[1190] Wenchao Chen, Guanqun Shen, Kaikai Chi, Shubin Zhang, and Xiaolong Chen. DRL based partial offloading for maximizing sum computation rate of FDMA-based wireless powered mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002675.

Duan:2022:FFG

[1191] Jian-Hui Duan, Wenzhong Li, Xiao Zhang, and Sanglu Lu. Forecasting fine-grained city-scale cellular traffic with sparse crowdsourced measurements. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002663.

Huang:2022:FEC

[1192] Shan Huang, Dezun Dong, Zejia Zhou, Hanyi Shi, Wenxiang Yang, and Xiangke Liao. FastCredit: Expediting credit-based congestion control in datacenters. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002481.

Mertens:2022:MFC

[1193] J. S. Mertens, L. Galluccio, and G. Morabito. MGM-4-FL: Combining federated learning and model gossiping in WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002584.

Wang:2022:UMS

[1194] Jiacheng Wang, Jianhui Zhang, Liming Liu, Xuzhao Zheng, Hanxiang Wang, and Zhigang Gao. Utility maximization for splittable task offloading in IoT edge network. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002705.

Qian:2022:DDS

[1195] Mimi Qian, Lin Cui, Xiaoquan Zhang, Fung Po Tso, and Yuhui Deng. Drops: Detecting silent packet drops on programmable data plane. Computer Networks (Amsterdam, Netherlands: 1999), 214(??):??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002742.

Jurdak:2022:EBB

[1196] Raja Jurdak, Juan M. Corchado, Jong Hyuk Park, Chintan Bhatt, and Kapal Dev. Editorial: Blockchain based sustainable, secure healthcare systems. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200278X.

Wang:2022:RAO

[1197] Zhifei Wang, Xiangming Wen, Zhaoming Lu, Wenpeng Jing, and Yujing Zhang. Random access optimization for initial access and seamless handover for 5g-satellite network. Computer Networks (Amsterdam, Netherlands: 1999), 214(??):??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002791.

Eltaief:2022:FCF

[1198] Hamdi Eltaief. Flex-CC: a flexible connected chains scheme for multicast source authentication in dynamic SDN environment. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002766.

Fu:2022:GIE

[1199] Juan juan Fu and Xing lan Zhang. Gradient importance enhancement based feature fusion intrusion detection technique. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002754.

Almasan:2022:EER

[1200] Paul Almasan, Shihan Xiao, Xiangle Cheng, Xiang Shi, Pere Barlet-Ros, and Albert Cabellos-Aparicio. Enero: Efficient real-time wan routing optimization with deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002717.

Dong:2022:SSI

[1201] Xiaodong Dong and Binlei Cai. Slardar: Scheduling information incomplete inter-datacenter deadline-aware coflows with a decentralized framework. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002808.

Wang:2022:ODO

[1202] Cen Wang, Noboru Yoshikane, Filippos Balasis, and Takehiro Tsuritani. Osdl: Dedicated optical slice provisioning in support of distributed deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002857.

Ribeiro:2022:DAE

[1203] Rafael Hengen Ribeiro, Arthur Selle Jacobs, Luciano Zembruzki, Ricardo Parizotto, Eder John Scheid, Alberto Egon Schaeffer-Filho, Lisandro Zambenedetti Granville, and Burkhard Stiller. A deterministic approach for extracting network security intents. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002389.

deSouza:2022:IDP

[1204] Cristiano Antonio de Souza, Carlos Becker Westphall, Renato Bobsin Machado, Leandro Loffi, Carla Merkle Westphall, and Guilherme Arthur Geronimo. Intrusion detection and prevention in fog based IoT environments: a systematic literature review. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002651.

Anonymous:2022:Oa

[1205] Anonymous. 9 October 2022. Computer Networks (Amsterdam, Netherlands: 1999), 215(??):??, October 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBn

[1206] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 215(??):??, October 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128622003802.

Alam:2022:ESB

[1207] Gowthul Alam MM, Jerald Nirmal Kumar S., Uma Mageswari R., and Michael Raj TF. An efficient SVM based DEHO classifier to detect DDoS attack in cloud computing environment. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002559.

Zhang:2022:PUT

[1208] Jiang Zhang, Lillian Clark, Matthew Clark, Konstantinos Psounis, and Peter Kairouz. Privacy-utility trades in crowdsourced signal map obfuscation. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002833.

Mouawad:2022:RSS

[1209] Mostafa Mouawad, Firmin Mah, and Zbigniew Dziong. Rrh-sector selection and load balancing based on MDP and dynamic RRH-Sector-BBU mapping in C-RAN. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002869.

Wu:2022:CMU

[1210] Fan Wu, Chaonong Xu, Yutong Zhu, and Chao Li. Complexity of minimum uplink power scheduling with delay bound for backbone-assisted pdnoma wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 215(??):??, October 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002870.

Li:2022:MOO

[1211] Jiahui Li, Geng Sun, Aimin Wang, Ming Lei, Shuang Liang, Hui Kang, and Yanheng Liu. A many-objective optimization charging scheme for wireless rechargeable sensor networks via mobile charging vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 215(??):??, October 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002882.

Zeng:2022:OMQ

[1212] Hongxin Zeng, Lin Cui, Fung Po Tso, and Zhen Zhang. Optimizing multipath QUIC transmission over heterogeneous paths. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002894.

Wang:2022:PPF

[1213] Weilong Wang, Yingjie Wang, Yan Huang, Chunxiao Mu, Zice Sun, Xiangrong Tong, and Zhipeng Cai. Privacy protection federated learning system based on blockchain and edge computing in mobile crowdsourcing. Computer Networks (Amsterdam, Netherlands: 1999), 215(??):??, October 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002936.

Tu:2022:RRN

[1214] Huaqing Tu, Gongming Zhao, Hongli Xu, Yangming Zhao, Yuhang Qiu, and Liusheng Huang. Rons: Robust network function services in clouds. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002961.

Leyva-Pupo:2022:DUP

[1215] Irian Leyva-Pupo, Cristina Cervelló-Pastor, Christos Anagnostopoulos, and Dimitrios P. Pezaros. Dynamic UPF placement and chaining reconfiguration in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 215(??):??, October 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002900.

Bringhenti:2022:ODF

[1216] Daniele Bringhenti and Fulvio Valenza. Optimizing distributed firewall reconfiguration transients. Computer Networks (Amsterdam, Netherlands: 1999), 215(??):??, October 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200281X.

Qu:2022:RAM

[1217] Bin Qu, Yan Bai, Yul Chu, Li e Wang, Feng Yu, and Xianxian Li. Resource allocation for MEC system with multi-users resource competition based on deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002778.

Robin:2022:PPS

[1218] Debobroto Das Robin and Javed I. Khan. P4te: Pisa switch based traffic engineering in fat-tree data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200295X.

Gedikli:2022:DRL

[1219] Ahmet Melih Gedikli, Mehmet Koseoglu, and Sevil Sen. Deep reinforcement learning based flexible preamble allocation for RAN slicing in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002912.

DiCicco:2022:OTR

[1220] Nicola Di Cicco, Federico Tonini, Valentina Cacchiani, and Carla Raffaelli. Optimization over time of reliable 5g-ran with network function migrations. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002985.

Costa:2022:MFC

[1221] Breno Costa, João Bachiega, Leonardo Rebouças Carvalho, Michel Rosa, and Aleteia Araujo. Monitoring fog computing: a review, taxonomy and open challenges. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002845.

Anonymous:2022:Ob

[1222] Anonymous. 24 October 2022. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBo

[1223] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128622004066.

Khawam:2022:CSG

[1224] Kinda Khawam, Hassan Fawaz, Samer Lahoud, Odalric-Ambrym Maillard, and Steven Martin. A channel selection game for multi-operator Lo-RaWAN deployments. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002821.

Oligeri:2022:GSD

[1225] Gabriele Oligeri, Savio Sciancalepore, Omar Adel Ibrahim, and Roberto Di Pietro. GPS spoofing detection via crowd-sourced information for connected vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003103.

Vanitha:2022:OPR

[1226] C. N. Vanitha, S. Malathy, Rajesh Kumar Dhanaraj, and Anand Nayyar. Optimized Pollard route deviation and route selection using Bayesian machine learning techniques in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003097.

Beduneau:2022:MPC

[1227] Guillaume Béduneau, Ghada Jaber, and Bertrand Ducourthial. A method for predicting its cooperative applications performances. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002614.

Paolucci:2022:LCS

[1228] Francesco Paolucci, Davide Scano, Piero Castoldi, and Emiliano De Paoli. Latency control in service chaining using p4-based data plane programmability. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003085.

Bhamare:2022:IBN

[1229] Deval Bhamare, Andreas Kassler, Jonathan Vestin, Mohammad Ali Khoshkholghi, Javid Taheri, Toktam Mahmoodi, Peter Öhlén, and Calin Curescu. IntOpt: In-band network telemetry optimization framework to monitor network slices using P4. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002973.

Lima:2022:GAS

[1230] Diogo Lima and Hugo Miranda. A geographical-aware state deployment service for fog computing. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002948.

Gharagezlou:2022:EEP

[1231] Abdolrasoul Sakhaei Gharagezlou, Mahdi Nangir, and Nima Imani. Energy efficient power allocation with joint antenna and user selection in massive MIMO systems. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003073.

Xie:2022:FFB

[1232] Shengxu Xie, Guyu Hu, Changyou Xing, Jiachen Zu, and Yaqun Liu.

Fint: Flexible in-band network telemetry method for data center network. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003115.

Niu:2022:DLO

[1233] Juncui Niu, Shubin Zhang, Kaikai Chi, Guanqun Shen, and Wei Gao. Deep learning for online computation offloading and resource allocation in NOMA. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003152.

Li:2022:TTS

[1234] Yanbin Li, Jiajie Zhu, Ming Tang, Shougang Ren, and Fusheng Wu. Tscl: a time-space crossing location for sidechannel leakage detection. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003176.

Elkael:2022:MBR

[1235] Maxime Elkael, Massinissa Ait Aba, Andrea Araldo, Hind Castel-Taleb, and Badii Jouaber. Monkey business: Reinforcement learning meets neighborhood search for virtual network embedding. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002924.

Sharma:2022:BME

[1236] Payal Sharma and B. R. Purushothama. Bp-mgkm: an efficient multi-group key management scheme based on bivariate polynomial. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003188.

Yan:2022:FLP

[1237] Lei Yan, Maode Ma, Dandan Li, Xiaohong Huang, Yan Ma, and Kun Xie. A flexible and lightweight privacypreserving handshake protocol based on DTLShps for IoT. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002730.

Pirayesh:2022:PHB

[1238] Jamshid Pirayesh, Alberto Giaretta, Mauro Conti, and Parviz Keshavarzi. A PLS-HECC-based device authentication and key agreement scheme for smart home networks. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200216X.

Febro:2022:SDD

[1239] Aldo Febro, Hannan Xiao, Joseph Spring, and Bruce Christianson. Synchronizing DDoS defense at network edge with P4, SDN, and Blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003310.

Naha:2022:MLR

[1240] Ranesh Naha, Saurabh Garg, Sudheer Kumar Battula, Muhammad Bilal Amin, and Dimitrios Georgakopoulos. Multiple linear regression-based energy-aware resource allocation in the Fog computing environment. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003164.

Castellanos:2022:MOO

[1241] German Castellanos, Simon De Gheselle, Luc Martens, Niels Kuster, Wout Joseph, Margot Deruyck, and Sven Kuehn. Multi-objective optimisation of human exposure for various 5G network topologies in Switzerland. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003231.

Sun:2022:ASW

[1242] Guoying Sun, Zhaoxin Zhang, Yanan Cheng, and Tingting Chai. Adaptive segmented webpage text based malicious website detection. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003140.

Herreria-Alonso:2022:ELS

[1243] Sergio Herrería-Alonso, Andrés Suárez-González, Miguel Rodríguez-Pérez, and Cándido López-García. Enhancing LoRaWAN scalability with Longest First Slotted CSMA. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200322X.

Chen:2022:DSC

[1244] Junmei Chen, Hao Chen, and Zongpeng Li. A double serial concatenated code using CRC-aided error correction for highly reliable communication. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003279.

Zhao:2022:CPA

[1245] Yubo Zhao and Jinyong Chang. Certificateless public auditing scheme with designated verifier and privacypreserving property in cloud storage. Computer Networks (Amsterdam, Netherlands: 1999), 216(??): ??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003322.

Miura:2022:PBD

[1246] Hideyoshi Miura, Kouji Hirata, and Takuji Tachibana. P4-based design of fast failure recovery for softwaredefined networks. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003358.

Sharma:2022:OEO

[1247] Abhilasha Sharma and Lalit Kumar Awasthi. Ob-EID: Obstacle aware event information dissemination for SDN enabled vehicular network. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003255.

Blaise:2022:GAD

[1248] Agathe Blaise, Mathieu Bouet, Vania Conan, and Stefano Secci. Group anomaly detection in mobile app usages: a spatiotemporal convex hull methodology. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003371.

Lv:2022:ANN

[1249] Yunxin Lv, Zixin Liu, Meihua Bi, Hao Chi, Yanrong Zhai, Yang Lu, and Zhengfeng Qian. Artificial neural network assisted polling scheme for central coordinated low-latency WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003541.

Li:2022:UNA

[1250] Kehong Li, Wengang Ma, Huawei Duan, Han Xie, Juanxiu Zhu, and Ruiqi Liu. Unbalanced network attack traffic detection based on feature extraction and GFDA-WGAN. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003401.

Khorasani:2022:NTD

[1251] Yaghoub Khorasani, Akbar Ghaffarpour Rahbar, and Mohammad Jafari-Beyrami. A novel two-dimensional metric for fragmentation evaluation in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200336X.

Batchu:2022:IAE

[1252] Raj Kumar Batchu and Hari Seetha.

An integrated approach explaining the detection of distributed denial of service attacks. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003334.

Restuccia:2022:GWA

[1253] Francesco Restuccia, Salvatore D’Oro, Amani Al-Shawabka, Bruno Costa Rendon, Kaushik Chowdhury, Stratis Ioannidis, and Tommaso Melodia. Generalized wireless adversarial deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003292.

Wang:2022:EEP

[1254] Yantong Wang and Vasilis Friderikos.

Energy-efficient proactive caching with multipath routing. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003346.

Gao:2022:RCS

[1255] Jing Gao, Lei Feng, Peng Yu, Fanqin Zhou, Zihao Wu, Xuesong Qiu, Jingchun Li, and Yifei Zhu. Resource consumption and security-aware multi-tenant service function chain deployment based on hypergraph matching. Computer Networks (Amsterdam, Netherlands: 1999), 216(??): ??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003516.

Silveira:2022:TCB

[1256] Lucas B. D. Silveira, Henrique C. de Resende, Cristiano B. Both, Johann M. Marquez-Barja, Bruno Silvestre, and Kleber V. Cardoso. Tutorial on communication between access networks and the 5G core. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003528.

Tedeschi:2022:SBC

[1257] Pietro Tedeschi, Savio Sciancalepore, and Roberto Di Pietro. Satellitebased communications security: a survey of threats, solutions, and research challenges. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200319X.

Wu:2022:SAB

[1258] Weiqi Wu, Xingfu Wang, Ammar Hawbani, Longzhi Yuan, and Wei Gong. A survey on ambient backscatter communications: Principles, systems, applications, and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003139.

Vargas-Arcila:2022:SDS

[1259] Angela Vargas-Arcila, Juan Carlos Corrales, Araceli Sanchis, and Alvaro Rendón. SOFI dataset: Symptom-fault relationship for IP-network. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003127.

Rajak:2022:EEM

[1260] Shaik Rajak, Poongundran Selvaprabhu, Sunil Chinnadurai, A. S. M. Sanwar Hosen, Aldosary Saad, and Amr Tolba. Energy efficient MIMO-NOMA aided IoT network in B5G communications. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003218.

Satpathy:2022:CEV

[1261] Anurag Satpathy, Manmath Narayan Sahoo, Arun Kumar Sangaiah, Chittaranjan Swain, and Sambit Bakshi.

CoMap: an efficient virtual network re-mapping strategy based on coalitional matching theory. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003206.

Zhang:2022:BEM

[1262] Jianfeng Zhang, Wensheng Zhang, and Jingdong Xu. Bandwidth-efficient multi-task AI inference with dynamic task importance for the Internet of Things in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003280.

Anonymous:2022:N

[1263] Anonymous. 9 November 2022. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBp

[1264] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128622004467.

Xiao:2022:SAA

[1265] Da Xiao, Shuo Chen, Wei Ni, Jie Zhang, Andrew Zhang, and Renping Liu. A sub-action aided deep reinforcement learning framework for latencysensitive network slicing. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003383.

Wang:2022:CUN

[1266] Kewei Wang, Changzhen Hu, and Chun Shan. Calculation of utility of network services based on state manifolds. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003267.

Zhang:2022:SCV

[1267] Lejun Zhang, Jinlong Wang, Weizheng Wang, Zilong Jin, Yansen Su, and Huiling Chen. Smart contract vulnerability detection combined with multi-objective detection. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003437.

Cavdar:2022:APS

[1268] Mustafa Can Çavdar, Ibrahim Korpeoglu, and Özgür Ulusoy. Application placement with shared monitoring points in multi-purpose IoT wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200353X.

Su:2022:SRS

[1269] Yishan Su, Yao Liu, Rong Fan, Luyu Li, Meng Han, and Hongliang Zhang. A secure relay selection scheme based on cooperative jamming for Underwater Acoustic Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003565.

Jiang:2022:AMA

[1270] Minghao Jiang, Zhen Li, Peipei Fu, Wei Cai, Mingxin Cui, Gang Xiong, and Gaopeng Gou. Accurate mobileapp fingerprinting using flow-level relationship with graph neural networks. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003577.

Rago:2022:TDS

[1271] Arcangela Rago, Sergio Martiradonna, Giuseppe Piro, Andrea Abrardo, and Gennaro Boggia. A tenant-driven slicing enforcement scheme based on Pervasive Intelligence in the Radio Access Network. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003413.

Kang:2022:FTR

[1272] Rui Kang, Fujun He, and Eiji Oki. Fault-tolerant resource allocation model for service function chains with joint diversity and redundancy. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003425.

Kolukisa:2022:DLA

[1273] Burak Kolukisa, Veli Can Yildirim, Bahadir Elmas, Cem Ayyildiz, and Vehbi Cagri Gungor. Deep learning approaches for vehicle type classification with 3-D magnetic sensor. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003668.

Chhea:2022:EEF

[1274] Kimchheang Chhea and Jung-Ryun Lee. Energy-efficient full-duplex D2D for SWIPT-empowered underlay cellular networks using a deep neural network. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003656.

Buccafurri:2022:IPT

[1275] Francesco Buccafurri, Vincenzo De Angelis, and Gianluca Lax. An integritypreserving technique for range queries over data streams in two-tier sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003619.

Mesodiakaki:2022:REE

[1276] A. Mesodiakaki, E. Zola, and A. Kassler. Robust and energy-efficient user association and traffic routing in B5G HetNets. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003553.

Alrubei:2022:HHB

[1277] Subhi Alrubei, Edward Ball, and Jonathan Rigelsford. HDPoA: Honesty-based distributed proof of authority via scalable work consensus protocol for IoT-blockchain applications. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003723.

Abdollahi:2022:OOE

[1278] Mehran Pourmohammad Abdollahi, Javad Musevi Niya, and Mahdi Nangir.

Optimizing the operator’s economy in multi-tier cellular networks with underlaid D2D communications. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003644.

Zhang:2022:ACB

[1279] Pengfei Zhang, Xiang Cheng, Sen Su, and Ning Wang. Area coveragebased worker recruitment under geo-indistinguishability. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003747.

Mountaser:2022:GMO

[1280] Ghizlane Mountaser, Enric Pardo, and Toktam Mahmoodi. Graphical modelling and optimization of RAN function split deployed through UAVs. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003309.

Yang:2022:DPA

[1281] Yi Yang, Fenglei Li, Xinzhe Zhang, Zhixin Liu, and Kit Yan Chan. Dynamic power allocation in cellular network based on multi-agent double deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003760.

Wang:2022:TAM

[1282] Xueyi Wang, Xingwei Wang, Dongkuo Wu, Lianbo Ma, and Min Huang. Truthful auction mechanisms for VNF chain provisioning and allocation across geo-distributed datacenters. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003693.

Kumari:2022:PQL

[1283] Swati Kumari, Maninder Singh, Raman Singh, and Hitesh Tewari. A post-quantum lattice based lightweight authentication and code-based hybrid encryption scheme for IoT devices. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200367X.

Li:2022:PBR

[1284] Yupeng Li, Mengjia Xia, Jingpu Duan, and Yang Chen. Pricing-based resource allocation in three-tier edge computing for social welfare maximization. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003589.

Mazziane:2022:ACM

[1285] Younes Ben Mazziane, Sara Alouf, and Giovanni Neglia. Analyzing count min sketch with conservative updates. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003607.

Lu:2022:FAM

[1286] Tingting Lu and Junfeng Wang.

F2DC: Android malware classification based on raw traffic and neural networks. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003632.

Hou:2022:NID

[1287] Tianhao Hou, Hongyan Xing, Xinyi Liang, Xin Su, and Zenghui Wang. Network intrusion detection based on DNA spatial information. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003620.

Violos:2022:IHA

[1288] John Violos, Stylianos Tsanakas, Theodoros Theodoropoulos, Aris Leivadeas, Konstantinos Tserpes, and Theodora Varvarigou. Intelligent horizontal autoscaling in edge computing using a double tower neural network. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003735.

Kuang:2022:PAE

[1289] Ye Kuang, Dandan Li, Xiaohong Huang, Mo Zhou, and Weidu Wang. PoiEvent: an approach to extract the persistent and destructive routing events. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003590.

Tran:2022:CMR

[1290] Phuong T. Tran, Ba Cao Nguyen, Tran Manh Hoang, Le The Dung, and Nguyen Van Vinh. Combining multi-RIS and relay for performance improvement of multi-user NOMA systems. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003875.

Duong:2022:FWH

[1291] Tho Minh Duong and Sungoh Kwon. A frame work of handover analysis for randomly deployed heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003851.

Ferriol-Galmes:2022:BDT

[1292] Miquel Ferriol-Galmés, José Suárez-Varela, Jordi Paillissé, Xiang Shi, Shihan Xiao, Xiangle Cheng, Pere Barlet-Ros, and Albert Cabellos-Aparicio. Building a Digital Twin for network optimization using graph neural networks. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003681.

Ren:2022:NMT

[1293] Junyu Ren and Tuanfa Qin. A novel multidimensional trust evaluation and fusion mechanism in fog-based Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003887.

Luglio:2022:SMB

[1294] M. Luglio, S. P. Romano, C. Roseti, and F. Zampognaro. Satellite multibeam multicast support for an efficient community-based CDN. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003863.

Shalini:2022:DDD

[1295] P. V. Shalini, V. Radha, and Sriram G. Sanjeevi. DOCUS-DDoS detection in SDN using modified CUSUM with flash traffic discrimination and mitigation. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003954.

Shariq:2022:EES

[1296] Mohd Shariq, Karan Singh, Chhagan Lal, Mauro Conti, and Tayyab Khan. ESRAS: an efficient and secure ultra-lightweight RFID authentication scheme for low-cost tags. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003942.

Barmpounakis:2022:ADQ

[1297] Sokratis Barmpounakis, Nikolaos Maroulis, Nikolaos Koursioumpas, Apostolos Kousaridas, Angeliki Kalamari, Panagiotis Kontopoulos, and Nancy Alonistioti. AI-driven, QoS prediction for V2X communications in beyond 5G systems. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003759.

Selim:2022:SBF

[1298] Mohamed Y. Selim and Ahmed E. Kamal. Self-backhauling failure mitigation using 5G new radio. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003899.

Masoudi-Sobhanzadeh:2022:RTI

[1299] Yosef Masoudi-Sobhanzadeh and Shabnam Emami-Moghaddam. A real-time IoT-based botnet detection method using a novel two-step feature selection technique and the support vector machine classifier. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003991.

Saad:2022:RTB

[1300] Muhammad Saad, Afsah Anwar, Ashar Ahmad, Hisham Alasmary, Murat Yuksel, and David Mohaisen. RouteChain: Towards blockchain-based secure and efficient BGP routing. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003966.

M:2022:SHO

[1301] Shyama M., Anju S. Pillai, and Alagan Anpalagan. Self-healing and optimal fault tolerant routing in wireless sensor networks using genetical swarm optimization. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003930.

Kaur:2022:RVI

[1302] Karamjeet Kaur, Veenu Mangat, and Krishan Kumar. A review on Virtualized Infrastructure Managers with management and orchestration features in NFV architecture. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003395.

Ramezanpour:2022:IZT

[1303] Keyvan Ramezanpour and Jithin Jagannath. Intelligent zero trust architecture for 5G/6G networks: Principles, challenges, and the role of machine learning in the context of O-RAN. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003929.

Akram:2022:EES

[1304] Waseem Akram, Khalid Mahmood, Xiong Li, Mazhar Sadiq, Zhihan Lv, and Shehzad Ashraf Chaudhry. An energy-efficient and secure identity based RFID authentication scheme for vehicular cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003711.

Taneja:2022:EAR

[1305] Ashu Taneja, Shalli Rani, Sahil Garg, Mohammad Mehedi Hassan, and Salman A. AlQahtani. Energy aware resource control mechanism for improved performance in future green 6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200370X.

Anonymous:2022:Da

[1306] Anonymous. 9 December 2022. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBq

[1307] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128622004753.

Xie:2022:IEA

[1308] Pengshou Xie, Haoxuan Yang, Tao Feng, and Yan Yan. Implementing efficient attribute encryption in IoV under cloud environments. Computer Networks (Amsterdam, Netherlands:

1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003978.

Sarraf-Maralaniyan:2022:MNM

[1309] Sarah Sarraf-Maralaniyan and Akbar Ghaffarpour Rahbar. MT^{3}A: a novel multicast routing, spectrum and modulation-level assignment in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004224.

Diarra:2022:RRA

[1310] Mamoutou Diarra, Walid Dabbous, Amine Ismail, Brice Tetu, and Thierry Turletti. RAPID: a RAN-aware performance enhancing proxy for high throughput low delay flows in MECenabled cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003917.

Gao:2022:CRS

[1311] Caina Gao, Jia Zhang, Linlin Guo, Ke Liu, Lili Meng, and Jiande Sun. Coordinated rate splitting and power allocation in energy-spectral efficiency tradeoff-based multicell networks. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200398X.

Li:2022:RNT

[1312] Wenhao Li, Xiao-Yu Zhang, Huaifeng Bao, Qiang Wang, and Zhaoxuan Li. Robust network traffic identification with graph matching. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004029.

Bella:2022:DIA

[1313] Giampaolo Bella, Pietro Biondi, Gianpiero Costantino, and Ilaria Matteucci. Designing and implementing an AUTOSAR-based Basic Software Module for enhanced security. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200411X.

Nandhini:2022:LEE

[1314] P. S. Nandhini, S. Kuppuswami, S. Malliga, and R. DeviPriya. A lightweight energy-efficient algorithm for mitigation and isolation of internal rank attackers in RPL based Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200425X.

Lopes:2022:PAT

[1315] Rafael Lopes, Denis Rosário, Eduardo Cerqueira, Helder Oliveira, and Sherali Zeadally. Priority-aware traffic routing and resource allocation mechanism for space-division multiplexing elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004236.

Ranjan:2022:TSD

[1316] Shashi Ranjan, Pranav Jha, Abhay Karandikar, and Prasanna Chaporkar. Two stage downlink scheduling for balancing QoS in multihop IAB networks. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004297.

Shan:2022:LBE

[1317] Minghui Shan, Sheng Zhang, Mingjun Xiao, and Yanchao Zhao. LENS: Bandwidth-efficient video analytics with adaptive super resolution. Computer Networks (Amsterdam, Netherlands: 1999), 218(??): ??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004261.

Wilhelmi:2022:AES

[1318] Francesc Wilhelmi, Lorenza Giupponi, and Paolo Dini. Analysis and evaluation of synchronous and asynchronous FLchain. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004248.

Ramineni:2022:PER

[1319] Chaturasan Ramineni, T. G. Venkatesh, and Lokesh Bommisetty. Performance evaluation of random access in narrow band Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004339.

Fedrizzi:2022:AGB

[1320] Riccardo Fedrizzi, Rasoul Behravesh, Cristina Costa, and Fabrizio Granelli. An adaptive GA-based slice provisioning method for vertical industries in 5G and beyond networks. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004315.

Ghosal:2022:SAS

[1321] Amrita Ghosal, Subir Halder, and Mauro Conti. Secure over-the-air software update for connected vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004285.

Tu:2022:BBT

[1322] Zhe Tu, Huachun Zhou, Kun Li, Haoxiang Song, and Yuzheng Yang. A blockchain-based trust and reputation model with dynamic evaluation mechanism for IoT. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004388.

Huang:2022:QDV

[1323] Guanglun Huang, Jianming Liu, Baoxian Zhang, and Cheng Li. Qualitydriven video streaming for ultradense OFDMA heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004327.

Uras:2022:MAR

[1324] Marco Uras, Enrico Ferrara, Raimondo Cossu, Antonio Liotta, and Luigi Atzori. MAC address de-randomization for WiFi device counting: Combining temporal- and content-based fingerprints. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004273.

Yadav:2022:IPS

[1325] Awaneesh Kumar Yadav, Manoj Misra, Pradumn Kumar Pandey, An Braeken, and Madhusanka Liyange. An improved and provably secure symmetric-key based 5G-AKA Protocol. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004340.

Ye:2022:MOT

[1326] Weidu Ye, Junzhou Luo, Wenjia Wu, Feng Shan, and Ming Yang. MU-TAA: an online trajectory optimization and task scheduling for UAV-aided edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200439X.

Tian:2022:RPB

[1327] Chuang Tian, Qi Jiang, Teng Li, Junwei Zhang, Ning Xi, and Jianfeng Ma. Reliable PUF-based mutual authentication protocol for UAVs towards multi-domain environment. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004558.

Hadi:2022:JRA

[1328] Majid Hadi and Reza Ghazizadeh.

Joint resource allocation, user clustering and 3-d location optimization in multi-UAV-enabled mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004546.

Shahid:2022:DRC

[1329] Syed Maaz Shahid and Sungoh Kwon. Distributed robust channel allocation for clustered cognitive radio-based IoT networks using graph theory. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004406.

Ou:2022:OCC

[1330] Wei Ou, Shiying Huang, Jingjing Zheng, Qionglu Zhang, Guang Zeng, and Wenbao Han. An overview on cross-chain: Mechanism, platforms, challenges and advances. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004121.

Samikwa:2022:AAR

[1331] Eric Samikwa, Antonio Di Maio, and Torsten Braun. ARES: Adaptive resource-aware split learning for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004145.

Koita:2022:GLS

[1332] Moussa Koïta, Youssouf M. Diagana, Oumar Y. Maïga, and Mamadou K. Traore. A generic learning simulation framework to assess security strategies in cyber-physical production systems. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004157.

Liu:2022:AGA

[1333] Jianmin Liu, Qi Wang, and Yongjun Xu. AR-GAIL: Adaptive routing protocol for FANETs using generative adversarial imitation learning. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004169.

Lyu:2022:CTE

[1334] Minzhao Lyu, Hassan Habibi Gharakheili, and Vijay Sivaraman. Classifying and tracking enterprise assets via dual-grained network behavioral analysis. Computer Networks (Amsterdam, Netherlands: 1999), 218(??): ??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004212.

Alani:2022:PCB

[1335] Mohammed M. Alani and Hissam Tawfik. PhishNot: a cloud-based machinelearning approach to phishing URL detection. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004418.

Alwan:2022:TSC

[1336] Ahmed A. Alwan, Allan J. Brimicombe, Mihaela Anca Ciupala, Seyed Ali Ghorashi, Andres Baravalle, and Paolo Falcarin. Timeseries clustering for sensor fault detection in large-scale cyber-physical systems. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004182.

Zhang:2022:DDP

[1337] Ziming Zhang, Xiaolong Xu, and Fu Xiao. 5GMEC-DP: Differentially private protection of trajectory data based on 5G-based mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004108.

Dandachi:2022:RCT

[1338] Ghina Dandachi, Sophie Cerf, Yassine Hadjadj-Aoul, Abdelkader Outtagarts, and Eric Rutten. A robust controltheory-based exploration strategy in deep reinforcement learning for virtual network embedding. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004005.

Parsa:2022:JPA

[1339] Ali Parsa, Neda Moghim, and Pouyan Salavati. Joint power allocation and MCS selection for energy-efficient link adaptation: a deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004200.

Zhang:2022:RDA

[1340] Junye Zhang, Peng Yu, Fanqin Zhou, Lei Feng, Wenjing Li, and Xuesong Qiu. Resource and delay aware finegrained service offloading in collaborative edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004170.

Azari:2022:ADL

[1341] Bahar Azari, Hai Cheng, Nasim Soltani, Haoqing Li, Yanyu Li, Mauro Belgiovine, Tales Imbiriba, Salvatore D’Oro, Tommaso Melodia, Yanzhi Wang, Pau Closas, Kaushik Chowdhury, and Deniz Erdoğmuş. Automated deep learning-based wide-band receiver. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004017.

Peng:2022:SOT

[1342] Yabin Peng, Caixia Liu, Shuxin Liu, Yuchen Liu, and Yiteng Wu. Smart-TRO: Optimizing topology robustness for Internet of Things via deep reinforcement learning with graph convolutional networks. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004194.

Rathee:2022:AIA

[1343] Geetanjali Rathee, Chaker Abdelaziz Kerrache, and Carlos T. Calafate. An ambient intelligence approach to provide secure and trusted Pub/Sub messaging systems in IoT environments. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004352.

Anonymous:2022:Db

[1344] Anonymous. 24 December 2022. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2022:EBr

[1345] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912862200528X.

Pinyoanuntapong:2022:ETN

[1346] Pinyarash Pinyoanuntapong, Prabhu Janakaraj, Ravikumar Balakrishnan, Minwoo Lee, Chen Chen, and Pu Wang. EdgeML: Towards networkaccelerated federated learning over wireless edge. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004303.

Ruchel:2022:ERS

[1347] Lucas V. Ruchel, Rogério C. Turchetti, and Edson T. de Camargo. Evaluation of the robustness of SDN controllers ONOS and ODL. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004376.

Debski:2022:RTS

[1348] Roman Debski and Rafał Drezewski.

Real-time surrogate-assisted preprocessing of streaming sensor data. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912862200456X.

Halgamuge:2022:ESP

[1349] Malka N. Halgamuge. Estimation of the success probability of a malicious attacker on blockchain-based edge network. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004364.

Mihailovic:2022:CMT

[1350] Andrej Mihailovic, Benyamin Abrishamchi, Mohammad Farhoudi, and A. Hamid Aghvami. A comprehensive multi-topology minimum set cover link-state routing approach for emerging random all-I access network topologies. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004522.

Li:2022:FFM

[1351] Zhao Li, Bigui Zhang, Chengyu Liu, Zhixian Chang, Kang G. Shin, and Zheng Yan. Framed Fidelity MAC: Losslessly packing multi-user transmissions in a virtual point-to-point framework. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004595.

Li:2022:HRN

[1352] Tengyao Li, Kaiyue Liu, Wangxin Feng, Chunfang Yang, and Xiangyang Luo. HeteroTiC: a robust network flow watermarking based on heterogeneous time channels. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004583.

Yi:2022:TPM

[1353] Yuzi Yi, Nafei Zhu, Jingsha He, Anca Delia Jurcut, and Bin Zhao. Toward pragmatic modeling of privacy information propagation in online social networks. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004637.

Latah:2022:CDC

[1354] Majd Latah and Kübra Kalkan. CWT-DPA: Component-wise waiting time for BC-enabled data plane authentication. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004571.

Tripathy:2022:EAI

[1355] Pujasuman Tripathy and P. M. Khilar. An ensemble approach for improving localization accuracy in wireless sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004613.

Huang:2022:LLW

[1356] Kaiwen Huang, Ronghui Hou, and Yingming Zeng. LWSBFT: Leaderless weakly synchronous BFT protocol. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004534.

Attanasio:2022:DSR

[1357] Giulia Attanasio, Claudio Fiandrino, Marco Fiore, Joerg Widmer, Norbert Ludant, Bastian Bloessl, Konstantinos Kousias, Özgü Alay, Lise Jacquot, and Razvan Stanica. In-depth study of RNTI management in mobile networks: Allocation strategies and implications on data trace analysis. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004625.

Lia:2022:NPD

[1358] Gianmarco Lia, Marica Amadeo, Giuseppe Ruggeri, Claudia Campolo, Antonella Molinaro, and Valeria Loscrì. In-network placement of delayconstrained computing tasks in a softwarized intelligent edge. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004662.

Enge:2022:OMA

[1359] Alexander Enge, Abylay Satybaldy, and Mariusz Nowostawski. An offline mobile access control system based on self-sovereign identity standards. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004686.

Hasselquist:2022:QTF

[1360] David Hasselquist, Christoffer Lindström, Nikita Korzhitskii, Niklas Carlsson, and Andrei Gurtov. QUIC throughput and fairness over dual connectivity. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004650.

Wang:2022:SBF

[1361] Hui Wang, Xiaowei Zhang, Ming Li, and Donghui Hu. SRRS: a blockchain fast propagation protocol based on non-Markovian process. Computer Networks (Amsterdam, Netherlands:

1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004698.

Sindjoung:2022:MAB

[1362] Miguel Landry Foko Sindjoung, Mthulisi Velempini, and Alain Bertrand Bomgni. A MEC architecture for a better quality of service in an Autonomous Vehicular Network. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004881.

Zheng:2022:IBN

[1363] Xiaoang Zheng, Aris Leivadeas, and Matthias Falkner. Intent Based Networking management with conflict detection and policy resolution in an enterprise network. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004911.

Hong:2022:RPN

[1364] Hsiang-Jen Hong, Wenjun Fan, Simeon Wuthier, Jinoh Kim, C. Edward Chow, Xiaobo Zhou, and Sang-Yoon Chang. Robust P2P networking connectivity estimation engine for permissionless Bitcoin cryptocurrency. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004704.

Bagnulo:2022:WLM

[1365] Marcelo Bagnulo and Alberto García-Martínez. When less is more: BBR versus LEDBAT++. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004947.

Busacca:2022:MMD

[1366] F. Busacca, L. Galluccio, and S. Palazzo. A marketplace model for drone-assisted edge computing in 5G scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200487X.

Mohammadi:2022:DSO

[1367] Marzieh Mohammadi, S. Ahmad Motamedi, and Saeed Sharifian. Drops on surface optimization (DSO): a new reinforcement learning based metaheuristic algorithm for virtual network functions placement in distributed cloud architectures. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004935.

Zou:2022:ECD

[1368] Hongcheng Zou, Jinshu Su, Ziling Wei, Shuhui Chen, and Baokang Zhao. An efficient cross-domain few-shot website fingerprinting attack with Brownian distance covariance. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004959.

Voulgaridis:2022:IDC

[1369] Konstantinos Voulgaridis, Thomas Lagkas, Constantinos Marios Angelopoulos, and Sotiris E. Nikoletseas. IoT and digital circular economy: Principles, applications, and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200490X.

Jagannath:2022:CSR

[1370] Anu Jagannath, Jithin Jagannath, and Prem Sagar Pattanshetty Vasanth Kumar. A comprehensive survey on radio frequency (RF) fingerprinting: Traditional approaches, deep learning, and open challenges. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004893.

Sahni:2022:DRS

[1371] Yuvraj Sahni, Jiannong Cao, Lei Yang, and Shengwei Wang. Distributed resource scheduling in edge computing: Problems, solutions, and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004649.

Lei:2022:QOM

[1372] Jianjun Lei, Lu Li, and Ying Wang. QoS-oriented media access control using reinforcement learning for next-generation WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004601.

Khoshkholghi:2022:EIS

[1373] Mohammad Ali Khoshkholghi and Toktam Mahmoodi. Edge intelligence for service function chain deployment in NFV-enabled networks. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004856.

Fan:2022:EEI

[1374] Linna Fan, Lin He, Enhuan Dong, Jiahai Yang, Chenglong Li, Jinlei Lin, and Zhiliang Wang. EvoIoT: an evolutionary IoT and non-IoT classification model in open environments. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004844.

Solomon:2022:SEL

[1375] Yossi Solomon, Osnat Mokryn, and Tsvi Kuflik. SDNSandbox — enabling learning-based innovation in provider networks. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004807.

Grasso:2022:HHL

[1376] Christian Grasso, Raoul Raftopoulos, Giovanni Schembra, and Salvatore Serrano. H-HOME: a learning framework of federated FANETs to provide edge computing to future delayconstrained IoT systems. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004832.

Guarino:2022:CCM

[1377] Idio Guarino, Giuseppe Aceto, Domenico Ciuonzo, Antonio Montieri, Valerio Persico, and Antonio Pescapè. Contextual counters and multimodal deep learning for activity-level traffic classification of mobile communication apps during COVID-19 pandemic. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004868.

Ayoub:2022:EAI

[1378] Omran Ayoub, Nicola Di Cicco, Fatima Ezzeddine, Federica Bruschetta, Roberto Rubino, Massimo Nardecchia, Michele Milano, Francesco Musumeci, Claudio Passera, and Massimo Tornatore. Explainable artificial intelligence in communication networks: a use case for failure identification in microwave networks. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200500X.

Navarro:2022:HRN

[1379] Jose M. Navarro, Alexis Huet, and Dario Rossi. Human readable network troubleshooting based on anomaly detection and feature scoring. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004819.

Ashraf:2022:FCV

[1380] M. Wasim Abbas Ashraf, Chuanhe Huang, Arif Raza, Kashif Sharif, Md Monjurul Karim, and Shidong Huang. Forwarding and caching in video streaming over ICSDN: a clean-slate publish–subscribe approach. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004674.

Anonymous:2023:Ja

[1381] Anonymous. January 2023. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBa

[1382] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005795.

Sharvari:2023:CCC

[1383] N. P. Sharvari, Dibakar Das, Jyotsna Bapat, and Debabrata Das. Connectivity and collision constrained opportunistic routing for emergency communication using UAV. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005023.

Avanzato:2023:DBA

[1384] Roberta Avanzato, Francesco Beritelli, and Corrado Rametta. An DL-based approach for packet error compensation using radio mobile network quality parameters in a rainfall scenario. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004972.

Baidas:2023:NCU

[1385] Mohammed W. Baidas, Ahmed M.

AbdelGhaffar, and Emad Alsusa. Network-coded uplink clustered NOMA relay networks: Models and performance comparisons. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004996.

Wan:2023:CPM

[1386] Zheng Wan and Xiaogang Dong. Computation power maximization for mobile edge computing enabled dense network. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004923.

Anedda:2023:SSC

[1387] Matteo Anedda, Mauro Fadda, Roberto Girau, Giovanni Pau, and Daniele Giusto. A social smart city for public and private mobility: a real case study. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004984.

Robles-Enciso:2023:MLG

[1388] Alberto Robles-Enciso and Antonio F. Skarmeta. A multi-layer guided reinforcement learning-based tasks offloading in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005102.

Foroughi:2023:AAD

[1389] Parisa Foroughi, Frank Brockners, and Jean-Louis Rougier. ADT: AIdriven network telemetry processing on routers. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005084.

Zhang:2023:LCV

[1390] Shengyu Zhang and Kwan L. Yeung. Location constrained virtual optical network embedding in spacedivision multiplexing elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005096.

Lira:2023:SAU

[1391] Clayton J. N. Lira, Raul C. Almeida, and Daniel A. R. Chaves. Spectrum allocation using multiparameter optimization in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005126.

Ma:2023:LDT

[1392] Dongchao Ma, Pengyu Wang, Lihua Song, Wenlong Chen, Li Ma, Mingwei Xu, and Laizhong Cui. A lightweight deployment of TD routing based on SD-WANs. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005205.

Liu:2023:JCZ

[1393] Zhixin Liu, Xi Chen, Yi Yang, Kit Yan Chan, and Yazhou Yuan. Joint cell zooming and sleeping strategy in ultra dense heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005163.

Qi:2023:LTB

[1394] Liuling Qi, Junfeng Tian, Mengjia Chai, and Hongyun Cai. LightPoW: a trust based time-constrained PoW for blockchain in Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200514X.

Xu:2023:CDM

[1395] Zhilin Xu, Hao Sun, and Weibin Han. A collaboration-driven mechanism for AI diagnose with multiple requesters under incomplete information. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005060.

Jiang:2023:AAP

[1396] Wenhao Jiang and Yuebin Bai. APGNN : Alarm propagation graph neural network for fault detection and alarm root cause analysis. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005199.

Venkateswararao:2023:NPC

[1397] Kuna Venkateswararao, Pravati Swain, Shashi Shekhar Jha, Iacovos Ioannou, and Andreas Pitsillides. A novel power consumption optimization framework in 5G heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005217.

Deng:2023:BSG

[1398] Hanqiang Deng, Jian Huang, Quan Liu, Cong Zhou, and Jialong Gao. BGSD: a SBERT and GAT-based service discovery framework for heterogeneous distributed IoT. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005229.

Wang:2023:CAF

[1399] Zhiyuan Wang, Hongli Xu, Yang Xu, Zhida Jiang, and Jianchun Liu. CoopFL: Accelerating federated learning with DNN partitioning and offloading in heterogeneous edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005242.

FeCastro:2023:UGU

[1400] Igor C. F.e Castro, Eduardo P. M.

Câmara, Marcos A. M. Vieira, and Luiz F. M. Vieira. UW-GRE: Underwater greedy geographic routing by network embedding. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005072.

Vrind:2023:FBA

[1401] Tushar Vrind, Chandan Kumar, Lalit Pathak, and Debabrata Das. Feedbackbased algorithm for aerial cell’s trajectory using deep learning for efficient performance in 6G. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005333.

Ariemma:2023:LLS

[1402] Lorenzo Ariemma, Alessandro Dell’Orco, Simone Liotta, Massimo Candela, and Giuseppe Di Battista. Long-lasting sequences of BGP updates. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005151.

Kazemi:2023:OSF

[1403] Seyed Mojtaba Kazemi, Shamsollah Ghanbari, Manochehr Kazemi, and Mohamed Othman. Optimum scheduling in fog computing using the Divisible Load Theory (DLT) with linear and nonlinear loads. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005175.

Darbandeh:2023:SSA

[1404] Foroozan Ghosairi Darbandeh and Masoumeh Safkhani. SAPWSN: a secure authentication protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005035.

Biomo:2023:NRP

[1405] Jean-Daniel Medjo Me Biomo, Thomas Kunz, and Marc St-Hilaire. A novel routing protocol for reducing packet delay with multi-beam antennas. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005138.

Esubonteng:2023:EIA

[1406] Paa Kwesi Esubonteng and Roberto Rojas-Cessa. Effect of the incident angle of a transmitting laser light on the coverage of a NLOS-FSO network. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005382.

Bonati:2023:OGA

[1407] Leonardo Bonati, Michele Polese, Salvatore D’Oro, Stefano Basagni, and Tommaso Melodia. OpenRAN Gym: AI/ML development, data collection, and testing for O-RAN on PAWR platforms. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005369.

Yan:2023:NMA

[1408] Yan Yan, Baoxian Zhang, and Cheng Li. A networked multi-agent reinforcement learning approach for cooperative FemtoCaching assisted wireless heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005473.

Sun:2023:VCB

[1409] Sheng Sun, Zengqi Zhang, Quyang Pan, Min Liu, and Zhongcheng Li. Vehicle-cluster-based opportunistic relays for data collection in intelligent transportation systems. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005436.

Xiong:2023:FTE

[1410] Runqun Xiong, Chuan Liang, Huajun Zhang, Xiangyu Xu, and Junzhou Luo. FlyingLoRa: Towards energy efficient data collection in UAVassisted LoRa networks. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200545X.

Mehmood:2023:IDA

[1411] Kashif Mehmood, Katina Kralevska, and David Palma. Intent-driven autonomous network and service management in future cellular networks: a structured literature review. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005114.

Sharma:2023:ISS

[1412] Rohit Sharma, Danda B. Rawat, Amiya Nayak, Sheng-Lung Peng, and Qin Xin. Introduction to the special section on survivability analysis of wireless networks with performance evaluation (VSI-networks survivability). Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005321.

Luxemburk:2023:FGT

[1413] Jan Luxemburk and Tomás Cejka.

Fine-grained TLS services classification with reject option. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005011.

Perdices:2023:WBP

[1414] Daniel Perdices, Jorge E. López de Vergara, Iván González, and Luis de Pedro. Web browsing privacy in the deep learning era: Beyond VPNs and encryption. Computer Networks (Amsterdam, Netherlands:

1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005059.

Ji:2023:AQA

[1415] Xiaolan Ji, Biao Han, Cao Xu, Congxi Song, and Jinshu Su. Adaptive QoS-aware multipath congestion control for live streaming. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005047.

Yu:2023:ESP

[1416] Jiayi Yu, Anfeng Liu, Neal N. Xiong, Shaobo Zhang, Tian Wang, and Mianxiong Dong. Employing social participants for timely data collection using pub/sub solutions in dynamic IoT systems. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005357.

Anonymous:2023:Fa

[1417] Anonymous. February 2023. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBb

[1418] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128623000129.

Nunez-Gomez:2023:HBS

[1419] Carlos Núñez-Gómez, Carmen Carrión, Blanca Caminero, and Francisco M. Delicado. S-HIDRA: a blockchain and SDN domain-based architecture to orchestrate fog computing environments. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005461.

Du:2023:EFL

[1420] Haizhou Du, Yijian Chen, Xiaojie Feng, Qiao Xiang, and Haoyu Liu. An efficient federated learning framework for multi-channeled mobile edge network with layered gradient compression. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005515.

Xu:2023:TTT

[1421] Luyang Xu, Haoyu Liu, Junping Song, Rui Li, Yahui Hu, Xu Zhou, and Paul Patras. TransMUSE: Transferable Traffic Prediction in MUlti-Service Edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005527.

Peng:2023:SPA

[1422] Chengzong Peng, Danyang Zheng, Yihan Zhong, and Xiaojun Cao. Offsite protection against service function forwarder failures in NFV. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005448.

Ibrahim:2023:DPS

[1423] Rita Ibrahim, Mohamad Assaad, and Berna Sayrac. A dynamic policy for selecting D2D mobile relays. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005394.

Salehi:2023:ADC

[1424] Shahab Salehi, Hamed Farbeh, and Alireza Rokhsari. An adaptive data coding scheme for energy consumption reduction in SDN-based Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200562X.

Waqas:2023:DSA

[1425] Muhammad Waqas, Shanshan Tu, Jialin Wan, Talha Mir, Hisham Alasmary, and Ghulam Abbas. Defense scheme against advanced persistent threats in mobile fog computing security. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005539.

Takahashi:2023:SMS

[1426] Tomoki Takahashi, Takehiro Sato, and Eiji Oki. Scheduling model for simultaneous update of multiple service function chains with state consistency. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005540.

Maier:2023:LMS

[1427] Markus Maier and Johanna Ullrich.

In the loop: a measurement study of persistent routing loops on the IPv4/ IPv6 Internet. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005345.

Gong:2023:ESA

[1428] Kai Gong, Dong Yang, Weiting Zhang, and Jie Ren. An efficient scheduling approach for multi-level industrial chain flows in time-sensitive networking. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005503.

Carvalho:2023:MPL

[1429] Rodrigo Carvalho, Noélia Correia, and Faroq Al-Tam. Mobility planning of LoRa gateways for edge storage of IoT data. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005552.

Liu:2023:OMU

[1430] Jiadi Liu, Songtao Guo, Quyuan Wang, Chengsheng Pan, and Li Yang. Optimal multi-user offloading with resources allocation in mobile edge cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005564.

Ramos:2023:WCP

[1431] Javier Ramos, José Luis García-Dorado, and Javier Aracil. Workforce capacity planning for proactive troubleshooting in the Network Operations Center. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005576.

Garrido-Hidalgo:2023:EOR

[1432] Celia Garrido-Hidalgo, Luis Roda-Sanchez, F. Javier Ramírez, Antonio Fernández-Caballero, and Teresa Olivares. Efficient online resource allocation in large-scale LoRaWAN networks: a multi-agent approach. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200559X.

Luglio:2023:FWT

[1433] M. Luglio, M. Quadrini, C. Roseti, and F. Zampognaro. A flexible Web traffic generator for the dimensioning of a 5G backhaul in NPN. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005655.

Roy:2023:APM

[1434] Prasanta Kumar Roy and Ansuman Bhattacharya. An anonymitypreserving mobile user authentication protocol for global roaming services. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005667.

Zhou:2023:SMA

[1435] Boqing Zhou, Sujun Li, Jianxin Wang, Yun Cheng, and Jie Wu. A secure model against mobile sink replication attacks in unattended sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005631. See

corrigendum [1765].

Zhang:2023:RMF

[1436] Shengyu Zhang and Kwan L. Yeung. Revisiting the modulation format selection problem in crosstalk-aware SDM-EONs. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005588.

Pattaranantakul:2023:SFC

[1437] Montida Pattaranantakul, Chalee Vorakulpipat, and Takeshi Takahashi. Service function chaining security survey: Addressing security challenges and threats. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005187.

Ramezanpour:2023:SPV

[1438] Keyvan Ramezanpour, Jithin Jagannath, and Anu Jagannath. Security and privacy vulnerabilities of 5G/ 6G and WiFi 6: Survey and research directions from a coexistence perspective. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005497.

Tuan:2023:UDD

[1439] Tong Anh Tuan, Nguyen Viet Anh, Tran Thi Luong, and Hoang Viet Long. UTL DGA22 — a dataset for DGA botnet detection and classification. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005424.

Giuliano:2023:MMS

[1440] Romeo Giuliano, Anna Maria Vegni, Valeria Loscrí, Eros Innocenti, Alessandro Vizzarri, and Franco Mazzenga. MuSLi: a multi sensor LiDAR detection for C-V2X networks. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005485.

Bao:2023:MMC

[1441] Yu Bao, Yonggang Li, Liang Zhao, Aijuan Zhang, and Yutao Wang. MCEN: Maximum cooperative equilibrium WSN based on greedy prediction to reduce opposite transmission. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005400.

Anonymous:2023:Fb

[1442] Anonymous. February 2023. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBc

[1443] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128623000336.

Yu:2023:CPR

[1444] Mengxin Yu, Yibo Pi, Aimin Tang, and Xudong Wang. Coordinated parallel resource allocation for integrated access and backhaul networks. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005679.

Xie:2023:YFB

[1445] Yi Xie, Xianliang Jiang, Guanghui Gong, Ziyi Jiang, Guang Jin, and Haiming Chen. Yinker: a flexible BBR to achieve the high-throughput and low-latency data transmission over Wi-Fi and 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005643.

Liu:2023:GEB

[1446] Xiuwen Liu and Yanjiao Chen. Group effect-based privacy-preserving data aggregation for mobile crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005412.

Bak:2023:CAW

[1447] Charn doh Bak and Seung-Jae Han.

Contention alleviation in WiFi networks by using light-weight machine learning model. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005680.

Shashin:2023:APS

[1448] Aleksei Shashin, Andrey Belogaev, Artem Krasilov, and Evgeny Khorov. Adaptive parameters selection for uplink grant-free URLLC transmission in 5G systems. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005618.

Seth:2023:DDD

[1449] Abhay Deep Seth, Santosh Biswas, and Amit Kumar Dhar. DAD-CNF: Diagnoser design for duplicate address detection threat using conjunctive normal form. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005734.

Zhou:2023:CFA

[1450] Jianer Zhou, Zengxie Ma, Weijian Tu, Xinyi Qiu, Jingpu Duan, Zhenyu Li, Qing Li, Xinyi Zhang, and Weichao Li. Cable: a framework for accelerating 5G UPF based on eBPF. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005692.

Tian:2023:TSE

[1451] Junfeng Tian, Jin Tian, and Hongwei Xu. TSBFT: a scalable and efficient leaderless Byzantine consensus for consortium blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005758.

Liu:2023:MEM

[1452] Yongbo Liu, Yongqiang Dong, Jun Shen, and Chong Feng. MAYA: Exploring multiform attributes of node to align YANG data models. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005722.

Li:2023:NSS

[1453] David Chunhu Li, Muhamad Rizka Maulana, and Li-Der Chou. NNSplit-SØREN: Supporting the model implementation of large neural networks in a programmable data plane. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005710.

Fu:2023:DCM

[1454] Yanming Fu, Xian Zhang, Xiaoqiong Qin, Qingwen Meng, and Bocheng Huang. Data collection of multiplayer cooperative game based on edge computing in mobile crowd sensing. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005850.

Benlloch-Caballero:2023:DDL

[1455] Pablo Benlloch-Caballero, Qi Wang, and Jose M. Alcaraz Calero. Distributed dual-layer autonomous closed loops for self-protection of 5G/6G IoT networks from distributed denial of service attacks. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005606.

Jha:2023:JRP

[1456] Rakesh Kumar Jha, Mittal K. Pedhadiya, Anutusha Dogra, Haneet Kour, and Puja. Joint resource and power allocation for 5G enabled D2D networking with NOMA. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005709.

Zhang:2023:RES

[1457] Le Zhang and Ye Du. Resilience enhancement scheme for gateway placement in space information networks. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005898.

Zhai:2023:MMS

[1458] Dong Zhai, Xiangru Meng, Zhenhua Yu, Hang Hu, and Yuan Liang. A migration method for service function chain based on failure prediction. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005886.

Augello:2023:CBU

[1459] Alessandro Augello, Pierluigi Gallo, Eleonora Riva Sanseverino, Gaetano Sciabica, and Giuseppe Sciumè. Certifying battery usage for V2G and second life with a blockchain-based framework. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000038.

Nandyala:2023:QQL

[1460] Chandra Sukanya Nandyala, Hee-Won Kim, and Ho-Shin Cho. QTAR: a Q-learning-based topology-aware routing protocol for underwater wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000075.

Hirsch:2023:DDG

[1461] Christian Hirsch, Luca Davoli, Radu Grosu, and Gianluigi Ferrari. Dyn-GATT: a dynamic GATT-based data synchronization protocol for BLE networks. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000051.

Wang:2023:CSA

[1462] Xueyi Wang, Xingwei Wang, Ying Shi, Dongkuo Wu, Lianbo Ma, and Min Huang. Core-selecting auctionbased mechanisms for service function chain provisioning and pricing in NFV markets. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000026.

deNeira:2023:DDS

[1463] Anderson Bergamini de Neira, Burak Kantarci, and Michele Nogueira. Distributed denial of service attack prediction: Challenges, open issues and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005874.

Farquhar:2023:MRF

[1464] Collin Farquhar, Swatantra Kafle, Kian Hamedani, Anu Jagannath, and Jithin Jagannath. Marconi–Rosenblatt framework for intelligent networks (MRiNet Gym): For rapid design and implementation of distributed multiagent reinforcement learning solutions for wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005230.

Yoshinaka:2023:DAL

[1465] Yutaro Yoshinaka, Junji Takemasa, Yuki Koizumi, and Toru Hasegawa.

Design and analysis of lightweight anonymity protocol for host- and ASlevel anonymity. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300004X.

Anonymous:2023:Ma

[1466] Anonymous. March 2023. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBd

[1467] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000634.

Bi:2023:BAB

[1468] Xiang Bi, Xiaokai Sun, Zengwei Lyu, Benhong Zhang, and Xing Wei. A back adjustment based dependent task offloading scheduling algorithm with fairness constraints in VEC networks. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005862.

Monfared:2023:DFD

[1469] Saleh Khalaj Monfared and Saeed Shokrollahi. DARVAN: a fully decentralized anonymous and reliable routing for VANets. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000063.

Sheng:2023:GDB

[1470] Siyuan Sheng, Qun Huang, and Patrick P. C. Lee. A general delta-based in-band network telemetry framework with extremely low bandwidth overhead. Computer Networks (Amsterdam, Netherlands: 1999), 223 (??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300018X.

Ju:2023:COT

[1471] Xiang Ju, Shengchao Su, Chaojie Xu, and Haoxuan Wang. Computation offloading and tasks scheduling for the Internet of Vehicles in edge computing: a deep reinforcement learningbased pointer network approach. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000178.

Xu:2023:CTO

[1472] Dahu Xu and Ding Xu. Cooperative task offloading and resource allocation for UAV-enabled mobile edge computing systems. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000191.

Li:2023:TDC

[1473] Shuyang Li, Gianluca Francini, and Enrico Magli. Temporal dynamics clustering for analyzing cell behavior in mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000233.

Truong:2023:SPO

[1474] Truong Van Truong and Anand Nayyar. System performance and optimization in NOMA mobile edge computing surveillance network using GA and PSO. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000208.

Zhu:2023:OTD

[1475] Konglin Zhu, Wentao Chen, Lei Jiao, Jiaxing Wang, Yuyang Peng, and Lin Zhang. Online training data acquisition for federated learning in cloud-edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000014.

Patel:2023:LLD

[1476] Chintan Patel and Nishant Doshi. LDA-2IoT: a level dependent authentication using two factor for IoT paradigm. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000257.

Aghapour:2023:TOR

[1477] Zahra Aghapour, Saeed Sharifian, and Hassan Taheri. Task offloading and resource allocation algorithm based on deep reinforcement learning for distributed AI execution tasks in IoT edge computing environments. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000221.

Qin:2023:SFC

[1478] Yudong Qin, Deke Guo, Lailong Luo, Jingyu Zhang, and Ming Xu. Service function chain migration with the long-term budget in dynamic networks. Computer Networks (Amsterdam, Netherlands: 1999), 223 (??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000087.

Guler:2023:BEC

[1479] Evrim Guler, Murat Karakus, and Suleyman Uludag. Blockchainenhanced cross-ISP spectrum assignment framework in SDONs: SpectrumChain. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000245.

Hoang:2023:SAC

[1480] Tran Manh Hoang, Le The Dung, Ba Cao Nguyen, Nguyen Van Vinh, and Gia Thien Luu. Secrecy analysis of cooperative NOMA-FDR systems with imperfect CSI and colluding eavesdroppers. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000397.

Shi:2023:AUD

[1481] Yao Shi, Emad Alsusa, and Mohammed W. Baidas. On the application of uplink/downlink decoupled access in heterogeneous mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000385.

Campolo:2023:SIE

[1482] Claudia Campolo, Giacomo Genovese, Gurtaj Singh, and Antonella Molinaro. Scalable and interoperable edge-based federated learning in IoT contexts. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300021X.

Anonymous:2023:Aa

[1483] Anonymous. April 2023. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBe

[1484] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001196.

Laroui:2023:SFC

[1485] Mohammed Laroui, Hatem Ibn Khedher, Hassine Moungla, and Hossam Afifi. Service function chains multiresource orchestration in virtual mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000270.

Chen:2023:AAQ

[1486] Xue Chen, Shiyuan Xu, Yibo Cao, Yunhua He, and Ke Xiao. AQRS: Anti-quantum ring signature scheme for secure epidemic control with blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000403.

Luconi:2023:IFM

[1487] Valerio Luconi and Alessio Vecchio.

Impact of the first months of war on routing and latency in Ukraine. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000415.

Zhou:2023:PLP

[1488] Menglun Zhou, Yifeng Zheng, Songlei Wang, Zhongyun Hua, Hejiao Huang, Yansong Gao, and Xiaohua Jia. PPTA: a location privacy-preserving and flexible task assignment service for spatial crowdsourcing. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000452.

Zhu:2023:IIL

[1489] Wenbin Zhu, Xiuli Ma, Yanliang Jin, and Rui Wang. ILETC: Incremental learning for encrypted traffic classification using generative replay and exemplar. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000476.

Shi:2023:JTP

[1490] KunLi Shi and Fei Wang. Joint task processing/offloading mode selection and resource-allocation for backscatter-aided and wireless-powered MEC. Computer Networks (Amsterdam, Netherlands: 1999), 224 (??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000294.

Bakhshi:2023:MPN

[1491] Bahador Bakhshi, Josep Mangues-Bafalluy, and Jorge Baranda. Multiprovider NFV network service delegation via average reward reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000567.

Yu:2023:LUL

[1492] Sungjin Yu, Joonyoung Lee, Anil Kumar Sutrala, Ashok Kumar Das, and Youngho Park. LAKA-UAV: Lightweight authentication and key agreement scheme for cloud-assisted unmanned aerial vehicle using blockchain in flying ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000579.

Raju:2023:DEA

[1493] Mekala Ratna Raju and Sai Krishna Mothku. Delay and energy aware task scheduling mechanism for fogenabled IoT applications: a reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000488.

Pirmoradian:2023:EEB

[1494] Fatemeh Pirmoradian, Masoumeh Safkhani, and Seyed Mohammad Dakhilalian. ECCPWS: an ECC-based protocol for WBAN systems. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000439.

Huang:2023:EDA

[1495] Ying Huang, Lang Li, Ying Guo, Yu Ou, and Xiantong Huang. An efficient differential analysis method based on deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000671.

Jiang:2023:BBF

[1496] Tongtong Jiang, Guowei Shen, Chun Guo, Yunhe Cui, and Bo Xie. BFLS: Blockchain and federated learning for sharing threat detection models as cyber threat intelligence. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300049X.

Liu:2023:PNSa

[1497] Jiayi Liu, Wenbin Yao, Chen Wang, and Qinghai Yang. Provisioning network slice for mobile content delivery in uncertain MEC environment. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000749.

Arribas:2023:OFC

[1498] Edgar Arribas, Vincenzo Mancuso, and Vicent Cholvi. Optimizing fairness in cellular networks with mobile drone relays. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000683.

Feng:2023:TPP

[1499] Zhenni Feng, Sijia Yu, and Yanmin Zhu. Towards personalized privacy preference aware data trading: a contract theory based approach. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000828.

Peng:2023:DQL

[1500] Bao Peng, Zhi Yao, Xin Liu, and Guofu Zhou. Deep Q-learning multiple networks based dynamic spectrum access with energy harvesting for green cognitive radio network. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000750.

Tokuda:2023:NSR

[1501] Kairi Tokuda, Takehiro Sato, and Eiji Oki. Network slice reconfiguration with deep reinforcement learning under variable number of service function chains. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000816.

Akram:2023:BBP

[1502] Muhammad Arslan Akram, Hira Ahmad, Adnan Noor Mian, Anca Delia Jurcut, and Saru Kumari. Blockchainbased privacy-preserving authentication protocol for UAV networks. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300083X.

Aghasi:2023:DAM

[1503] Ali Aghasi, Kamal Jamshidi, Ali Bohlooli, and Bahman Javadi. A decentralized adaptation of model-free Q-learning for thermal-aware energyefficient virtual machine placement in cloud data centers. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000695.

Attir:2023:LAM

[1504] Azeddine Attir, Farid Naït-Abdesselam, and Kamel Mohamed Faraoun. Lightweight anonymous and mutual authentication scheme for wireless body area networks. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000701.

Brahmi:2023:TFA

[1505] Saloua Brahmi and Mohand Yazid. Towards a fair allocation and effective utilization of resource units in multi-user WLANs-based OFDMA technology. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000841.

Yu:2023:CNO

[1506] Jiachao Yu, Chao Zhai, Yujun Li, and Lina Zheng. Cooperative non-orthogonal multiple access with D2D communication and opportunistic relaying in large-scale cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000762.

Diao:2023:EGE

[1507] Zulong Diao, Gaogang Xie, Xin Wang, Rui Ren, Xuying Meng, Guangxing Zhang, Kun Xie, and Mingyu Qiao. EC-GCN: a encrypted traffic classification framework based on multi-scale graph convolution networks. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000592.

Zhang:2023:DGA

[1508] Hengwei Zhang, Yan Mi, Xiaohu Liu, Yuchen Zhang, Jindong Wang, and Jinglei Tan. A differential game approach for real-time security defense decision in scale-free networks. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000804.

Ahmad:2023:DCN

[1509] Rami Ahmad, Matti Hämäläinen, Raniyah Wazirali, and Tarik Abu-Ain. Digital-care in next generation networks: Requirements and future directions. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000440.

Li:2023:ROT

[1510] Mao Li, Wei Liu, and Jing Lei. A review on orthogonal time-frequency space modulation: State-of-art, hotspots and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000427.

Bothra:2023:HCA

[1511] Priyanka Bothra, Raja Karmakar, Sanjukta Bhattacharya, and Sayantani De. How can applications of blockchain and artificial intelligence improve performance of Internet of Things? — A survey. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000798.

Ali:2023:SHA

[1512] Yasir Ali and Habib Ullah Khan. A survey on harnessing the applications of mobile computing in healthcare during the COVID-19 pandemic: Challenges and solutions. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000506.

Mekdad:2023:SSP

[1513] Yassine Mekdad, Ahmet Aris, Leonardo Babun, Abdeslam El Fergougui, Mauro Conti, Riccardo Lazzeretti, and A. Selcuk Uluagac. A survey on security and privacy issues of UAVs. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000713.

Longo:2023:DIA

[1514] Edoardo Longo and Alessandro E. C.

Redondi. Design and implementation of an advanced MQTT broker for distributed pub/sub scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000464.

Chen:2023:MMS

[1515] Ye Chen, Yingxu Lai, Zhaoyi Zhang, Hanmei Li, and Yuhang Wang. MDFD: a multi-source data fusion detection framework for Sybil attack detection in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000531.

Song:2023:DEA

[1516] Cha Hwan Song, Xin Zhe Khooi, Dinil Mon Divakaran, and Mun Choon Chan. DySO: Enhancing application offload efficiency on programmable switches. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300052X.

Gartner:2023:FIR

[1517] Christoph Gärtner, Amr Rizk, Boris Koldehofe, René Guillaume, Ralf Kundel, and Ralf Steinmetz. Fast incremental reconfiguration of dynamic timesensitive networks at runtime. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000518.

Bruhn:2023:PIT

[1518] Philipp Bruhn, Mirja Kühlewind, and Maciej Muehleisen. Performance and improvements of TCP CUBIC in lowdelay cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000543.

Toledo:2023:CFA

[1519] Kfir Toledo, David Breitgand, Dean Lorenz, and Isaac Keslassy. CloudPilot: Flow acceleration in the cloud. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000555.

Das:2023:CRM

[1520] Sree Krishna Das, Fatma Benkhelifa, Yao Sun, Hanaa Abumarshoud, Qammer H. Abbasi, Muhammad Ali Imran, and Lina Mohjazi. Comprehensive review on ML-based RISenhanced IoT systems: basics, research progress and future challenges. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000269.

Anonymous:2023:Ab

[1521] Anonymous. April 2023. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBf

[1522] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001536.

Jin:2023:ESP

[1523] Nansen Jin, Jinsong Gui, and Xinran Zhou. Equalizing service probability in UAV-assisted wireless powered mmWave networks for postdisaster rescue. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000890.

Chen:2023:BBA

[1524] Hao Chen, Hua Qin, Weimin Chen, Ni Li, Tao Wang, Jianxin He, Gelan Yang, and Yang Peng. BMS: Bandwidth-aware multi-interface scheduling for energy-efficient and delayconstrained gateway-to-device communications in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000907.

Iqbal:2023:UQA

[1525] Faheem Iqbal, Moneeb Gohar, Hanen Karamti, Walid Karamti, Seok-Joo Koh, and Jin-Ghoo Choi. Use of QUIC for AMQP in IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000853.

Hamnache:2023:JLB

[1526] Mohamed Hamnache, Rahim Kacimi, and André-Luc Beylot. Joint loadbalancing and power control strategy to maximize the data extraction rate of LoRaWAN networks. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000786.

Zhou:2023:CPD

[1527] Hongliang Zhou, Yifeng Zheng, Xiaohua Jia, and Jiangang Shu. Collaborative prediction and detection of DDoS attacks in edge computing: a deep learning-based approach with distributed SDN. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000877.

Ullah:2023:ITS

[1528] Farman Ullah, Muhammad Bilal, and Su-Kyung Yoon. Intelligent timeseries forecasting framework for nonlinear dynamic workload and resource prediction in cloud. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000981.

Mabrouk:2023:IDG

[1529] Abdelfettah Mabrouk and Assia Naja.

Intrusion detection game for ubiquitous security in vehicular networks: a signaling game based approach. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000944.

Li:2023:PRA

[1530] Ziyong Li, Yuxiang Hu, Le Tian, and Zhao Lv. Packet rank-aware active queue management for programmable flow scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000774.

Syed:2023:FCB

[1531] Naeem Firdous Syed, Mengmeng Ge, and Zubair Baig. Fog-cloud based intrusion detection system using Recurrent Neural Networks and feature selection for IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300107X.

G:2023:NNR

[1532] Keerthan Kumar T. G., Sourav Kanti Addya, Anurag Satpathy, and Shashidhar G. Koolagudi. NORD: NOde Ranking-based efficient virtual network embedding over single Domain substrate networks. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001068.

Markovitz:2023:DIR

[1533] Oren Markovitz and Michael Segal. Demand Island routing for LEO satellite constellations. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001007.

Deebak:2023:IDA

[1534] B. D. Deebak and Seong Oun Hwang. Intelligent drone-assisted robust lightweight multi-factor authentication for military zone surveillance in the 6G era. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001093.

Goscien:2023:TAS

[1535] Róza Goścień. Traffic-aware service relocation in software-defined and intentbased elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001056.

Valsamas:2023:VTB

[1536] Polychronis Valsamas, Sotiris Skaperas, Lefteris Mamatas, and Luis M. Contreras. Virtualization Technology Blending for resource-efficient edge clouds. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000919.

Hong:2023:ATE

[1537] Hsiang-Jen Hong, Sang-Yoon Chang, and Xiaobo Zhou. Auto-tune: an efficient autonomous multi-path payment routing algorithm for Payment Channel Networks. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001044.

Sun:2023:JSS

[1538] Jiayu Sun, Huiqiang Wang, Lili Nie, Guangsheng Feng, Zhibo Zhang, and Jingyao Liu. A joint strategy for service deployment and task offloading in satellite-terrestrial IoT. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001019.

Carpi:2023:EAR

[1539] Fabrizio Carpi, Marco Martalò, Luca Davoli, Antonio Cilfone, Yingjie Yu, Yi Wang, and Gianluigi Ferrari. Experimental analysis of RSSI-based localization algorithms with NLOS premitigation for IoT applications. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001081.

Chu:2023:ILD

[1540] Yi Chu, Paul Mitchell, David Grace, Jonathan Roberts, Dominic White, and Tautvydas Mickus. IRIS: a low duty cycle cross-layer protocol for long-range wireless sensor networks with low power budget. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001111.

Drainakis:2023:CFL

[1541] Georgios Drainakis, Panagiotis Pantazopoulos, Konstantinos V. Katsaros, Vasilis Sourlas, Angelos Amditis, and Dimitra I. Kaklamani. From centralized to federated learning: Exploring performance and end-to-end resource consumption. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001020.

Wang:2023:GQA

[1542] Hua Wang, Adrián Pérez Aguilar, Almudena Díaz Zayas, Germán Corrales Madueño, Changming Zhang, Nan Hao, and Xianbin Yu. A general QoE assessment framework for applications and services. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000865.

Vidal:2023:USM

[1543] José-Ramón Vidal, Luis Guijarro, and Vicent Pla. A user subscription model in mobile radio access networks with network slicing. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300110X.

Borgohain:2023:LDA

[1544] Ponjit Borgohain and Hiten Choudhury. A lightweight D2D authentication protocol for relay coverage scenario in 5G mobile network. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300124X.

Jiang:2023:SOR

[1545] Jinfang Jiang, Guangjie Han, and Chuan Lin. A survey on opportunistic routing protocols in the Internet of Underwater Things. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001032.

Jo:2023:SAE

[1546] Hyun-Chul Jo, Hyun-Wook Jin, and Joongheon Kim. Self-adaptive endto-end resource management for realtime monitoring in cyber-physical systems. Computer Networks (Amsterdam, Netherlands: 1999), 225 (??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001147.

Malekghaini:2023:DLE

[1547] Navid Malekghaini, Elham Akbari, Mohammad A. Salahuddin, Noura Limam, Raouf Boutaba, Bertrand Mathieu, Stephanie Moteau, and Stephane Tuffin. Deep learning for encrypted traffic classification in the face of data drift: an empirical study. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000932.

Jung:2023:TPO

[1548] Soyi Jung, Jae-Hyun Kim, David Mohaisen, and Joongheon Kim. Truthful and performance-optimal computation outsourcing for aerial surveillance platforms via learning-based auction. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000968.

Dai:2023:GGL

[1549] Jianbang Dai, Xiaolong Xu, and Fu Xiao. GLADS: a global-local attention data selection model for multimodal multitask encrypted traffic classification of IoT. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300097X.

Gupta:2023:FLU

[1550] Rajni Gupta and Juhi Gupta. Federated learning using game strategies: State-of-the-art and future trends. Computer Networks (Amsterdam, Netherlands: 1999), 225 (??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000956.

Anonymous:2023:Mb

[1551] Anonymous. May 2023. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBg

[1552] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300186X.

Akar:2023:PFS

[1553] Nail Akar and Ezhan Karasan. Is proportional fair scheduling suitable for age-sensitive traffic? Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001135.

Hassanpour:2023:PPE

[1554] Seyedeh Bahereh Hassanpour, Ahmad Khonsari, Masoumeh Moradian, and Seyed Pooya Shariatpanahi.

Privacy-preserving edge caching: a probabilistic approach. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000993.

Rengaraj:2023:SUS

[1555] Ramaprabha Rengaraj, Lavanya Dhamodharan Loganathan, and Gunaseelan Konganathan. Secure upgraded spatial modulation for wireless transmission systems. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001251.

Mushtaq:2023:KHA

[1556] Earum Mushtaq, Aneela Zameer, and Rubina Nasir. Knacks of a hybrid anomaly detection model using deep auto-encoder driven gated recurrent unit. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001263.

Zheng:2023:DCT

[1557] Jiaxing Zheng and Li Pan. Dominant coverage for target users at the lowest cost under competitive propagation in social networks. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300138X.

Huang:2023:RFF

[1558] Xiaohong Huang, Lu Han, Dandan Li, Kun Xie, and Yong Zhang. A reliable and fair federated learning mechanism for mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001238.

Singh:2023:OLA

[1559] Prabhjot Singh, Parulpreet Singh, Nitin Mittal, Urvinder Singh, and Supreet Singh. An optimum localization approach using hybrid TSN-MRA in 2D WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001275.

Fazio:2023:NPA

[1560] Peppino Fazio, Miralem Mehic, Miroslav Voznak, Floriano De Rango, and Mauro Tropea. A novel predictive approach for mobility activeness in mobile wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001342.

Alqarni:2023:OBO

[1561] Manal Alqarni, Asma Cherif, and Entisar Alkayyal. ODM-BCSA: an offloading decision-making framework based on binary cuckoo search algorithm for mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000920.

Jeon:2023:HNC

[1562] So-Eun Jeon, Sun-Jin Lee, and Il-Gu Lee. Hybrid in-network computing and distributed learning for large-scale data processing. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001317.

Liu:2023:PNSb

[1563] Xiaoxue Liu, Yichuan Wang, Yanping Li, and Hao Cao. PTAP: a novel secure privacy-preserving and traceable authentication protocol in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000889.

Zeng:2023:SSW

[1564] Zijie Zeng, Lin Cui, Mimi Qian, Zhen Zhang, and Kaimin Wei. A survey on sliding window sketch for network measurement. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300141X.

Castell-Uroz:2023:TLW

[1565] Ismael Castell-Uroz and Pere Barlet-Ros. TrackSign-labeled web tracking dataset. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001329.

Anonymous:2023:Mc

[1566] Anonymous. May 2023. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBh

[1567] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002098.

Hu:2023:WEB

[1568] Xinxin Hu, Hongchang Chen, Shuxin Liu, Haocong Jiang, Kai Wang, and Yahui Wang. Who are the evil backstage manipulators: Boosting graph attention networks against deep fraudsters. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001433.

Ugurlu:2023:APA

[1569] Onur Ugurlu. ADA-PC: an asynchronous distributed algorithm for minimizing pairwise connectivity in wireless multi-hop networks. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001482.

Stokkink:2023:WSA

[1570] Quinten Stokkink, Can Umut Ileri, Dick Epema, and Johan Pouwelse. Web3 Sybil avoidance using network latency. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001469.

Gupta:2023:ABV

[1571] Naveen Kumar Gupta, Rama Shankar Yadav, Rajendra Kumar Nagaria, Deepak Gupta, Achyut Mani Tripathi, and Om Jee Pandey. Anchor-based void detouring routing protocol in three dimensional IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001366.

Guo:2023:AES

[1572] Hui Guo, Rui chang Shi, Ping li Gu, Jia lu Li, and Shu long Wang. Allocating edge service resources to the up-offloaded vehicle tasks in ICV environment. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001603.

Alhamed:2023:PTC

[1573] Faris Alhamed, Davide Scano, Piero Castoldi, Juan Jose Vegas Olmos, Ilya Vershkov, Francesco Paolucci, and Filippo Cugini. P4 Telemetry collector. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300172X.

Zidic:2023:APP

[1574] Dinko Zidic, Toni Mastelic, Ivana Nizetic Kosovic, Mario Cagalj, and Josip Lorincz. Analyses of ping-pong handovers in real 4G telecommunication networks. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001445.

Yuan:2023:AAA

[1575] Jingling Yuan, Nana Wang, Siqi Cai, Mincheng Chen, and Xinping Li. AMLFN-AD: Adaptive multilevel integrated fusion attack detection framework for intelligent building systems. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001457.

Liu:2023:MNM

[1576] Jian Liu, Qingsai Xiao, Liling Xin, Qiuyun Wang, Yepeng Yao, and Zhengwei Jiang. M3F: a novel multisession and multi-protocol based malware traffic fingerprinting. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001688.

Vizziello:2023:IBC

[1577] Anna Vizziello, Maurizio Magarini, Pietro Savazzi, and Laura Galluccio. Intra-body communications for nervous system applications: Current technologies and future directions. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001639.

Zhang:2023:CDN

[1578] Chunrui Zhang, Gang Wang, Shen Wang, Dechen Zhan, and Mingyong Yin. Cross-domain network attack detection enabled by heterogeneous transfer learning. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001378.

Liu:2023:LBI

[1579] Wai-Xi Liu, Jun Cai, Ying-Hao Zhu, Jun-Ming Luo, and Jin Li. Load balancing inside programmable data planes based on network modeling prediction using a GNN with network behaviors. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001408.

Zhang:2023:DSB

[1580] Hang Zhang, Dongqi Yan, Yanxi Zhang, Jiamu Liu, and Mingwu Yao. Distributed synchronization based on model-free reinforcement learning in wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001159.

Lu:2023:NTP

[1581] Yufei Lu, Qian Ning, Linyu Huang, and Bingcai Chen. A network traffic prediction model based on reinforced staged feature interaction and fusion. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001640.

Shao:2023:MTD

[1582] Sujie Shao, Lili Su, Qinghang Zhang, Shuang Wu, Shaoyong Guo, and Feng Qi. Multi task dynamic edgeend computing collaboration for urban Internet of Vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001354.

Ji:2023:RLB

[1583] Pengshuo Ji, Jie Jia, Jian Chen, Liang Guo, An Du, and Xingwei Wang. Reinforcement learning based joint trajectory design and resource allocation for RIS-aided UAV multicast networks. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001421.

Sarah:2023:RAM

[1584] Annisa Sarah, Gianfranco Nencioni, and Md. Muhidul I. Khan. Resource allocation in multi-access edge computing for 5G-and-beyond networks. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001652.

Mathur:2023:SRB

[1585] Shikha Mathur, Anshuman Kalla, Gürkan Gür, Manoj Kumar Bohra, and Madhusanka Liyanage. A survey on role of blockchain for IoT: Applications and technical aspects. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001718.

Myneni:2023:USS

[1586] Sowmya Myneni, Kritshekhar Jha, Abdulhakim Sabur, Garima Agrawal, Yuli Deng, Ankur Chowdhary, and Dijiang Huang. Unraveled — a semisynthetic dataset for advanced persistent threats. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001330.

Anonymous:2023:ISIa

[1587] Anonymous. Introduction to the special issue on machine learning and artificial intelligence for the Internet of Things, 5G, and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001676.

Barakabitze:2023:CSN

[1588] Alcardo Alex Barakabitze and Ray Walshe. Corrigendum to “SDN and NFV for QoE-driven multimedia services delivery: the road towards 6G and beyond networks” COMPNW, Volume 214, September 2022, pp. 1–25, 109133. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001391. See

[1176].

Anonymous:2023:Jb

[1589] Anonymous. June 2023. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBi

[1590] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002347.

Guo:2023:DOR

[1591] Zhongfu Guo, Xinsheng Ji, Wei You, Mingyan Xu, Yu Zhao, Zhimo Cheng, and Deqiang Zhou. Delay optimal for reliability-guaranteed concurrent transmissions with raptor code in multi-access 6G edge network. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001615.

Zhang:2023:DAP

[1592] XiaoXu Zhang, Run hua Shi, WanPeng Guo, PengBo Wang, and WeiYang Ke. A dual auditing protocol for finegrained access control in the edgecloud-based smart home. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001809.

Boudermine:2023:DLB

[1593] Antoine Boudermine, Rida Khatoun, and Jean-Henri Choyer. Dynamic logicbased attack graph for risk assessment in complex computer systems. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001755.

Yang:2023:EID

[1594] Run Yang, Hui He, Yixiao Xu, Bangzhou Xin, Yulong Wang, Yue Qu, and Weizhe Zhang. Efficient intrusion detection toward IoT networks using cloud-edge collaboration. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300169X.

Wu:2023:CFU

[1595] Xiaocan Wu, He Huang, Yang Du, Yu-E Sun, and Shigang Chen. Coupon Filter: a universal and lightweight filter framework for more accurate data stream processing. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001937.

Jiang:2023:ZRM

[1596] Minghao Jiang, Mingxin Cui, Chang Liu, Gaopeng Gou, Gang Xiong, and Zhen Li. Zero-relabelling mobileapp identification over drifted encrypted network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001731.

Sun:2023:MDA

[1597] Yongxia Sun, Weijin Jiang, Ying Yang, Haoran Zhu, and Yirong Jiang. Multidomain authorization and decisionmaking method of access control in the edge environment. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001664.

Quezada:2023:RTB

[1598] Vicente Quezada, Fabian Astudillo-Salinas, Luis Tello-Oquendo, and Paul Bernal. Real-time bot infection detection system using DNS fingerprinting and machine-learning. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001706.

Liu:2023:TCT

[1599] Yaqun Liu, Jun Xie, Changyou Xing, and Shengxu Xie. Topology construction and topology adjustment in flying ad hoc networks for relay transmission. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001986.

Herrera:2023:PAC

[1600] Manuel Herrera, Manu Sasidharan, Stephen Cassidy, and Ajith Kumar Parlikad. Performance assessment of a communication infrastructure with redundant topology: a complex network approach. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001925.

Dursun:2023:GNB

[1601] Yunus Dursun, Muhammed Burak Goktas, and Zhiguo Ding. Green NOMA based MU-MIMO transmission for MEC in 6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001949.

Yang:2023:FSE

[1602] Chen Yang, Gang Xiong, Qing Zhang, Junzheng Shi, Gaopeng Gou, Zhen Li, and Chang Liu. Fewshot encrypted traffic classification via multi-task representation enhanced meta-learning. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001767.

Bringhenti:2023:NAS

[1603] Daniele Bringhenti, Riccardo Sisto, and Fulvio Valenza. A novel abstraction for security configuration in virtual networks. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001901.

Lin:2023:CCT

[1604] Po-Ching Lin, Wen-Hao Hsu, Ying-Dar Lin, Ren-Hung Hwang, Hsiao-Kuang Wu, Yuan-Cheng Lai, and Chung-Kuan Chen. Correlation of cyber threat intelligence with sightings for intelligence assessment and augmentation. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001810.

Nguyen:2023:PRA

[1605] Ba Cao Nguyen, Nguyen Thu Phuong, Xuan Nam Tran, and Pham Thanh Hiep. On performance of RIS-aided ground-to-air and air-to-ground communications in multi-user NOMA systems. Computer Networks (Amsterdam, Netherlands: 1999), 228 (??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001998.

Fei:2023:VEA

[1606] Teng Fei and Wenye Wang. The vulnerability and enhancement of AKA protocol for mobile authentication in LTE/5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001305.

Li:2023:TAE

[1607] Pan Li, Guiyan Liu, Songtao Guo, and Yue Zeng. Traffic-aware efficient consistency update in NFV-enabled software defined networking. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002001.

Zhang:2023:SFT

[1608] Qian Zhang, Shangping Wang, Duo Zhang, Jifang Wang, and Yinjuan Deng. Scalable and faulttolerant selection method of verification and accounting nodes for permissionless blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002025.

Yao:2023:OAM

[1609] Pei Yao, Longkun Guo, Peng Li, and Jiawei Lin. Optimal algorithm for minmax line barrier coverage with mobile sensors on 2-dimensional plane. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001627.

Marinho:2023:CEA

[1610] Ricardo Pagoto Marinho, Luiz F. M.

Vieira, Marcos A. M. Vieira, and Antonio A. F. Loureiro. CAIN: an energy-aware and intelligent increasing coverage area routing protocol for future 6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001780.

Roy:2023:GBR

[1611] Debashri Roy, Batool Salehi, Stella Banou, Subhramoy Mohanti, Guillem Reus-Muns, Mauro Belgiovine, Prashant Ganesh, Chris Dick, and Kaushik Chowdhury. Going beyond RF: a survey on how AI-enabled multimodal beamforming will shape the NextG standard. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001743.

Hasslinger:2023:OAM

[1612] Gerhard Hasslinger, Mahshid Okhovatzadeh, Konstantinos Ntougias, Frank Hasslinger, and Oliver Hohlfeld. An overview of analysis methods and evaluation results for caching strategies. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000282.

Harms:2023:GLC

[1613] Laura Harms, Christian Richter, and Olaf Landsiedel. Grace: Lowcost time-synchronized GPIO tracing for IoT testbeds. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001913.

Garg:2023:DRL

[1614] Sahil Garg, Jia Hu, Giancarlo Fortino, Laurence T. Yang, Mohsen Guizani, Xianjun Deng, and Danda B. Rawat. Deep reinforcement learning for nextgeneration IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002050.

Anonymous:2023:Jc

[1615] Anonymous. June 2023. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBj

[1616] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002645.

Jiang:2023:WGP

[1617] Jianguo Jiang, Shang Jiang, Yi Liu, Siye Wang, Yanfang Zhang, Yue Feng, and Ziwen Cao. Wi-Gait: Pushing the limits of robust passive personnel identification using Wi-Fi signals. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001962.

Li:2023:LCH

[1618] Junfeng Li, Dan Li, Huiyou Jiang, Du Lin, Jinkun Geng, Yukai Huang, K. K. Ramakrishnan, and Kai Zheng. Light: a compatible, highperformance and scalable user-level network stack. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002013.

Seyoum:2023:DLB

[1619] Yemane Teklay Seyoum, Syed Maaz Shahid, Eun Seon Cho, and Sungoh Kwon. Distributed load balancing algorithm considering QoS for next generation multi-RAT HetNets. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002037.

Ali:2023:RPB

[1620] Zainab H. Ali, Noha A. Sakr, Nora El-Rashidy, and Hesham A. Ali. A reliable position-based routing scheme for controlling excessive data dissemination in vehicular ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300230X.

Lau:2023:GOP

[1621] Wei Jian Lau, Joanne Mun-Yee Lim, Chun Yong Chong, Nee Shen Ho, and Thomas Wei Min Ooi. General outage probability model for UAV-to-UAV links in multi-UAV networks. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001974.

Yang:2023:DJD

[1622] Cheng Yang, Yangming Zhao, Gongming Zhao, and Hongli Xu. DFS: Joint data formatting and sparsification for efficient communication in Distributed Machine Learning. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002220.

Lastovicka:2023:POS

[1623] Martin Lastovicka, Martin Husák, Petr Velan, Tomás Jirsík, and Pavel Celeda. Passive operating system fingerprinting revisited: Evaluation and current challenges. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300227X.

Rivas:2023:RIG

[1624] Angel Esteban Labrador Rivas and Taufik Abrão. Raptor-IRSA grantfree random access protocol for smart grids applications. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002207.

Wang:2023:TRM

[1625] Jian Wang, Yuping Yan, and Guosheng Zhao. Task recommendation method combining multimodal cognition and collaboration in mobile crowdsensing systems. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002414.

Chkirbene:2023:DIL

[1626] Zina Chkirbene, Ridha Hamila, Aiman Erbad, Serkan Kiranyaz, and Nasser Al-Emadi. D2DLive: Iterative live video streaming algorithm for D2D networks. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001792.

Mohammadi:2023:REE

[1627] Razieh Mohammadi and Zahra Shirmohammadi. RLS^{2}: an energy efficient reinforcement learning-based sleep scheduling for energy harvesting WBANs. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002268.

Cheng:2023:SDR

[1628] Zhimo Cheng, Xinsheng Ji, Wei You, Yu Zhao, and Zhongfu Guo. SE-CHO: a deep reinforcement learningbased scheme for secure handover in mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002141.

Khalil:2023:DLB

[1629] Alvi Ataur Khalil, Mohamed Y. Selim, and Mohammad Ashiqur Rahman. Deep learning-based energy harvesting with intelligent deployment of RIS-assisted UAV-CFmMIMOs. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002293.

Nande:2023:QET

[1630] Swaraj Shekhar Nande, Marius Paul, Stefan Senk, Marian Ulbricht, Riccardo Bassoli, Frank H. P. Fitzek, and Holger Boche. Quantum enhanced time synchronisation for communication network. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002177.

Hu:2023:JTR

[1631] Bo Hu, Liangyu Chen, and Shanzhi Chen. Joint trajectory-resource optimization for UAV-enabled uplink communication networks with wireless backhaul. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002244.

Zhou:2023:FSW

[1632] Qiang Zhou, Liangmin Wang, Huijuan Zhu, and Tong Lu. Fewshot website fingerprinting attack with cluster adaptation. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002256.

Javanmardi:2023:SPI

[1633] Saeed Javanmardi, Mohammad Shojafar, Reza Mohammadi, Mamoun Alazab, and Antonio M. Caruso. An SDN perspective IoT-Fog security: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001779.

Kress:2023:COS

[1634] Fabian Kreß, Vladimir Sidorenko, Patrick Schmidt, Julian Hoefer, Tim Hotfilter, Iris Walter, Tanja Harbaum, and Jürgen Becker. CN-NParted: an open source framework for efficient Convolutional Neural Network inference partitioning in embedded systems. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002049.

Almstedt:2023:CRA

[1635] Lennart Almstedt, Kai Bleeke, Mohammad Mahhouk, Leander Jehl, Rüdiger Kapitza, and Lars Wolf. ContractBox: Realizing accountable data sharing on the edge using a small scale blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300213X.

Anonymous:2023:Jd

[1636] Anonymous. July 2023. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBk

[1637] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002773.

Kim:2023:RRM

[1638] Sungwook Kim. Radio resource management scheme in radar and communication spectral coexistence platform. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128623002189.

Kurt:2023:SRA

[1639] Isil Kurt, Fatih Alagöz, and Mete Akgün. Scalable RFID authentication protocol based on physically unclonable functions. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002384.

Garroppo:2023:DMI

[1640] Rosario Giuseppe Garroppo and Maria Grazia Scutellà. Design model of an IEEE 802.11ad infrastructure for TSN-based industrial applications. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002165.

Shu:2023:NCM

[1641] Zhaogang Shu, Haoxian Feng, Tarik Taleb, and Zhifang Zhang. A novel combinatorial multi-armed bandit game to identify online the changing top-K flows in softwaredefined networks. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002281.

Fotiou:2023:SVC

[1642] Nikos Fotiou, Yannis Thomas, Vasilios A. Siris, George Xylomenos, and George C. Polyzos. Selfverifiable content using decentralized identifiers. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300244X.

Tang:2023:RRB

[1643] Tong Tang, Yingxu Lai, and Yipeng Wang. Relational reasoning-based approach for network protocol reverse engineering. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002426.

Li:2023:DEC

[1644] Pengmiao Li, Yuchao Zhang, Huahai Zhang, Wendong Wang, Ke Xu, and Zhili Zhang. A delayed eviction caching replacement strategy with unified standard for edge servers. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002396.

Kayathri:2023:SDL

[1645] T. L. Kayathri, N. Kumaresan, and R. Vijayabhasker. SDBGPChain: a decentralized low complexity framework to detect and prevent the BGPattacks using SDN with smart contract based Dendrimer tree blockchain. Computer Networks (Amsterdam, Netherlands:

1999), 230(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002451.

Kowalski:2023:TMR

[1646] Mikołaj Kowalski and Wojciech Mazurczyk. Toward the mutual routing security in wide area networks: a scoping review of current threats and countermeasures. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002232.

Mazloum:2023:SRT

[1647] Ali Mazloum, Elie Kfoury, Jose Gomez, and Jorge Crichigno. A survey on rerouting techniques with P4 programmable data plane switches. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002402.

Qin:2023:HCD

[1648] Xingsheng Qin, Frank Jiang, Mingcan Cen, and Robin Doss. Hybrid cyber defense strategies using Honey-X: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002219.

Anonymous:2023:ISIb

[1649] Anonymous. Introduction to special issue on selected papers from IFIP networking conference 2022. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002517.

Cerquitelli:2023:MLE

[1650] Tania Cerquitelli, Michela Meo, Marilia Curado, Lea Skorin-Kapov, and Eirini Eleni Tsiropoulou. Machine learning empowered computer networks. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002529.

Anonymous:2023:Je

[1651] Anonymous. July 2023. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBl

[1652] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003043.

Zhao:2023:AMP

[1653] Zhe Zhao, Lu Gao, Weitao Pan, Zhiliang Qiu, Xu Guo, Ya Gao, and Ling Zheng. Access mechanism for period flows of non-deterministic end systems for time-sensitive networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002505.

Yin:2023:JDR

[1654] Yabo Yin, Chuanhe Huang, Neal N. Xiong, Dong-Fang Wu, and Shidong Huang. Joint dynamic routing and resource allocation in satelliteterrestrial integrated networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002682.

Bozkaya:2023:DTA

[1655] Elif Bozkaya. Digital twin-assisted and mobility-aware service migration in Mobile Edge Computing. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002438.

Troia:2023:PCP

[1656] Sebastian Troia, Marco Savi, Giulia Nava, Ligia Maria Moreira Zorello, Thomas Schneider, and Guido Maier. Performance characterization and profiling of chained CPU-bound Virtual Network Functions. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002608.

Milanez:2023:VVB

[1657] Guilherme A. S. Milanez, Marcos A. M. Vieira, Luiz F. M. Vieira, and José Augusto M. Nacif. VariBan: a variable bandwidth channel allocation algorithm for IEEE 802.15.4ebased networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002190.

Yang:2023:MME

[1658] Huan Yang, Sheng Sun, Min Liu, Qiuping Zhang, and Yuwei Wang. MJOA-MU: End-to-edge collaborative computation for DNN inference based on model uploading. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002463.

Wang:2023:IFB

[1659] Yang Wang, Yawen Chen, Zhaoming Lu, and Xiangming Wen. Intelligent flying-beamformer for hybrid mmWave systems: a deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002554.

Leon:2023:MLB

[1660] Juan Pablo Astudillo León, Luis J. de la Cruz Llopis, and Francisco J. Rico-Novella. A machine learning based Distributed Congestion Control Protocol for multi-hop wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300258X.

Almasaeid:2023:MCS

[1661] Hisham M. Almasaeid. Minimum cost spectrum allocation with QoS guarantees in multi-interface multi-hop dynamic spectrum access networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002694.

Sahin:2023:CCP

[1662] Mehmet Emin Sahin and Mehmet Demirci. ConPoolUBF: Connection pooling and updatable Bloom filter based SYN flood defense in programmable data planes. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002475.

Wu:2023:PCP

[1663] Hua Wu, Ziling Shao, Fuhao Yang, Guang Cheng, Xiaoyan Hu, Jing Ren, and Wei Wang. PD-CPS: a practical scheme for detecting covert port scans in high-speed networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002700.

Moorthy:2023:SUN

[1664] Sabarish Krishna Moorthy, Nicholas Mastronarde, Scott Pudlewski, Elizabeth Serena Bentley, and Zhangyu Guan. Swarm UAV networking with collaborative beamforming and automated ESN learning in the presence of unknown blockages. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002499.

Zhu:2023:RFD

[1665] Mengfei Zhu and Eiji Oki. Robust function deployment against uncertain recovery time in different protection types with workloaddependent failure probability. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002712.

Baena:2023:MEK

[1666] Carlos Baena, O. S. Peñaherrera-Pulla, Raquel Barco, and Sergio Fortes. Measuring and estimating Key Quality Indicators in Cloud Gaming services. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002530.

Attarian:2023:EWF

[1667] Reyhane Attarian and Alireza Keshavarz-Haddad. Effective website fingerprinting attack based on the first packet direction only. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002566.

Birabwa:2023:MAD

[1668] Denise Joanitah Birabwa, Daniel Ramotsoela, and Neco Ventura. Multiagent deep reinforcement learning for user association and resource allocation in integrated terrestrial and non-terrestrial networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002724.

Ma:2023:UAC

[1669] Bo Ma, Heng Kuang, Shiyuan Liu, and Chuanhuang Li. UAV assisted cellular network traffic offloading: Joint swarm, 3D deployment, and user allocation optimization based on a data-aware method. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002578.

Hu:2023:MTC

[1670] Chia-Cheng Hu, Wen-Wu Liu, and Jeng-Shyang Pan. Minimizing traffic cost of content distribution and storage allocation in cloud radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002815.

Li:2023:MAA

[1671] Peng Li, Junzuo Lai, Dehua Zhou, Ye Yang, Wei Wu, and Junbin Fang. Multi-authority anonymous authentication with public accountability for incentive-based applications. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002736.

Alnajar:2023:TIF

[1672] Omar Alnajar and Ahmed Barnawi. Tactile internet of federated things: Toward fine-grained design of FL-based architecture to meet TIoT demands. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001573.

Alatram:2023:DMI

[1673] Alaa Alatram, Leslie F. Sikos, Mike Johnstone, Patryk Szewczyk, and James Jin Kang. DoS/DDoS-MQTT-IoT: a dataset for evaluating intrusions in IoT networks using the MQTT protocol. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002542.

Chang:2023:TNA

[1674] Hyunseok Chang, Fang Hao, Murali Kodialam, T. V. Lakshman, Sarit Mukherjee, and Matteo Varvello. Towards network-assisted publish– subscribe over wide area networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128623001470.

Anonymous:2023:Ac

[1675] Anonymous. August 2023. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBm

[1676] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003316.

Mu:2023:DRL

[1677] Liting Mu, Bin Ge, Chenxing Xia, and Cai Wu. Deep reinforcement learning based adaptive threshold multi-tasks offloading approach in MEC. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002487.

Wang:2023:ECR

[1678] You-Chiun Wang and Chia-Wei Chou.

Efficient coordination of radio frames to mitigate cross-link interference in 5G D-TDD systems. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002852.

Garcia-Martin:2023:ECA

[1679] Juan Pablo García-Martín and Antonio Torralba. Energy consumption analytical modeling of NB-IoT devices for diverse IoT applications. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003006.

Biswas:2023:MRP

[1680] Kamanashis Biswas, Vallipuram Muthukkumarasamy, Mohammad Jabed Morshed Chowdhury, Xin-Wen Wu, and Kalvinder Singh. A multipath routing protocol for secure energy efficient communication in Wireless Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002876.

Zhao:2023:FSL

[1681] Guozheng Zhao, Kaiqiang Lin, and Tong Hao. A feasibility study of LoRaWAN-based wireless underground sensor networks for underground monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 232 (??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002967.

Jung:2023:OFS

[1682] Sehun Jung and Hyang-Won Lee. Optimization framework for splitting DNN inference jobs over computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002591.

Zeba:2023:MPH

[1683] Ousmane Zeba, Shunsuke Saruwatari, and Takashi Watanabe. MAC protocols for high capacity and increased reliability in bistatic backscatter networks for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002153.

Duong:2023:CMA

[1684] Tho Minh Duong, Seung-Ho Hyun, and Sungoh Kwon. Channel modeling and achievable data rate for 1-D molecular communication in bounded environments. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002979.

Dung:2023:EED

[1685] Le The Dung and Taejoon Kim. Exploring the effects of directional antennas on the secure connectivity and hop count of secure multi-hop adhoc wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002906.

Chen:2023:ASM

[1686] Jie Chen, Jiaxu Niu, Hao Lei, Li Lin, and Yunhao Ling. Adaptively secure multi-authority attribute-based broadcast encryption in fog computing. Computer Networks (Amsterdam, Netherlands: 1999), 232 (??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300289X.

Wang:2023:CCB

[1687] Zhujun Wang, Lejun Zhang, Ran Guo, Guopeng Wang, Jing Qiu, Shen Su, Yuan Liu, Guangxia Xu, and Zhihong Tian. A covert channel over blockchain based on label tree without long waiting times. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002888.

Abubakar:2023:CTA

[1688] Attai Ibrahim Abubakar, Michael S. Mollel, Oluwakayode Onireti, Metin Ozturk, Iftikhar Ahmad, Syed Muhammad Asad, Yusuf Sambo, Ahmed Zoha, Sajjad Hussain, and Muhammad Ali Imran. Coverage and throughput analysis of an energy efficient UAV base station positioning scheme. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002992.

Dao:2023:ONI

[1689] Thi-Nga Dao, Duc Van Le, and Xuan Nam Tran. Optimal network intrusion detection assignment in multi-level IoT systems. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002918.

Ali:2023:SUV

[1690] Mohammad Furqan Ali and Dushantha Nalin K. Jayakody. SIMO-Underwater Visible Light Communication (UVLC) system. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001950. See

corrigendum [1764].

Guda:2023:AAO

[1691] Sruthi Guda and Sreenivasa Rao Duggirala. Adaptive Aquila optimized relay selection and power allocation in CR-NOMA model. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002931.

Valente:2023:OBC

[1692] Francesco Valente, Vincenzo Eramo, and Francesco G. Lavacca. Optimal bandwidth and computing resource allocation in low earth orbit satellite constellation for earth observation applications. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002943.

Yigit:2023:NFT

[1693] Beytüllah Yigit, Gürkan Gür, Fatih Alagöz, and Bernhard Tellenbach. Network fingerprinting via timing attacks and defense in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002955.

Dong:2023:EEE

[1694] Yongpeng Dong, Shiyou Qian, Wanghua Shi, Jian Cao, and Guangtao Xue. EEM: an elastic event matching framework for content-based publish/subscribe systems. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002827.

Anonymous:2023:S

[1695] Anonymous. September 2023. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBn

[1696] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128623003742.

Ren:2023:CAD

[1697] Xiangyu Ren, Lin Cai, Pu Yang, and Jiequ Ji. Congestion-aware delayguaranteed scheduling and routing with renewal optimization. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003080.

Zhang:2023:CRT

[1698] Menghan Zhang, Xianliang Jiang, Guang Jin, Penghui Li, and Haiming Chen. CapRadar: Real-time adaptive bandwidth prediction for dynamic wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003109.

Bagnulo:2023:DIV

[1699] Marcelo Bagnulo, Alberto García-Martínez, Anna Maria Mandalari, Praveen Balasubramanian, Daniel Havey, and Gabriel Montenegro. Design, implementation and validation of a receiver-driven less-than-best-effort transport. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002864.

Zhou:2023:EPL

[1700] Yuwen Zhou, Bangbang Ren, Junjie Xie, Lailong Luo, Deke Guo, and Xiaobo Zhou. Enable the proactively load-balanced control plane for SDN via intelligent switch-to-controller selection strategy. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003122.

Zhang:2023:EES

[1701] Shengyu Zhang and Kwan L. Yeung. Efficient embedding of service function chains in space-division multiplexing elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003146.

Alonso:2023:WSP

[1702] Rodney Martinez Alonso, David Plets, Luc Martens, Wout Joseph, Ernesto Fontes Pupo, and Glauco Guillen Nieto. White spaces pattern finding and inference based on machine learning for multi-frequency spectrum footprints. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300316X.

Ruiz-Villafranca:2023:MII

[1703] Sergio Ruiz-Villafranca, José Roldán-Gómez, Javier Carrillo-Mondéjar, Juan Manuel Castelo Gómez, and José Miguel Villalón. A MEC-IIoT intelligent threat detector based on machine learning boosted tree algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003134.

Lyu:2023:PPE

[1704] Minzhao Lyu, Hassan Habibi Gharakheili, and Vijay Sivaraman. PEDDA: Practical and effective detection of distributed attacks on enterprise networks via progressive multi-stage inference. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003183.

Yao:2023:UFS

[1705] Mingwu Yao, Jiamu Liu, Jing Du, Dongqi Yan, Yanxi Zhang, Wei Liu, and Anthony Man-Cho So. A unified flow scheduling method for time sensitive networks. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300292X.

Dogan:2023:SSO

[1706] Orkun Dogan and Kübra Kalkan. SOREC: Self-organizing and resource efficient clustered blockchain network. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003389.

Wong:2023:FGH

[1707] KaKei Wong and Lin Cui. Finegrained HTTP/3 prioritization via reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003250.

Fang:2023:RRA

[1708] Jin Fang, Gongming Zhao, Hongli Xu, Huaqing Tu, and Haibo Wang. Reveal: Robustness-aware VNF placement and request scheduling in edge clouds. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003274.

Li:2023:DAW

[1709] Yinlong Li, Siyao Cheng, Hao Zhang, and Jie Liu. Dynamic adaptive workload offloading strategy in mobile edge computing networks. Computer Networks (Amsterdam, Netherlands:

1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003237.

Lekssays:2023:MBB

[1710] Ahmed Lekssays, Barbara Carminati, and Elena Ferrari. Mal-Con: a blockchain-based malware containment framework for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002980.

Li:2023:PEC

[1711] Chen Li, Xiaoyu Wang, Tongyu Zong, Houwei Cao, and Yong Liu. Predictive edge caching through deep mining of sequential patterns in user content retrievals. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003110.

Wang:2023:TUS

[1712] Jian Wang, Jia Liu, and Guosheng Zhao. Trusted user selection for fusion of multimodal cognition in selforganizing pattern of mobile crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003481.

Naqvi:2023:IID

[1713] Haidlir Achmad Naqvi, Muhammad Hafizhuddin Hilman, and Bayu Anggorojati. Implementability improvement of deep reinforcement learning based congestion control in cellular network. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003195.

Satoh:2023:AFT

[1714] Daisuke Satoh. Analysis for flow termination in the decentralized and centralized pre-congestion notification architecture. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003213.

Jiang:2023:PPS

[1715] Wanchun Jiang, Renfu Yao, Kaiqin Liao, Yulong Yan, Jiawei Huang, Weiping Wang, and Jianxin Wang. Practical periodic strategy for 40/100 Gbps energy efficient Ethernet. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003468.

Megias:2023:EBT

[1716] Carlos Megías, Víctor Vázquez, Eduardo Ros, Mauro Cappelli, and Javier Díaz. Ethernet-based timing system for accelerator facilities: the IFMIF-DONES case. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003420.

Liu:2023:SSS

[1717] Jing Liu, Jianguo Xie, Biao Liu, Liwei Xu, Xiaoming Chen, and Huamin Feng. SED-SGC: a scalable, efficient, and distributed secure group communication scheme based on superlattice PUF. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003432.

Oliveira:2023:TGM

[1718] Afonso Oliveira and Teresa Vazão. Towards green machine learning for resource allocation in beyond 5G RAN slicing. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003225.

Sedrati:2023:IGS

[1719] Anass Sedrati, Abdellatif Mezrioui, and Aafaf Ouaddah. IoT-Gov: a structured framework for Internet of Things governance. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300347X.

Liu:2023:DFF

[1720] Zunyi Liu, Dian Shen, Jiaang Bao, Fang Dong, and Jiong You. Deep-MetricCorr: Fast flow correlation for data center networks with deep metric learning. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003493.

Amador:2023:SIP

[1721] Oscar Amador, Ignacio Soto, Maria Calderon, and Manuel Urueña. Studying and improving the performance of ETSI ITS contention-based forwarding (CBF) in urban and highway scenarios: S-FoT+. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003444.

Li:2023:PBM

[1722] Xinhao Li, Yingjie Wang, Bingyi Xie, Lingkang Meng, Zhaowei Liu, Xiangrong Tong, Ao Zhou, and Zhipeng Cai. Planning-based mobile crowdsourcing bidirectional multi-stage online task assignment. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003249.

Parra-Espin:2023:SBA

[1723] José Antonio Parra-Espín, Rafael Marín-López, Gabriel López-Millán, Fernando Pereñíguez-García, and Oscar Canovas. SDN-based automated rekey of IPsec security associations: Design and practical validations. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300350X.

Damsgaard:2023:ACB

[1724] Hans Jakob Damsgaard, Aleksandr Ometov, Md Munjure Mowla, Adam Flizikowski, and Jari Nurmi. Approximate computing in B5G and 6G wireless systems: a survey and future outlook. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003171.

Li:2023:CSD

[1725] Qing Li, He Huang, Ruoyu Li, Jianhui Lv, Zhenhui Yuan, Lianbo Ma, Yi Han, and Yong Jiang. A comprehensive survey on DDoS defense systems: New trends and challenges. Computer Networks (Amsterdam, Netherlands:

1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003407.

N:2023:SSC

[1726] Pruthvi C. N., H. S. Vimala, and Shreyas J. A systematic survey on content caching in ICN and ICN-IoT: Challenges, approaches and strategies. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003419.

Anonymous:2023:O

[1727] Anonymous. October 2023. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBo

[1728] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300419X.

Sharara:2023:MEC

[1729] Mahdi Sharara, Francesca Fossati, Sahar Hoteit, Véronique Vèque, and Francesca Bassi. Minimizing energy consumption by joint radio and computing resource allocation in Cloud-RAN. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003158.

Cao:2023:RLB

[1730] Shaohua Cao, Di Liu, Congcong Dai, Chengqi Wang, Yansheng Yang, Weishan Zhang, and Danyang Zheng. Reinforcement learning based tasks offloading in vehicular edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003390.

Talau:2023:ITP

[1731] Marcos Talau, Thiago A. Herek, Mauro Fonseca, and Emilio C. G. Wille. Improving TCP performance over a common IoT scenario using the Early Window Tailoring method. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003201.

Xu:2023:LAD

[1732] Chenglin Xu, Cheng Xu, Bo Li, Siqi Li, and Tao Li. Load-aware dynamic controller placement based on deep reinforcement learning in SDNenabled mobile cloud-edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003456.

Rezaei:2023:ESE

[1733] Zahra Rezaei and Behrouz Shahgholi Ghahfarokhi. Energy and spectrum efficient cell switch-off with channel and power allocation in ultradense networks: a deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003572.

Liu:2023:ERT

[1734] Libin Liu, Jingzong Li, Hong Xu, Kaiwen Xue, and Jason Chun Xue. Efficient real-time video conferencing with adaptive frame delivery. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003638.

Wang:2023:MCE

[1735] Yang Wang, Qian Hu, Long Nguyen, and Maryam Jalalitabar. Minimumcost embedding of virtual networks: an iterative decomposition approach. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003523.

Liu:2023:SDN

[1736] Jiayi Liu, Xiangjie Shi, Yizhi Huang, and Qinghai Yang. Selective and ondemand network measurement with SRv6 and INT. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003596.

Fu:2023:PPM

[1737] Yanming Fu, Bocheng Huang, Xiao Liu, Jiayuan Chen, and Shenglin Lu. Privacy-preserving mobile crowd sensing task assignment with Stackelberg game. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003626.

Zeng:2023:JOM

[1738] Yaoping Zeng, Dongyang Lu, and Jianbo Du. Joint optimized multiuser access and UAV deployments based on heterogeneous revenue in IoT network. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300364X.

Allegretta:2023:CSC

[1739] Mauro Allegretta, Giuseppe Siracusano, Roberto Gonzalez, and Marco Gramaglia. Are crowd-sourced CTI datasets ready for supporting anticybercrime intelligence? Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003651.

K:2023:KSI

[1740] Venkatakrishna Rao. K. and C. Ravindranath Chowdary. K++ Shell: Influence maximization in multilayer networks using community detection. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003614.

Chen:2023:MMC

[1741] Luyao Chen, Yong Tang, Jingwen Xia, Siyang Chen, Chengyu Zheng, Hai Lin, and Wenyong Wang. Multi-MEC collaboration for VR video transmission: Architecture and cache algorithm design. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003092.

Chen:2023:OAM

[1742] Tingjun Chen, Prasanthi Maddala, Panagiotis Skrimponis, Jakub Kolodziejski, Abhishek Adhikari, Hang Hu, Zhihui Gao, Arun Paidimarri, Alberto Valdes-Garcia, Myung Lee, Sundeep Rangan, Gil Zussman, and Ivan Seskar. Open-access millimeter-wave software-defined radios in the PAWR COSMOS testbed: Design, deployment, and experimentation. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003675.

Yu:2023:CJA

[1743] Kan Yu, Jiguo Yu, Zhiyong Feng, and Min Guo. Cooperative jamming aided securing wireless communications without CSI of eavesdroppers. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003511.

Fernandez-Veiga:2023:IRS

[1744] Manuel Fernández-Veiga, M. E. Sousa-Vieira, Ana Fernández-Vilas, and Rebeca P. Díaz-Redondo. Irregular repetition slotted Aloha with multiuser detection: a density evolution analysis. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003663.

Kar:2023:DPD

[1745] Snigdhaswin Kar, Prabodh Mishra, and Kuang-Ching Wang. Dynamic packet duplication for reliable low latency communication under mobility in 5G NR-DC networks. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003687.

Boutiba:2023:ORR

[1746] Karim Boutiba, Miloud Bagaa, and Adlen Ksentini. Optimal radio resource management in 5G NR featuring network slicing. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003821.

Li:2023:EET

[1747] Bin Li, Wenshuai Liu, Wancheng Xie, and Xiaohui Li. Energy-efficient task offloading and trajectory planning in UAV-enabled mobile edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003857.

Zhu:2023:DEE

[1748] Rongxin Zhu, Azzedine Boukerche, Xiangdang Huang, and Qiuling Yang. DESLR: Energy-efficient and secure layered routing based on channel-aware trust model for UASNs. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003845.

Antevski:2023:ABC

[1749] Kiril Antevski and Carlos J. Bernardos. Applying blockchain consensus mechanisms to network service federation: Analysis and performance evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003584.

Minmin:2023:TEC

[1750] Huang Minmin, Yuan Lingyun, Pan Xue, and Zhou Chuan. Trusted edge and cross-domain privacy enhancement model under multi-blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003262.

Onsu:2023:HCM

[1751] Murat Arda Onsu, Burak Kantarci, and Azzedine Boukerche. How to cope with malicious federated learning clients: an unsupervised learningbased approach. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003833.

Qu:2023:ETO

[1752] Xiaofeng Qu and Huiqiang Wang. Emergency task offloading strategy based on cloud-edge-end collaboration for smart factories. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003602.

Shen:2023:JTO

[1753] Bo Shen, Qi Gu, and Gang Yang. Joint task offloading and UAVs deployment for UAV-assisted mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003882.

Phu:2023:DSA

[1754] Anh Tuan Phu, Bo Li, Faheem Ullah, Tanvir Ul Huque, Ranesh Naha, Muhammad Ali Babar, and Hung Nguyen. Defending SDN against packet injection attacks using deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003808.

Luo:2023:IAD

[1755] Feng Luo, Zitong Wang, and Baoyin Zhang. Impact analysis and detection of time-delay attacks in time-sensitive networking. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300381X.

Su:2023:IDL

[1756] Lina Su and Zongpeng Li. Incentivedriven long-term optimization for hierarchical federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003894.

Qu:2023:HNS

[1757] Lianwei Qu, Yong Wang, Jing Yang, and Meng Zhao. A heterogeneous network structure publishing security framework based on cloud-edge collaboration. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003924.

Park:2023:TTC

[1758] Kibeom Park, Sangmo Sung, Hokeun Kim, and Jae il Jung. Technology trends and challenges in SDN and service assurance for end-to-end network slicing. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003535.

Ergun:2023:SHN

[1759] Serap Ergun, Ibrahim Sammour, and Gerard Chalhoub. A survey on how network simulators serve reinforcement learning in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003791.

TedT:2023:HDI

[1760] Ted “Taekyoung” Kwon, Junghwan Song, Heeyoung Jung, Selin Chun, Hyunwoo Lee, Minhyeok Kang, Minkyung Park, and Eunsang Cho. How to decentralize the Internet: a focus on data consolidation and user privacy. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003560.

Nguyen:2023:OAC

[1761] Tien-Tung Nguyen and Anh-Vinh Nguyen. Outage analysis of cognitive inspired NOMA networks in the presence of imperfect SIC, CCI, and non-ID fading channels. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003547.

Moure-Garrido:2023:RTD

[1762] Marta Moure-Garrido, Celeste Campo, and Carlos Garcia-Rubio. Real time detection of malicious DoH traffic using statistical analysis. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003559.

Manzoni:2023:EPS

[1763] Pietro Manzoni, Claudio Palazzi, Flavia Delicato, Erika Rosas, and Spyridon Mastorakis. Editorial: Pub/ sub solutions for interoperable and dynamic IoT systems. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003705.

Ali:2023:CSU

[1764] Mohammad Furqan Ali and Dushantha Nalin K. Jayakody. Corrigendum to “SIMO-Underwater Visible Light Communication (UVLC) system” Computer Networks. 2023 May 9:109750. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003870. See

[1690].

Zhou:2023:CSM

[1765] Boqing Zhou, Sujun Li, Jianxin Wang, Yun Cheng, and Jie Wu. Corrigendum to “A secure model against mobile sink replication attacks in unattended sensor networks” COMPNW, Volume 221, February 2023, 109529. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004000. See

[1435].

Pei:2023:RNP

[1766] Jiaming Pei, Kaiyang Zhong, Mian Ahmad Jan, and Jinhai Li. Retraction notice to ‘Personalized Federated Learning Framework for Network Traffic Anomaly Detection’. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003900. See

[1076].

Anonymous:2023:Na

[1767] Anonymous. November 2023. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBp

[1768] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128623004711.

Souza:2023:RDS

[1769] Otavio A. de O. Souza, Caio A.

Caldeira, and Olga Goussevskaia. Randomized distributed self-adjusting tree networks. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003869.

Ma:2023:GGB

[1770] Zhaorui Ma, Shicheng Zhang, Na Li, Tianao Li, Xinhao Hu, Hao Feng, Qinglei Zhou, Fenlin Liu, Xiaowen Quan, Hongjian Wang, Guangwu Hu, Shubo Zhang, Yaqi Zhai, Shuaibin Chen, and Shuaiwei Zhang. GraphNEI: a GNNbased network entity identification method for IP geolocation. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003912.

Campazas-Vega:2023:AIS

[1771] Adrián Campazas-Vega, Ignacio Samuel Crespo-Martínez, Ángel Manuel Guerrero-Higueras, Claudia Álvarez-Aparicio, Vicente Matellán, and Camino Fernández-Llamas. Analyzing the influence of the sampling rate in the detection of malicious traffic on flow data. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003961.

Araujo:2023:CCS

[1772] Guilherme Araujo, Maycon Peixoto, and Leobino Sampaio. A comprehensive and configurable simulation environment for supporting vehicular named-data networking applications. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003948.

Chen:2023:CRS

[1773] Junmei Chen, Zongpeng Li, Guang Fang, Yeqiao Hou, and Xianglong Li. A comprehensive repair scheme for distributed storage systems. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003997.

Perifanis:2023:FLB

[1774] Vasileios Perifanis, Nikolaos Pavlidis, Remous-Aris Koutsiamanis, and Pavlos S. Efraimidis. Federated learning for 5G base station traffic forecasting. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300395X.

Kumar:2023:OTA

[1775] G. V. Pradeep Kumar, V. V. Satyanarayana Tallapragada, and N. Alivelu Manga. Optimized transmit antenna selection and self-attention based convolutional resource allocation model for massive MIMO technology. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003936.

Alkwai:2023:TCM

[1776] Lamia Alkwai, Abdelfettah Belghith, Achraf Gazdar, and Saad AlAhmadi. Transparent consumer mobility management in named data networking under the push communication mode. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003985.

Abbasi:2023:FTS

[1777] Felor Beikzadeh Abbasi, Ali Rezaee, Sahar Adabi, and Ali Movaghar. Fault-tolerant scheduling of graphbased loads on fog/cloud environments with multi-level queues and LSTMbased workload prediction. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004097.

Wang:2023:MSD

[1778] Kaiyu Wang, Zhiying Tu, Zhenzhou Ji, and Shufan He. Multistage data synchronization for public blockchain in complex network environment. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003973.

Cao:2023:NAI

[1779] Chenhong Cao, Miaoling Dai, Bonan Shen, Guobing Zou, and Wei Dong. Neural adaptive IoT streaming analytics with RL-Adapt. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003699.

Naderi:2023:HLE

[1780] Mohammad Naderi, Jacob Chakareski, and Mohammad Ghanbari. Hierarchical Q-learning-enabled neutrosophic AHP scheme in candidate relay set size adaption in vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004139.

Xu:2023:FLM

[1781] Yuwei Xu, Jie Cao, Kehui Song, Qiao Xiang, and Guang Cheng. FastTraffic: a lightweight method for encrypted traffic fast classification. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004103. See

corrigendum [1903].

Rafique:2023:SFC

[1782] Wajid Rafique, Abdelhakim Senhaji Hafid, and Soumaya Cherkaoui. Soft-Caching: a framework for caching node selection and routing in Software-Defined Information Centric Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004115.

Wang:2023:RTI

[1783] Shiyu Wang, Wenxiang Xu, and Yiwen Liu. Res-TranBiLSTM: an intelligent approach for intrusion detection in the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004279.

Min:2023:RLB

[1784] Junhong Min, Yongjun Kim, Moonbeom Kim, Jeongyeup Paek, and Ramesh Govindan. Reinforcement learning based routing for time-aware shaper scheduling in time-sensitive networks. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004280.

Dong:2023:BBD

[1785] Xiaodong Dong and Binlei Cai. Balanar: Balancing deadline guarantee and Jain’s fairness for inter-datacenter transfers. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004334.

He:2023:RLM

[1786] Rui He, Bangbang Ren, Junjie Xie, Deke Guo, Yuwen Zhou, Laiping Zhao, and Yong Li. A reinforcement learning method for scheduling service function chains with multi-resource constraints. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004309.

Liu:2023:MDG

[1787] Lingling Liu, Aimin Wang, Geng Sun, and Jiahui Li. Maximizing data gathering and energy efficiency in UAV-assisted IoT: a multi-objective optimization approach. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004310.

Wang:2023:CED

[1788] Han Wang, Zhangguo Tang, Huanzhou Li, Jian Zhang, Shuangcheng Li, and Junfeng Wang. CI GRU: an efficient DGA botnet classification model based on an attention recurrence plot. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004371.

Liu:2023:RLB

[1789] Ning Liu, Chunbo Jia, Bingnan Hou, Changsheng Hou, Yingwen Chen, and Zhiping Cai. 6Search: a reinforcement learning-based traceroute approach for efficient IPv6 topology discovery. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004322.

Wang:2023:PDR

[1790] Ting Wang, Xi Fan, Kai Cheng, Xiao Du, Haibin Cai, and Yang Wang. Parameterized deep reinforcement learning with hybrid action space for energy efficient data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004346.

Erazo-Agredo:2023:ERE

[1791] Cristian C. Erazo-Agredo, Luis Diez, Ramón Agüero, Mario Garza-Fabre, and Javier Rubio-Loyola. Enabling realistic experimentation of disaggregated RAN: Design, implementation, and validation of a modular multisplit eNodeB. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004383.

Lin:2023:NOD

[1792] Xu Lin, Chuchu Liu, Lailong Luo, Deke Guo, and Ming Xu. NEST: Optimal deploying DAG-SFCs to maximize the flows wholly served in the network edge. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004401.

Hidouri:2023:QIL

[1793] Abdelhak Hidouri, Haifa Touati, Mohamed Hadded, Nasreddine Hajlaoui, Paul Muhlethaler, and Samia Bouzefrane. Q-ICAN: a Q-learning based cache pollution attack mitigation approach for named data networking. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004437.

Wang:2023:DHO

[1794] Xiang Wang, Yang Liu, Jiahui Ma, Yunling Dong, Jiaxing Chen, and Zhihua Liu. A DV-Hop optimization localization algorithm based on topological structure similarity in three-dimensional wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004589.

Shao:2023:BAC

[1795] Huishuang Shao, Yurong Xia, and Changhao Piao. Blockchain-assisted certificateless signcryption for vehicleto-vehicle communication in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128623004140.

Zhu:2023:DRL

[1796] Xiaojuan Zhu, Tianhao Zhang, Jinwei Zhang, Bao Zhao, Shunxiang Zhang, and Cai Wu. Deep reinforcement learning-based edge computing offloading algorithm for softwaredefined IoT. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004516.

Jin:2023:MAN

[1797] Zilong Jin, Jun Qian, Zhixiang Kong, and Chengsheng Pan. A mobility aware network traffic prediction model based on dynamic graph attention spatio-temporal network. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004267.

Kumar:2023:CSH

[1798] E. Praveen Kumar and S. Priyanka. A comprehensive survey on hardwareassisted malware analysis and primitive techniques. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004127.

Ullah:2023:SBE

[1799] Syed Sajid Ullah, Vladimir Oleshchuk, and Harsha S. Gardiyawasam Pussewalage. A survey on blockchain envisioned attribute based access control for Internet of Things: Overview, comparative analysis, and open research challenges. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004395.

Nguyen:2023:TWU

[1800] Tien-Tung Nguyen, Anh-Vinh Nguyen, Tan-Loc Nguyen, and Van-Hung Le. Two-way UAV-aided systems with short packet communications. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004024.

Leon:2023:EPL

[1801] Juan Pablo Astudillo León, Carlos Lester Duenas Santos, Ahmad Mohamad Mezher, Julián Cárdenas Barrera, Julian Meng, and Eduardo Castillo Guerra. Exploring the potential, limitations, and future directions of wireless technologies in smart grid networks: a comparative analysis. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004012.

Bine:2023:NAC

[1802] Lailla M. S. Bine, Azzedine Boukerche, Linnyer B. Ruiz, and Antonio A. F. Loureiro. A novel ant colony-inspired coverage path planning for Internet of Drones. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004085.

Garcia-Luna-Aceves:2023:ANC

[1803] J. J. Garcia-Luna-Aceves and Dylan Cirimelli-Low. ALOHA-NUI: a collision-free version of ALOHA using a neighborhood-understood index. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004048.

Santana:2023:ALR

[1804] Juan Ramón Santana, Pablo Sotres, Jesús Pérez, Luis Sánchez, Jorge Lanza, and Luis Muñoz. Assessing LoRaWAN radio propagation for smart parking service: an experimental study. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004073.

Mukhtaruzzaman:2023:SDF

[1805] Mohammad Mukhtaruzzaman and Mohammed Atiquzzaman. Stable dynamic feedback-based predictive clustering protocol for vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300405X.

Belda:2023:DSS

[1806] Román Belda, Pau Arce, Juan Carlos Guerri, and Ismael de Fez. A DASH server-side delay-based representation switching solution to improve the quality of experience for lowlatency live video streaming. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004061.

Batalla:2023:SAR

[1807] Jordi Mongay Batalla, Sławomir Sujecki, Jan M. Kelner, Piotr Śliwka, and Dariusz Zmysłowski. On studying active radio measurements estimating the mobile network quality of service for the Regulatory Authority’s purposes. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004255.

Xu:2023:IBQ

[1808] Yaqi Xu, Yan Shi, Yuming Ge, Shanzhi Chen, and Longxiang Wang. Informerbased QoS prediction for V2X communication: a method with verification using reality field test data. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004036.

Anonymous:2023:Nb

[1809] Anonymous. November 2023. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBq

[1810] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128623005121.

Dahanayaka:2023:ROS

[1811] Thilini Dahanayaka, Yasod Ginige, Yi Huang, Guillaume Jourjon, and Suranga Seneviratne. Robust openset classification for encrypted traffic fingerprinting. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300436X.

deOliveira:2023:FNN

[1812] Jonathas A. de Oliveira, Vinícius P. Gonçalves, Rodolfo I. Meneguette, Rafael T. de Sousa, Daniel L. Guidoni, José C. M. Oliveira, and Geraldo P. Rocha Filho. F-NIDS — a network intrusion detection system based on federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004553.

Camargo:2023:DSR

[1813] Juan Sebastian Camargo, Estefanía Coronado, Wilson Ramirez, Daniel Camps, Sergi Sánchez Deutsch, Jordi Pérez-Romero, Angelos Antonopoulos, Oscar Trullols-Cruces, Sergio Gonzalez-Diaz, Borja Otura, and Giovanni Rigazzi. Dynamic slicing reconfiguration for virtualized 5G networks using ML forecasting of computing capacity. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004462.

Xu:2023:STO

[1814] Jing Xu, Jiarun Tang, Yuze Zou, Ruikai Wen, Wei Liu, and Jianhua He. Sum throughput optimization of wireless powered IRS-assisted multi-user MISO system. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004292.

Zafar:2023:DRB

[1815] Waseeq Ul Islam Zafar, Muhammad Atif Ur Rehman, Farhana Jabeen, Rehmat Ullah, Ghulam Abbas, and Abid Khan. Decentralized receiver-based link stability-aware forwarding scheme for NDN-based VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004413.

Hirayama:2023:SBV

[1816] Takahiro Hirayama, Masahiro Jibiki, and Ved P. Kafle. Salience-based VNF placement method for evolving multiple network slices. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004474.

Wang:2023:EVP

[1817] Ying Wang, Jiang Liu, Mingwei Cui, Weihong Wu, and Tao Huang. ED-VNE: a profit-oriented VNE optimization scheme of energy and delay in 5G SlaaS. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004486.

Khah:2023:DML

[1818] Sahar Ahmadi Khah, Ali Barati, and Hamid Barati. A dynamic and multi-level key management method in wireless sensor networks (WSNs). Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004425.

Lv:2023:ADO

[1819] Sicai Lv, Chao Wang, Zibo Wang, Shuo Wang, Bailing Wang, and Yongzheng Zhang. AAE-DSVDD: a one-class classification model for VPN traffic identification. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004358.

Liu:2023:PWP

[1820] Pei Liu, Bing Qian, Qi Sun, and Longgang Zhao. Prompt-WNQA: a prompt-based complex question answering for wireless network over knowledge graph. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004590.

Amiri-Zarandi:2023:SFL

[1821] Mohammad Amiri-Zarandi, Rozita A.

Dara, and Xiaodong Lin. SIDS: a federated learning approach for intrusion detection in IoT using Social Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004504.

Mohammadzad:2023:MMA

[1822] Maryam Mohammadzad, Jaber Karimpour, and Farnaz Mahan. MAGD: Minimal attack graph generation dynamically in cyber security. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004498.

Wang:2023:CCM

[1823] Zhendong Wang, Liwei Shao, Shuxin Yang, Junling Wang, and Dahai Li. CRLM: a cooperative model based on reinforcement learning and metaheuristic algorithms of routing protocols in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004644.

Lu:2023:EUL

[1824] Yuan Lu, Guoyuan Yuan, Yang Bai, Dezun Dong, and Renjie Zhou. EagerCC: an ultra-low latency congestion control mechanism in datacenter networks. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004541.

Lu:2023:URU

[1825] Qiuyu Lu, June Li, Kai Yuan, Kaipei Liu, Ming Ni, and Jianbo Luo. UDP-RT: a UDP-based reliable transmission scheme for power WAPS. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004577.

Cotter:2023:CAP

[1826] Jamie Cotter, Ignacio Castiñeiras, Donna O’Shea, and Victor Cionca. A comparative analysis of proactive and reactive methods for privacy-aware interleaved DNN offloading. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004449.

Xiao:2023:DPB

[1827] Guangbing Xiao, Haibo Zhang, Zhiyi Huang, and Yawen Chen. Decentralized piggybacking-based dissemination of Cooperative Awareness Messages in vehicular ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004772.

Amin:2023:PAD

[1828] Md. Ruhul Amin, Md. Shohrab Hossain, and Mohammed Atiquzzaman. Performance analysis of distributed coordination function with early collision detection in in-band full-duplex wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004620.

Liang:2023:MAM

[1829] Yu Liang and Sheng Zhang. Mobilityaware multi-user service placement and resource allocation in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004656.

Liu:2023:BBI

[1830] Zhenpeng Liu, Shuo Wang, and Yi Liu. Blockchain-based integrity auditing for shared data in cloud storage with file prediction. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004851.

Awad:2023:EED

[1831] Sumaya D. Awad, A. Sali, Mohanad M. Al-Wani, Ali M. Al-Saegh, J. S. Mandeep, and R. S. A. Raja Abdullah. End-to-end DVB-S2X system design with DL-based channel estimation over satellite fading channels at Ka-band. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300467X.

Garcia-Rodriguez:2023:PPA

[1832] Jesús García-Rodríguez and Antonio Skarmeta. A privacy-preserving attribute-based framework for IoT identity lifecycle management. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912862300484X.

Okhovvat:2023:JTE

[1833] Morteza Okhovvat, Mohammad Taghi Kheirabadi, Mohammad Reza Okhovvat, and Ali Nodehi. Joint time and energy-optimal approach to allocate task to actors in wireless sensor actor networks. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004632.

Wang:2023:LZK

[1834] Zhangquan Wang, Jiaxuan Huang, Kelei Miao, Xiaowen Lv, Yourong Chen, Bing Su, Liyuan Liu, and Meng Han. Lightweight zero-knowledge authentication scheme for IoT embedded devices. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004668.

Li:2023:PSS

[1835] Man Li, Shuangxing Deng, Huachun Zhou, and Yajuan Qin. A path selection scheme for detecting malicious behavior based on deep reinforcement learning in SDN/NFVenabled network. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004796.

Li:2023:GDA

[1836] Kun Li, Huachun Zhou, Zhe Tu, Ouyang Liu, and Hongke Zhang. AT-GCN: a DDoS attack path tracing system based on attack traceability knowledge base and GCN. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004814.

Tian:2023:PEB

[1837] He Tian, Kaihong Guo, Ran Zhang, and Shiliang Shao. Prediction of evolution behavior of Internet bottleneck delay based on improved logistic equation. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004863.

Chaudhary:2023:DAI

[1838] Shubhankar Chaudhary and Pramod Kumar Mishra. DDoS attacks in Industrial IoT: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004607.

M:2023:DRL

[1839] Jesús F. Cevallos M., Alessandra Rizzardi, Sabrina Sicari, and Alberto Coen Porisini. Deep reinforcement learning for intrusion detection in Internet of Things: Best practices, lessons learnt, and open challenges. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004619.

McManus:2023:DTE

[1840] Maxwell McManus, Yuqing Cui, Josh (Zhaoxi) Zhang, Jiangqi Hu, Sabarish Krishna Moorthy, Nicholas Mastronarde, Elizabeth Serena Bentley, Michael Medley, and Zhangyu Guan. Digital twin-enabled domain adaptation for zero-touch UAV networks: Survey and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128623004450.

Banotra:2023:EHS

[1841] Atul Banotra, Sarbani Ghose, Deepak Mishra, and Sudhakar Modem. Energy harvesting in self-sustainable IoT devices and applications based on cross-layer architecture design: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004565.

Khan:2023:RAT

[1842] Fátima Khan, Cristina Hervella, Luis Diez, Fátima Fernández, Néstor J. Hernández Marcano, Rune Hylsberg Jacobsen, and Ramón Agüero. Realistic assessment of transport protocols performance over LEO-based communications. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300453X.

deMattos:2023:SWS

[1843] Ekler Paulino de Mattos, Augusto C. S. A. Domingues, Fabrício A. Silva, Heitor S. Ramos, and Antonio A. F. Loureiro. Slicing who slices: Anonymization quality evaluation on deployment, privacy, and utility in mix-zones. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004528.

Anonymous:2023:D

[1844] Anonymous. December 2023. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2023:EBr

[1845] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128623005716.

Mahboob:2023:CVC

[1846] Shadab Mahboob, Koushik Kar, Jacob Chakareski, and Md Ibrahim Alam. CLoSER: Video caching in small-cell edge networks with local content sharing. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004784.

Omri:2023:EDN

[1847] Aymen Omri, Javier Hernandez Fernandez, and Roberto Di Pietro. Extending device noise measurement capacity for OFDM-based PLC systems: Design, implementation, and onfield validation. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004838.

Fernandez:2023:SAH

[1848] Yaime Fernández, Javier E. Soto, Sofía Vera, Yasmany Prieto, Cecilia Hernández, and Miguel Figueroa. A streaming algorithm and hardware accelerator to estimate the empirical entropy of network flows. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004802.

Leyva-Pupo:2023:ISU

[1849] Irian Leyva-Pupo and Cristina Cervelló-Pastor. An intelligent scheduling for 5G user plane function placement and chaining reconfiguration. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004826.

Liang:2023:LNB

[1850] Weibiao Liang, Lin You, and Gengran Hu. LRS PKI: a novel blockchainbased PKI framework using linkable ring signatures. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004887.

Wang:2023:MKS

[1851] Miaomiao Wang, Lanlan Rui, Siya Xu, Zhipeng Gao, Huiyong Liu, and Shaoyong Guo. A multi-keyword searchable encryption sensitive data trusted sharing scheme in multi-user scenario. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004905.

Arat:2023:NMV

[1852] Ferhat Arat and Sedat Akleylek. A new method for vulnerability and risk assessment of IoT. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004917.

Bozorgchenani:2023:NMO

[1853] Arash Bozorgchenani, Charilaos C.

Zarakovitis, Su Fong Chien, Tiew On Ting, Qiang Ni, and Wissam Mallouli. Novel modeling and optimization for joint Cybersecurity-vs-QoS Intrusion Detection Mechanisms in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004966.

Sartayeva:2023:SIP

[1854] Yerkezhan Sartayeva, Henry C. B.

Chan, Yik Him Ho, and Peter H. J. Chong. A survey of indoor positioning systems based on a six-layer model. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004875.

Zhang:2023:ESL

[1855] Yuandi Zhang, Jiawangnan Lu, Hongtao Zhang, Ziyi Huang, César Briso-Rodríguez, and Lei Zhang. Experimental study on low-altitude UAVto-ground propagation characteristics in campus environment. Computer Networks (Amsterdam, Netherlands:

1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005005.

S:2023:DED

[1856] Ahmad Salehi S., Runchao Han, Carsten Rudolph, and Marthie Grobler. DACP: Enforcing a dynamic access control policy in cross-domain environments. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004942.

Wang:2023:BEF

[1857] Houtian Wang, Taotao Wang, Long Shi, Naijin Liu, and Shengli Zhang. A blockchain-empowered framework for decentralized trust management in Internet of Battlefield Things. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004930.

Liu:2023:OIS

[1858] Yujie Liu, Shangping Wang, Duo Zhang, Qian Zhang, and Jifang Wang. Optimal incentive strategy in blockchain-based mobile crowdsensing using game theory. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300498X.

Dong:2023:SSS

[1859] Enhuan Dong, Peng Gao, Yuan Yang, Mingwei Xu, Xiaoming Fu, and Jiahai Yang. SmartSBD: Smart shared bottleneck detection for efficient multipath congestion control over heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004929.

Wang:2023:GGB

[1860] Xiao Wang and Xudong Wang. Graph-PowerNet: Graph-based power consumption profiling for mobile phone applications. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005017.

Chang:2023:DES

[1861] Wei-Teng Chang and Ben-Jye Chang.

Dynamic extended sigmoid-based scheduling with virtualized RB allocation for maximizing frequency numerology efficiency in 5G/B5G NR networks. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300508X.

Zhu:2023:SAE

[1862] Yan Zhu and Dawei Lu. Satelliteassisted edge computing management based on deep reinforcement learning in industrial Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004954.

Rochman:2023:CAC

[1863] Muhammad Iqbal Rochman, Vanlin Sathya, Damian Fernandez, Norlen Nunez, Ahmed S. Ibrahim, William Payne, and Monisha Ghosh. A comprehensive analysis of the coverage and performance of 4G and 5G deployments. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005054.

Zhu:2023:FFF

[1864] Zhengbin Zhu, Hong Yu, Qinrang Liu, Dongpei Liu, and Bo Mei. FFRLI: Fast fault recovery scheme based on link importance for data plane in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005078.

Ahmad:2023:ISR

[1865] Muneeb Ahmad, Tanzeela Shakeel, and Soo Young Shin. Image super resolution based channel estimation for future wireless communication. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005029.

Ullah:2023:TIT

[1866] Safi Ullah, Jawad Ahmad, Muazzam A.

Khan, Mohammed S. Alshehri, Wadii Boulila, Anis Koubaa, Sana Ullah Jan, and M Munawwar Iqbal Ch. TNN-IDS: Transformer neural network-based intrusion detection system for MQTTenabled IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005170.

Luo:2023:NCC

[1867] Pengcheng Luo, Yuan Liu, Zekun Wang, Jian Chu, and Genke Yang. A novel Congestion Control algorithm based on inverse reinforcement learning with parallel training. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005169.

Zhang:2023:GSE

[1868] Quanwei Zhang, Qingjun Xiao, and Yuexiao Cai. A generic sketch for estimating super-spreaders and per-flow cardinality distribution in high-speed data streams. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005042.

Ahmadi:2023:SAE

[1869] Mojtaba Ahmadi, Sahar Kianian, and Zahra Shirmohammadi. Social-aware energy management in D2D communications. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005066.

Abdelhakam:2023:CCT

[1870] Mostafa M. Abdelhakam, Mahmoud M. Elmesalawy, Ibrahim I. Ibrahim, and Samir G. Sayed. Collaborative CoMP and trajectory optimization for energy minimization in multi-UAV-assisted IoT networks with QoS guarantee. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005194.

Liu:2023:JMT

[1871] Jingyao Liu, Guangsheng Feng, Zhibo Zhang, Liying Zheng, Huiqiang Wang, and Jyri Hämäläinen. Joint mixedtimescale optimization of content caching and delivery policy in NOMAbased vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005200.

Wu:2023:MAC

[1872] Lisheng Wu, Xiaoming You, and Sheng Liu. Multi-ant colony algorithm based on cooperative game and dynamic path tracking. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005224.

Robitzsch:2023:PAM

[1873] Sebastian Robitzsch, Marco Centenaro, Nicola di Pietro, Luis Cordeiro, André S. Gomes, Peter Sanders, and Arif Ishaq. Prospects on the adoption of a microservice-based architecture in 5G systems and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005030.

Hajar:2023:RMA

[1874] Muhammad Shadi Hajar, Harsha Kumara Kalutarage, and M. Omar Al-Kadri. 3R: a reliable multi agent reinforcement learning based routing protocol for wireless medical sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005182.

Wu:2023:LVS

[1875] Qiyu Wu, Fucai Zhou, Jian Xu, and Da Feng. Lightweight and verifiable secure aggregation for multi-dimensional data in edge-enhanced IoT. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005248.

Moorthy:2023:OHS

[1876] Sabarish Krishna Moorthy, Nicholas Mastronarde, Elizabeth Serena Bentley, Michael Medley, and Zhangyu Guan. OSWireless: Hiding specification complexity for zero-touch software-defined wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128623005212.

Ren:2023:NNA

[1877] Qingqing Ren, Shuyong Zhu, Lu Lu, Zhiqiang Li, Guangyu Zhao, and Yujun Zhang. NetShield: an in-network architecture against Byzantine failures in distributed deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005261.

Chen:2023:ELS

[1878] Ciyuan Chen, Junzhou Luo, Zhuqing Xu, Runqun Xiong, Dian Shen, and Zhimeng Yin. Enabling large-scale lowpower LoRa data transmission via multiple mobile LoRa gateways. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005285.

Kurihara:2023:ODA

[1879] Jun Kurihara, Toshiaki Tanaka, and Takeshi Kubo. μODNS: a distributed approach to DNS anonymization with collusion resistance. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005236.

Sahin:2023:NBS

[1880] Emre Sahin, Mustafa Alper Akkas, and Orhan Dagdeviren. Nanonetworkbased search and rescue operations in debris areas. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005273.

Wang:2023:STW

[1881] Wei Wang, Weiping Zhao, Dingcheng Yang, An Li, Yapeng Wang, and Lin Xiao. Secure two-way transmission via untrusted UAV relay: Joint path design and slot-pairing strategy. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005352.

Ebrahim:2023:PAL

[1882] Maad Ebrahim and Abdelhakim Hafid. Privacy-aware load balancing in fog networks: a reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005406.

Gan:2023:STD

[1883] Chenquan Gan, Anqi Liu, Qingyi Zhu, Ye Zhu, Yong Xiang, and Jun Liu. Social tie-driven coupling propagation of user awareness and information in device-to-device communications. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005327.

Xiao:2023:FRP

[1884] Qingjun Xiao, Yifei Li, and Yeke Wu. Finding recently persistent flows in high-speed packet streams based on cuckoo filter. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300542X.

Bolla:2023:MGS

[1885] Raffaele Bolla, Roberto Bruschi, Chiara Lombardo, Alireza Mohammadpour, Riccardo Trivisonno, and Wint Yi Poe. A 5G multi-gNodeB simulator for ultra-reliable 0.5–100 GHz communication in indoor Industry 4.0 environments. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005480.

Takeda:2023:LPM

[1886] Kenta Takeda, Takehiro Sato, Bijoy Chand Chatterjee, and Eiji Oki. Lightpath provisioning model considering crosstalk-derived fragmentation in spectrally-spatially elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005443.

Dangi:2023:FSB

[1887] Ramraj Dangi and Praveen Lalwani.

Feature selection based machine learning models for 5G network slicing approximation. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005388.

Liu:2023:FJC

[1888] Su Liu, Peiyuan Guan, Jiong Yu, and Amir Taherkordi. FedSSC: Joint client selection and resource management for communication-efficient federated vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005455.

Awad:2023:MRP

[1889] Mirna Awad, Aris Leivadeas, and Abir Awad. Multi-resource predictive workload consolidation approach in virtualized environments. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005339.

Quan:2023:FOC

[1890] Yu xuan Quan, Yu ning Dong, Yang Xiang, Shan shan Chen, Zai jian Wang, and Jiong Jin. Fast online classification of network traffic using new feature-embedded hierarchical structure. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005510.

Tan:2023:NRO

[1891] Haining Tan, Tao Ye, Sadaqat ur Rehman, Obaid ur Rehman, Shanshan Tu, and Jawad Ahmad. A novel routing optimization strategy based on reinforcement learning in perception layer networks. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005509.

Chaudhary:2023:EIC

[1892] Pankaj Chaudhary, Neminath Hubballi, and Sameer G. Kulkarni. eN-Cache: Improving content delivery with cooperative caching in Named Data Networking. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005492.

Yang:2023:LPH

[1893] Xu Yang, Jiahe Yu, Saiyu Qi, Qiuhao Wang, Jianfeng Wang, Yanan Qiao, and Yong Qi. Less payment and higher efficiency: a verifiable, fair and forward-secure range query scheme using blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005534.

Li:2023:ABA

[1894] Peng Li, Junzuo Lai, Ye Yang, Meng Sun, Chi Wu, Wei Wu, and Xiaowei Yuan. Attribute-based anonymous credential: Delegation, traceability, and revocation. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005315.

Zaballos:2023:UTT

[1895] Agustín Zaballos, Adrià Mallorquí, and Joan Navarro. Unboxing trustworthiness through quantum internet. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300539X.

Li:2023:DAM

[1896] Xia Li, Xiaoli Wang, and Yuping Wang. DRL assisted multi-objective algorithm for multicast scheduling in elastic optical network. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005364.

Karmakar:2023:TAO

[1897] Kallol Krishna Karmakar, Vijay Varadharajan, Michael Hitchens, Uday Tupakula, and Prajna Sariputra. A trust-aware openflow switching framework for software defined networks (SDN). Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005546.

Chen:2023:VYB

[1898] Yitong Chen, Chaoqin Gan, Shibao Wu, Xiaoqi Wang, and Yixin Chen. (Vehicle’s YB)-based LPB prediction model and anti-LPB (communication link)-reconstruction strategy in the (VLC/RF)-V2X network. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005571.

Hu:2023:FGE

[1899] Xiaoyan Hu, Zhuozhuo Shu, Zhongqi Tong, Guang Cheng, Ruidong Li, and Hua Wu. Fine-grained Ethereum behavior identification via encrypted traffic analysis with serialized backward inference. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005558.

Taleb:2023:AMB

[1900] Tarik Taleb, Chafika Benzaïd, Rami Akrem Addad, and Konstantinos Samdanis. AI/ML for beyond 5G systems: Concepts, technology enablers and solutions. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004899.

Gomez:2023:SNS

[1901] Jose Gomez, Elie F. Kfoury, Jorge Crichigno, and Gautam Srivastava. A survey on network simulators, emulators, and testbeds used for research and education. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004991.

Jawad:2023:CSB

[1902] Aqeel Thamer Jawad, Rihab Maaloul, and Lamia Chaari. A comprehensive survey on 6G and beyond: Enabling technologies, opportunities of machine learning and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005303.

Xu:2023:CFL

[1903] Yuwei Xu, Jie Cao, Kehui Song, Qiao Xiang, and Guang Cheng. Corrigendum to “FastTraffic: a lightweight method for encrypted traffic fast classification” [Computer Networks, Volume 235, November 2023, 109965]. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005431. See

[1781].

Zhang:2023:CVN

[1904] Guoan Zhang, Wei Cao, and Xiaohui Gu. Cooperative vehicular networks over Nakagami-m fading: Joint power control and spectrum scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004978.

Gonzalez:2023:QEM

[1905] N. González, M. Solera, F. Ruiz, C. Gijón, and M. Toril. A quality of experience model for live video in first-person-view drone control in cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005340.

Anonymous:2024:Ja

[1906] Anonymous. January 2024. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2024:EBa

[1907] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005960.

Premalatha:2024:OEE

[1908] B. Premalatha and P. Prakasam. Optimal energy-efficient resource allocation and fault tolerance scheme for task offloading in IoT-FoG computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 238 (??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300525X.

Wang:2024:ETP

[1909] Ke Wang, Xiaojuan Ma, Heng Kang, Zheng Lyu, Baorui Feng, Wenliang Lin, Zhongliang Deng, and Yun Zou. An efficient topology partitioning algorithm for system-level parallel simulation of mega satellite constellation communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005479.

Kashif:2024:EEP

[1910] Muhammad Kashif and Kubra Kalkan. EPIoT: Enhanced privacy preservation based blockchain mechanism for Internet-of-Things. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005522.

Al-hammadi:2024:CCO

[1911] Ikhlas Al-hammadi, Mingchu Li, Sardar M. N. Islam, and Esmail Al-Mosharea. Collaborative computation offloading for scheduling emergency tasks in SDN-based mobile edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005467.

Wu:2024:QBR

[1912] Yutong Wu, Jianyue Zhu, Xiao Chen, Yu Zhang, Yao Shi, and Yaqin Xie. QoS-based resource allocation for uplink NOMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005297.

Alam:2024:JES

[1913] Amjad Alam, Kamran Ali, Ramona Trestian, Purav Shah, and Glenford Mapp. Joint energy and spectral optimization in heterogeneous vehicular network. Computer Networks (Amsterdam, Netherlands: 1999), 238 (??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300556X.

Yang:2024:RBO

[1914] Sijin Yang, Lei Zhuang, Julong Lan, Jianhui Zhang, and Bingkui Li. Reusebased online joint routing and scheduling optimization mechanism in deterministic networks. Computer Networks (Amsterdam, Netherlands: 1999), 238 (??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005625.

Gu:2024:TRT

[1915] Xiaolin Gu, Wenjia Wu, Yusen Zhou, Aibo Song, Ming Yang, Zhen Ling, and Junzhou Luo. TEA-RFFI: Temperature adjusted radio frequency fingerprint-based smartphone identification. Computer Networks (Amsterdam, Netherlands: 1999), 238 (??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005601.

Kankane:2024:CRW

[1916] Bhawna Kankane, Sandeep Sharma, and Rajesh Mishra. κ-coverage reliability for wireless multihop network incorporating boundary effect. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005649.

Zhang:2024:PIP

[1917] Ze Zhang, Chingfang Hsu, Man Ho Au, Lein Harn, Jianqun Cui, Zhe Xia, and Zhuo Zhao. PRLAP-IoD: a PUF-based robust and lightweight authentication protocol for Internet of Drones. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005637.

Dou:2024:MIC

[1918] Songshi Dou, Li Qi, and Zehua Guo.

Mitigating the impact of controller failures on QoS robustness for softwaredefined wide area networks. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005418.

Wang:2024:MCM

[1919] Rui Wang, Gang Tian, and Shi Ying.

MicroCM: a cloud monitoring architecture for microservice invocation. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005662.

Yang:2024:PPA

[1920] Xiaodu Yang, Sijie Ruan, Yinliang Yue, and Bo Sun. PETNet: Plaintextaware encrypted traffic detection network for identifying Cobalt Strike HTTPS traffics. Computer Networks (Amsterdam, Netherlands: 1999), 238 (??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005650.

Lin:2024:SPC

[1921] Hsin-Tsung Lin and Pi-Chung Wang. Scalable packet classification based on rule categorization and crossproducting. Computer Networks (Amsterdam, Netherlands: 1999), 238 (??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005613.

Islam:2024:PIP

[1922] SK Hafizul Islam, Krittibas Parai, and Daya Sagar Gupta. PF-IBDA: Provably secure and pairing-free identitybased deniable authentication protocol for MANET environments. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005583.

Le:2024:ACU

[1923] Siqi Le, Yingxu Lai, Yipeng Wang, and Huijie He. An adaptive classification and updating method for unknown network traffic in open environments. Computer Networks (Amsterdam, Netherlands: 1999), 238 (??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005595.

Chakour:2024:UQP

[1924] Imane Chakour, Sajida Mhammedi, Cherki Daoui, and Mohamed Baslam. Unlocking QoS potential: Integrating IoT services and Monte Carlo control for heterogeneous IoT device management in gateways. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005790.

Kumar:2024:SCB

[1925] Neeraj Kumar and Rifaqat Ali. A smart contract-based robotic surgery authentication system for healthcare using 6G-Tactile Internet. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005789.

Xiong:2024:EAL

[1926] Bing Xiong, Jing Wu, Qiaorong Huang, Jinyuan Zhao, Qiang Tang, Jin Zhang, Kun Yang, and Keqin Li. Elastically accelerating lookup on virtual SDN flow tables for softwaredefined cloud gateways. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005376.

Anonymous:2024:Fa

[1927] Anonymous. February 2024. Computer Networks (Amsterdam, Netherlands: 1999), 239(??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2024:EBb

[1928] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 239(??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128624000094.

Ramos:2024:ETC

[1929] Daniel Ramos, Oscar Esparza, Jorge Mata-Díaz, and Juanjo Alins. Evaluation of TCP congestion control algorithms with traffic control policies in a PEP-based geosynchronous satellite scenario. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005765.

Duan:2024:TLB

[1930] Zhonglei Duan and Ting Wang. Towards learning-based energy-efficient online coordinated virtual network embedding framework. Computer Networks (Amsterdam, Netherlands: 1999), 239(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005844.

Sousa:2024:FSF

[1931] Bruno Sousa and Carolina Gonçalves. FedAAA-SDN: Federated authentication, authorization and accounting in SDN controllers. Computer Networks (Amsterdam, Netherlands: 1999), 239(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005753.

Chen:2024:GNN

[1932] Hongyang Chen, Pengfei Chen, Benran Wang, Xian Yu, Xiaofan Chen, Dandan Ma, and Zibin Zheng. Graph neural network based robust anomaly detection at service level in SDN driven microservice system. Computer Networks (Amsterdam, Netherlands: 1999), 239(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005807.

Kita:2024:UAA

[1933] Nozomi Kita and Eiji Oki. Unavailabilityaware allocation of backup resources considering failures of virtual and physical machines. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005868.

Yu:2024:EEN

[1934] Ya-Ju Yu and Shan-Yao Lo. Energyefficient non-anchor channel allocation in NB-IoT cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 239(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300590X.

Cao:2024:WTL

[1935] Tuo Cao, Qinhui Wang, Yuhan Zhang, Zhuzhong Qian, Yue Zeng, Mingtao Ji, Hesheng Sun, and Baoliu Ye. Walking on two legs: Joint service placement and computation configuration for provisioning containerized services at edges. Computer Networks (Amsterdam, Netherlands: 1999), 239(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005893.

Tu:2024:PDD

[1936] Huaqing Tu, Gongming Zhao, Hongli Xu, and Chunming Qiao. Programmable device deployment for efficient network function offloading. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623006084.

Imputato:2024:BWF

[1937] Pasquale Imputato, Stefano Avallone, Malcolm Smith, David Nunez, and Boris Bellalta. Beyond Wi-Fi 7: Spatial reuse through multi-AP coordination. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623006059.

Doriguzzi-Corin:2024:IPL

[1938] Roberto Doriguzzi-Corin, Luis Augusto Dias Knob, Luca Mendozzi, Domenico Siracusa, and Marco Savi. Introducing packet-level analysis in programmable data planes to advance Network Intrusion Detection. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623006072.

Mohammed:2024:CSR

[1939] Ahmed Burhan Mohammed, Lamia Chaari Fourati, and Ahmed M. Fakhrudeen. Comprehensive systematic review of intelligent approaches in UAV-based intrusion detection, blockchain, and network security. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005856.

Bine:2024:CID

[1940] Lailla M. S. Bine, Azzedine Boukerche, Linnyer B. Ruiz, and Antonio A. F. Loureiro. Connecting Internet of Drones and urban computing: Methods, protocols and applications. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005819.

Cen:2024:RED

[1941] Mingcan Cen, Frank Jiang, Xingsheng Qin, Qinghong Jiang, and Robin Doss. Ransomware early detection: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005832.

Bimpas:2024:LPC

[1942] Athanasios Bimpas, John Violos, Aris Leivadeas, and Iraklis Varlamis. Leveraging pervasive computing for ambient intelligence: a survey on recent advancements, applications and open challenges. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623006011.

Trabelsi:2024:IMB

[1943] Nessrine Trabelsi, Lamia Chaari Fourati, and Chung Shue Chen. Interference management in 5G and beyond networks: a comprehensive survey. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623006047.

Gonzalez-Soto:2024:DCM

[1944] Martín González-Soto, Rebeca P. Díaz-Redondo, Manuel Fernández-Veiga, Bruno Fernández-Castro, and Ana Fernández-Vilas. Decentralized and collaborative machine learning framework for IoT. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005820.

Anonymous:2024:Fb

[1945] Anonymous. February 2024. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2024:EBc

[1946] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624000598.

Griner:2024:BMD

[1947] Chen Griner, Chen Avin, and Gil Einziger. Beyond matchings: Dynamic multi-hop topology for demandaware datacenters. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005881.

Firouzjaee:2024:UPA

[1948] Mostafa Taghizade Firouzjaee, Kamal Jamshidi, Neda Moghim, and Sachin Shetty. User preference-aware content caching strategy for video delivery in cache-enabled IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300587X.

Li:2024:DLB

[1949] Shuyang Li, Enrico Magli, Gianluca Francini, and Giorgio Ghinamo. Deep learning based prediction of traffic peaks in mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623006126.

Pathan:2024:PBE

[1950] Md. Naimul Pathan, Maisha Muntaha, Selina Sharmin, Sajeeb Saha, Md. Ashraf Uddin, Fernaz Narin Nur, and Sunil Aryal. Priority based energy and load aware routing algorithms for SDN enabled data center network. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623006114.

Chang:2024:BSG

[1951] Sang-Yoon Chang, Arijet Sarker, Simeon Wuthier, Jinoh Kim, Jonghyun Kim, and Xiaobo Zhou. Base station gateway to secure user channel access at the first hop edge. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623006102.

Zhang:2024:RBR

[1952] Yiguo Zhang, Wei Wang, and Feiyang Shi. Reputation-based Raft-Poa layered consensus protocol converging UAV network. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000021.

Koumar:2024:NEI

[1953] Josef Koumar, Karel Hynek, Jaroslav Pesek, and Tomás Cejka. Net-TiSA: Extended IP flow with timeseries features for universal bandwidthconstrained high-speed network traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005923.

Zhai:2024:NEH

[1954] Chao Zhai, Jiachao Yu, Kun Du, Hao Dai, Xinhua Wang, and Yujun Li. Nonlinear energy harvesting based alternate cooperative nonorthogonal multiple access with adaptive interference cancellation. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400001X.

Arif:2024:CJA

[1955] Mohammad Arif and Abdul Shakoor. Clustered jamming and antenna beamwidth fluctuations for UAV-assisted cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000033.

Seifhosseini:2024:MOC

[1956] Seyyedamin Seifhosseini, Mirsaeid Hosseini Shirvani, and Yaser Ramzanpoor. Multi-objective cost-aware bagof-tasks scheduling optimization model for IoT applications running on heterogeneous fog environment. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623006060.

Younes:2024:HDL

[1957] Osama S. Younes. A hybrid deep learning model for detecting DDoS flooding attacks in SIP-based systems. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005911.

Wadatkar:2024:JMO

[1958] Prachi Vinod Wadatkar, Rosario Giuseppe Garroppo, Gianfranco Nencioni, and Marco Volpi. Joint multi-objective MEH selection and traffic path computation in 5G-MEC systems. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623006138.

Hu:2024:INT

[1959] Xueyan Hu, Wei Liu, and Hua Huo.

An intelligent network traffic prediction method based on Butterworth filter and CNN-LSTM. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000045.

Song:2024:IIA

[1960] Junda Song, Jiajin Liu, Peixuan Gao, Guyue Liu, and H. Jonathan Chao. Inversion impact of approximate PIFO to Start-Time Fair Queueing. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623006096.

Magoula:2024:SCI

[1961] Lina Magoula, Nikolas Koursioumpas, Theodora Panagea, Nancy Alonistioti, Chaima Ghribi, and Joshua Shakya. SIM+: a comprehensive implementation-agnostic information model assisting AI-driven optimization for beyond 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000227.

Pi:2024:MCI

[1962] Yibo Pi, Sugih Jamin, and Yichen Wei. Measuring congestion-induced performance imbalance in Internet load balancing at scale. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000215.

Qiao:2024:BBM

[1963] Yuncheng Qiao, Qiujun Lan, Shiyu Jia, Xianhua Kuang, Qingyue Xiong, and Chaoqun Ma. Blockchain-based multiinstitution collaborative credit evaluation modeling method. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000197.

Cai:2024:ACL

[1964] Yinan Cai, Wei Li, Xiangxu Meng, Wenqi Zheng, Chuhao Chen, and Zhuoxuan Liang. Adaptive contrastive learning based network latency prediction in 5G URLLC scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000173.

Garcia-Martinez:2024:BCM

[1965] Alberto García-Martínez, Pedro Rodrigues Torres, and Marcelo Bagnulo. BGP convergence in an MRAI-free Internet. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400015X.

Zheng:2024:NLL

[1966] Xiao Zheng, Yan Liu, Shuangping Zhan, Yao Xin, and Yi Wang. A novel low-latency scheduling approach of TSN for multi-link rate networking. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000161.

Hara:2024:PKU

[1967] Takanori Hara and Masahiro Sasabe.

Practicality of in-kernel/user-space packet processing empowered by lightweight neural network and decision tree. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000203.

Anonymous:2024:Ma

[1968] Anonymous. March 2024. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2024:EBd

[1969] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000963.

Liu:2024:QAB

[1970] Ying Liu, Jiawang Zhi, Xiaoyu Xia, Yuzheng Han, Changsheng Zhang, and Bin Zhang. QoE-aware budgeted edge data caching online: a primal-dual approach. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000276.

Chen:2024:MAD

[1971] Lin Chen, Weijia Wang, Yichuan Song, Jihong Yu, Kehao Wang, Shurong Zhang, Weihua Yang, and Celimuge Wu. Meet in the air: Distributed neighbor discovery in 3D networks with directional transceivers. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000331.

Song:2024:QIS

[1972] Jiahui Song. Quantifying the importance of specific types of higher-order structures in complex networks. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400032X.

Ji:2024:IPB

[1973] Mingtao Ji, Chenwei Su, Yitao Fan, Yibo Jin, Zhuzhong Qian, Yuting Yan, Yu Chen, Tuo Cao, Sheng Zhang, and Baoliu Ye. INTaaS: Provisioning in-band network telemetry as a service via online learning. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000434.

Pieska:2024:LDC

[1974] Marcus Pieska, Alexander Rabitsch, Anna Brunstrom, Andreas Kassler, Markus Amend, and Eckard Bogenfeld. Low-delay cost-aware multipath scheduling over dynamic links for access traffic steering, switching, and splitting. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000185.

Moazzeni:2024:VTN

[1975] Shadi Moazzeni, Konstantinos Katsaros, Nasim Ferdosian, Konstantinos Antonakoglou, Mark Rouse, Dritan Kaleshi, Adriana Fernández-Fernández, Miguel Catalan-Cid, Constantinos Vrontos, Reza Nejabati, and Dimitra Simeonidou. 5G-VIOS: Towards next generation intelligent interdomain network service orchestration and resource optimisation. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000343.

BenMazziane:2024:TML

[1976] Younes Ben Mazziane, Sara Alouf, Giovanni Neglia, and Daniel S. Menasche. TTL model for an LRU-based similarity caching policy. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000380.

Fan:2024:DRN

[1977] Chenyu Fan, Xiaoning Zhang, Yangming Zhao, Yexiao He, and Yi Yang. Dynamic relay node selection and routing for cloud-native Software Defined WANs. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000513.

Chong:2024:MPU

[1978] Kah Meng Chong and Amizah Malip. May the privacy be with us: Correlated differential privacy in location data for ITS. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400046X.

Lei:2024:RES

[1979] Wenli Lei, Yang Lei, Hongbo Guo, and Kun Jia. Research on energysaving relay selection algorithm based on olive forwarding area in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000318.

Pei:2024:EER

[1980] Xinglong Pei, Penghao Sun, Yuxiang Hu, Dan Li, Bo Chen, and Le Tian. Enabling efficient routing for traffic engineering in SDN with Deep Reinforcement Learning. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000525.

Gohar:2024:OCM

[1981] Ali Gohar and Gianfranco Nencioni.

An online cost minimization of the slice broker based on deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000306.

Singh:2024:PPQ

[1982] Nisha Singh, Mukesh Kumar, and Bhaskar Biswas. PQCLP: Parameterized quantum circuit based link prediction in dynamic networks. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000422.

Yuan:2024:HEW

[1983] Xiaoyun Yuan, Tengyao Li, Lingling Li, Ruixiang Li, Zhiquan Wang, and Xiangyang Luo. HSWF: Enhancing website fingerprinting attacks on Tor to address real world distribution mismatch. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000495.

Thomas:2024:DTA

[1984] Yannis Thomas, Stavros Toumpis, and Nikolaos Smyrnioudis. Digital twin approach to estimating and utilizing the capacity region of wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000458.

Kumari:2024:HCP

[1985] Deepa Kumari, Abhirath Singh Parmar, Harshvadhan Sunil Goyal, Kushal Mishra, and Subhrakanta Panda. HealthRec-Chain: Patient-centric blockchain enabled IPFS for privacy preserving scalable health data. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000550.

Wang:2024:EEM

[1986] Zhongyu Wang, Yashuai Cao, Zheng Chang, Tiejun Lv, and Wei Ni. Energy efficiency maximization in UAV communication networks with nonlinear energy harvesting. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000549.

Ali:2024:IZT

[1987] Belal Ali, Mark A. Gregory, Shuo Li, and Omar Amjad Dib. Implementing zero trust security with dual fuzzy methodology for trust-aware authentication and task offloading in Multiaccess Edge Computing. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400029X.

Yarinezhad:2024:NSA

[1988] Ramin Yarinezhad and Eylem Ekici. A novel scheduling algorithm for LTE on unlicensed bands to ensure fair coexistence with Wi-Fi. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000641.

Gulzari:2024:CAM

[1989] Usman Ali Gulzari, Zoran Salcic, Waqar Farooq, Sheraz Anjum, Sarzamin Khan, Muhammad Sajid, and Frank Sill Torres. Comparative analysis of 2D mesh topologies with additional communication links for onchip networks. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000252.

Liu:2024:ARR

[1990] Yantao Liu and Na Qin. Achievable rate regions for fractional routing. Computer Networks (Amsterdam, Netherlands: 1999), 241 (??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000537.

Lin:2024:PPB

[1991] Yuming Lin, Youjia Jiang, You Li, and Ya Zhou. Privacy-preserving batchbased task assignment over spatial crowdsourcing platforms. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000288.

Kim:2024:ESS

[1992] Jinwoo Kim, Minjae Seo, Seungsoo Lee, Jaehyun Nam, Vinod Yegneswaran, Phillip Porras, Guofei Gu, and Seungwon Shin. Enhancing security in SDN: Systematizing attacks and defenses from a penetration perspective. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000355.

Zormati:2024:RAR

[1993] Mohamed Ali Zormati, Hicham Lakhlef, and Sofiane Ouni. Review and analysis of recent advances in intelligent network softwarization for the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000471.

Lowes:2024:PPA

[1994] Gavin James Lowes and Jeffrey Neasham. PADAL — Passive Acoustic Detection and Localisation: Low energy underwater wireless vessel tracking network. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000483.

Teglasy:2024:AUA

[1995] Bálint Z. Téglásy, Emil Wengle, John R. Potter, and Sokratis Katsikas. Authentication of underwater assets. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000239.

Busacca:2024:SLO

[1996] Fabio Busacca, Stefano Mangione, Sergio Palazzo, Francesco Restuccia, and Ilenia Tinnirello. SDR-LoRa, an opensource, full-fledged implementation of LoRa on Software-Defined-Radios: Design and potential exploitation. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000264.

Rossanese:2024:DVS

[1997] Marco Rossanese, Placido Mursia, Andres Garcia-Saavedra, Vincenzo Sciancalepore, Arash Asadi, and Xavier Costa-Perez. Design and validation of scalable reconfigurable intelligent surfaces. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000409.

Kousias:2024:EPA

[1998] Konstantinos Kousias, Mohammad Rajiullah, Giuseppe Caso, Ozgu Alay, Anna Brunstrom, Usman Ali, Luca De Nardis, Marco Neri, and Maria-Gabriella Di Benedetto. Empirical performance analysis and ML-based modeling of 5G non-standalone networks. Computer Networks (Amsterdam, Netherlands: 1999), 241 (??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000392.

Wen:2024:RRB

[1999] Tian Wen, Hanqing Zhang, Han Zhang, Huixin Wu, Danxin Wang, Xiuwen Liu, Weishan Zhang, Yuwei Wang, and Shaohua Cao. RTIFed: a Reputation based Triple-step Incentive mechanism for energy-aware Federated learning over battery-constricted devices. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000240.

Liu:2024:ESR

[2000] Dongquan Liu, Zhengyan Zhou, Dong Zhang, Kaiwei Guo, Yanni Wu, and Chunming Wu. Efficient service reconfiguration with partial virtual network function migration. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000379.

Miano:2024:ANA

[2001] Sebastiano Miano, Giuseppe Lettieri, Gianni Antichi, and Gregorio Procissi. Accelerating network analytics with an on-NIC streaming engine. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400063X.

Ozaki:2024:MSM

[2002] Shintaro Ozaki, Takehiro Sato, and Eiji Oki. Multi-source multicast service chaining model guaranteeing reliability of network services. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000367.

Anonymous:2024:Aa

[2003] Anonymous. April 2024. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2024:EBe

[2004] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400149X.

Amend:2024:CEM

[2005] Markus Amend and Veselin Rakocevic. Cost-efficient multipath scheduling of video-on-demand traffic for the 5G ATSSS splitting function. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000501.

Zhao:2024:LAU

[2006] Xueyuan Zhao, Zhuoran Qi, and Dario Pompili. Link adaptation in Underwater Wireless Optical Communications based on deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000653.

Mohamed:2024:SFC

[2007] Ramy Mohamed, Marios Avgeris, Aris Leivadeas, Ioannis Lambadaris, John Chinneck, Todd Morris, and Petar Djukic. Service function chain network planning through offline, online and infeasibility restoration techniques. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000732.

Zhu:2024:LGJ

[2008] Yan Zhu and Dawei Lu. Latency guaranteed joint optimal traffic intelligent scheduling in large-scale satellite networks. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000689.

Livadariu:2024:TSC

[2009] Ioana Livadariu, Ahmed Elmokashfi, and Georgios Smaragdakis. Tracking submarine cables in the wild. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000665.

Huang:2024:RBT

[2010] Chung-Ming Huang and Chia-Han Hsieh. Registered-backoff-time (RBT) based channel access with grouping control for the trigger RAW mode of IEEE 802.11ah IoT network. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000410.

Vega:2024:MLC

[2011] Christian Vega, Elie F. Kfoury, Jose Gomez, Jorge E. Pezoa, Miguel Figueroa, and Jorge Crichigno. Machine learning controller for data rate management in science DMZ networks. Computer Networks (Amsterdam, Netherlands: 1999), 242 (??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000690.

Wei:2024:CCA

[2012] Zhenchun Wei, Yang Zhao, Zengwei Lyu, Xiaohui Yuan, Yu Zhang, and Lin Feng. Cooperative caching algorithm for mobile edge networks based on multi-agent meta reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000793.

Dai:2024:MHR

[2013] Jun Dai, Xinbin Li, Song Han, Zhixin Liu, Haihong Zhao, and Lei Yan. Multi-hop relay selection for underwater acoustic sensor networks: a dynamic combinatorial multi-armed bandit learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000744.

Feng:2024:MDC

[2014] Huifen Feng, Weiting Zhang, Ying Liu, Chuan Zhang, Chenhao Ying, Jian Jin, and Zhenzhen Jiao. Multi-domain collaborative two-level DDoS detection via hybrid deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000835.

Ahmed:2024:MAC

[2015] Imran Ahmed, Eiji Oki, and Bijoy Chand Chatterjee. Modeling and analysis of crosstalk-avoided and crosstalk-aware approaches in spectrally-spatially elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 242 (??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000884.

Sari:2024:UCB

[2016] T. Tolga Sari and Gökhan Seçinti. Using centrality based topology control for FANET survivability against jamming. Computer Networks (Amsterdam, Netherlands: 1999), 242 (??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000823.

Raj:2024:SME

[2017] Raina Raj. Stochastic modeling of enhancing energy efficiency for HAP in HAP-LEO systems using Markov regenerative process. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000859.

Baganal-Krishna:2024:FLA

[2018] Nehal Baganal-Krishna, Ralf Lübben, Eirini Liotou, Konstantinos V. Katsaros, and Amr Rizk. A federated learning approach to QoS forecasting in cellular vehicular communications: Approaches and empirical evidence. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000719.

Scott:2024:MPD

[2019] Ben A. Scott, Michael N. Johnstone, Patryk Szewczyk, and Steven Richardson. Matrix profile data mining for BGP anomaly detection. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000896.

Yang:2024:BBC

[2020] Huan Yang, Yajun Guo, and Yimin Guo. Blockchain-based cloud-fog collaborative smart home authentication scheme. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000720.

Luo:2024:UAF

[2021] Xijian Luo, Jun Xie, Liqin Xiong, Zhen Wang, and Yaqun Liu. UAVassisted fair communications for multipair users: a multi-agent deep reinforcement learning method. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001099.

K:2024:PAI

[2022] Aswini K. and Surendar M. Performance analyses of intelligent reflecting surface aided downlink multiuser rate-splitting multiple access system for 6G applications. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001038.

Zhen:2024:GSC

[2023] Zixian Zhen, Xiangfu Zhao, Jinkai Zhang, Yichen Wang, and Haiyue Chen. DA-GNN: a smart contract vulnerability detection method based on dual attention graph neural network. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000707.

Liu:2024:JUG

[2024] Huanlin Liu, Xilin Yuan, Yong Chen, Yang Hu, Huiling Zhou, and Wenjie Song. Joint user grouping and resource allocation scheme based on user satisfaction management for indoor multicolor VLC-NOMA network. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000446.

Cai:2024:SSA

[2025] Binlei Cai, Xiaoli Wang, Bin Wang, Meihong Yang, Ying Guo, and Qin Guo. A self-stabilizing and autoprovisioning orchestration for microservices in edge-cloud continuum. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001117.

Liu:2024:QAT

[2026] Jingxian Liu, Yitian Wang, Duotao Pan, and Decheng Yuan. QoS-aware task offloading and resource allocation optimization in vehicular edge computing networks via MADDPG. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001142.

Soler:2024:PPK

[2027] David Soler, Carlos Dafonte, Manuel Fernández-Veiga, Ana Fernández Vilas, and Francisco J. Nóvoa. A privacy-preserving key transmission protocol to distribute QRNG keys using zk-SNARKs. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000914.

Yang:2024:CBP

[2028] Zhimei Yang, Changgen Peng, Chongyi Zhong, and Yangyang Long. Consortium blockchain private key protection scheme based on rational secret sharing and blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000926.

Mayahi:2024:IEI

[2029] Meysam Mayahi, Valeria Loscrí, and Anna Maria Vegni. INVISIBLE 2.0: an enhanced interference-based handover technique for visible light communications in vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001129.

Du:2024:BPS

[2030] Zhiqiang Du, Liangxin Liu, Muhong Huang, Yanfang Fu, and Wendong Zhang. Bulwark: a proofof-stake protocol with strong consistency and liveness. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400077X.

Chen:2024:SMW

[2031] Weilin Chen, Wei Yang, and Wei Gong. A survey of millimeter wave backscatter communication systems. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000677.

Zhang:2024:AIE

[2032] Shunliang Zhang, Dali Zhu, and Yinlong Liu. Artificial intelligence empowered physical layer security for 6G: State-of-the-art, challenges, and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000872.

Robin:2024:OSP

[2033] Debobroto Das Robin and Javed I.

Khan. An open-source P4_{16} compiler backend for reconfigurable matchaction table switches: Making networking innovation accessible. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000781.

Miao:2024:SEC

[2034] Junfeng Miao, Zhaoshun Wang, Mei Wang, Sahil Garg, M. Shamim Hossain, and Joel J. P. C. Rodrigues. Secure and efficient communication approaches for Industry 5.0 in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000768.

Alvarez-Horcajo:2024:MDP

[2035] Joaquin Alvarez-Horcajo, Isaias Martinez-Yelmo, Elisa Rojas, Juan A. Carral, and Victoria Noci-Luna. MuHoW: Distributed protocol for resource sharing in collaborative edge-computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 242 (??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000756.

Wang:2024:CDS

[2036] Yuantao Wang, Zhaogang Shu, Shuwu Chen, Jiaxiang Lin, and Zhenchang Zhang. A cost and demand sensitive adjustment algorithm for service function chain in data center network. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000860.

Yu:2024:HLB

[2037] Hao Yu and Qiang Zhang. Hybrid learning based service migration for cost minimization with deadlines in multi-user mobile edge computing systems. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000811.

Polverini:2024:RBN

[2038] Marco Polverini, Stefania Sardellitti, Sergio Barbarossa, Antonio Cianfrani, Paolo Di Lorenzo, and Marco Listanti. Reducing the In band Network Telemetry overhead through the spatial sampling: Theory and experimental results. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001014.

Shakeri:2024:REC

[2039] Ali Shakeri, Tianliang Zhang, Shunmugapriya Ramanathan, Miguel Razo, Marco Tacca, and Andrea Fumagalli. Reliable edge-to-core optical networks: an optimal algorithm for maximal path diversity. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001002.

Pelekanakis:2024:CEC

[2040] Konstantinos Pelekanakis, Pietro Paglierani, Alberto Alvarez, and João Alves. Comparison of error correction codes via optimal channel replay of high north underwater acoustic channels. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001026.

Rzepka:2024:PER

[2041] Karol Rzepka, Przemysław Szary, Krzysztof Cabaj, and Wojciech Mazurczyk. Performance evaluation of Raspberry Pi 4 and STM32 Nucleo boards for security-related operations in IoT environments. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000847.

Alaez:2024:DTM

[2042] D. Aláez, P. Lopez-Iturri, M. Celaya-Echarri, L. Azpilicueta, F. Falcone, J. Villadangos, and J. J. Astrain. Digital twin modeling of open category UAV radio communications: a case study. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001087.

Anonymous:2024:Ab

[2043] Anonymous. April 2024. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2024:EBf

[2044] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001981.

Guerra:2024:FLR

[2045] Renata Rojas Guerra, Anna Vizziello, Pietro Savazzi, Emanuele Goldoni, and Paolo Gamba. Forecasting LoRaWAN RSSI using weather parameters: a comparative study of ARIMA, artificial intelligence and hybrid approaches. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000902.

Malandrino:2024:EIR

[2046] Francesco Malandrino, Alessandro Nordio, and Carla Fabiana Chiasserini. Eavesdropping with intelligent reflective surfaces: Nearoptimal configuration cycling. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001166.

Liao:2024:CFB

[2047] Zhuofan Liao, Wenqiang Deng, Shiming He, and Qiang Tang. Collaborative Filtering-based Fast Delay-aware algorithm for joint VNF deployment and migration in edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001324.

Li:2024:LBC

[2048] Zhiyuan Li and Xiaoping Xu. L2-BiTCN-CNN: Spatio-temporal features fusion-based multi-classification model for various internet applications identification. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001300.

Li:2024:TTL

[2049] Xia Li and Yuping Wang. TABDeep: a two-level action branch architecturebased deep reinforcement learning for distributed sub-tree scheduling of online multicast sessions in EON. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001208.

ElGhafour:2024:SCV

[2050] Mohamed G. Abd El Ghafour, Ahmed H. Abd El-Malek, Ola E. Hassan, and Mohammed Abo-Zahhad. Securing cooperative vehicular networks amid obstructing vehicles and mixed fading channels. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001233.

Zhao:2024:DIF

[2051] Zhe Zhao, Zhiliang Qiu, Weitao Pan, Hui Li, Ling Zheng, and Ya Gao. Design and implementation of a frame preemption model without guard bands for time-sensitive networking. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001178.

Valko:2024:RCE

[2052] Danila Valko and Daniel Kudenko.

Reducing CO_{2} emissions in a peerto-peer distributed payment network: Does geography matter in the lightning network? Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001294.

Buccafurri:2024:HDT

[2053] Francesco Buccafurri, Vincenzo De Angelis, Maria Francesca Idone, and Cecilia Labrini. A hierarchical distributed trusted location service achieving location k-anonymity against the global observer. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001336.

Gu:2024:LGH

[2054] Zheyuan Gu, Gaopeng Gou, Chang Liu, Chen Yang, Xiyuan Zhang, Zhen Li, and Gang Xiong. Let gambling hide nowhere: Detecting illegal mobile gambling apps via heterogeneous graphbased encrypted traffic analysis. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001105.

Kilic:2024:THP

[2055] Ahmet Kilic. TLS-handshake for Plug and Charge in vehicular communications. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001130.

Chen:2024:RSC

[2056] Shiyou Chen, Lanlan Rui, Zhipeng Gao, Yang Yang, Xuesong Qiu, and Shaoyong Guo. Resource sharing for collaborative edge learning: a privacyaware incentive mechanism combined with demand prediction. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001348.

Haj-Hassan:2024:DMP

[2057] Ali Haj-Hassan, Youcef Imine, Antoine Gallais, and Bruno Quoitin. Detecting malicious proxy nodes during IoT network joining phase. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001403.

Song:2024:EEF

[2058] Yu Song, Shilong Zhang, and Shubin Wang. An energy efficient fusing data gathering protocol in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001373.

He:2024:BBC

[2059] Yunhua He, Shuang Wang, Zhihao Zhou, Ke Xiao, Anke Xie, and Bin Wu. A blockchain-based carbon emission security accounting scheme. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001361.

Jin:2024:HIH

[2060] Yanliang Jin, Jiahao Zhou, and Yuan Gao. HSGAN-IoT: a hierarchical semi-supervised generative adversarial networks for IoT device classification. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001312.

Chintapalli:2024:EED

[2061] Venkatarami Reddy Chintapalli, Rajat Partani, Bheemarjuna Reddy Tamma, and Siva Ram Murthy C. Energy efficient and delay aware deployment of parallelized service function chains in NFV-based networks. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400121X.

Batham:2024:ICF

[2062] Deepak Batham and Vandana Vikas Thakare. An improved cost functionbased class of service provisioning scheme for elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 243 (??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001154.

Jovanovic:2024:RIB

[2063] Zorka Jovanovic, Zhe Hou, Kamanashis Biswas, and Vallipuram Muthukkumarasamy. Robust integration of blockchain and explainable federated learning for automated credit scoring. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400135X.

Zhou:2024:TRP

[2064] Hongliang Zhou, Yifeng Zheng, and Xiaohua Jia. Towards robust and privacy-preserving federated learning in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001531.

Bi:2024:CSB

[2065] Meihua Bi, Chenghao Yan, Rui Lin, Yuanyuan Hong, and Jun Li. Client scheduling and bandwidth slicing for multiple federated learning tasks over multiple passive optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001415.

Liu:2024:ECO

[2066] Liang Liu, Wuping Mao, Wenwei Li, Jie Duan, Guanyu Liu, and Bingchuan Guo. Edge computing offloading strategy for space-air-ground integrated network based on game theory. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001634.

Larsen:2024:EMN

[2067] Line M. P. Larsen, Henrik L. Christiansen, Sarah Ruepp, and Michael S. Berger. The evolution of mobile network operations: a comprehensive analysis of open RAN adoption. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001245.

ElRajab:2024:ZTN

[2068] Mirna El Rajab, Li Yang, and Abdallah Shami. Zero-touch networks: Towards next-generation network automation. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001269.

Ruiz:2024:STU

[2069] Olga Ruiz, Juan Sánchez-González, Jordi Pérez-Romero, Oriol Sallent, and Irene Vilà. Space and time user distribution measurements dataset in a university campus. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001610.

Ardizzon:2024:RBA

[2070] Francesco Ardizzon, Paolo Casari, and Stefano Tomasin. A RNN-based approach to physical layer authentication in underwater acoustic networks with mobile devices. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001439.

Maier:2024:WSN

[2071] Guido Maier, Antonino Albanese, Michele Ciavotta, Nicola Ciulli, Stefano Giordano, Elisa Giusti, Alfredo Salvatore, and Giovanni Schembra. WatchEDGE: Smart networking for distributed AI-based environmental control. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400080X.

Gallego-Madrid:2024:MLP

[2072] Jorge Gallego-Madrid, Irene Bru-Santa, Alvaro Ruiz-Rodenas, Ramon Sanchez-Iborra, and Antonio Skarmeta. Machine learning-powered traffic processing in commodity hardware with eBPF. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001270.

Vila:2024:REB

[2073] I. Vilà, O. Sallent, and J. Pérez-Romero. Relay-empowered beyond 5G radio access networks with edge computing capabilities. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001191.

Botta:2024:AOS

[2074] Alessio Botta, Roberto Canonico, Annalisa Navarro, Giovanni Stanco, and Giorgio Ventre. Adaptive overlay selection at the SD-WAN edges: a reinforcement learning approach with networked agents. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001427.

Wu:2024:ABD

[2075] Yunyi Wu, Kai Zhang, and Yongbing Zhang. An accelerated Benders decomposition approach for virtual base station formation in stochastic Cloud-RANs. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001609.

Gudepu:2024:DHF

[2076] Venkateswarlu Gudepu, Venkatarami Reddy Chintapalli, Piero Castoldi, Luca Valcarenghi, Bheemarjuna Reddy Tamma, and Koteswararao Kondepu. The drift handling framework for open radio access networks: an experimental evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001221.

DiMauro:2024:HLS

[2077] Mario Di Mauro, Giovanni Galatro, Fabio Postiglione, Wei Song, and Antonio Liotta. Hybrid learning strategies for multivariate time series forecasting of network quality metrics. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400118X.

Peng:2024:NTP

[2078] Yufei Peng, Yingya Guo, Run Hao, and Chengzhe Xu. Network traffic prediction with Attention-based Spatial-Temporal Graph Network. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001282.

Anonymous:2024:Mb

[2079] Anonymous. May 2024. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2024:EBg

[2080] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002433.

Marin:2024:QML

[2081] Andrea Marin, Marco Ajmone Marsan, Michela Meo, and Matteo Sereno. Queuing models of links carrying streaming and elastic services. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001385.

Liang:2024:PPM

[2082] Junjie Liang, Lei Zhang, Can Bu, Guangyu Yang, Hao Wu, and Aiguo Song. Plug-and-play multi-dimensional attention module for accurate Human Activity Recognition. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001701.

Zhu:2024:ECC

[2083] Sifeng Zhu, Xiaohua Tian, Hao Chen, Hai Zhu, and Rui Qiao. Edge collaborative caching solution based on improved NSGA II algorithm in Internet of Vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001397.

Li:2024:ORA

[2084] Yue Li, Yanjun Li, Yuzhe Chen, Jiahui Tong, Xianzhong Tian, and Kaikai Chi. Online resolution adaptation and resource allocation for edgeassisted video analytics. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001749.

Faticanti:2024:OOC

[2085] Francescomaria Faticanti and Giovanni Neglia. Optimistic online caching for batched requests. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001737.

Mortagua:2024:EAI

[2086] Duarte Mortágua, André Zúquete, and Paulo Salvador. Enhancing 802.1X authentication with identity providers using EAP-OAUTH and OAuth 2.0. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001695.

Zhang:2024:TNT

[2087] Zixu Zhang, Guangsheng Yu, Caijun Sun, Xu Wang, Ying Wang, Ming Zhang, Wei Ni, Ren Ping Liu, Andrew Reeves, and Nektarios Georgalas. TbDd: a new trustbased, DRL-driven framework for blockchain sharding in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001750.

Baidas:2024:NSR

[2088] Mohammed W. Baidas, Ahmed M. AbdelGhaffar, and Emad Alsusa. Network sum-rate maximization for network-coded clustered uplink NOMA networks with SWIPT-enabled relays. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128624001725.

Bilir:2024:HCH

[2089] Emin Bilir, Murtaza Cicioglu, and Ali Çalhan. Hybrid cell handover strategy for O-RAN-based campus networks. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001798.

Neto:2024:FFL

[2090] Helio N. Cunha Neto, Jernej Hribar, Ivana Dusparic, Natalia C. Fernandes, and Diogo M. F. Mattos. FedSBS: Federated-learning participant-selection method for Intrusion Detection Systems. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400183X.

Toorchi:2024:DRL

[2091] Niloofar Toorchi, Weiqiang Lyu, Linsheng He, Jiamiao Zhao, Iftikhar Rasheed, and Fei Hu. Deep reinforcement learning enhanced skeleton based pipe routing for highthroughput transmission in flying adhoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001622.

Hosseini:2024:EEM

[2092] Soheil Hosseini, Ignacio de Miguel, Noemí Merayo, Ramón de la Rosa, Rubén M. Lorenzo, and Ramón J. Durán Barroso. Energy efficient multipath routing in space division multiplexed elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001816.

Kumar:2024:BEA

[2093] Neeraj Kumar and Rifaqat Ali.

Blockchain-enabled authentication framework for Maritime Transportation System empowered by 6G-IoT. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001853.

Du:2024:OTT

[2094] An Du, Jie Jia, Jian Chen, Liang Guo, and Xingwei Wang. Online twotimescale service placement for timesensitive applications in MEC-assisted network: a TMAGRL approach. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001713.

Mlika:2024:BST

[2095] Fatma Mlika, Wafa Karoui, and Lotfi Ben Romdhane. Blockchain solutions for trustworthy decentralization in social networks. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001683.

Hakiri:2024:CSD

[2096] Akram Hakiri, Aniruddha Gokhale, Sadok Ben Yahia, and Nedra Mellouli. A comprehensive survey on digital twin for future networks and emerging Internet of Things industry. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001828.

Bingol:2024:WQD

[2097] Gülnaziye Bingöl, Simone Porcu, Alessandro Floris, and Luigi Atzori. WebRTC-QoE: a dataset of QoE assessment of subjective scores, network impairments, and facial and speech features. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001889.

deAlmeida:2024:DDE

[2098] Leandro C. de Almeida, Washington Rodrigo Dias da Silva, Thiago C. Tavares, Rafael Pasquini, Chrysa Papagianni, and Fábio L. Verdi. DESiRED — Dynamic, Enhanced, and Smart iRED: a P4-AQM with deep reinforcement learning and in-band network telemetry. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001580.

Bellin:2024:MBA

[2099] Arturo Bellin, Fabrizio Granelli, and Daniele Munaretto. A measurementbased approach to analyze the power consumption of the softwarized 5G core. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001440.

Wengle:2024:EAJ

[2100] Emil Wengle, Elias Strandell Erstorp, Viktor Lidström, Damiano Varagnolo, and Hefeng Dong. Experimental assessment of a JANUSbased consensus protocol. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001774.

delRio:2024:MGS

[2101] Alberto del Rio, Javier Serrano, David Jimenez, Luis M. Contreras, and Federico Alvarez. Multisite gaming streaming optimization over virtualized 5G environment using deep reinforcement learning techniques. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400166X.

Zhao:2024:MSB

[2102] Baosen Zhao, Wanghong Yang, Wenji Du, Yongmao Ren, Jianan Sun, Qinghua Wu, and Xu Zhou. A multipath scheduler based on crosslayer information for low-delay applications in 5G edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001658.

Shirsath:2024:STS

[2103] Vaishali A. Shirsath, Madhav M. Chandane, Chhagan Lal, and Mauro Conti. SYNTROPY: TCP SYN DDoS attack detection for Software Defined Network based on Rényi entropy. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001592.

Iqbal:2024:IQO

[2104] Muhammad Shahid Iqbal and Chien Chen. Instant queue occupancy used for automatic traffic scheduling in data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001786.

Feraudo:2024:DDB

[2105] Angelo Feraudo, Nicolò Romandini, Carlo Mazzocca, Rebecca Montanari, and Paolo Bellavista. DIVA: a DID-based reputation system for secure transmission in VANETs using IOTA. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001646.

Anonymous:2024:Mc

[2106] Anonymous. May 2024. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2024:EBh

[2107] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002718.

Hu:2024:RSA

[2108] Jianwei Hu, Kaiqi Huang, Genqing Bian, and Yanpeng Cui. Redact4Trace: a solution for auditing the data and tracing the users in the redactable blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001920.

Liu:2024:MTE

[2109] Libin Liu and Xiuting Xu. Marvel: Towards efficient federated learning on IoT devices. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400207X.

Deng:2024:TOO

[2110] Tao Deng, Zhanwei Yu, and Di Yuan. Task offloading optimization in mobile edge computing under uncertain processing cycles and intermittent communications. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001919.

Zhang:2024:RDT

[2111] Xuhui Zhang, Qing Li, Feixue Han, and Yong Jiang. A receiver-driven transport protocol using differentiated algorithms for differential congestion in datacenters. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001890.

Peng:2024:LLI

[2112] Min Peng, Xianxin Fu, Haiyang Zhao, Yu Wang, and Caihong Kai. LiKey: Location-independent keystroke recognition on numeric keypads using WiFi signal. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001865.

Stergiopoulos:2024:SPP

[2113] George Stergiopoulos, Panagiotis Kotzanikolaou, Konstantinos Adamos, and Lilian Mitrou. Scaling private proximity testing protocols for geofenced information exchange: a metropolitan-wide case study. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002135.

Ge:2024:HIP

[2114] Fei Ge and Liansheng Tan. HITLinQ:

Improving path capacity through hybrid information-theoretic link scheduling in multi-hop wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624002147.

Han:2024:GDE

[2115] Xinbo Han, Guizhong Xu, Meng Zhang, Zheng Yang, Ziyang Yu, Weiqing Huang, and Chen Meng. DE-GNN: Dual embedding with graph neural network for fine-grained encrypted traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002044.

Garima:2024:DLB

[2116] Garima, Vivekanand Jha, and Rakesh Kumar Singh. A deep learning based dynamic bandwidth allocation method for XG-PON based mobile fronthaul for CRAN. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001762.

Pei:2024:PRE

[2117] Hongmei Pei, Peng Yang, Weihao Li, Miao Du, and Zhongjian Hu. Proxy re-encryption for secure data sharing with blockchain in Internet of Medical Things. Computer Networks (Amsterdam, Netherlands:

1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002056.

Zhang:2024:ABA

[2118] Jingci Zhang, Jun Zheng, Zheng Zhang, Tian Chen, Yu an Tan, Quanxin Zhang, and Yuanzhang Li. ATT&CK-based Advanced Persistent Threat attacks risk propagation assessment model for zero trust networks. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002081.

Liu:2024:AMA

[2119] Xiaoying Liu, Xiaodong Li, Kechen Zheng, and Jia Liu. AoI minimization of ambient backscatterassisted EH-CRN with cooperative spectrum sensing. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002214.

Li:2024:MNA

[2120] Beibei Li and Wei Hu. A multinode attack scheme based on community partitioning in large scale infrastructure networks. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002184.

Cerasuolo:2024:MNA

[2121] Francesco Cerasuolo, Alfredo Nascita, Giampaolo Bovenzi, Giuseppe Aceto, Domenico Ciuonzo, Antonio Pescapè, and Dario Rossi. MEMENTO: a novel approach for class incremental learning of encrypted traffic. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002068.

Kim:2024:HAO

[2122] Sungwook Kim. Hierarchical aerial offload computing algorithm based on the Stackelberg-evolutionary game model. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128624001804.

Guo:2024:CEL

[2123] Feng Guo, Shidong Sun, Junjie Hu, Ning Zhang, and Zhiqiang Lv. CIPO: Efficient, lightweight and programmable packet scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001877.

Han:2024:CAU

[2124] Xiaoli Han and Bo Hu. A congestionaware user-cell association scheme to assist with traffic offloading in a threetier heterogeneous network. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002202.

Wang:2024:SIB

[2125] Kaiyu Wang, Guangsheng Feng, Zhenzhou Ji, Zhiying Tu, and Shufan He. SP-PoR: Improve blockchain performance by semi-parallel processing transactions. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002263.

Wang:2024:RLB

[2126] Yi Wang, Yu Deng, Ling Kang, Fang Jiang, and Fulin Jiang. Reinforcement learning-based energy efficiency optimization for RIS-assisted UAV hybrid uplink and downlink system. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002226.

Lavernelle:2024:ASI

[2127] Jean de Bonfils Lavernelle, Pierre-François Bonnefoi, Benoît Gonzalvo, and Damien Sauveron. Assessment of spatial isolation in Jailhouse: Towards a generic approach. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002342.

Li:2024:QQR

[2128] Junhao Li, Qiang Nong, and Ziyu Liu. QBDD: Quantum-resistant blockchain-assisted deep data deduplication protocol for vehicular crowdsensing system. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002251.

Xie:2024:OTO

[2129] Mande Xie, Xiangquan Su, Hao Sun, and Guoping Zhang. Online task offloading algorithm based on multiobjective optimization caching strategy. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128624002329.

Rashid:2024:PUD

[2130] Fariza Rashid, Ben Doyle, Soyeon Caren Han, and Suranga Seneviratne. Phishing URL detection generalisation using Unsupervised Domain Adaptation. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128624002305.

Wang:2024:CII

[2131] Xiangbin Wang, Qingjun Yuan, Yongjuan Wang, Gaopeng Gou, Chunxiang Gu, Gang Yu, and Gang Xiong. Combine intra- and interflow: a multimodal encrypted traffic classification model driven by diverse features. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002354.

Zhang:2024:RPP

[2132] Fan Zhang, Hui Huang, Zhixiong Chen, and Zhenjie Huang. Robust and privacy-preserving federated learning with distributed additive encryption against poisoning attacks. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002159.

Feng:2024:GGL

[2133] Pengbin Feng, Dawei Wei, Qiaoyang Li, Qin Wang, Youbing Hu, Ning Xi, and Jianfeng Ma. GlareShell: Graph learning-based PHP webshell detection for web server of industrial internet. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400238X.

Dao:2024:RNT

[2134] Nhu-Ngoc Dao, Ngo Hoang Tu, Trong-Dai Hoang, Tri-Hai Nguyen, Luong Vuong Nguyen, Kyungchun Lee, Laihyuk Park, Woongsoo Na, and Sungrae Cho. A review on new technologies in 3GPP standards for 5G access and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002020.

Kazmi:2024:SFL

[2135] Syed Hussain Ali Kazmi, Faizan Qamar, Rosilah Hassan, Kashif Nisar, and Mohammed Azmi Al-Betar. Security of federated learning in 6G era: a review on conceptual techniques and software platforms used for research and analysis. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001907.

Wilhelmi:2024:IDB

[2136] Francesc Wilhelmi, Nima Afraz, Elia Guerra, and Paolo Dini. The implications of decentralization in blockchained federated learning: Evaluating the impact of model staleness and inconsistencies. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001932.

Guo:2024:MCO

[2137] Yinzhi Guo, Xiaolong Xu, and Fu Xiao.

MADRLOM: a computation offloading mechanism for software-defined cloudedge computing power network. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001841.

Centofanti:2024:IPC

[2138] Carlo Centofanti, José Santos, Venkateswarlu Gudepu, and Koteswararao Kondepu. Impact of power consumption in containerized clouds: a comprehensive analysis of open-source power measurement tools. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002032.

Kvasic:2024:DRC

[2139] Igor Kvasić, Derek Orbaugh Antillon, ula Na, Christopher Walker, Iain Anderson, and Nikola Mišković. Diver-robot communication dataset for underwater hand gesture recognition. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912862400224X.

Zorello:2024:BBO

[2140] Ligia Maria Moreira Zorello, Laurens Bliek, Sebastian Troia, Guido Maier, and Sicco Verwer. Black-box optimization for anticipated baseband-function placement in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002160.